A Complete Bibliography of ACM Transactions on Sensor Networks

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

22 May 2024 Version 1.51

Title word cross-reference

2 [BY19, CWY^{+}15, TJZ^{+}13]. 3 [Amm16, BY19, NXW^{+}22, SNK^{+}22, TJZ^{+}13, TGG^{+}19, WWL^{+}16, WJD16, XYW^{+}22, YRB^{+}17]. ^{2} [XWC^{+}23]. _{2} [AAHS18]. α [ZH05]. k [Amm13, Amm16, Amm23, SCWC13]. μ [RHS20]. × [GDM22].

-Coverage [Amm23, Amm13, SCWC13]. -Covered [Amm16]. -D [BY19]. -lifetime [ZH05]. -Mote [CWY^{+}15].

1 [SMS22]. 19 [AAJ^{+}23, CC23].

2 [BNN^{+}20, XDX^{+}14].

3-Conversion [ZZG^{+}24].

5 [BDP24]. 5.0 [YYC^{+}19]. 5G [CWK^{+}22, DTY^{+}22, GXQ^{+}22, MLS^{+}22, SE23, SJP^{+}22, YQLD22]. 5G/B5G [MLS^{+}22].

6D [XWL24].

802.15.4 [PEFSV13, PFJ13]. 802.15.4e [TDD^{+}19]. 802.15.4m [BAP^{+}17].

A-MAC [DDHC^{+}12]. A-MCI [GZK^{+}23]. Abnormality [GZK^{+}23]. Abstraction [JJ15, RKJ09]. AC [MKFD^{+}23]. AC-DC [MKFD^{+}23]. Accelerating [CS17]. Accelerations [ZHL^{+}15]. Access [GLG^{+}23, LGLD23, SBSD18, ZHJ^{+}20, PFJ13, RDR07].

Accuracy [LHX^{+}21, BHA^{+}13]. Accurate [AHK16, COS19, CLX^{+}21, PKC^{+}18, VTY18, ZLW^{+}15, ZW24]. ACES [FBAG20]. Achieving [VHC^{+}09, WC13, ZGHZ12]. Acoustic [CK09, GYNY16, LYL^{+}24, LWH^{+}22, SDW^{+}23, WSC^{+}23, WLX^{+}23, ZW24, GAJ^{+}06, KVI^{+}13, SHY13]. Acoustic-based [SDW^{+}23]. Acoustical [CSLJ23, MKK^{+}13]. acquisition [AAA06]. Across [HPS^{+}18, SPK^{+}10]. Activation [MNLZ18, BCL^{+}12, HR13, JKK08]. Active [ALS23, MGS^{+}15, IW14]. Activities [KSR^{+}20, ZZY^{+}23]. Activity [DBC^{+}24, LPW^{+}23, OXZ^{+}23, Pha16, WL23, WHYC19, YYSL08, dLM14]. Actor [WHST16]. Acts [HL17]. aCtuation [WWB^{+}19]. Actuator [CS23, GRE^{+}07, PCR13, ZVPS10]. Ad [CS17, CS18, JYC^{+}24, VDV16, CVY09, DRC06, KPK12, LYG^{+}13, NJS05, PR10, SS13]. ad-hoc [CVY09, SS13]. AdaMEC [PLW^{+}24]. Adaptation [HL17, ZWG24, BCL^{+}12, CUdVY13, EMBP12, SPK14, XTZ08]. Adapting [JJ15]. Adaptive [AKSM15, BCMY22, CRZ^{+}20, HF17, HKG^{+}19, KCE^{+}20, LDZ13, LMZ^{+}16, LC14b, LHX16, SGM08, SCWC13, YTR^{+}22, ZCLJ14, ZTZX23, KLJ12, KRJ09, LPW^{+}23, PDMJ10, PLW^{+}24, QM13, YH13]. Addressing [ZWL^{+}24]. Adjustable [FLS^{+}14, MZW^{+}19]. Advanced [AH14, MDB^{+}23, ZYZ^{+}19]. Advances [SYL^{+}22]. Adversarial [LXYT24, SYT22, XLG^{+}22, Yan22, ZWG24]. Aerial [HWS^{+}20]. After [HBW^{+}18]. After-sales [HBW^{+}18]. Against [MY24, TDD^{+}19, CKHP19, HMG^{+}24, LPV^{+}09, LLH22, LWCJ14, NLD08, SYT22, WWZ^{+}21, WC09, WC12, XBWX13, ZSJN07]. Age [SCLG24, YMY^{+}23]. Age-of-Information [YMY^{+}23]. agent [JR08]. Agents [SHWW20]. AGgregation [YS07, ARWK19, BYD^{+}15, CCC^{+}21, CDR08, HMLJ17, HLN^{+}11, LCC^{+}17, LYY24, PNL^{+}22, SCL^{+}14, WWZ^{+}21, XAKV15, ZSZ20, CCMT09, CC11, CNMH08, ELR08, Kal10, KLJ12, MS09, NGSA08, ZJX10]. Agile [WYC^{+}24]. agnostic [LGLD23]. Agreement [MLX^{+}24, YLSZ19]. Ahead [RS19]. Ahead-of-time [RS19]. ahoi [RHS20]. AI [CWK^{+}22, GXQ^{+}22, LLH22, RRA22]. AI-Based [LLH22, RRA22]. AI-enabled [CWK^{+}22]. Aided [QWC^{+}22, WLW^{+}23, XQL^{+}24]. Aids [YYL^{+}23]. AIoMT [ALS23]. AIoT [LYZ^{+}24]. Air [ALNT22, CML^{+}21, LWL^{+}21, PKS^{+}23, XXW^{+}24, YXG^{+}19, SNY^{+}24]. AirContour [YXG^{+}19]. Akte [SDW^{+}23]. Akte-Liquid [SDW^{+}23]. Algorithm [AH20, CS17, GSM^{+}22, LWX^{+}21, PNL^{+}22, TBS^{+}24, XWC^{+}23, ZWWL23, CNMH08, CVY09, FKMS06, KLC13]. algorithmic [Su07]. Algorithms [GSGA23, TJLK14, WJD16, BLWY06, CKL^{+}09, Dji10, MAG13, NEKK12, ZSG09]. Alignment [WZZ^{+}23, WCLD23]. Alive [BR15]. Allocation [HCL15, MSAJ18, TZZ22, WCW^{+}23, YM14, ZWWL23, ZGX^{+}16, SC12]. Ambient [ZZH^{+}23]. AMOC [ZHJ^{+}20]. Amongst [MSAJ18]. Analyses [ZZW^{+}23a]. Analysis [BBD^{+}23, BAP^{+}17, BQB^{+}11, CPL^{+}20, CML^{+}21, DIE14, FC18, GKRW17, GZJE23, JYC^{+}24, LCC10, MYH^{+}24, MB16, PS17, RDR07, XYW^{+}22, ZJZ12, CKL^{+}09, JTS09, JKS^{+}10, PFJ13, WKA14, ZK07, ZBA07]. Analytic [LPR09]. Analytics [BIMD19, FPA^{+}20, LZGX23, NJL24, RKLM23]. Analyze [MSK^{+}23]. Analyzing [LM10a, LM10b]. Anchor [CWY24, TJZ^{+}13]. anchor-free [TJZ^{+}13]. Android [GLL^{+}24, ZLB^{+}23]. Angle [GZJE23, BGJ09]. Anisotropic [ZLW^{+}15, LH09]. Annealing [YTZ^{+}23]. anomalies [RBLP09]. Anomaly [DD11, HWT^{+}22, IPMGL18, LYF^{+}23, PC10, dLM14]. anonymity [YSZC13]. Anonymous [MLX^{+}24, SLS^{+}22, YBY^{+}24]. Antenna [HXZ23a]. Antennas [YTB^{+}14, ZJZ12]. Anyone [LXYT24]. Anytime [LXYT24]. AoA [PAYL22]. App [GLL^{+}24, YZZD23]. Appliance [NZM21]. Application [JAC19, KKRR15, MDB^{+}23, YBY^{+}24, YCL^{+}19, LHRM09, WZL08, IBS^{+}10]. Application-specific [IBS^{+}10]. Applications [BASM16, DLG^{+}21, LLX^{+}22, RFB^{+}14, TJLK14, WJZ21, XZL^{+}20, ZHL^{+}15, ACG^{+}13, CHN^{+}13, CCJ08, LM10a, LM10b, LS10, SPK^{+}10, ZSG09]. Applied [BDP24]. Applying [GZK^{+}23, YPW^{+}13]. Apportionment [WCV^{+}18, WJ21]. Approach [Amm23, CLX^{+}21, FSTH24, GHZ^{+}22, KPRH14, LDGG21, LTZ^{+}24, MCLM20, MG24, PNL^{+}22, SBCF20, SZ19, SCD^{+}24, SGB15, TCN^{+}17, WYY^{+}19, WLLZ24, YHC^{+}24, ABM13, EGG13, HM07b, IR12, KBD14, LS10, NJS05, Su07, VAC13, WWLX13, XRH^{+}13, ZLGG10]. approaches [EFI^{+}10]. Approximate [CG18, LCC^{+}17]. approximately [Kal10]. Approximation [Dji10]. ApproxNet [XKW^{+}22]. APs [YYXL22]. AQMon [XXW^{+}24]. AQuaMoHo [PKS^{+}23]. Aquatic [WTX^{+}16]. Arbitrary [ÁKSW22]. Architecture [HBW^{+}18, LLDZ23, PGG^{+}10]. Area [DSH16, DGS16, Hau14, LFNS14, LWKZ22, MSAJ18, RHD17, SBK22, WQH^{+}22, XDL^{+}24, ZZX^{+}20, CJS11, HM07b, HR13, KNSM14, LYG^{+}13, LCD22, YSM08]. Arms [LJLW19]. Arrival [GZJE23]. Arrivals [JZL^{+}19]. Artificial [LCF^{+}22, MGN22, QXZZ22]. AS-MAC [QM13]. as-rigid-as-possible [ZLGG10]. As-You-Go [GCAK17]. Assessment [BAP^{+}17, KR18]. Assignment [GSM^{+}22, MKM^{+}20, LWH^{+}06, RJL^{+}10, TP07]. Assignments [HBKP14]. Assisted [DGS16, TZZ22, XJL^{+}23, DPB19, LLZ^{+}22, LYY24, SDYC22, SCLG24, SNY^{+}24, WLZ13]. association [WL14]. Assurance [WRYL11]. AsTAR [YTR^{+}22]. Asymmetric [KLC^{+}16]. asymmetry [SAZ10, ZK07]. Asymptotic [VMS10]. Asynchronous [CZMM23, ELR08, HY07, LLL14, FSTH24, WLD10]. ATPC [LMZ^{+}16]. ATPP [YZZD23]. Attack [CD21, GJT^{+}22, LTZ^{+}24, TDD^{+}19, WWZ^{+}21, Yan22]. Attack-aware [GJT^{+}22]. Attacks [CKHP19, CPL^{+}20, HAH22, HLZ^{+}24, HMG^{+}24, LLH22, LWCJ14, MY24, MB16, SBCF20, SE23, TDZ^{+}22, CKL^{+}09, LPV^{+}09, NZR10, NLD08, PX13, XWDN12, ZSJN07]. Attention [ALS23]. Attestation [KBD13]. audio [LCH^{+}09]. Auditing [TCN^{+}17]. Augment [ZWG24]. Augmentation [LLDZ23, LYST23]. Augmented [CYD^{+}24, LLZ^{+}22, ZYL^{+}24, SPK14]. Authenticated [MLX^{+}24, YLSZ19]. Authentication [CLJ^{+}23, LHZZ20, LTDZ22, LLDZ23, LWJ^{+}23, NLD08, WDLN09, XWDN12, ZSJN07]. authenticity [ADF12]. Authority [AKC^{+}18]. Auto [KRP15, LLDZ23, LYF^{+}23, RKRP17]. Auto-Encoder [LYF^{+}23]. AutoCalib [BTR^{+}18]. Automated [NLH^{+}19]. Automatic [BTR^{+}18, FBAG20, NZH^{+}23]. Automatically [SWH^{+}24]. Autonomous [CS23, SAK^{+}19]. AutoWitness [GPL^{+}12]. AUVs [RHS20]. Availability [LGXC23, ZGH^{+}21, ADF12]. Availability-aware [LGXC23]. Average [CG18]. Averages [Kou18]. Aviation [BBD^{+}23]. AviSense [BBD^{+}23]. Avoidance [XJL^{+}23, WEC11]. Aware [ARWK19, BIMD19, CS23, EA15, GSM^{+}22, MCLM20, RBS16, TNBG18, XXHL16, XZL^{+}20, XKW^{+}22, YXFL17, ZZZ^{+}20, ZZW^{+}23b, COS19, CCC^{+}21, DLD09, FS13, GAJ^{+}06, GJT^{+}22, HR13, LDG^{+}21, LGXC23, LCC10, MKM^{+}20, SDYC22, ZGH^{+}21, CYD^{+}24, HBLR05]. Awareness [SPI^{+}24].

B5G [MLS^{+}22]. Backscatter [SSL^{+}22, WYC^{+}24, ZLZ21]. Balancing [KKP18, LP08, LKA10]. Band [CSLJ23, GTL19, SCS22, ZZW^{+}23b]. Bands [SMS22]. Bandstitched [PKC^{+}18]. bandwidth [CHN^{+}13, CRW07, EMBP12]. bandwidth-constrained [CRW07]. Barometer [DSA^{+}20]. BaroSense [DSA^{+}20]. Barrier [FLS^{+}14, ZHT^{+}23, CLX09]. Barycentric [PWS^{+}23]. Base [YHC^{+}24, SH09]. Based [AH14, BWP^{+}24, BNN^{+}20, CKHP19, CZX^{+}22, CC23, DWF^{+}23, EY14, FHST22, FSTH23, FLCH23, GAMW22, GCAK17, GZJE23, HMLJ17, HSL^{+}15, HKW^{+}24, JAC19, KGBS18, KGDC22, KLC^{+}16, Kou18, KRP15, LWZ24, LHHW24, LWJ^{+}23, LWCJ14, LWX^{+}21, LYF^{+}23, MDC17, MNLZ18, NGBB14, RRA22, RKRP17, SBK22, SMR^{+}14, SNC^{+}23, SLC^{+}22, SUZK19, SCD^{+}24, SZG^{+}15, TZZ22, WJD16, WTX^{+}16, WZZ^{+}21, WL23, WLZ23, WLLZ24, WMT^{+}19, XXW^{+}24, XCT^{+}16, XYW^{+}22, XWW^{+}20, XJR^{+}17, XDM^{+}21, YSK^{+}15, YZZD23, YRB^{+}17, YTZ^{+}23, ZZZ^{+}22, ZCZL22, ZLB^{+}23, ZSZ20, ZLL^{+}22, Amm23, AAA06, BLWY06, BJW^{+}22, CLSW12, CTWG24, DBC^{+}24, EMBP12, GCRB12, GBS08, GWS^{+}24, GZZ^{+}23, HSGW21, HM07a, HLZ^{+}24, HCXT09, JHU^{+}13, JYC^{+}24, KBD14, KKK08, KPS12, KAS^{+}10, LWG09, LHZZ20, LL21, LTDZ22, LDS^{+}22, LLW^{+}23, LLDZ23, LXYT24, LND08, LHX^{+}21, MLZ^{+}24, MDM^{+}20, MG24, MS12, NEKK12, NJS05, NLH^{+}19, OXZ^{+}23, PDMJ10, RS19, SW22, SGM08, SCL^{+}19]. based [SDW^{+}23, TJZ^{+}13, TXC^{+}13, TBL07, VG10, VAC13, WYY^{+}19, WZLM21, WJY^{+}24, WHYC19, YQLD22, YH13, YXG^{+}19, YYL^{+}23, ZKS10, ZJX10, ZLZ21, ZZ21, ZWWL23, ZDS^{+}21, ZBA07, ZWG24, LLH22, BHA^{+}13]. bases [JLYG13]. Bats [DML^{+}16]. Battery [CKHP19, HKG^{+}19, SCL^{+}19, WXG^{+}24, ZLGL19, ZLGL20]. Battery-Free [ZLGL19, SCL^{+}19, WXG^{+}24, ZLGL20]. Batteryless [BAHS24]. Bayesian [BT18, NP12, ORRJ12, WB17]. Beam [WCLD23]. Beamforming [SNY^{+}24, FLJ^{+}13]. Beamforming-assisted [SNY^{+}24]. Beams [TCB^{+}14]. BEANet [XDL^{+}24]. Bed [AJH^{+}20]. Behave [FSTH24]. Behavior [CPSS23, GZK^{+}23, HL17, KGBS18, LLW^{+}23, LZGX23, NDM^{+}13, SYX^{+}23, YTZ^{+}23, ZZW^{+}23a, ZGH^{+}21]. Behavior-aware [ZGH^{+}21]. Behavior-based [LLW^{+}23]. Behavior-oriented [NDM^{+}13]. Behaviors [KSR^{+}20, MDB^{+}23]. Behaviour [MSK^{+}23]. belief [WL14]. belts [CLX09]. benchmark [LDH06]. benefits [JSBN^{+}12]. between [ÁKSW22, FLFW13]. Beyond [CWK^{+}22, QXZZ22, YJWL13]. Bi [JAC19]. Bi-dimensional [JAC19]. BikeGPS [CT19]. BikeNet [EML^{+}09]. Bikes [CT19]. BiLSTM [ZWWL23]. BiLSTM-based [ZWWL23]. Bin [YRB^{+}17]. Bin-Based [YRB^{+}17]. Binary [BQB^{+}11, LMP14, SKM^{+}11, SMMS09, WBS10]. biological [KAH^{+}10]. Biometric [WWZ24]. Bit [HCL15]. Bitrate [ZTZX23]. Blame [GLL^{+}24]. BLE [BDP24, XDL^{+}24]. Blind [BY19, KGDC22]. Blinder [YA24]. BLITZ [SDBT19]. block [LDH06]. Blockchain [HKW^{+}24, LWZ24, SNC^{+}23, TBS^{+}24]. Blockchain-Based [LWZ24, SNC^{+}23]. Blood [ZYC^{+}23]. BLOW [WWL^{+}16]. Blueprints [LSW14]. Bluetooth [YYC^{+}19]. Body [AJH^{+}20, DSH16, DGS16, Hau14, MSAJ18, RHD17, LYG^{+}13, VG10]. bogus [XWDN12]. BOND [MCGZ21]. Boosting [HXZ23a]. both [HTW07]. Bottleneck [MCGZ21]. bound [ZH05]. Boundaries [Sch15]. Boundary [CS17, CS18, SSGM10, ZBA07]. Bounds [Bra07, JTE20, MCW^{+}16]. breach [CRW07]. Breaking [TDZ^{+}22, YYXL22]. BreatheBand [GYG^{+}23]. Brick [FC18]. Bridging [ZWWZ20]. Bringing [IHGS15]. Broadcast [XCC^{+}15, ZCZL22, ZLGL19, JROH09, NLD08, SGM08, WDLN09, XWDN12]. broadcasting [HM07a]. Buffer [WJZ21]. buffering [LCC10]. Bug [SCD^{+}24]. bugs [KLA^{+}14]. Building [DCD24, ECPC14, FPA^{+}20, KOD^{+}14, LCM21, SCL^{+}14, YXG^{+}19]. Buildings [ABC^{+}18, CHSA18, HBW^{+}18, WCV^{+}18, ZWWZ20]. BuildSense [COS19]. BuildSys’17 [NJZ18]. Built [AKC^{+}18]. bulk [GCRB12]. Bundling [ZZ23]. Bytecode [RS19]. Byzantine [ZJZ24b]. Byzantine-Robust [ZJZ24b].

C4IoT [GDM22]. cache [PA05]. Caching [XFZ^{+}21, ZTZX23]. CAG [YS07]. Calibrating [KNSM14]. Calibration [ALNT22, BTR^{+}18, CML^{+}21, DRC06, TXY^{+}13]. CAMA [DRW^{+}14]. Camera [BTR^{+}18, DSZ^{+}24, GLQ^{+}22, HLL^{+}23, TAT14, TMAP14, CHN^{+}13, DRC06, ES12, ELYR14, IW14, KNSM14, MCT14, SPK14, ST12, WL14, WC13]. Cameras [DXC^{+}21, YRB^{+}17, EGG13]. Campaigns [DD11]. Can [LSW14]. cane [HBC^{+}09]. canonical [TP07]. Canyons [CT19]. capabilities [Bra07]. capacitor [ZGHZ12]. capacitor-driven [ZGHZ12]. Capacity [BIST18, HR13, LFW^{+}19, XDL^{+}24, ZJZ12]. Capacity- [HR13]. CapNet [SSL^{+}19]. Capping [SSL^{+}19]. Capture [DRW^{+}14, MDC17]. Cardiac [WWZ24]. Carpooling [ZHZ^{+}16]. Carrier [BBEM^{+}24, GLG^{+}23]. Carrier-Sense [GLG^{+}23]. Carriers [SDZZ24]. Carries [ZHJ^{+}20]. Cascaded [RSK^{+}21]. Case [COP^{+}16, ZGJ^{+}22, IV12, JKS^{+}10, MRM09]. Casual [WTC22]. Catching [GSW09]. CATS [ZGX^{+}16]. CDS [FKMS06]. Cell [CZX^{+}22, MLS^{+}22, JHU^{+}13]. Cell-based [JHU^{+}13]. Cells [WXG^{+}24]. Cellular [BRR^{+}18, SJP^{+}22, SDZZ24, TDZ^{+}22, ZZX^{+}20]. Center [LWL^{+}21, SSL^{+}19]. Centers [CTW^{+}15]. Centric [HCL15, LCM21, XDX^{+}14, CUdVY13, LCH^{+}09, YSM08]. certification [GSL10]. Chain [PK20, YBY^{+}24]. Chaining [XZL^{+}20]. Chains [LGXC23]. Challenges [AAJ^{+}23, GSGA23, RDP16, RGB^{+}17]. Channel [KR18, LHHW24, NK15, RRA22, TNBG18, WZLM21, SC12, XTZ08, Yan22]. Channels [CSLJ23, GM14, LWH^{+}22, WQH^{+}22, VMS10, WWXY13]. Characterization [ZZX^{+}20]. Charge [SCG^{+}15, ZZZ^{+}20]. Charge-Aware [ZZZ^{+}20]. Charger [WXD^{+}23, YRM^{+}24]. Chargers [WTX^{+}23]. Charging [CKHP19, KJD^{+}23, LDC^{+}19, LXR^{+}16, LWX^{+}21, MZW^{+}19, WCW^{+}23, WYD^{+}22, WXD^{+}23, YWD^{+}21, ZWY21, ZZW^{+}23b]. Checking [GZK^{+}23, KA13]. Chest [CC23]. Chief [Liu21]. Child [CJL^{+}20]. Children [YRB^{+}17]. Chipnet [SSL^{+}22]. Chromophore [BNN^{+}20]. ciphers [LDH06]. Circuits [ZJZ^{+}24a]. Cities [XXW^{+}24]. City [SDZZ24, WJ21, XFZ^{+}21]. City-wide [WJ21]. Class [LTZ^{+}24, GZZ^{+}23]. Classification [AJH^{+}20, BBD^{+}23, LTZ^{+}24, PSR^{+}22, RSK^{+}21, XKW^{+}22, XWL24, YRB^{+}17]. classifying [BNG12]. Clear [KR18]. Client [LGLD23, ZWL^{+}24]. Client-agnostic [LGLD23]. Clients [XKW^{+}22]. Clock [JTE20, VTY18]. clocks [SSC^{+}10]. Clothing [SZX17]. Cloud [LDS^{+}22, MYW^{+}24, MLS^{+}22, NJL24, QWC^{+}22, XWL24, LLW^{+}23]. Cloud-Aided [QWC^{+}22]. Cloud-Edge [NJL24]. Cloud-Edge-based [LDS^{+}22]. CloudNavi [TGG^{+}19]. Clouds [TGG^{+}19, TTBH14]. Cluster [AH20, KKK08, NGBB14, HM07a, JKS^{+}10]. Cluster-based [KKK08, HM07a].

Cluster-tree [AH20, JKS^{+}10]. Clustered [RRA22, MZWT10, YS07]. ClusterFL [OXZ^{+}23]. Clustering [FSTH24, FLCH23, LHX^{+}21, OXZ^{+}23, MB09]. Clustering-based [LHX^{+}21, OXZ^{+}23]. CMAC [LFS09]. CNN [LTDZ22]. CNN-based [LTDZ22]. CO [AAHS18]. coal [LL09]. coalition [VAC13]. Code [DCBL15, PBM11, QM13]. codebook [ZLZ21]. coded [ME21]. Codes [DML^{+}16, LCD22, JJ15]. Coding [EA15, JAC19, VRSR15, WKYH17, DVS^{+}14, KAAF13, MB09, WZL08]. Coding-Aware [EA15]. Coexistence [DSH16]. Coexisting [MSAJ18]. COFlood [CZMM23]. Cognitive [SMW23, ZSLL23]. CoHop [WZLM21]. Cold [SMZ^{+}17]. Cold-Start [SMZ^{+}17]. Collaboration [LLW^{+}23, MYW^{+}24, PCPK14, SWYW21, WTH^{+}23, ZCZL22]. Collaborative [CRZ^{+}20, GSL10, HM07a, KQ14, LLZ^{+}22, LWY^{+}21, NJL24, WYY^{+}19]. Collaboratively [LSW14]. Collection [DDA11, HLN^{+}11, JJ15, LCLY22, WBS14, YB17, ZZW^{+}23a, ZLGL20, GFJ^{+}13, JHU^{+}13, LKA10, Su07, WZL08]. collision [CCC^{+}21]. Collisions [WZZ^{+}23]. Combating [CWY24]. Combinable [PLW^{+}24]. Combinatorial [TCB^{+}14, RR09, Su07]. ComFor [Amm16]. Commercial [WCV^{+}18, ZZX^{+}20]. Commodity [SYX^{+}23]. Communicate [SLS^{+}22]. Communication [ÁKSW22, BY19, CSA06, CD21, CSLJ23, DGS16, EY14, FM15, GM14, GHZ^{+}22, Hau14, HBW^{+}18, HWF^{+}24, LCJ^{+}23, MSK^{+}23, ME21, PK20, PCA^{+}23, RRA22, RHS20, SJP^{+}22, SBSD18, SMS22, SCS22, SDBT19, ZGJ^{+}22, ZJZ^{+}24a, ZDS^{+}21, ZZW^{+}23b, KGGK11, KAR^{+}14, LJY^{+}10, PDMJ10, XLZ^{+}07]. communication-efficient [KGGK11]. Communication-Topology-preserving [HWF^{+}24]. Communications [SE23, WWFX11, WLS^{+}16, ZLZ21, SYL09]. Communities [SBSD18]. compact [SZG13]. Comparative [MPRS16, MPC^{+}10, RBD13]. Compensation [BNN^{+}20, WJZ21, XXHL16, SC12]. Compilation [RS19]. Complete [XTXW22]. Complex [CS18, LFNS14, TJLK14, WHYC19, LWG09]. Complex-Valued [WHYC19]. Complexity [VRSR15, GJNC^{+}14, KLA^{+}14, MB09]. Complexity-Constrained [VRSR15]. Component [AH14]. Component-Based [AH14]. Components [ZWW^{+}23, TLRE13]. Composite [Amm16]. Composition [FM15]. Comprehensive [PCA^{+}23, PGY^{+}24, SYL^{+}22, WXD^{+}23]. Compressed [CTWG24]. Compression [AKSM15, AH14, JAC19, LL16, RBD13, TCN^{+}17, WB17, ZMVR14, HM07a, KLJ12, PKG08]. Compressive [CGB^{+}19, CZC^{+}24, EA15, XAKV15, ZLL^{+}22]. compromise [DLD09, PX13]. compromises [SZZC08]. Compromising [LHX^{+}21]. Computation [SHWW20, ZWWL23]. Computational [Amm23, XRS10]. Computer [CZC^{+}24, IW14]. Computing [ELR^{+}22, HMG^{+}24, LDG^{+}21, LLX^{+}22, LGXC23, LTL^{+}24, LLH22, MLS^{+}22, PLW^{+}24, QXZZ22, SMW23, SHWW20, TZZ22, XQL^{+}24, ZLX^{+}24, Dji10]. concave [WX08]. Concealed [ARWK19]. Concept [WZL08]. Concepts [BASM16]. Concurrency [LCH^{+}19b, LCH^{+}20]. Concurrent [BBEM^{+}24, CZMM23, CP20, LCJ^{+}23, WYC^{+}24, XHZG22]. condition [TBL07]. condition-based [TBL07]. Conditioning [CA22]. conditions [FT06]. Confident [DTY^{+}22]. Configuration [FBAG20, JZX^{+}20, WLW^{+}23, WWXY13, XWZ^{+}05, XLZ^{+}07]. conflicting [WKA14]. Congestion [DSA^{+}20, KKK08, WEC11]. Connected [GCAK17, MDB^{+}23, SBSD18, XWC^{+}23, YTB^{+}14, ZDG09]. Connecting [SWH^{+}24]. Connectivity [BGMP15, ENPNF13, LWG09, TJZ^{+}13, WJD16, YJL^{+}22, CJS11, HTW07, XWZ^{+}05]. Connectivity-Based [WJD16, LWG09, TJZ^{+}13]. Consensus [RBS16, TBS^{+}24]. Consensus-Aware [RBS16]. conservation [XWZ^{+}05, YPW^{+}13]. conserving [HLTC06, PA05]. Considering [PZOZ21, ZZPW23]. Consistency [JM16]. constant [FT06, LHRM09]. Constrained [ÁKSW22, DBOD^{+}16, LDC^{+}19, VRSR15, ZMVR14, BJW^{+}22, CSA06, CRW07, RS19]. Constraints [RD16, YWD^{+}21, GCBL06]. Constructing [PSB^{+}14]. Construction [SCL^{+}19, WWL^{+}16, WJD16, PR10]. Consuming [LLH22]. Consumption [JZX^{+}20, LP08]. Contactless [LWJ^{+}23, LJLW19, SYX^{+}23]. Containing [XWDN12]. Contamination [PK19]. Content [XFZ^{+}21, XKW^{+}22]. Contention [XKW^{+}22, DIE14, RDR07, ZJX10]. Contention-Aware [XKW^{+}22]. contention-based [ZJX10]. Context [BIMD19, KSR^{+}20, PLW^{+}24, YXFL17, ZZW^{+}23b]. Context-adaptive [PLW^{+}24]. Context-Aware [BIMD19, YXFL17, ZZW^{+}23b]. Contextual [LJW^{+}21]. Continuous [LHZZ20, LTDZ22, LLDZ23, LYL^{+}24, NJL24, JHU^{+}13, WZL08]. Contour [YXG^{+}19, SCWC13]. Contour-based [YXG^{+}19]. contract [GDM22]. Control [DCD24, GTL19, HL17, JZL^{+}19, KCE^{+}20, KPCB20, LWL^{+}21, LYZ^{+}24, LMZ^{+}16, PK20, WCPC20, ZLW^{+}24, IW14, KKK08, KRJ09, LSW06, NC10, OBB^{+}13, SG10, WWLX13, ZCLJ14]. Controlled [KSMH13, PG10]. Controlling [BIST18]. Convenient [CWS^{+}22]. convergent [LFS09]. Conversion [ZZG^{+}24]. Convex [CS18, TJLK14]. Convolution [LLW^{+}23]. Convolutional [CC23, LHZZ20]. cooled [LWL^{+}21]. Cooperation [CT19, HWS^{+}20].

Cooperative

[BIMD19, DSH16, DGS16, Lam15, LK09, NK14, RRA22, ZGX^{+}16, HZX^{+}24, SYL09]. coordinate [DABNR10]. Coordinated [YYXL22]. coordinates [CA06]. Core [GZZ^{+}23]. CoRec [LLW^{+}23]. Correction [JTE20, KRP15, RKRP17, KLC13]. Correlated [HCL15, WKYH17, GNDC08, JP06]. Correlation [SUZK19, WZLM21, PKG08]. Correlation-based [WZLM21]. Correlations [LWY^{+}21, JKK08, YS07]. Cost [CWS^{+}22, COS19, CML^{+}21, LFL^{+}19, TAT14, WXD^{+}23, ZLX^{+}24, ALNT22, ODCP13, PKS^{+}23]. Cost-aware [COS19]. COTS [HZX^{+}24, SYX^{+}23, WSC^{+}23]. CoUAS [HWS^{+}20]. count [NEKK12]. Countermeasure [HXZ23b, TDZ^{+}22]. Countersniper [LNV^{+}05]. Counterstrategy [CPL^{+}20]. Counting [CG18]. Counts [HCL15]. Cov [Amm16]. Cov-ComFor [Amm16]. cover [ZDG09]. Coverage [Amm23, CRW07, DTY^{+}22, DSZ^{+}24, FLS^{+}14, GM14, KQ12, Lam15, LFNS14, MZWT10, MCT14, MAG13, SAK^{+}19, SCL^{+}19, YJL^{+}22, YTB^{+}14, Amm13, Bra07, CGVC06, CLX09, CLH^{+}13, CGD12, ENPNF13, HLTC06, HTW07, LP06, MRM09, SCWC13, WC13, WLZ13, XWZ^{+}05, YYM^{+}10, YLL13]. coverage-preserving [HLTC06]. Covered [Amm16]. COVID [AAJ^{+}23, CC23]. COVID-19 [AAJ^{+}23, CC23]. CPS [JYC^{+}24, LTZ^{+}24]. CPU [JCZ^{+}22]. Crashes [GLL^{+}24]. created [MPC^{+}10]. Credential [YLSZ19]. criteria [MCT14]. Critical [CJS11, CML^{+}21, CWK^{+}22, GXQ^{+}22, PSB^{+}14, TYGW15]. CRONOS [SZ19]. Crop [LWLT24]. Cross [CD21, GHZ^{+}22, KPRH14, LCD22, SMS22, SCS22, WXL^{+}19, YBY^{+}24, ZGJ^{+}22, ZZY^{+}23]. Cross-chain [YBY^{+}24]. Cross-labelling [ZZY^{+}23]. Cross-Layer [KPRH14, LCD22]. Cross-Technology [CD21, GHZ^{+}22, WXL^{+}19, ZGJ^{+}22, SMS22, SCS22]. Crowd [HSL^{+}15, MJS^{+}19, SLC^{+}22, SML18, XLO^{+}23, ZZ21, ZZ23]. Crowd-sensed [SLC^{+}22]. Crowd-Sensing [SML18, XLO^{+}23]. crowded [KQ12]. CrowdLoc [BRR^{+}18]. Crowds [BRR^{+}18]. Crowdsensing [CGB^{+}19, Kou18, LLZ^{+}20, RGB^{+}17, RFS^{+}19, TGG^{+}17, WYY^{+}19, WLZ23, WJY^{+}24, ZLL^{+}22, ZGH^{+}21]. Crowdsourcer [LLZ^{+}20]. Crowdsourcing [DSA^{+}20, LWZ24, MKM^{+}20, PZOZ21]. CSI [LWJ^{+}23, WHYC19]. CTP [GFJ^{+}13]. Current [AAJ^{+}23, AMTH^{+}17, BJR15]. Curve [WWL^{+}16, WJD16]. cuts [SST08]. Cyber [HLZ^{+}24, KSR^{+}20, LSX24, LTZ^{+}24, SJH^{+}18, SDX^{+}20, WLLZ24]. Cyber-Physical [LSX24, SJH^{+}18, SDX^{+}20, WLLZ24, HLZ^{+}24]. Cycle [CZMM23, GLS^{+}14, Pha16, XCC^{+}15, PEFSV13, SPK14, WWLX13]. Cycled [Amm16, BGMP15, LCH^{+}19b, SSC^{+}10, YH13]. Cycling [LLL14, NK15, ZZZ^{+}20, JCC^{+}13, LCJ^{+}23]. cyclist [EML^{+}09].

D [Amm16, TJZ^{+}13, BY19, NXW^{+}22, SNK^{+}22, TJZ^{+}13, TGG^{+}19, WWL^{+}16, WJD16, XYW^{+}22, YRB^{+}17]. D- [Amm16]. D/ [TJZ^{+}13]. D2D [WYY^{+}19]. DAG [LTL^{+}24]. DAML [ZSZ20]. Data [ALS23, ARWK19, AAHS18, ADF12, BYD^{+}15, CTWG24, CTW^{+}15, DD11, DDA11, EA15, FSTH23, GJT^{+}22, GZZ^{+}14, GZZ^{+}23, HMLJ17, HBKP14, HLN^{+}11, HL17, HCL15, HKW^{+}24, JZL^{+}19, KYM17, LDDL24, LWL^{+}21, LLX^{+}14, LZGX23, LWCJ14, LC14a, LLZ^{+}20, LCM21, LCLY22, LYST23, MY24, MKFD^{+}23, PNL^{+}22, PSB^{+}14, PSR^{+}22, SSL^{+}19, SJH^{+}18, SZ19, SCL^{+}14, SLC^{+}22, SDZZ24, SXD^{+}15, SG11, SWYW21, TCN^{+}17, TDZ^{+}22, WRYL11, WWZ^{+}21, WCW^{+}23, WJY^{+}24, WJ21, WBS14, XAKV15, XQL^{+}24, XWL24, YMY^{+}23, YB17, YHC^{+}24, ZZW^{+}23a, ZCZL22, ZZY^{+}23, ZGX^{+}16, ZSZ20, ZLL^{+}22, ZZG^{+}24, ZLGL20, Amm13, AAA06, CDGC12, CCMT09, CC11, CNMH08, CGD12, CUdVY13, FLJ^{+}13, GCBL06, GNDC08, JHU^{+}13, JP06, Kal10, KBD13, KLJ12, KLA^{+}14, KVI^{+}13, LM10a, LM10b, LKA10, LK09, MDC^{+}09, NRC^{+}09, NP12, NDM^{+}13, ORRJ12, PA05, PH10, RKW^{+}06, SG10, TXY^{+}13, TJWK13, WL14, WZL08, WLD10, ZKS10, ZJX10]. data [ZSJN07]. Data-Anomaly [DD11]. Data-Centric [HCL15, LCM21, CUdVY13]. Data-Driven [PSR^{+}22, WCW^{+}23, LC14a, WJ21, ZZG^{+}24]. Data-plane [TDZ^{+}22]. data-rate [LM10a, LM10b]. Dataset [MG24]. datasets [SGG10]. DC [MKFD^{+}23]. DCS [CUdVY13]. DDoS [HMG^{+}24]. Deadline [YWD^{+}21]. Dealing [NZR10]. Decentralized [HLTC06, KRJ09, VDV16]. Decode [ZDS^{+}21]. Decoding [WZZ^{+}23, XTXW22]. Decomposition [AAHS18, SDYC22]. Dedicated [LZN19]. Deep [ALS23, BNPR20, CLX^{+}21, CTWG24, DCD24, FLCH23, JGK^{+}23, Kun22, LWL^{+}21, LTDZ22, LLW^{+}23, LYF^{+}23, LYST23, MDB^{+}23, RKLM23, SYT22, XFZ^{+}21, YZZD23]. DeepHeart [CLX^{+}21]. DeepMTD [SYT22]. Defending [LWCJ14, XTZ08]. Defense [MY24, SYT22]. DeFFusion [LTDZ22]. Delay [DBOD^{+}16, KPK12, PS17, VRSR15, WXL^{+}19, WTX^{+}23, WWLX13]. delays [LWSL12]. Delivered [ZZC^{+}23]. Delivery [DLD^{+}23, KLC^{+}16, PSB^{+}14, WXL^{+}19, ZZG^{+}24, PH10]. demand [DLD^{+}23, KPB^{+}08]. Democratizing [AKC^{+}18]. Demodulation [XTXW22]. Dense [YJL^{+}22, NEKK12]. denser [JSBN^{+}12]. density [CJS11]. Dependable [TNBG18, WRYL11]. dependent [CLJ^{+}23]. Depleting [CPL^{+}20]. deployed [Amm13]. Deploying [ZHT^{+}23, GRE^{+}07]. Deployment [DLD09, DTY^{+}22, GSGA23, GCAK17, LYZ^{+}24, PLW^{+}24, WXD^{+}23, XWW^{+}23, XWC^{+}23, DEM^{+}12, JSBN^{+}12, KC14, LN05, MPS10, OBB^{+}13, RR09, SCWC13]. Deployment-aware [DLD09]. Deposit [LWZ24]. Deposit-Free [LWZ24]. deprivation [SZZC08]. Depth [GLL^{+}24, YRB^{+}17]. derived [KLC13]. Design [BR15, CPP^{+}17, CSLJ23, DEM^{+}12, FC18, GKRW17, GZJE23, HBC^{+}09, LYZ^{+}24, LCJ^{+}23, LCH^{+}09, OBB^{+}13, ODCP13, PDP^{+}17, RFB^{+}14, XDX^{+}14, ZWY21, CK09, TBL07, ZSG09]. Designing [COP^{+}16, SBSD18]. designs [RR09]. Detect [HLZ^{+}24]. Detecting [GZZ^{+}14, LGLD23, SST08, WLX^{+}23, YRB^{+}17]. Detection [AJH^{+}20, ARWK19, BBD^{+}23, BNPR20, CLL^{+}23, CS17, CS18, CA22, DD11, DSA^{+}20, GZK^{+}23, GZZ^{+}23, HZX^{+}24, HSL^{+}15, HWT^{+}22, IPMGL18, LZZ^{+}15, LDS^{+}22, LDL^{+}24, LYF^{+}23, LTZ^{+}24, MLS^{+}22, MNLZ18, NXW^{+}22, PTDD16, Sch15, SCD^{+}24, SDČ10, TCC^{+}23, WHQ^{+}23, WNM^{+}24, XWW^{+}20, ZLB^{+}23, ZYC^{+}23, Bra07, CGVC06, KBD14, KC14, KPK12, LPR09, NP12, PC10, TXC^{+}13, TTBH14, WEC11, WRS10, ZDW^{+}10, dLM14, SGG10]. detector [GAJ^{+}06]. determine [RMB^{+}10]. Determining [IPMGL18]. Deterministic [BDO14, BQB^{+}11, SC15, SB16]. Developing [SMR^{+}14, GRE^{+}07]. Development [DLG^{+}21, ODCP13]. Device [JCZ^{+}22, LZGX23, ME21, WHQ^{+}23, ZW24, ZYL^{+}24, SWYW21]. Device-free [WHQ^{+}23, ZW24]. Device-to-device [ME21]. Devices [BAHS24, GDM22, HPS^{+}18, JZX^{+}20, LDG^{+}21, LWX^{+}21, MDM^{+}20, RS19, SDX^{+}20, SSL^{+}22, SWH^{+}24, WSC^{+}23, XWW^{+}20, XJR^{+}17, ZZH^{+}23, ZJZ^{+}24a, KNSM14, MKK^{+}13]. Diagnosis [CC23, YSK^{+}15]. Diagnostic [SEZA13]. Diagram [MLZ^{+}24]. Diary [FSSR15]. DICTUM [WWB^{+}19]. differences [XRS10]. Differentiating [KR18]. Differently [FSTH24]. diffusion [Gel07, NGSA08]. Digital [GXQ^{+}22, LCF^{+}22, ZLX^{+}24]. Digraphs [KKRR15]. Dimensional [Amm16, JAC19]. Dimensioning [JKS^{+}10]. Dimensions [ALY^{+}23]. Dimming [ZMXM24]. Direct [Den09]. Directed [JROH09, EFI^{+}10, LYST23]. Directional [YTB^{+}14, ZJZ12]. Directions [AAJ^{+}23, AMTH^{+}17]. Discovery [MJS^{+}19, WJY^{+}24, ZHL^{+}15, ZGH^{+}21, ZVPS10]. Discrete [KKP18]. Disease [TCC^{+}23]. DISH [TDD^{+}19]. Disjoint [HSD16]. disk [FKMS06]. Disorder [ALS23]. Dispatching [MCLM20]. Disruptions [MCLW23]. Disruptive [PS17, SXD^{+}15]. dissemination [FLJ^{+}13]. Distance [HMLJ17, ZWW^{+}23, KASD09, SS13, YJWL13]. Distance-Based [HMLJ17]. distance-sensitive [KASD09]. distances [XRS10]. distortion [GCBL06, VMS10]. Distributed [AH20, AHK16, BYD^{+}15, BJR15, BIST18, CVY09, CPH06, DRC06, GSGA23, HTW07, JJ15, KJD^{+}23, LED20, LWSL12, LH09, LWCJ14, SZG13, SGB15, VRSR15, WL14, WBS10, WWL^{+}16, YM14, YLL13, ZLL^{+}22, ABM13, CNMH08, ELYR14, FS13, FKMS06, GJNC^{+}14, KC14, KASD09, PG09, TMAP14, WC09, WC12, ZVPS10, ZSJ06, TDD^{+}19, WWB^{+}19]. Distribution [CTW^{+}15, PK19, SPK^{+}10, ZW05]. distributions [SZG13]. Districts [ZZX^{+}20]. Diversities [HXZ23a, XHZG22]. diversity [KAR^{+}14]. Division [ZYZ^{+}19]. DMCP [KJD^{+}23]. DNN [JYB^{+}21, PLW^{+}24]. DOA [BY19]. DOA/Symbols [BY19]. Does [RSK^{+}21]. Domain [ZWG24]. Dominating [SCL^{+}19]. Don’t [HXZ23a]. Doorway [GKRW17]. Doppler [KAS^{+}10]. Downtime [SXD^{+}15]. Downward [IIPK20, KLC^{+}16, KJP^{+}15]. DPIVE [ZLD^{+}24]. DQN [YTZ^{+}23]. Drift [KRP15, RKRP17]. Driven [PK19, PSR^{+}22, SZ19, WCW^{+}23, JLZL19, LC14a, SPK^{+}10, SLC^{+}22, WJ21, ZZG^{+}24, ZGHZ12]. Driver [CLL^{+}23, ZGH^{+}21]. Drivers [XWW^{+}20]. Driving [BNPR20, LYF^{+}23, WLX^{+}23]. Drone [SCD^{+}24]. Drone-Based [SCD^{+}24]. Drones [SAK^{+}19, SPI^{+}24]. droplet [LCC^{+}13]. Drowsiness [CLL^{+}23, XWW^{+}20]. DrunkWalk [CRZ^{+}20]. DSME [ÁKSW22]. DSME-LoRa [ÁKSW22]. Dual [XWL24, ZLZ21]. Dual-codebook-based [ZLZ21]. DualMOP [KJP^{+}15]. During [CGB^{+}19, JYB^{+}21]. Duty [Amm16, BGMP15, CZMM23, GLS^{+}14, LLL14, LCH^{+}19b, LCJ^{+}23, PEFSV13, Pha16, XCC^{+}15, ZZZ^{+}20, JCC^{+}13, SSC^{+}10, SPK14, WWLX13, YH13]. Duty-Cycle [GLS^{+}14, Pha16, PEFSV13, WWLX13]. Duty-Cycled [Amm16, BGMP15, LCH^{+}19b, SSC^{+}10, YH13]. Duty-Cycling [LLL14, LCJ^{+}23]. DutyCon [WWLX13]. dWatch [XWW^{+}20]. Dynamic [AHK16, CQDW21, DD11, FM15, GM14, GDM22, Lam15, LDG^{+}21, MDM^{+}20, ME21, NC10, RKW^{+}06, SBSD18, SGB15, WRYL11, WB17, WJZ21, YLSZ19, ZKS10, ZYZ^{+}19, ZLW^{+}24, IR12, KBD14, WWLX13]. Dynamically [PLW^{+}24, SML18].

Each [CWS^{+}22]. Early [JYB^{+}21]. earthquake [TXC^{+}13]. EATU [HWT^{+}22]. Eavesdropping [LHHW24, PX13]. EchoSensor [LDL^{+}24]. Economic [MKFD^{+}23, ELYR14]. Economical [ZZW^{+}23a]. ECPC [SXD^{+}15]. ECRLoRa [MYW^{+}24]. ECT [WXL^{+}19]. eDeepSave [JYB^{+}21]. Edge [BWP^{+}24, HMG^{+}24, JYB^{+}21, LLZ^{+}22, LDG^{+}21, LDS^{+}22, LLX^{+}22, LGXC23, LLW^{+}23, LTL^{+}24, LYY24, LLH22, MYW^{+}24, MLX^{+}24, MLS^{+}22, ME21, NJL24, PLW^{+}24, SHWW20, SDYC22, SCLG24, TZZ22, XYJ^{+}23, XZL^{+}20, XFZ^{+}21, XQL^{+}24, YMY^{+}23, ZZW^{+}23a, ZCZL22, ZTZX23, ZZPW23, ZLX^{+}24]. Edge-assisted [LLZ^{+}22, LYY24, SDYC22, SCLG24]. Edge-Cloud [MYW^{+}24, MLS^{+}22, LLW^{+}23]. Edge-coded [ME21]. Edge-Computing-Supported [SHWW20]. Editor [Liu21]. Editor-in-Chief [Liu21]. Editorial [LSX24, Liu21]. Effect [DRW^{+}14, MDC17, ZJZ^{+}24a]. Efficiency [LFW^{+}19, PAYL22, XCC^{+}15, FLFW13, SYL09, VAC13, WIF^{+}11]. Efficient [Amm16, BAHS24, CCMT09, CA22, DRW^{+}14, DCBL15, DML^{+}16, EA15, FSTH23, GLG^{+}23, GNDC08, HSGW21, HBKP14, HMG^{+}24, IIPK20, KLC^{+}16, LED20, LLW^{+}23, LWZ24, LWM^{+}21, LHX^{+}21, MCLM20, NGBB14, NZLH15, NZH^{+}23, PBM11, PCPK14, QWC^{+}22, RRA22, SDBT19, TFL^{+}24, TBS^{+}24, VPB^{+}20, WTX^{+}16, WLS^{+}16, WMT^{+}19, XLG^{+}22, XXHL16, YB17, ZSKH08, AH20, CNMH08, CLH^{+}13, CGD12, DDHC^{+}12, FLJ^{+}13, GCRB12, GCBL06, GFJ^{+}13, HKL^{+}06, HWT^{+}22, JCC^{+}13, KPB^{+}08, KGGK11, KW09, LPV^{+}09, LDZ13, LWY^{+}21, LFS09, MP10, NLH^{+}19, QXZZ22, SDYC22, Su07, SNY^{+}24, TJWK13, TBL07, VG10, WEC11, WBS10, WLD10, WLW^{+}20, WYC^{+}24, XDL^{+}24, ZLZ21, ZLGL20, ELR08, ZSJ06]. EGM [XLG^{+}22]. EH [AMAT^{+}18]. EH-WSNs [AMAT^{+}18]. eigenvector [CLS12]. Electric [WCW^{+}23]. Electrical [VTY18]. Electromagnetic [LTY18]. Elements [DDA11]. elephants [GSW09]. Eliminating [WCLD23]. Elliptical [RBLP09]. Embedded [CBSA18, DCBL15, JZX^{+}20, XKW^{+}22, IV12, LJY^{+}10, MKK^{+}13, SSC^{+}10]. Embedding [WL23]. Embedding-Based [WL23]. Emerging [CPSS23]. EMG [DWF^{+}23]. Emotion [JLZL19, SMZ^{+}17]. Emotion-driven [JLZL19]. Empirical [DGS16, GKRW17, YJL^{+}22, SDTL10].

Empowered [KCE^{+}20]. Emstar [GRE^{+}07]. Emulation [HSSS17, ZGJ^{+}22]. Enable [HWS^{+}20]. Enabled [DSH16, KOD^{+}14, CWK^{+}22, GXQ^{+}22, SUR^{+}23, WWZ^{+}21, SNC^{+}23]. Enabling [CWS^{+}22, DXC^{+}21, HWF^{+}24, LJW^{+}21, MNLZ18, PHKK17, SMS22, SCS22, SSL^{+}22]. Encode [WKYH17]. Encoder [LYF^{+}23]. Encoding [SMS22]. encrypted [CCMT09]. Encryption [FHST22, FSTH23, TCN^{+}17, ZCZL22]. End [MSK^{+}23, YSK^{+}15, YA24, WWLX13]. End-Point [MSK^{+}23]. End-to-end [YA24, WWLX13]. Energy [AMAT^{+}18, AH20, Amm16, BAHS24, BDO14, BASM16, CBSA18, CKHP19, CCC^{+}21, CPL^{+}20, DBOD^{+}16, DML^{+}16, EA15, ECPC14, FLJ^{+}13, FBAG20, GSM^{+}22, HSSS17, HWT^{+}22, JZL^{+}19, JGK^{+}23, JCC^{+}13, KOD^{+}14, KLC^{+}16, KPB^{+}08, KW09, LPV^{+}09, LED20, LLL14, LWY^{+}21, LWM^{+}21, LFW^{+}19, MDM^{+}20, MZKC23, NZLH15, NZM21, PA05, QXZZ22, SPI^{+}24, SPK^{+}10, SDYC22, SCLG24, SNY^{+}24, SDBT19, TCN^{+}17, TJWK13, TBL07, VAC13, WEC11, WLD10, WTX^{+}16, WCV^{+}18, WJ21, XCC^{+}15, XXHL16, XDL^{+}24, YTR^{+}22, YXFL17, YB17, ZLYW19, ZZZ^{+}20, ZLZ21, ZGCL23, ZPL^{+}24, ZWY21, ZMVR14, ABM13, CNMH08, CLH^{+}13, CGD12, FLFW13, GAJ^{+}06, HKL^{+}06, HLTC06, HR13, Kal10, LP08, LDZ13, LFS09, SYL09, SGM08, SS13, Su07, SC12, WBS10, WIF^{+}11, XWZ^{+}05, YPW^{+}13, ZGHZ12, MGS^{+}15]. Energy-Aware [GSM^{+}22, GAJ^{+}06, HR13]. Energy-collision-aware [CCC^{+}21]. Energy-conserving [PA05, HLTC06]. Energy-Delay [DBOD^{+}16]. Energy-Depleting [CPL^{+}20]. Energy-driven [SPK^{+}10]. Energy-Efficient [Amm16, DML^{+}16, EA15, KLC^{+}16, LED20, LWM^{+}21, NZLH15, SDBT19, WTX^{+}16, XXHL16, YB17, AH20, FLJ^{+}13, HWT^{+}22, JCC^{+}13, KPB^{+}08, KW09, LPV^{+}09, LWY^{+}21, QXZZ22, SDYC22, SNY^{+}24, TJWK13, TBL07, WEC11, WLD10, XDL^{+}24, ZLZ21, CNMH08, CLH^{+}13, CGD12, HKL^{+}06, LDZ13, LFS09, WBS10]. Energy-Fairness [LLL14]. Energy-Harvesting [AMAT^{+}18, JZL^{+}19, CCC^{+}21, MDM^{+}20, SCLG24, MGS^{+}15]. Energy-Optimal [BDO14]. Energy-Saving [YXFL17, JGK^{+}23, SGM08]. Enhanced [SJH^{+}18, ZYZ^{+}19, ZZC^{+}23]. Enhancement [GXQ^{+}22]. Enhancements [MLS^{+}22]. Enhancing [BHA^{+}13, LZGX23, PAYL22, WHYC19]. EnHANTs [MGS^{+}15]. Enlargement [PTDD16]. Ensemble [LTZ^{+}24]. ensuring [HTW07]. Entropy [RKRP17]. Entropy-Based [RKRP17]. EnviroMic [LCH^{+}09]. Environment [AKC^{+}18, JYB^{+}21, LFNS14, WTX^{+}16, GRE^{+}07]. Environmental [CTWG24, DD11, Kou18, ACG^{+}13, IBS^{+}10, ORRJ12]. Environments [GM14, GKRW17, HSSS17, MNLZ18, WLX^{+}23, XCT^{+}16, YJL^{+}22, KMS^{+}10, WX08]. epidemic [DLD09]. equal [MPC^{+}10]. equally [NCV10]. Equipment [XDL^{+}24]. Erasure [DML^{+}16]. Erasure-Resilient [VRSR15]. Error [PPM15, SNK^{+}22, VRSR15, AAA06]. error-based [AAA06]. Error/Erasure [VRSR15]. Error/Erasure-Resilient [VRSR15]. Errors [GZZ^{+}14, GHZ^{+}22]. establishment [HM07b]. Estimating [GLQ^{+}22, Kou18]. Estimation [BY19, CLX^{+}21, DSA^{+}20, KYM17, KRP15, SMR^{+}14, WWL15, XWL24, ZGJ^{+}22, BKM^{+}12, CK09, FS13, KQ12, LWSL12, SAZ10, SC12, VMS10, WLW12]. Estimation-Based [KRP15]. Estimator [WZZ^{+}21]. Euclidean [CLS12, KA13]. Evaluating [CZC^{+}24]. Evaluation [ALNT22, DWF^{+}23, LYZ^{+}24, XLO^{+}23, HBC^{+}09, KA13, LPR09, LCH^{+}09, ODCP13, RBD13, SCWC13]. Event [CA22, ES12, IPMGL18, SDBT19, WJZ21, ZHCA17, KPK12]. Event-Triggered [SDBT19]. events [YYM^{+}10]. Every [HCL15]. Everywhere [Kal10]. Evolution [CQDW21, KKRR15, PCR13]. Evolvable [HAH22]. Evolving [GDM22]. Examples [SYT22, XLG^{+}22]. Execution [MDM^{+}20]. Exercise [MNLZ18]. Exergames [COP^{+}16]. Exit [JYB^{+}21]. experience [EML^{+}09]. Experiences [BASM16, CPP^{+}17, LGTL19, OBB^{+}13]. Experimental [BDP24, PG09]. Experimentation [MGS^{+}15]. Exploiting [BNN^{+}20, CWY24, LCH^{+}19b, LWH^{+}22, LCD22, SSL^{+}19, VTY18, WXL^{+}19, ZMXM24]. Exploring [DCD24, MCGZ21, WQH^{+}22]. exponents [VMS10]. exposure [Dji10]. Extending [CWY^{+}15, HKG^{+}19]. Extraction [GZZ^{+}23, PCPK14, ZZH^{+}23].

Face [LHX^{+}21, SUR^{+}23, HBLR05]. Face-Aware [HBLR05]. Factories [LYZ^{+}24]. Facts [LGTL19]. Fading [GM14]. Failure [BCMY22, KBD14]. Fair [LDC^{+}19]. Fairness [LLL14]. Fall [WHQ^{+}23]. False [MY24, CDGC12, ZSJN07]. False-Data [MY24]. Familiarity [PZOZ21]. FAR [HBLR05]. Fast [BLGS19, MZW^{+}19, PKC^{+}18, WCLD23]. Fault [COS19, CHSA18, JTE20, LMP14, LDS^{+}22, NRC^{+}09, NP12]. Fault-Tolerant [LMP14, COS19]. faults [SGG10]. Faulty [GZZ^{+}14]. Feasibility [BAP^{+}17]. Feature [FLCH23, LTDZ22, LYY24]. Features [HLZ^{+}24, LC14a]. Features-based [HLZ^{+}24]. Federated [FHST22, FSTH24, MG24, OXZ^{+}23, SDYC22, WTH^{+}23, YA24, ZWWL23, ZWL^{+}24, ZJZ24b]. FedSuper [ZJZ24b]. Fi [XYJ^{+}23, ZZZ^{+}22]. Fidelity [CTW^{+}15]. Field [LLH22, ZYZ^{+}19, Dji10, MRM09, WLZ13, WLW12, XRH^{+}13, ZW05, ZSG09]. Fields [TJLK14]. Filling [WWL^{+}16, WJD16]. filtering [CDGC12]. Filters [TCB^{+}14]. Finding [CHPP23]. Fine [GYG^{+}23, LDL^{+}24, XXW^{+}24, YYL^{+}23, ZTZX23, ZMXM24, MB16]. Fine-Grained [YYL^{+}23, GYG^{+}23, LDL^{+}24, XXW^{+}24, ZTZX23, ZMXM24, MB16]. Fingerprint [GWS^{+}24]. Fingerprint-based [GWS^{+}24]. Fingerprinting [BRR^{+}18, HLZ^{+}24, JCZ^{+}22, LDGG21, WTC22]. Fingerprints [KK15, LXY^{+}22, LGLD23]. finite [ENPNF13]. Firmware [SNY^{+}24]. FIRST [RFS^{+}19]. Fit [RSK^{+}21]. fitting [LPW^{+}23]. Flash [LLX^{+}14]. Flash-Optimized [LLX^{+}14]. flat [CK13]. Fleet [WCW^{+}23]. Fleet-Oriented [WCW^{+}23]. Flexibility [BSI^{+}15]. Flexible [BGP^{+}23, WYD^{+}22]. Floating [ZLW^{+}24]. Flood [IIPK20]. Flooding [BLGS19, CZMM23]. Floor [WHQ^{+}23]. FLoRa [SNY^{+}24]. Flow [PK19, SZG^{+}15, XQL^{+}24, YHC^{+}24, KPS12]. Flow-Based [SZG^{+}15]. Flow-Time [XQL^{+}24]. FlowerCast [TFL^{+}24]. Flux [SML18]. Flying [CPP^{+}17]. Fog [BIMD19]. Follower [XDM^{+}21]. Following [WPL^{+}16]. Food [PK20]. Footprinting [WJ21]. Footprints [WCV^{+}18, ZZX^{+}20]. Force [EFI^{+}10]. Force-directed [EFI^{+}10]. Forecasting [CTW^{+}15, FWF^{+}23, LL21]. Forests [DPB19]. ForETaxi [WCW^{+}23]. Forged [TDZ^{+}22]. formation [VAC13]. Forward [KKRR15]. Forward-Secure [KKRR15]. Forwarding [Amm16, Den09, LCH^{+}19b, WBS14, HCXT09, LFS09, SGM08]. Framework [Amm16, DBOD^{+}16, FM15, GDM22, HBKP14, HWT^{+}22, LLW^{+}23, LZN19, MY24, NK14, NZLH15, PLW^{+}24, RFS^{+}19, SJH^{+}18, SLC^{+}22, SDYC22, SUZK19, VPB^{+}20, WTH^{+}23, ZLB^{+}23, CA06, CC11, CGD12, GBS08, HZGS05, KBD13, KT11, MS09, SPK14]. Free [LWL^{+}21, LWZ24, Sch15, WHST16, ZLW^{+}15, ZLGL19, GJT^{+}22, HCXT09, SCL^{+}19, SSL^{+}22, TJZ^{+}13, WXG^{+}24, WTC22, WHQ^{+}23, ZW24, ZLGL20]. Free-cooled [LWL^{+}21]. Frequency [BBEM^{+}24, GWS^{+}24, LWCJ14, ACG^{+}13]. Frequency-Based [LWCJ14]. Frequent [WTH^{+}23]. ftTRACK [LMP14]. Full [DSZ^{+}24, SCL^{+}19, WC13]. full-view [WC13]. Fully [XWC^{+}23]. Function [LGXC23]. Fusion [HPS^{+}18, HBKP14, LTDZ22, LWLT24, MCW^{+}16, TXC^{+}13, XWL24, ZW24, ZDW^{+}10, RKW^{+}06, TXY^{+}13]. Fusion-based [TXC^{+}13]. FusionTrack [ZW24]. Future [AAJ^{+}23, AMTH^{+}17, RKW^{+}06]. Fuzzy [YRB^{+}17].

Gains [IPMGL18]. Gait [XYW^{+}22, XJR^{+}17, ZZZ^{+}22, XJR^{+}17]. Gait-Based [XJR^{+}17, ZZZ^{+}22]. Gait-Key [XJR^{+}17]. GaitSense [ZZZ^{+}22]. GaitTracker [XYW^{+}22]. Game [CPL^{+}20, DSH16, DBOD^{+}16, LLH22, YMY^{+}23, YHC^{+}24, ABM13, VAC13, YLL13]. Game-Theoretic [CPL^{+}20, VAC13]. Garment [LPW^{+}23]. Gated [FLCH23]. Gathering [EA15, HCL15, YMY^{+}23, Amm13, CGD12, GCBL06, GNDC08, Kal10, WLD10]. Gauss [KLC13]. Gaussian [ORRJ12, WZZ^{+}21]. General [LZN19, CLX09]. Generalized [WL23]. Generate [KVS23]. Generation [LWH^{+}22, PKC^{+}18, WXG^{+}24, XJR^{+}17, ELYR14]. Generative [XLG^{+}22]. Generic [LZZ^{+}15, ZHL^{+}15, ZWW^{+}23]. Genus [WJD16]. Geographic [LFL^{+}19, WS14, ZSKH08]. Geographical [LYF^{+}23]. Geomagnetic [WTC22]. Geomagnetism [WMT^{+}19]. geometric [ABM06, NEKK12]. Geometry [Amm23, XRS10]. Geometry-based [Amm23]. Geospatial [KRP15]. Gesture [XYJ^{+}23, YXG^{+}19]. GHz [SMS22, SCS22].

GINSENG [OBB^{+}13]. Global

[QNN^{+}22, ZWW^{+}23]. Go

[GCAK17, SYOY12]. goals [LHRM09]. Gossip [SZG11]. GPART [ZWW^{+}23]. GPFS [LL21]. GPIO [JZX^{+}20]. GPS [CT19, FSSR15, GPL^{+}12, JCC^{+}13]. gradient [HCXT09]. gradient-based [HCXT09]. Grained [MB16, YYL^{+}23, GYG^{+}23, LDL^{+}24, XXW^{+}24, ZTZX23, ZMXM24]. Graph [DTW^{+}23, LDDL24, LL21, WYY^{+}19, ELYR14, NEKK12, ZBA07]. Graph-based [LL21, WYY^{+}19]. Graphical [WZZ^{+}21]. Graphs [CHPP23, ZWW^{+}23, FKMS06]. Grayspaces [BAP^{+}17]. greedy [KT11]. Green [SBSD18]. Greenifying [ABC^{+}18]. GreenLocs [NZLH15]. Greentooth [BAHS24]. Grid [LDS^{+}22, VTY18, WWZ^{+}21, RR09]. grid-group [RR09]. Grids [KKP18, MY24]. Group [LND08, MLX^{+}24, CLS12, MPS10, RR09]. Group-based [LND08]. grouping [RKJ09]. Growth [LWLT24]. Guarantee [SCL^{+}19]. Guaranteed [WS14]. guaranteeing [CLX09]. guarantees [WWLX13]. Guidance [GZK^{+}23]. guided [BJW^{+}22].

H [CRZ^{+}20]. H-DrunkWalk [CRZ^{+}20]. Hand [CLJ^{+}23]. Hand-dependent [CLJ^{+}23]. handover [ELYR14]. Handovers [JYB^{+}21]. HAR [ZWG24]. Harmonium [PKC^{+}18]. Harmony [YMY^{+}23]. Harvesting [AMAT^{+}18, BASM16, FBAG20, HSSS17, JZL^{+}19, YTR^{+}22, ZZZ^{+}20, ZPL^{+}24, ZWY21, CCC^{+}21, MDM^{+}20, SCLG24, MGS^{+}15]. Hazards [PDP^{+}17]. HDACS [XAKV15]. healing [PMST12]. Health [BWCW14, DBC^{+}24]. Healthcare [AAJ^{+}23, GZZ^{+}23, SUR^{+}23, SMW23, SNC^{+}23]. Heart [CLX^{+}21]. Heartbeat [KAH^{+}10]. Heat [SZX17]. Heterogeneity [ZZZ^{+}20, ZWL^{+}24, Amm13]. Heterogeneous [CRZ^{+}20, ELR^{+}22, LWY^{+}21, LFW^{+}19, MG24, SGB15, SWYW21, TYGW15, BCL^{+}12, GRE^{+}07, LP06, LPR09, LSW06, RKJ09]. Heterogeneous-device [SWYW21]. Hidden [MCGZ21, LCC^{+}13]. Hierarchical [ALS23, FSTH23, FLCH23, SZG11, XAKV15, IV12, LDZ13]. High [CTW^{+}15, KKP18, MNLZ18, PDP^{+}17, PCPK14, RKRP17, WJD16, XDL^{+}24, YSK^{+}15, ACG^{+}13, GBS08]. High- [RKRP17]. High-capacity [XDL^{+}24]. High-End [YSK^{+}15]. High-Fidelity [CTW^{+}15]. high-frequency [ACG^{+}13]. High-Level [PDP^{+}17]. High-Mobility [MNLZ18]. High-Rate [PCPK14]. Histograms [CG18]. Hoc [CS17, CS18, JYC^{+}24, VDV16, CVY09, DRC06, KPK12, LYG^{+}13, NJS05, PR10, SZ19, SS13]. Hoc-based [JYC^{+}24]. Holistic [DCD24, DLG^{+}21, LCC^{+}17, SPI^{+}24]. Home [HPS^{+}18, LL21, LDL^{+}24, LSW14]. homogeneous [MPS10]. Homomorphic [FHST22]. Hop [DGS16, GTL19, NEKK12, WXD^{+}23, ZSLL23, ZSJN07]. hop-by-hop [ZSJN07]. hop-count-based [NEKK12]. Hopping [TNBG18, WZLM21]. HP [LYZ^{+}24]. Human [Hau14, LL21, LPW^{+}23, OXZ^{+}23, WNM^{+}24, YXFL17, ZZZ^{+}22, ZZY^{+}23, ZHJ^{+}20, YSM08]. human-centric [YSM08]. Human-related [ZHJ^{+}20]. humans [GJNC^{+}14]. hUmidity [WWB^{+}19]. Hunting [XWW^{+}23]. Hunting-style [XWW^{+}23]. HVAC [ABC^{+}18]. Hybrid [AKSM15, MSK^{+}23, MKFD^{+}23, PSR^{+}22, ZLYW19, ES12, HBC^{+}09, PFJ13]. hygrometer [PKS^{+}23]. Hypergraph [WJY^{+}24]. Hypergraph-based [WJY^{+}24]. Hypothesis [BWP^{+}24, AAA06]. Hypothesis-Based [BWP^{+}24].

i-Sample [ZWG24]. ID [CYD^{+}24, FHST22, FSTH23]. ID-Aware [CYD^{+}24]. ID-Based [FHST22, FSTH23].

IdealVolting [KBW16]. Identification [CWS^{+}22, CRY^{+}10, GWS^{+}24, HPS^{+}18, HZX^{+}24, HSL^{+}15, KGBS18, NZLH15, PWS^{+}23, SDW^{+}23, WLW^{+}20, WWZ24, YYL^{+}23, ZZZ^{+}22, ZHJ^{+}20]. Identifying [CJL^{+}20]. iDiary [FSSR15]. IEEE [BAP^{+}17, PEFSV13, PFJ13, RDR07, TDD^{+}19]. IIoT [HWT^{+}22, QWC^{+}22]. Image [LLZ^{+}22, NLH^{+}19, XWL24]. Image-based [NLH^{+}19]. imagers [KAH^{+}10]. Images [CC23, LDGG21, XXW^{+}24]. IMeP [ZZC^{+}23]. IMF [XWC^{+}23]. Impact [Amm13, BBEM^{+}24, MCLW23, NCV10, PKG08]. Impedance [ZZC^{+}23]. Implementation [CSLJ23, XTXW22, GAJ^{+}06, LCH^{+}09, TBL07]. Implementing [MWS08]. Improve [KSR^{+}20]. Improved [RS19, SS13, YTZ^{+}23, ZMXM24, FKMS06]. improvement [ZJZ12]. Improving [DTY^{+}22, KCPC13, LN05, MDC17, SJP^{+}22]. Imputation [CTWG24]. In-Air [YXG^{+}19]. In-Band [CSLJ23, ZZW^{+}23b]. In-Bed [AJH^{+}20]. In-Depth [GLL^{+}24]. In-Network [BJR15, ELR08, KBD13]. In-situ [WLW12, WWL15]. Inaudible [LWH^{+}22]. Incentive [LLZ^{+}20, RDP16, YCL^{+}19, ZZ21, ZZ23]. Incidents [MSB17]. Incremental [PPM15, PBM11]. independent [WHQ^{+}23]. Indexing [LLX^{+}14, HZGS05]. Individual [MSK^{+}23]. Indoor [KVS23, LZZ^{+}15, LJW^{+}21, NZLH15, NLH^{+}19, PKC^{+}18, TAT14, TGG^{+}17, TGG^{+}19, WMT^{+}19, XCT^{+}16, XDM^{+}21]. Indoor-Outdoor [TGG^{+}17]. Indoor/Outdoor [LZZ^{+}15]. Induction [JCZ^{+}22]. Industrial [CS23, HLZ^{+}24, LYZ^{+}24, ZSLL23]. inequality [YJWL13]. inertia [YPW^{+}13]. Inertial [MNLZ18, XYW^{+}22]. Inexpensive [RHS20]. Inference [BWP^{+}24, DLD^{+}23, JYB^{+}21, SUZK19]. InferLoc [BWP^{+}24]. Inferring [SZX17].

Information [CDGC12, DTY^{+}22, GLQ^{+}22, HLN^{+}11, LTZ^{+}24, RGB^{+}17, RFS^{+}19, SCLG24, YMY^{+}23, BKS13, BGJ09, KVI^{+}13, MS09, ORRJ12, SSGM10, Su07]. information-seeking [KVI^{+}13]. Information-theoretic [CDGC12]. informative [KGGK11]. Infrastructure [COS19, MWS08]. Infrastructures [CWK^{+}22, GXQ^{+}22]. Ingestion [ZZM^{+}22]. initialization [LYG^{+}13]. initiated [DDHC^{+}12]. Injection [MY24, ZSJN07]. InPhase [SW22]. input [FLCH23]. insertion [XWDN12]. Insider [HLZ^{+}24]. Inspired [HL17]. Inspiring [YMY^{+}23]. Instant [ZZG^{+}24]. instantiation [ZCLJ14]. Insulation [SZX17]. Integrated [WLLZ24, XWZ^{+}05, YHC^{+}24, HKL^{+}06]. Integrity [IPMGL18, MKFD^{+}23, WRYL11, GBS08]. Intelligence [LCF^{+}22, MGN22, QXZZ22, XYJ^{+}23]. Intelligent [GZZ^{+}23, HL17, SPI^{+}24, SWYW21, ZZM^{+}22, ZPL^{+}24, ZDS^{+}21]. Intensity [CLJ^{+}23, XCT^{+}16]. Intensity-Based [XCT^{+}16]. Interaction [CYD^{+}24, PHKK17, SSC^{+}10]. Interactions [CJL^{+}20, SDX^{+}20]. Interactive [COP^{+}16, KLA^{+}14]. Intercepting [BH21]. Interference [BBEM^{+}24, MSAJ18, TNBG18, BNG12, XTZ08, ZCLJ14]. Interference-Aware [TNBG18]. Interleaved [ZSJN07]. Intermittent [MDM^{+}20]. Internet [AAJ^{+}23, BJW^{+}22, CQDW21, CPSS23, JGK^{+}23, LLW^{+}23, MDB^{+}23, MGS^{+}19, SMW23, SLS^{+}22, YTR^{+}22, YMY^{+}23, YTZ^{+}23, ZZW^{+}23a, ZLYW19, ZDS^{+}21]. interpolation [LS10]. Interpretable [TCC^{+}23]. interrelational [RKJ09]. Interval [SBK22]. Intervals [ZGX^{+}16]. Introduction [CPSS23, CWK^{+}22, LWKZ22, MGN22, NJZ18, QXZZ22, SMW23, Zha05]. Intrusion [LDL^{+}24]. Intrusive [NZM21, WNM^{+}24, ZYC^{+}23]. Inverted [ABC^{+}18]. Involution [YMY^{+}23]. Involved [ZWWZ20]. IODetector [LZZ^{+}15]. IoMT [ZLB^{+}23]. IONavi [TGG^{+}17]. IoT [ÁKSW22, CZX^{+}22, DTY^{+}22, DLG^{+}21, DTW^{+}23, FSTH24, GDM22, GZZ^{+}23, HBW^{+}18, KCE^{+}20, KGDC22, LDG^{+}21, LLX^{+}22, LZGX23, LCH^{+}19b, LCM21, MSK^{+}23, SBCF20, SUR^{+}23, SNC^{+}23, SHWW20, SWH^{+}24, SWYW21, TDZ^{+}22, WXL^{+}19, WWZ^{+}21, WZZ^{+}21, WTH^{+}23, XJL^{+}23, XZL^{+}20, YBY^{+}24, YJL^{+}22, YYC^{+}19, ZPL^{+}24]. IoT-based [GZZ^{+}23]. IoT-Empowered [KCE^{+}20]. IoT-Enabled [SNC^{+}23, SUR^{+}23, WWZ^{+}21]. IoV [XFZ^{+}21]. IR [TAT14]. Irregular [WWZ24, CK13]. Irregularity [MLZ^{+}24, ZHKS06]. Irrigation [WWB^{+}19, WCPC20]. iSelf [SMZ^{+}17]. iSleep [CPX^{+}20]. Issue [LWKZ22, LSX24, MGN22, NJZ18, SMW23]. Item [QWC^{+}22]. Itemsets [WTH^{+}23].

Jamming [CD21, CPL^{+}20, HXZ23b, TDD^{+}19, LPV^{+}09, SDČ10]. Joint [Amm13, BWP^{+}24, BY19, KSR^{+}20, KPCB20, TCN^{+}17, TZZ22, WLW^{+}23]. JVM [RS19].

Kamada [CS17]. Kawai [CS17]. kernel [NJS05]. kernel-based [NJS05]. Key [KKRR15, LWH^{+}22, MPS10, MLX^{+}24, PCPK14, RR09, WXG^{+}24, XJR^{+}17, YLSZ19, ZZH^{+}23, HM07b, LYG^{+}13, LN05, LND08, MWS08, TP07, WDLN09, XJR^{+}17]. knowledge [LN05]. Known [LGTL19].

Labeling [NZH^{+}23, SMZ^{+}17]. labelling [ZZY^{+}23]. Landmark [NZH^{+}23]. Lane [BNPR20]. LaPS [DPB19]. Large [LGTL19, LXR^{+}16, MCGZ21, MYH^{+}24, NJL24, SBK22, SSL^{+}22, TJLK14, VRSR15, WCW^{+}23, WS14, ZHZ^{+}16, CJS11, CDR08, HBLR05, HM07b, KSMH13, KPB^{+}08, LWG09, MB09, PCR13, PH10, TJZ^{+}13, ZH05, ZSJ06]. Large-Scale [LXR^{+}16, SBK22, TJLK14, VRSR15, WCW^{+}23, WS14, ZHZ^{+}16, LGTL19, MCGZ21, MYH^{+}24, NJL24, SSL^{+}22, CDR08, HBLR05, HM07b, KSMH13, KPB^{+}08, LWG09, MB09, PCR13, PH10, TJZ^{+}13, ZSJ06]. Latency [BYD^{+}15, CCC^{+}21, PNL^{+}22, SDBT19, XCC^{+}15, ZLGL20, LP08, WRS10]. Latency-efficient [ZLGL20]. Latent [LWY^{+}21]. Layer [BBEM^{+}24, KPRH14, LCM21, DDHC^{+}12, HWT^{+}11, LPV^{+}09, LFS09, LCD22]. Layers [KPRH14]. LeaD [ZDS^{+}21]. Leader [XDM^{+}21]. Leader-Follower [XDM^{+}21]. Leakage [PK19]. Leaked [LHHW24]. LEAP [ZSJ06]. Learn [ZDS^{+}21]. Learning [ALS23, BT18, CLX^{+}21, CQDW21, CPL^{+}20, DCD24, FHST22, FSTH24, FBAG20, GAMW22, JGK^{+}23, Kun22, LWL^{+}21, LL21, LYY24, LXYT24, LWY^{+}21, LC14b, LWX^{+}21, MLZ^{+}24, MLX^{+}24, MDB^{+}23, MY24, MG24, NJL24, OXZ^{+}23, RKLM23, SDYC22, SMZ^{+}17, WLZ23, WLLZ24, WTH^{+}23, Yan22, YA24, ZZ21, ZWWL23, ZZY^{+}23, ZWL^{+}24, ZSZ20, ZJZ24b, NJS05]. Learning-Based [GAMW22, LWX^{+}21, WLLZ24, LXYT24, ZZ21]. Least [SZZC08]. Leds [TAT14]. length [QM13]. Lesion [GZZ^{+}23]. LesionTalk [GZZ^{+}23]. less [YHC^{+}24]. Level [PDP^{+}17, VDV16, ZGJ^{+}22, CT19, CRY^{+}10, CK13, TXY^{+}13, KBD13]. Levels [SZX17, ZLD^{+}24]. Leveraging [BIMD19, CLL^{+}23, Hau14, LS10, WLLZ24, YS07]. Lexicographic [YM14]. LiDAR [DPB19]. LiDAR-assisted [DPB19]. Lifelogging [JLZL19]. Lifetime [QNN^{+}22, RD16, SCL^{+}14, ZSLL23, DD09, IR12, JTS09, LHRM09, LKA10, WRS10, YLL13, ZH05]. lifetime-maximized [YLL13]. LiFi [ZMXM24]. Light [CLJ^{+}23, XCT^{+}16]. Lighting [KCE^{+}20]. Lightweight [SC15, WS14]. like [AH20]. likelihood [WKA14]. Limit [YYXL22, ZCZ^{+}23]. Limited [LTZ^{+}24]. Limits [LCH^{+}20]. Linear [JAC19, PWS^{+}23]. Link [LC14b, MB16, PS17, TFL^{+}24, ZGJ^{+}22, BKM^{+}12, DDHC^{+}12, KCPC13, LPV^{+}09, LC14a, SAZ10]. link-layer [LPV^{+}09]. Links [CD21, CWY24, PS17, WKYH17, ZK07, ZSKH08]. LIPAuth [CLJ^{+}23]. LIPS [XCT^{+}16]. Liquid [SDW^{+}23, SDW^{+}23]. Liquidity [MYH^{+}24]. Listening [LCJ^{+}23]. LMAC [GLG^{+}23]. LMS [PPM15]. Load [KKP18, NZM21, ZZC^{+}23, LKA10]. Local [LTZ^{+}24, BGJ09]. Localisation [BCMY22]. Localizability [PWS^{+}23]. Localization [AHK16, BWP^{+}24, BGJ09, CWY24, EY14, GYNY16, KVI^{+}13, LXY^{+}22, LXYT24, LDGG21, NLH^{+}19, PKC^{+}18, PWS^{+}23, RHS20, SNK^{+}22, SW22, SLC^{+}22, WMT^{+}19, ZLW^{+}15, ZCZ^{+}23, ZBA07, dOEC^{+}23, BLWY06, CKL^{+}09, CVY09, CPH06, CLS12, EFI^{+}10, JR08, JCC^{+}13, KQ14, KMS^{+}10, LP05, LWG09, LK09, LH09, NEKK12, NJS05, PG09, TJZ^{+}13, WX08, XBWX13, XRS10, YJWL13, ZLGG10, ZGT11]. Localized [LSW06, MS12, PR10, PKS^{+}23]. Localizing [ALY^{+}23, CT19, SCG^{+}15, ZYZ^{+}19, ST12]. Locate [LXYT24]. Locating [GPL^{+}12]. Location [LYL^{+}24, PZOZ21, Sch15, TAT14, TYGW15, YQLD22, ZLD^{+}24, GSL10, SSGM10]. Location-based [YQLD22]. Location-Free [Sch15]. Locations [LSW14, KGGK11]. logical [CA06]. Logistics [NXW^{+}22]. Long [ÁKSW22, Pha16, XDX^{+}14, VHC^{+}09, ZGHZ12]. Long-Range [Pha16, ÁKSW22]. Long-Term [XDX^{+}14, VHC^{+}09, ZGHZ12]. longitudinal [KPS12]. Loose [LPW^{+}23]. Loose-fitting [LPW^{+}23]. LoRa [ÁKSW22, GLG^{+}23, HXZ23a, HXZ23b, LGTL19, LDGG21, MYW^{+}24, SMS22, SCS22, SYL^{+}22, SNY^{+}24, WZZ^{+}23, XHZG22, XTXW22, ZLW^{+}24]. LoRaWAN [GJT^{+}22, HAH22]. Loss [MB16, CK13].

Lossless [LL16]. Lossy [HSD16, KPCB20, LL16, ZMVR14, ZSKH08]. Low [ALNT22, BYD^{+}15, BLGS19, CWS^{+}22, CT19, CML^{+}21, DRW^{+}14, DRC17, GLS^{+}14, GJNC^{+}14, HSD16, KPCB20, LWKZ22, LFL^{+}19, LCH^{+}20, LCJ^{+}23, LCD22, MB09, MYW^{+}24, ME21, PKS^{+}23, RKRP17, RHS20, SBK22, SDBT19, TAT14, WZLM21, WQH^{+}22, WS14, XWW^{+}20, XCC^{+}15, CHN^{+}13, CRY^{+}10, DDHC^{+}12, IV12, LM10a, LM10b, MDC^{+}09, ODCP13, PH10, SDTL10, ZK07]. low-bandwidth [CHN^{+}13]. Low-complexity [GJNC^{+}14, MB09]. Low-Cost [CWS^{+}22, CML^{+}21, LFL^{+}19, TAT14, ALNT22, PKS^{+}23, ODCP13]. Low-Duty-Cycle [XCC^{+}15]. Low-Latency [BYD^{+}15]. Low-level [CT19, CRY^{+}10]. Low-Power [BLGS19, DRW^{+}14, DRC17, HSD16, KPCB20, SBK22, XWW^{+}20, LCJ^{+}23, LCD22, ME21, RHS20, WZLM21, WQH^{+}22, DDHC^{+}12, IV12, ODCP13, PH10, SDTL10, ZK07]. Low-Precision [RKRP17]. Low-Stretch-Guaranteed [WS14]. Lower [KPRH14]. LP [GSM^{+}22]. LR [LED20]. LR-WPANs [LED20]. LSAB [PAYL22]. LT [JJ15].

MAC [DBOD^{+}16, DDHC^{+}12, GCRB12, GAMW22, HF17, LM10a, LM10b, LPV^{+}09, LFS09, LHX16, NGBB14, QM13, RDR07, SC15, YH13]. Machine [HCL15, Yan22, ZSZ20]. Machine-Learning [Yan22]. Machine-to-Machine [HCL15]. macroscopic [KLC13]. Magnetic [JCZ^{+}22, LHHW24, ZZW^{+}23b, ZZC^{+}23]. Maintaining [LXR^{+}16]. Maintenance [CHSA18, HBW^{+}18, SB16, TBL07]. Malicious [ARWK19, WWZ^{+}21]. Malware [ZLB^{+}23]. Management [ECPC14, KOD^{+}14, LCH^{+}19a, SBCF20, TAT14, ZLYW19, ZHJ^{+}20, JLYG13, LYG^{+}13, NDM^{+}13, WECC07]. Managing [PCR13, SHY13]. Maneuver [LYF^{+}23]. Manipulation [SBCF20]. Map [LSW14]. Mapping [LCC^{+}13, MZKC23, EML^{+}09]. Maps [KVS23]. Marked [YZZD23]. Markov [KCPC13]. Massive [BY19]. Matching [ZZC^{+}23]. Material [SYX^{+}23]. Matrices [YB17]. MAV [CRZ^{+}20]. Max [YM14, YSM08]. Max-Min [YM14]. Maximal [ZWW^{+}23]. Maximization [QNN^{+}22]. maximized [YLL13]. Maximizing [ZGX^{+}16, IR12]. Maximum [DSZ^{+}24, RKRP17, SCL^{+}14, WKA14, NP12]. MC [XDX^{+}14]. MCI [GZK^{+}23]. MCRT [WWFX11]. MDF [Den09]. Mean [LLH22]. Measure [LJLW19, IR12]. Measurement [BNN^{+}20, CZX^{+}22, DXL^{+}15, GCAK17, LGTL19, WWL15, XYW^{+}22]. Measurement-Based [CZX^{+}22]. Measurements [SUZK19, YJWL13]. Measuring [CLX09]. MEC [YTZ^{+}23, ZWWL23]. Mechanism [XLO^{+}23, YCL^{+}19, ZZ21, ZZ23]. Mechanisms [BIST18, LLZ^{+}20, RDP16, SZX17, ZSJ06]. Medical [JGK^{+}23, SMW23, NDM^{+}13]. medium [Gel07]. meeting [LHRM09]. Memento [JLZL19]. Mental [ALS23]. Mesh [BDP24, YYC^{+}19]. Metaheuristics [PSR^{+}22]. Method [FLCH23, GYNY16, MLZ^{+}24, WL23, WLZ23, AAA06, XRS10]. Methods [ZZZ^{+}20, CDR08, KKP^{+}07, SGG10]. metric [DRC06]. Metrics [RFB^{+}14, ZLB^{+}23, SS13]. mice [GSW09]. micro [JC12]. micro-solar [JC12]. Microgrids [MKFD^{+}23]. Microphone [ZJZ^{+}24a]. Middleware [ZYZ^{+}19]. Milestones [YYC^{+}19]. Millimeter [BY19, YPZ^{+}17, ZCZ^{+}23]. Millimeter-wave [ZCZ^{+}23]. MIMO [BY19, KGDC22, NK14, YYXL22, ZZW^{+}23b, ZZC^{+}23]. Min [YM14]. mine [LL09]. Minimal [COS19, GLQ^{+}22, WTX^{+}23]. Minimalistic [CPP^{+}17]. Minimization [SNK^{+}22, XQL^{+}24, ZLX^{+}24]. Minimizing [PNL^{+}22]. Minimum [CCC^{+}21, WWXY13, XLZ^{+}07, XCC^{+}15, ZHT^{+}23, Dji10, FKMS06, Kal10]. Mining [WWZ^{+}21, WTH^{+}23, KLA^{+}14]. Miscontrol [PTDD16]. Miss [HXZ23a]. Missing [WLW^{+}20]. mission [EMBP12, RJL^{+}10]. mission-oriented [EMBP12]. Mitigate [SE23]. Mitigating [NLD08]. Mitigation [CD21, HAH22, MSAJ18]. Mixed [Lam15]. Mixing [KKRR15]. mmWave [HZX^{+}24, JYC^{+}24, WCLD23, WNM^{+}24]. mobicast [HBLR05]. Mobile [AHK16, CYD^{+}24, CGB^{+}19, CS17, DRC17, DDA11, GSGA23, HMG^{+}24, JYB^{+}21, KCE^{+}20, KJD^{+}23, Kou18, LLZ^{+}22, LLX^{+}22, LGXC23, LTL^{+}24, LXR^{+}16, LWX^{+}21, MKM^{+}20, MLS^{+}22, PLW^{+}24, RD16, RGB^{+}17, RFS^{+}19, SML18, SZG^{+}15, TZZ22, TGG^{+}17, VDV16, WPL^{+}16, WYY^{+}19, WTX^{+}23, WLZ23, WJY^{+}24, WHST16, XLO^{+}23, XWW^{+}20, XZL^{+}20, XQL^{+}24, YWD^{+}21, YZZD23, ZHL^{+}15, ZZ21, ZZ23, ZLX^{+}24, ZYL^{+}24, ZLL^{+}22, dOEC^{+}23, Bra07, CSA06, EML^{+}09, FLFW13, KKP^{+}07, KNSM14, KAS^{+}10, LCC^{+}13, RMB^{+}10, SZZC08, WRS10, WLZ13]. Mobility [Hau14, MNLZ18, NGBB14, ZWWZ20, Amm13]. Mode [MSK^{+}23, XDM^{+}21]. Model [GZK^{+}23, LWLT24, LYST23, MZW^{+}19, MG24, RBS16, SLC^{+}22, TCC^{+}23, XLG^{+}22, YXG^{+}19, ZWWZ20, DIE14, Gel07, KT11, KLC13, KA13, MS09, TP07, ZCLJ14]. model-derived [KLC13]. Model-driven [SLC^{+}22]. Modeling [DRW^{+}14, ECPC14, JP06, KGBS18, PFJ13, PS17, RRA22, WRS10, ZZW^{+}23a, BJW^{+}22, CDGC12, CK13, DLD09, KA13, NP12, SYOY12, WWB^{+}19]. Modelling [KSR^{+}20]. Models [ALNT22, DD11, WZZ^{+}21, ZHKS06, ZWG24, Bra07, KCPC13, NEKK12, SG08, JTS09].

Modern [IHGS15]. Modes

[KJP^{+}15, RMB^{+}10]. Modulation [SBK22]. Modules [JCZ^{+}22]. Moisture [WWL15, WLW12]. Monitor [BCMY22, GYG^{+}23]. Monitoring [AMTH^{+}17, BWCW14, BGP^{+}23, COS19, CPX^{+}20, CTWG24, CML^{+}21, DD11, DBC^{+}24, DML^{+}16, DSZ^{+}24, NZM21, PK19, SZG^{+}15, TPM^{+}17, WTX^{+}16, XDX^{+}14, XXW^{+}24, YPZ^{+}17, ZHCA17, ZZM^{+}22, ACG^{+}13, DEM^{+}12, GSW09, HBC^{+}09, IBS^{+}10, LL09, OBB^{+}13, YYM^{+}10]. Mortar [FPA^{+}20]. Mote [CWY^{+}15]. motifs [dLM14]. Motion [AJH^{+}20, HWF^{+}24, ZW24]. Motions [YXFL17]. Motivating [LLZ^{+}20]. Movement [ZHJ^{+}20, WIF^{+}11]. Moving [DSZ^{+}24, SYT22, WC09, WC12]. MQTT [FSTH23]. MSEva [DWF^{+}23]. MU [YYXL22]. MU-MIMO [YYXL22]. Mules [SG11, KVI^{+}13, SG10]. Multi [CYD^{+}24, ELR^{+}22, FLCH23, GTL19, GZZ^{+}23, HLZ^{+}24, HKW^{+}24, LLX^{+}22, LTZ^{+}24, RSK^{+}21, SZ19, WSC^{+}23, WZZ^{+}21, WXD^{+}23, XZL^{+}20, YWD^{+}21, YYL^{+}23, ZSLL23, ZZG^{+}24, MCT14]. multi-camera [MCT14]. Multi-Class [LTZ^{+}24, GZZ^{+}23]. Multi-Hop [GTL19, WXD^{+}23, ZSLL23]. Multi-input [FLCH23]. Multi-Node [LTZ^{+}24, YWD^{+}21]. Multi-Object [YYL^{+}23]. Multi-Parameter [ELR^{+}22]. Multi-physical [HLZ^{+}24]. Multi-scale [RSK^{+}21]. Multi-Sensor [SZ19, ZZG^{+}24]. Multi-sharding [HKW^{+}24]. Multi-source [LLX^{+}22]. Multi-task [WZZ^{+}21]. Multi-Tier [XZL^{+}20]. Multi-User [CYD^{+}24, WSC^{+}23]. Multicamera [dLM14, GJNC^{+}14]. Multicast [LFW^{+}19, TFL^{+}24]. Multichannel [WWFX11, WLS^{+}16, GCRB12]. Multichannels [MDC17]. Multicluster [LCH^{+}19a]. multicriteria [SS13]. multidimensional [CPH06]. multigroup [HM07b]. multihop [ADF12, Gel07, KW09, PDMJ10, VMS10, Den09]. Multihop/Direct [Den09]. Multilevel [LZAH^{+}15, KCPC13]. Multimedia [GAMW22, DIE14]. Multimodal [LYY24, LWLT24, ZZY^{+}23, ZZPW23]. Multimode [XDX^{+}14]. multiobjective [WC12]. Multipath [HSD16, SHY13, YH13]. Multiple [BWCW14, BQB^{+}11, GLG^{+}23, KJP^{+}15, LXR^{+}16, MCW^{+}16, SHWW20, SDZZ24, SKM^{+}11, WTX^{+}23, EGG13, PFJ13]. Multiple-Target [SKM^{+}11]. Multiplication [Yan22]. multiquery [ZKS10]. Multireceiver [FHST22]. Multiresolution [SZG11]. multiroot [ZKS10]. MultiSense [ZZY^{+}23]. Multisensor [KCE^{+}20]. Multiswimmer [COP^{+}16]. Multitask [HBKP14]. Muscle [MNLZ18]. Musculoskeletal [DWF^{+}23]. MuSiC [GZJE23]. MuSiC-Based [GZJE23]. Mutual [CWS^{+}22]. MyoVibe [MNLZ18].

Nanosensor [ZHCA17]. NAS [Kun22]. Natural [LTY18]. Navigate [DXC^{+}21]. Navigation [CRZ^{+}20, LR05, TGG^{+}17, TGG^{+}19, XDM^{+}21, KAS^{+}10]. NB [CZX^{+}22, YJL^{+}22]. NB-IoT [CZX^{+}22, YJL^{+}22]. Near [BCMY22, JKK08, LKA10, SB16]. Near-lifetime-optimal [LKA10]. Near-Optimal [SB16, JKK08]. Necessary [WKYH17]. Neighbor [ZHL^{+}15]. Neighborhood [JM16]. Neighbour [HSD16]. Neighbour-Disjoint [HSD16]. nest [KAH^{+}10]. Net [KKP18]. Net-Load [KKP18]. Nets [SCD^{+}24]. Network [BJR15, BH21, BASM16, BGP^{+}23, BQB^{+}11, CC23, CS17, DRC17, EA15, GZK^{+}23, JTE20, KOD^{+}14, KAAF13, KGDC22, KK15, KJP^{+}15, LCH^{+}19a, LZAH^{+}15, LFL^{+}19, MPRS16, PHKK17, QNN^{+}22, RRA22, Sch15, SSL^{+}22, TPM^{+}17, VPB^{+}20, VDV16, WKYH17, WB17, WZZ^{+}21, WHST16, XFZ^{+}21, XDL^{+}24, YHC^{+}24, ZSLL23, ZZG^{+}24, BLWY06, BNG12, CK09, CSA06, CRY^{+}10, CLS12, DEM^{+}12, ELR08, EGG13, ES12, GAJ^{+}06, HKL^{+}06, HBC^{+}09, HTW07, HR13, IBS^{+}10, KBD13, KT11, KVI^{+}13, KASD09, KNSM14, LP08, LPV^{+}09, LCH^{+}09, MCT14, NJS05, NRC^{+}09, NP12, ORRJ12, TLRE13, TBL07, WZL08, ZLGG10, ZSG09, ZGT11, ZGHZ12]. Network-Level [VDV16]. Networked [DCBL15, GM14, MGS^{+}15, MZKC23, MKK^{+}13, ZCLJ14]. Networking [BAHS24, CBSA18, CKHP19, CQDW21, LCM21, ZPL^{+}24, ZMVR14]. Networks [AAJ^{+}23, AMTH^{+}17, AMAT^{+}18, AKSM15, Amm16, Amm23, AH14, AHK16, BYD^{+}15, BGMP15, BWP^{+}24, BAP^{+}17, BCMY22, BNPR20, BIMD19, BLGS19, BSI^{+}15, BR15, CZMM23, CBSA18, CZX^{+}22, CCC^{+}21, CS23, CS18, DPB19, DRW^{+}14, DDA11, DSH16, DGS16, DTW^{+}23, DBOD^{+}16, DML^{+}16, EA15, EY14, GAMW22, GLS^{+}14, GSGA23, GCAK17, GTL19, GZZ^{+}14, HF17, HMLJ17, HSGW21, HBKP14, Hau14, HSD16, HCL15, HWF^{+}24, HKW^{+}24, IPMGL18, JJ15, JM16, KYM17, KPRH14, KJD^{+}23, KLC^{+}16, KPCB20, KKRR15, KRP15, Lam15, LMP14, LCH^{+}19a, LLL14, LL16, LCC^{+}17, LHZZ20, LWKZ22, LLW^{+}23, LXR^{+}16, LZAH^{+}15, LMZ^{+}16, LWM^{+}21, LWCJ14, LHX16, LCH^{+}19b, LZN19, LFW^{+}19, LCH^{+}20, LCF^{+}22, LCD22, MCGZ21, MB16, MSB17, MLS^{+}22, MGN22, MSAJ18, NGBB14, NK15, NK14, PK19, PCA^{+}23, PPM15, PDP^{+}17, PTDD16, PS17, PNL^{+}22, PSB^{+}14, PSR^{+}22, PCPK14, QNN^{+}22, RFB^{+}14, RBS16, RHD17]. Networks [RHS20, RD16, SNK^{+}22, SSL^{+}19, SBCF20, SBK22, SZG11, SCL^{+}14, SB16, SCL^{+}19, SCLG24, SXD^{+}15, SGB15, SG11, SNY^{+}24, SZG^{+}15, TJLK14, TCN^{+}17, TFL^{+}24, TNBG18, TYGW15, TDD^{+}19, VPB^{+}20, VRSR15, VDV16, WWFX11, WPL^{+}16, WB17, WYY^{+}19, WXL^{+}19, WZLM21, WQH^{+}22, WCW^{+}23, WS14, WBS14, WLS^{+}16, XDX^{+}14, XWW^{+}23, XWC^{+}23, XCC^{+}15, XXHL16, XZL^{+}20, YM14, YRM^{+}24, YTB^{+}14, YB17, YHC^{+}24, ZHCA17, ZZW^{+}23a, ZLW^{+}15, ZHZ^{+}16, ZLZ21, ZTZX23, ZSLL23, ZGCL23, ZLW^{+}24, ZWY21, ZLGL19, ZLGL20, dOEC^{+}23, Amm13, ADF12, BKM^{+}12, BCL^{+}12, BKS13, BHA^{+}13, Bra07, BGJ09, CJS11, CA06, CDGC12, CGVC06, CYS^{+}10, CCMT09, CC11, CLSW12, CNMH08, CLH^{+}13, CHN^{+}13, CRW07, CVY09, CDR08, CGD12, CK13, CPH06, CCJ08, DLD09, Den09, DRC06, DD09, DABNR10, DIE14, ELR08, ENPNF13, ELYR14, EMBP12, FLJ^{+}13, FT06, FLFW13, GCRB12, GSW09, GBS08, GSL10, GRE^{+}07, GFJ^{+}13]. networks [GNDC08, HZGS05, HM07a, HWT^{+}11, HTC^{+}10, HY07, HBLR05, HLTC06, HM07b, HCXT09, IW14, IR12, IV12, JKK08, JC12, JHU^{+}13, JLYG13, JP06, JKS^{+}10, JROH09, Kal10, KBD14, KXTZ09, KKP^{+}07, KC14, KQ12, KQ14, KKK08, KPK12, KLJ12, KAAF13, KLA^{+}14, KRJ09, KSMH13, KPB^{+}08, KW09, KAR^{+}14, KMS^{+}10, KA13, LDH06, LP05, LP06, LPR09, LWG09, LKA10, LR05, LSW06, LL09, LDZ13, LYG^{+}13, LWSL12, LS10, LH09, LCC10, LN05, LWH^{+}06, LND08, LFS09, MZWT10, MB09, MWS08, MS09, MPS10, MDC^{+}09, MP10, MS12, MPC^{+}10, MAG13, NGSA08, NEKK12, NLD08, NC10, ODCP13, PDMJ10, PG10, PGG^{+}10, PBM11, PEFSV13, PG09, PC10, PKG08, PR10, PMST12, PCR13, PA05, PH10, QM13, RBLP09, RKW^{+}06, RBD13, RJL^{+}10, RR09, SYL09, SAZ10, SZG13, SSGM10, SGM08, SPK^{+}10, SCWC13, SH09, SPK14, ST12, SS13]. networks [SST08, SYOY12, SZZC08, SDČ10, Su07, SG08, SG10, SC12, SEZA13, TP07, TJZ^{+}13, TXC^{+}13, TXY^{+}13, TJWK13, TMAP14, TYD^{+}07, VMS10, VG10, VAC13, WECC07, WEC11, WL14, WZL07, WZL08, WDLN09, WBS10, WLD10, WRS10, WC13, WWLX13, WWXY13, XBWX13, XWZ^{+}05, XLZ^{+}07, XWDN12, XTZ08, XRH^{+}13, YSZC13, YS07, YVS07, ZSKH08, ZH05, ZKS10, ZJX10, ZJZ12, ZVPS10, ZHKS06, ZDG09, ZSJ06, ZSJN07, ZDW^{+}10]. Neural [BNPR20, CC23, DTW^{+}23, LHZZ20, LLW^{+}23, LLDZ23]. NLOS [CWY24]. Node [ARWK19, BCMY22, CWY^{+}15, CPP^{+}17, CS18, GSGA23, LTZ^{+}24, MLZ^{+}24, MB16, PWS^{+}23, YSK^{+}15, YJL^{+}22, CVY09, CPH06, DLD09, JTS09, LK09, PX13, YWD^{+}21]. Nodes [ÁKSW22, DTY^{+}22, ELR^{+}22, GZZ^{+}14, KBW16, MCGZ21, HR13, MPS10, SSC^{+}10]. noisy [YJWL13]. Nomadic [XJL^{+}23]. Non [BT18, CS18, DSH16, HZX^{+}24, WNM^{+}24, ZYC^{+}23, KNSM14]. Non-Bayesian [BT18]. Non-Convex [CS18]. Non-Cooperative [DSH16, HZX^{+}24]. Non-intrusive [WNM^{+}24, ZYC^{+}23]. non-overlapping [KNSM14]. Nondeterministic [XLO^{+}23]. nonhomogeneous [MRM09]. Nonlinear [MZW^{+}19, LK09]. Nonlinearities [PPM15, LWSL12]. Nonlinearity [ZJZ^{+}24a]. nonuniform [KC14]. Novel [SBK22, SCD^{+}24, YLSZ19, ZLB^{+}23, CGD12]. Num2vec [FWF^{+}23]. Number [ZHT^{+}23]. Numeric [FWF^{+}23].

O [XWC^{+}23]. Obfuscation [ZLD^{+}24].

Object

[DSZ^{+}24, EGG13, HPS^{+}18, LYY24, LJLW19, XKW^{+}22, YYL^{+}23, ABM06, KASD09]. Objectives [BWCW14]. Objects [BQB^{+}11, NXW^{+}22]. Oblivious [KCE^{+}20]. Observation [BT18]. observations [WKA14]. observer [CSA06]. Obstacle [ZVPS10]. Obstacles [TCB^{+}14, XJL^{+}23, YRM^{+}24]. occlusions [EGG13]. Occupancy [AAHS18, ECPC14]. Occupant [HPS^{+}18]. occurring [LWSL12]. OFDM [KGDC22]. off [FLFW13, WRS10].

Offloading [BJW^{+}22, JGK^{+}23, SHWW20, TZZ22, YTZ^{+}23, ZWWL23]. Offset [BBEM^{+}24]. Oilfield [MYH^{+}24]. Older [ABC^{+}18]. On-demand [DLD^{+}23, KPB^{+}08]. On-Object [HPS^{+}18]. Once [LXYT24]. One [ABC^{+}18, GTL19, RSK^{+}21, SAZ10]. One-Hop [GTL19]. one-way [SAZ10]. Online [CGB^{+}19, IW14, LL21, LC14b, LCLY22, MKM^{+}20, SE23, MCT14]. OPCIO [JZX^{+}20]. Open [FPA^{+}20, WLW^{+}20]. OpenCarrier [YYXL22]. Operation [HKG^{+}19, MSK^{+}23, RFB^{+}14, ZGHZ12]. Opportunistic [CZMM23, GLS^{+}14, HSGW21, LCH^{+}19b, LFL^{+}19, MSAJ18, WYY^{+}19, WBS14]. OPTI [DLD^{+}23]. OPTICS [WCPC20]. Optimal [BGMP15, BDO14, DSH16, HBKP14, JZL^{+}19, JR08, KC14, KYM17, KKP18, LWH^{+}06, MGS^{+}19, SB16, SH09, SZG^{+}15, WC09, WC12, WLW12, WYD^{+}22, YM14, YHC^{+}24, JKK08, Kal10, KPK12, LKA10, SC12, ZW05]. Optimally [LP08]. Optimization [CZX^{+}22, CGB^{+}19, DBOD^{+}16, KPRH14, LCD22, PDP^{+}17, YMY^{+}23, ZZPW23, ZSLL23, ZWL^{+}24, ZZC^{+}23, ABM13, CSA06, PEFSV13]. Optimize [SCLG24]. Optimized [CC23, Lam15, LLX^{+}14, MZKC23, MB09]. OPTimizing [WCPC20, DCBL15, HWT^{+}11, JZX^{+}20, RD16, RFS^{+}19, TLRE13, WIF^{+}11, WXD^{+}23, XCC^{+}15]. Orchards [SCD^{+}24]. Orchestration [LDS^{+}22]. Order [DLD^{+}23, WJZ21]. organized [KSMH13]. organizing [CNMH08]. Oriented [WCW^{+}23, WYD^{+}22, YCL^{+}19, EMBP12, NDM^{+}13, ZGCL23]. Orienteering [SCD^{+}24]. Other [CWS^{+}22]. Our [LJLW19]. Out-of-Band [GTL19]. Out-of-order [WJZ21]. outages [GPL^{+}12]. Outdoor [CML^{+}21, LZZ^{+}15, LDGG21, PKS^{+}23, TGG^{+}17, KMS^{+}10]. outlier [YJWL13]. outliers [XBWX13]. Over-the-air [SNY^{+}24]. overcomplete [JLYG13]. overhearing [JROH09]. Overlapping [WQH^{+}22, KNSM14, WWXY13]. Overload [WECC07]. Own [LSW14].

P2P [MSK^{+}23]. Packages [NXW^{+}22].

Packet

[KLC^{+}16, MYW^{+}24, MB16, WXL^{+}19, Gel07, LFS09, PX13, XWDN12, KBD13]. Packet-Level [KBD13]. Packet-Loss [MB16]. Packets [HXZ23a]. pairwise [HM07b]. Paradigm [LCJ^{+}23]. Parallel [WZZ^{+}23, ZZW^{+}23b]. Parameter [DBOD^{+}16, ELR^{+}22]. Parameters [Kou18, HWT^{+}11]. Paring [ZYL^{+}24]. Parking [RKLM23, ZGH^{+}21]. Parkinson [TCC^{+}23]. Partial [WZL08, WLZ23, CJS11]. Partially [WQH^{+}22]. Participant [CGB^{+}19, WLZ23, YCL^{+}19]. Participants [MG24]. Participatory [RDP16]. Partitioning [LYF^{+}23, TJLK14, ZWW^{+}23, HM07b]. Passive [CWY^{+}15]. Path [DSA^{+}20, MRM09, SCL^{+}14, SG11, CSA06, CK13]. path-constrained [CSA06]. Paths [TCB^{+}14, Dji10]. Patients [GZK^{+}23]. Patterns [CLJ^{+}23, KGBS18, BNG12]. Payload [SMS22]. PC [KPCB20]. PC-RPL [KPCB20]. PCube [XHZG22]. PDA [HLN^{+}11]. PDGes [TCC^{+}23]. Pedometer [WTC22]. Pedometer-free [WTC22]. Penetration [KKP18]. Performance [BBEM^{+}24, BAP^{+}17, KA13, LZAH^{+}15, MDC17, PDP^{+}17, ZMVR14, CKL^{+}09, ODCP13, WZL08]. period [RDR07]. Periodic [HMLJ17, SE23, YYM^{+}10]. periodical [CLSW12]. Perishable [PK20]. Perpetually [LXR^{+}16]. Persistence [SXD^{+}15]. Person [KGBS18]. Personalization [MG24]. Personalization-based [MG24]. Personalized [YA24, ZLD^{+}24]. Perspective [LZAH^{+}15]. Perspectives [MKFD^{+}23]. perturbation [ZGT11]. Phase [SW22]. Phase-based [SW22]. phased [WLZ23]. Phases [MZW^{+}19]. Phenomena [AHK16, TTBH14]. phenomenon [HR13]. Phones [YXFL17, RMB^{+}10]. Photographing [YXFL17]. PHY [HXZ23b, XTXW22]. Physical [BBEM^{+}24, KSR^{+}20, LSX24, SJH^{+}18, SDX^{+}20, WLLZ24, XJL^{+}23, ZGJ^{+}22, HWT^{+}11, HLZ^{+}24, YSM08]. Physical-Assisted [XJL^{+}23]. physical-layer [HWT^{+}11]. Physical-Level [ZGJ^{+}22]. Physics [LYST23]. Physics-directed [LYST23]. Physiological [VG10]. Pigs [DBC^{+}24]. PigSense [DBC^{+}24]. PIP [GCRB12]. Pipelines [PK19, LCC^{+}13]. Pixel [ALY^{+}23]. PLA [KBD13]. Place [NZLH15]. Placement [BCMY22, BWCW14, DPB19, DXL^{+}15, MLZ^{+}24, WYD^{+}22, XZL^{+}20, YRM^{+}24, ZZPW23, GCBL06, JR08, PA05, SH09, WC09, WC12, WLW12]. Placements [ZLX^{+}24, KGGK11]. Placing [LFNS14]. Planar [Amm23]. plane [TDZ^{+}22]. Planes [GTL19]. Planning [HWF^{+}24, PZOZ21, SG11, WLW^{+}23, WIF^{+}11]. Platform [CPP^{+}17, LPW^{+}23, SML18, CHN^{+}13]. Platforms [LLX^{+}14, SWYW21]. Point [MSK^{+}23, TGG^{+}19, XWL24, YZZD23, CRY^{+}10]. Points [LGLD23]. PolarScheduler [ZLW^{+}24]. Policies [BIST18, JKK08]. policy [MS12]. policy-based [MS12]. Portable [FPA^{+}20]. Pose [LL21, XWL24]. POSE.R [HSGW21]. position [CK09]. Positioning [GZJE23, PTDD16, XCT^{+}16, YQLD22]. Positive [CKHP19]. Possible [TCB^{+}14, ZLGG10]. Post [SZ19]. Post-hoc [SZ19]. posteriori [NP12]. potential [XRH^{+}13]. Power [BLGS19, CKHP19, DRW^{+}14, DRC17, GCBL06, HSD16, JZX^{+}20, KLC^{+}16, KPCB20, KR18, LDC^{+}19, LWKZ22, LMZ^{+}16, LCH^{+}20, MGS^{+}19, SSL^{+}19, SBK22, TPM^{+}17, WLW^{+}23, XWW^{+}20, YSK^{+}15, ZZC^{+}23, CSA06, DDHC^{+}12, IV12, JC12, KT11, LCC10, LCJ^{+}23, LCD22, MDC^{+}09, ME21, ODCP13, PH10, RHS20, SSC^{+}10, SDTL10, WWXY13, WZLM21, WQH^{+}22, XLZ^{+}07, ZK07]. power-aware [LCC10]. Power-Based [KLC^{+}16, YSK^{+}15]. Power-Delivered-to-Load [ZZC^{+}23]. Power-efficient [GCBL06]. Power-Positive [CKHP19]. Powered [YM14, ZHCA17, ZLZ21, ZPL^{+}24, RKLM23]. Powerline [LTY18]. PPG [CLX^{+}21]. Practical [CLSW12, GLL^{+}24, SMR^{+}14, YRM^{+}24, ZSZ20, JC12]. Practice [ZWWZ20, KXTZ09]. Pre [FWF^{+}23, WBS14]. Pre-Forwarding [WBS14]. Pre-Training [FWF^{+}23]. Precision [RKRP17]. Predicting [MCLW23]. Prediction [AAHS18, BJR15, ECPC14, FLCH23, HSGW21, JAC19, KSR^{+}20, LWLT24, LC14b, YZZD23, ZZG^{+}24, AAA06, ELR08, ES12, LC14a, SYOY12, LDDL24]. Prediction-based [HSGW21]. predictive [SPK14]. predistribution [HM07b, LN05, LND08, MPS10, RR09, TP07]. Preparation [DLD^{+}23]. Preprocess [LLZ^{+}20]. Presence [GM14, YRB^{+}17, EGG13]. Preservation [SNC^{+}23, YHC^{+}24]. Preserving [HLN^{+}11, MJS^{+}19, SJH^{+}18, SXD^{+}15, CC11, HLTC06, HWF^{+}24, LHX^{+}21, WWZ^{+}21]. Pressure [ZYC^{+}23]. prevalence [SGG10]. Prevention [MSB17]. Price [ZZ21]. Primitive [SC15]. Principal [AH14]. prioritized [DIE14]. Privacy [HLN^{+}11, HLL^{+}23, LZGX23, LHX^{+}21, MJS^{+}19, SJH^{+}18, SNC^{+}23, SDYC22, WWZ^{+}21, WTH^{+}23, YQLD22, YBY^{+}24, YA24, ZLD^{+}24, CYS^{+}10, CC11, KXTZ09, PX13]. Privacy-aware [SDYC22]. Privacy-Preserving [HLN^{+}11, MJS^{+}19, SJH^{+}18, LHX^{+}21, WWZ^{+}21, CC11]. privilege [SZZC08]. Proactive [XJL^{+}23]. Probabilistic [GZK^{+}23, GHZ^{+}22, KGDC22]. probability [SGM08]. probability-based [SGM08]. Probing [NK15]. Problem [GYNY16, WZL07]. problems [CRW07]. Processes [YZZD23, ORRJ12]. Processing [VPB^{+}20, XQL^{+}24, ORRJ12, SPK^{+}10, ZKS10]. Processor [FC18, SSL^{+}22]. Processor-free [SSL^{+}22]. Profit [CGB^{+}19]. Programming [SG08, BLWY06, IR12]. Progressive [Kun22]. Progressively [DVS^{+}14]. projection [LK09]. propagation [WL14]. Properties [GLQ^{+}22, MZWT10]. Property [JLYG13, GPL^{+}12]. proportional [YYM^{+}10]. proportional-share [YYM^{+}10]. Prospect [SBCF20]. Protect [CKHP19]. Protection [FSTH23, YQLD22, Yan22, YBY^{+}24, YA24, WZL07]. Protocol [GAMW22, HF17, KPRH14, KJD^{+}23, LHX16, WS14, XJR^{+}17, YLSZ19, YBY^{+}24, ZSZ20, GFJ^{+}13, HCXT09, LFS09, PDMJ10, PG10, PFJ13, ZCLJ14]. Protocols [MDC17, ME21, NGBB14, HLTC06, HTW07, LM10a, LM10b, LPV^{+}09, LR05, YH13]. Prototyping [MGS^{+}15, LJY^{+}10]. provably [CCMT09]. Provenance [WB17]. providing [LHRM09]. Provision [LGXC23]. Provisioning [LLX^{+}22, LCLY22, SGB15]. Proximity [LJW^{+}21, SKM^{+}11, SMMS09]. Proxy [FHST22, ZCZL22]. Public [BDP24, MWS08, WDLN09]. public-key [MWS08]. Publishing [SJH^{+}18]. Pulse [PKC^{+}18]. purposeful [Amm13]. Push [ZCZ^{+}23]. Pushing [LCH^{+}20]. PV [KKP18].

Q [MLZ^{+}24]. Q-Learning [MLZ^{+}24]. QA [MCLM20]. QA-Share [MCLM20]. QoE [LDG^{+}21]. QoE-aware [LDG^{+}21]. QoS [MCLM20, Pha16, RHD17, RD16, XZL^{+}20].

QoS-Aware [MCLM20, XZL^{+}20]. Quality [AMTH^{+}17, ALNT22, CPX^{+}20, CML^{+}21, DXL^{+}15, LYZ^{+}24, LC14b, MKM^{+}20, PKS^{+}23, RGB^{+}17, RFS^{+}19, SJP^{+}22, SGB15, XXW^{+}24, YYM^{+}10, YCL^{+}19, ZGJ^{+}22, BKM^{+}12, BKS13, CLX09, LHRM09, LC14a, MCT14]. Quality-aware [MKM^{+}20]. Quality-of-Service [SGB15]. Quality-Oriented [YCL^{+}19]. Quantitative [WZLM21]. Quantization [SC12]. Quarantine [ZHT^{+}23]. quasi [NCV10]. quasi-equally [NCV10]. Queec [LDG^{+}21]. Query [CYS^{+}10, FC18, VPB^{+}20].

Radar [HZX^{+}24, RSK^{+}21, ZCZ^{+}23]. Radiated [JCZ^{+}22]. Radiation [LTY18, LDC^{+}19]. radii [ZDG09]. Radio [BKM^{+}12, GWS^{+}24, KAR^{+}14, MLZ^{+}24, MGS^{+}19, WHYC19, ZSLL23, GPL^{+}12, JCC^{+}13, ODCP13, XTZ08, ZHKS06]. Radio-based [WHYC19]. radioactive [CRY^{+}10]. Radios [PHKK17, WCLD23]. Radius [BGMP15, BCL^{+}12]. radon [JLYG13]. Raft [TBS^{+}24]. Rail [MCLW23]. Random [JZL^{+}19, KKRR15, YB17, CGD12, CUdVY13, Gel07, HY07, NEKK12, NZR10, ZW05]. Randomization [SE23]. randomly [LWSL12]. Range [CWY^{+}15, Pha16, WHST16, ZLW^{+}15, ÁKSW22, PR10]. Range-Extending [CWY^{+}15]. Range-Free [WHST16, ZLW^{+}15]. Ranges [FLS^{+}14]. Ranging [CP20, SW22, JCC^{+}13, MKK^{+}13]. Rapid [DLG^{+}21, LJY^{+}10]. RaPTEX [LJY^{+}10]. Rate [CLX^{+}21, JZL^{+}19, PCPK14, YM14, LM10a, LM10b, LWH^{+}06, PG10]. Rate-controlled [PG10]. Rateless [LCD22]. ray [CC23]. RCRT [PG10]. Re [FHST22, ZCZL22]. Re-Encryption [FHST22, ZCZL22]. REACH [CWY^{+}15]. Reactive [CD21, SDČ10]. Read [CWS^{+}22]. Real [BBD^{+}23, BCMY22, DRC17, GKRW17, KPCB20, LJW^{+}21, MZKC23, ORRJ12, WWFX11, WLLZ24, WHYC19, XYJ^{+}23, XRH^{+}13, ZJX10, ZZM^{+}22, ZYC^{+}23, LWH^{+}06, SGG10, SHY13, WWXY13]. Real-Time [DRC17, MZKC23, WWFX11, XYJ^{+}23, BBD^{+}23, BCMY22, LJW^{+}21, ORRJ12, WLLZ24, XRH^{+}13, ZJX10, ZZM^{+}22, ZYC^{+}23, LWH^{+}06, WWXY13]. Real-World [GKRW17, SGG10]. Realistic [HSSS17, SAK^{+}19]. Reality [CYD^{+}24, LLZ^{+}22, ZYL^{+}24]. Receiver [HF17, DDHC^{+}12]. receiver-initiated [DDHC^{+}12]. Receiver-Synchronized [HF17]. Reception [HXZ23a, XHZG22]. Rechargeable [KJD^{+}23, LXR^{+}16, QNN^{+}22, SCG^{+}15, WTX^{+}23, WYD^{+}22, YRM^{+}24, ZGCL23, ZHT^{+}23, JKK08]. Recognition [LLZ^{+}22, LPW^{+}23, LHX^{+}21, OXZ^{+}23, SUR^{+}23, SYX^{+}23, WL23, WHYC19, XYJ^{+}23, YXG^{+}19, SSGM10, YYSL08]. Recommendation [LLW^{+}23]. Recommendations [dOEC^{+}23]. Reconfigurable [SML18, TLRE13]. Reconfiguration [HKG^{+}19, KKP^{+}07, SGB15]. reconstruction [NCV10]. Recovery [MYW^{+}24, PKC^{+}18, PX13]. Recruitment [XLO^{+}23]. Recurrent [FLCH23]. redistribution [TJWK13]. Reducing [WXL^{+}19]. Redundancy [CGVC06, LS10]. Redundant [ZWW^{+}23]. reference [ABM06]. refined [DVS^{+}14]. Reflection [EY14]. Regionalized [ZLD^{+}24]. Regions [SMR^{+}14]. Regressive [Kun22]. Regressive/Progressive [Kun22]. Regulations [Pha16]. Regulator [HSL^{+}15]. Rehabilitation [DWF^{+}23]. Reinforcement [DCD24, FBAG20, GAMW22, JGK^{+}23, LWL^{+}21, LWX^{+}21, RKLM23]. Reinforcing [MKFD^{+}23]. rekeying [CLSW12]. Related [RFB^{+}14, ZHJ^{+}20]. Relay [DGS16, GCAK17, MLZ^{+}24, NK15].

Relay-Assisted [DGS16]. Relays

[GSM^{+}22]. Reliability [JYC^{+}24, KYM17, KBD13]. Reliability-Security [JYC^{+}24]. Reliable [CLL^{+}23, DRC17, KLC^{+}16, KBW16, LED20, MP10, MZKC23, PH10, SNY^{+}24, XWW^{+}20, GFJ^{+}13, KAAF13, KAR^{+}14, PG10, IIPK20]. Relocatable [DCBL15]. Relocation [WHST16]. Remote [YSK^{+}15]. Renewable [MKFD^{+}23]. Repeatable [HSSS17]. replication [CUdVY13]. report [FLFW13]. Representations [FWF^{+}23, SZG11]. Representative [CHPP23]. reproduction [HR13]. reprogramming [KPB^{+}08, KW09, MP10, TLRE13]. Reputation [GBS08]. Reputation-based [GBS08]. Research [AAJ^{+}23, AMTH^{+}17, RDP16, RGB^{+}17]. Reservoirs [DXL^{+}15]. Residential [TPM^{+}17]. Residual [XFZ^{+}21]. Resilience [IPMGL18, JTE20]. Resiliency [CWK^{+}22, MLS^{+}22]. Resilient [CLJ^{+}23, DTY^{+}22, HSGW21, KMS^{+}10, SC15, SJP^{+}22, VRSR15]. Resistance [Yan22]. Resolutions [GLL^{+}24]. Resource [BJW^{+}22, HBKP14, HCL15, LDS^{+}22, LLH22, LCLY22, NLH^{+}19, RS19, TZZ22, VPB^{+}20, WCW^{+}23, ZWWL23, ZTZX23, NDM^{+}13]. Resource-constrained [BJW^{+}22, RS19]. Resource-Consuming [LLH22]. Resource-efficient [NLH^{+}19]. Respiration [GYG^{+}23, WSC^{+}23]. Respiratory [WLX^{+}23]. Response [MSB17, ZZPW23]. Results [ENPNF13, PG09]. Rethinking [HLL^{+}23]. Retrieving [SDZZ24]. Reuse [BT18]. Review [AMAT^{+}18, KOD^{+}14, WNM^{+}24]. Revolving [NXW^{+}22]. REWIMO [DRC17]. RF [BBEM^{+}24, GWS^{+}24, KVS23, KAS^{+}10, SMR^{+}14, SCL^{+}19, ZHJ^{+}20]. RF-AMOC [ZHJ^{+}20]. RF-based [SCL^{+}19]. RF-TESI [GWS^{+}24]. RFID [NXW^{+}22, WLW^{+}20, YYL^{+}23, ZHJ^{+}20]. RFID-based [YYL^{+}23]. RFIDs [ALY^{+}23, SYX^{+}23]. RFSense [SMR^{+}14]. Rigid [ZWW^{+}23, ZLGG10]. Risks [HLL^{+}23]. River [BGP^{+}23]. RLC [LWX^{+}21]. RNNs [RSK^{+}21]. Road [DSA^{+}20, SMR^{+}14, SMR^{+}14]. Road-RFSense [SMR^{+}14]. Robin [SC15]. Robots [LFNS14, TAT14, WTX^{+}16]. Robust [BAHS24, CQDW21, GYG^{+}23, KGGK11, LXY^{+}22, LFL^{+}19, MY24, MGS^{+}19, PPM15, PKC^{+}18, PG09, XBWX13, XWL24, ZJZ24b, DABNR10, GFJ^{+}13, NGSA08, LP05]. Robustness [SPI^{+}24, CKL^{+}09]. Rogue [LGLD23]. Room [ABC^{+}18, AAHS18, LYL^{+}24, WSC^{+}23]. Room-Scale [WSC^{+}23, LYL^{+}24]. rooms [YPW^{+}13]. Round [SC15]. Route [IIPK20, ZZG^{+}24]. Routing [ARWK19, GLS^{+}14, HWF^{+}24, KPCB20, KJP^{+}15, LFL^{+}19, WS14, XJL^{+}23, BGJ09, CA06, IV12, KT11, KLC13, KSMH13, LP08, PKG08, SZG13, TYD^{+}07, XRH^{+}13, YH13, ZSKH08, HBLR05]. Routing-Aware [ARWK19]. RPL [IIPK20, KPCB20, KJP^{+}15]. RSA [CLSW12]. RSSI [BHA^{+}13]. RSSI-based [BHA^{+}13]. RT [LCH^{+}19a]. RT-WiFi [LCH^{+}19a]. Rulers [LJLW19]. rules [ZDW^{+}10].

S [GDM22]. Safety [BSI^{+}15]. sales [HBW^{+}18]. Sample [ZWG24]. Sampling [BNG12, CHPP23, WWL15, ZGX^{+}16, ACG^{+}13, GSW09, KRJ09, LS10, LWH^{+}06, WLD10]. sampling-interpolation [LS10]. SARA [BCL^{+}12]. Satellite [LDGG21]. SateLoc [LDGG21]. Saturation [PPM15]. Saving [JYB^{+}21, YXFL17, JGK^{+}23, SGM08]. Scalable [AAHS18, CA06, WWL^{+}16, WZZ^{+}21, WCV^{+}18, GCRB12, GJNC^{+}14]. Scalar [Yan22]. Scale [BTR^{+}18, GLL^{+}24, LXR^{+}16, SBK22, SDZZ24, TJLK14, VRSR15, WSC^{+}23, WCW^{+}23, WS14, ZHZ^{+}16, ZZX^{+}20, CDR08, HBLR05, HM07b, KSMH13, KPB^{+}08, LWG09, LYL^{+}24, LGTL19, MCGZ21, MYH^{+}24, MB09, NJL24, PCR13, PH10, RSK^{+}21, SSL^{+}22, TJZ^{+}13, ZSJ06, WCPC20]. Scaling [LFW^{+}19, XHZG22, CPH06]. SCANet [LHZZ20]. Scanning [NXW^{+}22, WCLD23]. Schedule [SE23]. Schedules [PSB^{+}14]. Scheduling [AH20, BYD^{+}15, CCC^{+}21, CS23, ELR^{+}22, KYM17, LED20, LTL^{+}24, MZW^{+}19, TYGW15, WLW^{+}23, WWL15, WYD^{+}22, YWD^{+}21, YTR^{+}22, ZTZX23, ZGCL23, ZGX^{+}16, ZLGL19, ZLGL20, CNMH08, FS13, LDZ13, SG10, TYD^{+}07, YYM^{+}10]. Scheme [FSTH23, LHX^{+}21, MLX^{+}24, SLS^{+}22, YXFL17, YRM^{+}24, ZLD^{+}24, CLSW12, KLJ12, KT11, RR09, WDLN09]. Schemes [AH14, MSK^{+}23, ZMVR14, CDGC12, LCC10]. SDCN [LCM21]. SDN [PSR^{+}22]. SDP [GYNY16]. Seamless [ÁKSW22]. Search [LLDZ23, YSM08]. Search-based [LLDZ23]. Searchable [FSSR15]. SearchAuth [LLDZ23]. SecEG [HMG^{+}24]. Secret [LWH^{+}22, PCPK14, XJR^{+}17]. Section [CPSS23, CWK^{+}22, QXZZ22]. Secure [AAJ^{+}23, DABNR10, HM07b, HKW^{+}24, HMG^{+}24, KKRR15, LYG^{+}13, PTDD16, QWC^{+}22, QXZZ22, SUR^{+}23, SLS^{+}22, SNY^{+}24, TBS^{+}24, VTY18, WRYL11, ZYL^{+}24, ZSZ20, CCMT09]. Securing [SDX^{+}20]. Security [CZC^{+}24, GDM22, HAH22, JYC^{+}24, LSX24, LTZ^{+}24, MS09, MSB17, PDP^{+}17, WLLZ24, ZCZL22, CC11, CKL^{+}09, VG10, ZSJ06]. Security-by-contract [GDM22]. seed [TP07]. seeking [KVI^{+}13]. Segmentation [LYY24, YYSL08]. Segmenting [ABM06, ZSG09]. Seidel [KLC13]. Selection [CZX^{+}22, CGB^{+}19, MGS^{+}19, NK15, WLZ23, ZWL^{+}24, MCT14, NP12, TMAP14]. Selective [TDD^{+}19, NZR10]. Self [BR15, HL17, PMST12, ST12, ZHCA17, ZWY21, CNMH08, KSMH13, WZL07]. Self-Adaptation [HL17]. Self-healing [PMST12]. Self-localizing [ST12]. self-organized [KSMH13]. self-organizing [CNMH08]. Self-Powered [ZHCA17]. self-protection [WZL07]. Self-Sufficient [BR15]. Self-sustainable [ZWY21]. Semi [FSTH24, NZM21]. Semi-asynchronous [FSTH24]. Semi-supervised [NZM21]. Semidefinite [BLWY06]. SEMON [ZHCA17]. SenCS [LJW^{+}21]. Sense [GLG^{+}23]. SenseCode [KAAF13]. sensed [SLC^{+}22]. SenseLens [CA22]. Sensing [BIMD19, CTWG24, CZC^{+}24, FWF^{+}23, GLQ^{+}22, HSGW21, HSL^{+}15, LDL^{+}24, LWY^{+}21, LZN19, LJLW19, LCM21, MJS^{+}19, PK20, PKS^{+}23, RDP16, SMR^{+}14, SML18, SUZK19, SYT22, SDBT19, WWL15, WLX^{+}23, WNM^{+}24, XLO^{+}23, XAKV15, YSK^{+}15, YA24, YCL^{+}19, ZZ21, ZZ23, ZZY^{+}23, ZLL^{+}22, EML^{+}09, KPS12, NDM^{+}13, PDMJ10, SPK14, WKA14, WLW12, ZCLJ14]. Sensing-Based [SMR^{+}14]. sensitive [KASD09, TFL^{+}24, WJZ21]. Sensor [AMTH^{+}17, AMAT^{+}18, AKSM15, Amm16, Amm23, AH14, AHK16, AAHS18, ALNT22, BYD^{+}15, BGMP15, BWP^{+}24, BCL^{+}12, BAP^{+}17, BCMY22, BIMD19, BASM16, BWCW14, BSI^{+}15, BR15, BGP^{+}23, BQB^{+}11, COS19, CHPP23, CWY^{+}15, CTW^{+}15, CPP^{+}17, CCC^{+}21, CS23, CML^{+}21, CLS12, DPB19, DDA11, DBOD^{+}16, DML^{+}16, DXL^{+}15, EA15, ELR^{+}22, EY14, GZK^{+}23, GAMW22, GLS^{+}14, GSGA23, GLQ^{+}22, GTL19, GZZ^{+}14, HF17, HPS^{+}18, HMLJ17, HSGW21, HBKP14, IPMGL18, JJ15, JM16, JTS09, KPRH14, KJD^{+}23, KOD^{+}14, KKRR15, KK15, KBW16, KRP15, Lam15, LMP14, LLX^{+}14, LLL14, LL16, LCC^{+}17, LHZZ20, LXR^{+}16, LZAH^{+}15, LMZ^{+}16, LWM^{+}21, LHX16, LZN19, LFW^{+}19, LYST23, LCF^{+}22, MCGZ21, MB16, MSB17, MPRS16, MNLZ18, MGN22, MCW^{+}16, NGBB14, NK15, NK14, NRC^{+}09, NP12, PK19, PCA^{+}23, PPM15, PHKK17, PDP^{+}17, PTDD16, PNL^{+}22, PX13, PSB^{+}14, PCPK14, QNN^{+}22, RFB^{+}14]. Sensor [RBS16, RHS20, RD16, RJL^{+}10, SNK^{+}22, SSL^{+}19, SZG11, SZ19, SCL^{+}14, SGG10, SB16, SCL^{+}19, SCLG24, SXD^{+}15, SGB15, SG11, SZG^{+}15, TJLK14, TPM^{+}17, TFL^{+}24, TNBG18, TYGW15, TCB^{+}14, VPB^{+}20, VRSR15, WX08, WRYL11, WWFX11, WPL^{+}16, WB17, WS14, WBS14, WLS^{+}16, WHST16, WYD^{+}22, XDX^{+}14, XWW^{+}23, XWC^{+}23, XCC^{+}15, XXHL16, XWL24, YM14, YJL^{+}22, YRM^{+}24, YB17, YHC^{+}24, ZLW^{+}15, ZSLL23, ZGCL23, ZWY21, ZGT11, ZLGL19, ZLGL20, ZMVR14, dOEC^{+}23, Amm13, AAA06, ADF12, BKM^{+}12, BKS13, BLWY06, BHA^{+}13, BNG12, BGJ09, CJS11, CA06, CDGC12, CGVC06, CYS^{+}10, CCMT09, CK09, CSA06, CC11, CLSW12, CNMH08, CLH^{+}13, CHN^{+}13, CRW07, CRY^{+}10, CDR08, CGD12, CK13, CPH06, CCJ08, DLD09, Den09, DD09, Dji10, DABNR10, DIE14, DEM^{+}12, ELR08, EFI^{+}10, EGG13, ENPNF13, EMBP12, FLJ^{+}13, FS13, FLFW13, GCRB12, GSW09, GBS08, GCBL06, GSL10, GRE^{+}07, GFJ^{+}13]. sensor [GAJ^{+}06, GNDC08, HZGS05, HKL^{+}06, HM07a, HWT^{+}11, HBC^{+}09, HTC^{+}10, HY07, HBLR05, HLTC06, HTW07, HM07b, HCXT09, HR13, IR12, IBS^{+}10, JKK08, JC12, JHU^{+}13, JLYG13, JP06, JSBN^{+}12, JR08, JKS^{+}10, JROH09, Kal10, KBD13, KBD14, KXTZ09, KKP^{+}07, KC14, KQ12, KQ14, KKK08, KPK12, KLJ12, KT11, KAAF13, KLA^{+}14, KRJ09, KVI^{+}13, KSMH13, KPB^{+}08, KGGK11, KASD09, KW09, KAS^{+}10, KAR^{+}14, KMS^{+}10, KA13, LP08, LCC^{+}13, LDH06, LPV^{+}09, LP05, LP06, LPR09, LWG09, LKA10, LR05, LSW06, LL09, LDZ13, LWSL12, LS10, LH09, LCC10, LN05, LWH^{+}06, LND08, LFS09, LCH^{+}09, MZWT10, MB09, MWS08, MRM09, MS09, MPS10, MDC^{+}09, MP10, MS12, MKK^{+}13, MPC^{+}10, MAG13, NGSA08, NEKK12, NJS05, NZR10, NLD08, NC10, NCV10, ODCP13, ORRJ12, PDMJ10, PG10, PGG^{+}10, PBM11, PEFSV13, PG09, PC10, PKG08, PMST12, PCR13]. sensor [PA05, PH10, QM13, RBLP09, RKW^{+}06, RBD13, RR09, SYL09, SAZ10, SZG13, SSGM10, SSC^{+}10, SGM08, SPK^{+}10, SCWC13, SH09, SST08, SYOY12, SZZC08, SDČ10, Su07, SG08, SG10, SC12, SEZA13, TP07, TLRE13, TJZ^{+}13, TXC^{+}13, TXY^{+}13, TJWK13, TBL07, TYD^{+}07, VMS10, VG10, VAC13, WECC07, WEC11, WZL07, WZL08, WDLN09, WBS10, WLD10, WRS10, WIF^{+}11, WC13, WWLX13, WLZ13, WWXY13, WLW12, XBWX13, XWZ^{+}05, XLZ^{+}07, XWDN12, XTZ08, XRH^{+}13, YH13, YSZC13, YYM^{+}10, YS07, YVS07, ZSKH08, ZH05, ZKS10, ZLGG10, ZJX10, ZJZ12, ZVPS10, ZHKS06, ZDG09, ZZG^{+}24, ZSJ06, ZSJN07, ZSG09, ZDW^{+}10]. Sensor-Actuator [CS23, GRE^{+}07]. Sensor-Based [MNLZ18, LHZZ20]. Sensor-mission [RJL^{+}10]. SensorFly [CPP^{+}17]. Sensorial [LDDL24]. Sensorless [ZHCA17]. Sensornets [IHGS15]. Sensors [DSZ^{+}24, FLS^{+}14, FBAG20, KCE^{+}20, LFNS14, LWY^{+}21, LSW14, Pha16, RKRP17, SCG^{+}15, SKM^{+}11, ZLYW19, Bra07, CLX09, DVS^{+}14, KC14, KAH^{+}10, RKJ09, SMMS09, WC09, WC12, ZW05, ZBA07]. SensorScope [IBS^{+}10]. Sensory [LCM21]. Separation [BNN^{+}20, KGDC22]. sequence [KBD14]. sequence-based [KBD14]. Series [AAHS18, FWF^{+}23, LLX^{+}14, CHPP23]. SeRLoc [LP05]. Server [ZZPW23]. Service [JGK^{+}23, LZZ^{+}15, LLX^{+}22, LGXC23, SJP^{+}22, SGB15, TGG^{+}17, TGG^{+}19, XZL^{+}20, ZHZ^{+}16, KASD09]. Services [FM15, MLX^{+}24, YQLD22]. Sets [SCL^{+}19]. SGF [HCXT09]. SGX [YQLD22]. Shape [KGBS18, LWG09]. sharding [HKW^{+}24]. share [YYM^{+}10, MCLM20]. Shared [CT19, LWH^{+}22, Pha16, VPB^{+}20, XJR^{+}17]. Sharing [HBW^{+}18, HKW^{+}24, MCLM20, ZGX^{+}16, ZKS10, ZGHZ12]. shift [KAS^{+}10]. shift-based [KAS^{+}10]. Shopping [SYX^{+}23]. short [WDLN09]. short-term [WDLN09]. Shortest [SCL^{+}14]. ShortPK [WDLN09]. Shot [WL23]. Should [GLL^{+}24]. SHuffling [TDD^{+}19]. Side [LHHW24, Yan22]. Side-Channel [LHHW24, Yan22]. Sifting [YJWL13]. Sign [WNM^{+}24, YPZ^{+}17]. Signal [CA22, JAC19, ZW24, CKL^{+}09, NCV10, SPK^{+}10]. Signaling [TDZ^{+}22]. Signals [BBD^{+}23, CLX^{+}21, DWF^{+}23, FSSR15, GYG^{+}23, JCZ^{+}22, KVS23, LHHW24, WWZ24]. signature [CLSW12]. Silence [YSK^{+}15]. Similar [SDZZ24]. Similarities [CHPP23]. Similarity [LJW^{+}21]. Simple [LSW14, FKMS06]. Simulated [YTZ^{+}23]. simulation [KCPC13]. Simulators [MPRS16]. Single [KJP^{+}15]. sink [SZZC08]. Sinks [RD16]. situ [TLRE13, WLW12, WWL15]. Size [LJLW19, RSK^{+}21]. Sizing [WJZ21]. Skeletal [XYW^{+}22]. Sleep [CPX^{+}20, NK15, YPZ^{+}17, NC10]. Sleep-Wake [NK15]. Sleeping [MLS^{+}22, HY07, YH13]. Slotted [TNBG18]. Smart [CHSA18, CWK^{+}22, DTY^{+}22, DCD24, GXQ^{+}22, HPS^{+}18, HBW^{+}18, KCE^{+}20, KYM17, KKP18, LL21, LDS^{+}22, LDL^{+}24, LPW^{+}23, LSW14, MY24, NZM21, PK20, SBSD18, SMW23, WWZ^{+}21, XXW^{+}24, XFZ^{+}21, YXFL17, ZZH^{+}23, ZJZ^{+}24a, CHN^{+}13, ELYR14, ST12, TMAP14, WL14]. Smartphone [BNN^{+}20, CPX^{+}20, GWS^{+}24, XDM^{+}21, HSL^{+}15, PHKK17, WTX^{+}16]. Smartphone-Based [BNN^{+}20, XDM^{+}21, HSL^{+}15, WTX^{+}16]. Smartphones [BNPR20, SJP^{+}22, SDW^{+}23, TCC^{+}23, ZYC^{+}23, SMZ^{+}17]. SmartRoad [HSL^{+}15]. smoothness [MCT14]. snapshot [JHU^{+}13]. SNR [MYW^{+}24]. Social [BT18, CA22, MKFD^{+}23, SDX^{+}20, WKA14]. Social-Economic [MKFD^{+}23]. Socially [DSH16]. Socio [ELYR14]. Socio-economic [ELYR14]. Sociopsychological [RBS16]. SOCP [GYNY16]. Soft [BT18]. Software [DCBL15, PHKK17, GRE^{+}07, PCR13]. Soil [WWL15, WLW12]. Solar [BJR15, BIST18, RKLM23, WXG^{+}24, YM14, JC12]. Solar-Powered [YM14, RKLM23]. SolarKey [WXG^{+}24]. Solution [WLLZ24, XDL^{+}24, YH13]. Solutions [HBKP14, VG10, ZHKS06]. SonicDoor [KGBS18]. Sounds [ZZH^{+}23]. Source [GYNY16, KGDC22, LLX^{+}22, MB09, PX13, YSZC13]. source-optimized [MB09]. sources [CRY^{+}10]. Space [GKRW17, WWL^{+}16, WJD16, WCLD23, ABM06]. spaced [NCV10]. spanner [PR10]. spanners [SS13]. Sparse [BWP^{+}24, WJY^{+}24, WWL15, YB17, Kal10, KVI^{+}13, GSW09]. sparsely [Amm13]. Spatial [FLCH23, Kou18, LXY^{+}22, LWZ24, PZOZ21, SZG11, ZLB^{+}23, JKK08, PKG08, SZG13, YS07]. Spatial-Feature-based [FLCH23]. Spatial-Temporal [LXY^{+}22]. spatially [JP06]. Spatio [CUdVY13, PAYL22, LKA10]. Spatio-temporal [CUdVY13, PAYL22, LKA10]. Spatiotemporal [DD11, XFZ^{+}21]. Speakers [LHHW24]. Special [CPSS23, CWK^{+}22, LWKZ22, LSX24, MGN22, NJZ18, QXZZ22, SMW23]. Specific [LYST23, IBS^{+}10]. spectral [LS10]. Spectrum [LZN19, MSAJ18, SBSD18, WYC^{+}24]. Spectrum-efficient [WYC^{+}24]. Speech [HL17]. Speed [SG10, WTC22]. SpO [BNN^{+}20]. Spray [WYC^{+}24]. spread [DLD09]. spreading [QM13]. SPRED [LDDL24]. stability [PFJ13]. Stable [LZAH^{+}15]. Stack [KPRH14, RS19].

Stack-based [RS19]. STARR [CUdVY13]. STARR-DCS [CUdVY13]. Start [SMZ^{+}17]. state [HCXT09, LWSL12]. state-free [HCXT09]. Static [HWF^{+}24, LWM^{+}21, Den09, LN05]. Station [YHC^{+}24, SH09]. Station-less [YHC^{+}24]. Statistical [PC10, IR12, KA13]. statistically [YSZC13]. Staying [BR15]. Stealthy [BH21]. Steganographic [CSLJ23]. Steiner [SB16]. Stochastic [LP06, KT11, PG09, YYM^{+}10]. stolen [GPL^{+}12]. Stone [KGDC22]. Storage [LLX^{+}14, LWCJ14, WRYL11, ZLL^{+}22, CUdVY13, LCH^{+}09, MDC^{+}09, ZGHZ12]. storage-centric [LCH^{+}09]. Strategies [LWM^{+}21]. Strategy [HMG^{+}24, WLW^{+}23, YTZ^{+}23]. Stream [KYM17, XQL^{+}24, LHZZ20]. Streams [MYH^{+}24]. Street [CT19]. strength [CKL^{+}09]. Stretch [WS14]. Strip [LFL^{+}19]. strong [YSZC13]. Structural [BWCW14, DBC^{+}24, ACG^{+}13]. Structure [NXW^{+}22, SJP^{+}22, GCBL06]. structures [ABM06]. sTube [HBW^{+}18]. Studies [DXL^{+}15]. Study [BDP24, COP^{+}16, DGS16, GLL^{+}24, LGTL19, MPRS16, YJL^{+}22, KPS12, MPC^{+}10, SDTL10, YPW^{+}13]. style [XWW^{+}23]. Sub [SMS22]. Sub-1 [SMS22]. Subject [LPW^{+}23, MLZ^{+}24, LWSL12]. Subject-adaptive [LPW^{+}23]. Submodular [ZWL^{+}24]. Subsets [CHPP23]. Sufficient [BR15]. summarization [dLM14]. Summary [PCA^{+}23, PGY^{+}24]. Superposition [MZW^{+}19]. supervised [NZM21]. Supervision [ZJZ24b]. Supervisory [YBY^{+}24]. Supplied [ZLYW19]. Supply [PK20]. Support [IIPK20, NGBB14]. Supported [SHWW20]. Supporting [KJP^{+}15]. Surface [CK13, EY14, WJD16]. Surface-level [CK13]. Surface-Reflection-Based [EY14]. Surveillance [DXC^{+}21, HLL^{+}23, TYGW15, GAJ^{+}06, HKL^{+}06, VHC^{+}09]. Survey [CML^{+}21, DDA11, DTW^{+}23, GSGA23, HAH22, LDH06, LWM^{+}21, RHD17, RDP16, RGB^{+}17, SYL^{+}22, YYC^{+}19, dOEC^{+}23, BKM^{+}12, RBD13, SG08]. Survivability [TYGW15]. Survivability-Heterogeneous [TYGW15]. Sustainability [KYM17]. Sustainable [YTR^{+}22, DEM^{+}12, ZWY21]. Swarm [CRZ^{+}20]. Switching [BT18]. Symbols [BY19]. SymListener [WLX^{+}23]. Symptoms [WLX^{+}23]. sync [YVS07]. Synchronization [BDO14, GJT^{+}22, JTE20, SZ19, VTY18, VDV16, XXHL16, CLS12, SSC^{+}10, YVS07]. Synchronization-free [GJT^{+}22]. Synchronized [HF17]. Synchronous [LHX16, MDC17]. Synopsis [NGSA08]. System [AJH^{+}20, BBD^{+}23, BR15, CPX^{+}20, CTW^{+}15, CC23, CSLJ23, CA22, DWF^{+}23, DLG^{+}21, DBC^{+}24, FWF^{+}23, GZJE23, GYG^{+}23, HKG^{+}19, JLZL19, KCE^{+}20, KGBS18, LYZ^{+}24, LL21, LHHW24, LWJ^{+}23, MYH^{+}24, MSB17, NZM21, OXZ^{+}23, RKLM23, SUR^{+}23, SMR^{+}14, SNC^{+}23, TXY^{+}13, WLW^{+}20, WYC^{+}24, WCV^{+}18, WJ21, XXW^{+}24, XCT^{+}16, XWW^{+}20, XKW^{+}22, YZZD23, ZZPW23, ZZC^{+}23, ZGH^{+}21, ACG^{+}13, DABNR10, EML^{+}09, HKL^{+}06, LNV^{+}05, OBB^{+}13, ODCP13]. System-level [TXY^{+}13]. Systematic [HAH22]. Systems [BY19, CZC^{+}24, DCBL15, GKRW17, HLZ^{+}24, HWS^{+}20, JZL^{+}19, KOD^{+}14, MLX^{+}24, MJS^{+}19, MZKC23, MCLW23, NXW^{+}22, PAYL22, RFS^{+}19, SJH^{+}18, SBSD18, SZG^{+}15, SDBT19, YSK^{+}15, YA24, YYL^{+}23, ZZZ^{+}20, ZPL^{+}24, LJY^{+}10, NZR10, NDM^{+}13].

Tag [CWS^{+}22, WLW^{+}20, ZHJ^{+}20]. TagFocus [YYL^{+}23]. Tagged [NXW^{+}22]. Tags [CWS^{+}22, MGS^{+}15]. Taking [PGY^{+}24]. Tamera [SYX^{+}23]. Taming [GHZ^{+}22, WWZ24]. Target [LMP14, SAK^{+}19, SMMS09, SKM^{+}11, SYT22, Bra07, LPR09, MS12, WBS10, WRS10, YLL13, ZDW^{+}10]. Targets [WPL^{+}16, KQ12, WC09, WC12]. TARS [HF17]. TAS [LHX16]. TAS-MAC [LHX16]. Task [BJW^{+}22, MDM^{+}20, MKM^{+}20, MZKC23, PZOZ21, YTR^{+}22, ZZ23, ZGCL23, WZZ^{+}21]. Task-based [MDM^{+}20]. Task-oriented [ZGCL23]. Tasks [ZGX^{+}16, IW14]. Taxi [MCLM20, WCW^{+}23]. Taxi-Sharing [MCLM20]. Taxicab [ZHZ^{+}16]. TDMA [AH20, GCRB12, NGBB14]. TDMA-Based [NGBB14, GCRB12]. Team [LFNS14]. Technique [HMLJ17, YS07]. Techniques [IHGS15, dOEC^{+}23, KLA^{+}14, MKK^{+}13]. Technologies [CPSS23, WNM^{+}24]. Technology [CD21, GHZ^{+}22, WXL^{+}19, ZGJ^{+}22, SMS22, SCS22]. Temperature [CTW^{+}15, GWS^{+}24, XXHL16]. Temperature-Aware [XXHL16]. Temperatures [BGP^{+}23]. TempMesh [BGP^{+}23]. Temporal [KXTZ09, LDDL24, LLX^{+}14, LL16, LXY^{+}22, LC14b, YZZD23, ZLB^{+}23, CUdVY13, LKA10, PAYL22, YS07]. Tenet [PGG^{+}10]. Term [XDX^{+}14, VHC^{+}09, WDLN09, ZGHZ12]. Terra [BSI^{+}15]. terrain [CK13]. TESI [GWS^{+}24]. Testbed [FPA^{+}20]. Testing [IHGS15, AAA06]. Text [FSSR15]. Text-Searchable [FSSR15]. TG [LDDL24]. TG-SPRED [LDDL24]. Their [LSW14, HAH22]. Theoretic [CPL^{+}20, SBCF20, CDGC12, VAC13]. Theory [DBOD^{+}16, NEKK12, YHC^{+}24, ZWWZ20, ABM13, CCJ08, DLD09, JC12, ZBA07, KXTZ09, PG09]. Thermal [FS13, YPW^{+}13]. Thermal-aware [FS13]. Thermo [PKS^{+}23]. Thermo-hygrometer [PKS^{+}23]. Things [YMY^{+}23, AAJ^{+}23, BJW^{+}22, CQDW21, JGK^{+}23, MGS^{+}19, SMW23, SLS^{+}22, YTR^{+}22, ZZW^{+}23a, ZLYW19, ZDS^{+}21]. Threat [BJW^{+}22].

Threat-modeling-guided [BJW^{+}22]. Three [Amm16]. Three-Dimensional [Amm16]. threshold [ZDW^{+}10]. Throughput [ZMXM24, FT06]. Tier [XZL^{+}20]. Tiered [WHST16, PGG^{+}10]. Tight [YVS07]. Time [ABC^{+}18, AAHS18, CHPP23, DLD^{+}23, DRC17, FWF^{+}23, FLCH23, GM14, LLX^{+}14, MZKC23, Pha16, PSB^{+}14, SBK22, SCG^{+}15, TFL^{+}24, TNBG18, WWFX11, WLW^{+}20, WJZ21, XYJ^{+}23, XXHL16, XQL^{+}24, ZZPW23, BBD^{+}23, BCMY22, Gel07, HZGS05, LJW^{+}21, LWSL12, LWH^{+}06, NC10, ORRJ12, RS19, VMS10, WWXY13, WLLZ24, XRH^{+}13, YVS07, ZJX10, ZZM^{+}22, ZYC^{+}23]. Time-Critical [PSB^{+}14]. Time-efficient [WLW^{+}20]. Time-Interval [SBK22]. Time-sensitive [TFL^{+}24, WJZ21]. Time-Series [LLX^{+}14, CHPP23]. Time-Slotted [TNBG18]. Time-Varying [GM14, VMS10]. Timely [XQL^{+}24]. Timestamping [GJT^{+}22]. Timestamps [LTY18]. Timing [SE23, TXC^{+}13]. Tiny [YVS07]. Tiny-sync [YVS07]. TinyLink [DLG^{+}21]. toad [HBC^{+}09]. TOC [SCG^{+}15]. Tolerant [LMP14, COS19]. tolerating [GPL^{+}12, SZZC08]. Tones [SHY13]. tool [LJY^{+}10]. tools [JTS09]. topologies [NCV10]. Topology [CQDW21, HWF^{+}24, KPCB20, LFL^{+}19, RFB^{+}14, LSW06]. Topology-Related [RFB^{+}14]. Touchscreen [CJL^{+}20]. Trace [LYL^{+}24, YYSL08]. Traceability [QWC^{+}22]. tracing [SEZA13]. trackability [CCJ08]. Tracking [BQB^{+}11, GKRW17, LMP14, LYL^{+}24, PAYL22, SYX^{+}23, SKM^{+}11, WSC^{+}23, WPL^{+}16, WCV^{+}18, XYW^{+}22, YXFL17, ZYZ^{+}19, ZW24, BHA^{+}13, EGG13, GJNC^{+}14, GPL^{+}12, KASD09, KAS^{+}10, MS12, SMMS09, TMAP14, TTBH14, WBS10]. Trade [FLFW13, ZZX^{+}20, WRS10]. Trade-off [FLFW13, WRS10]. Tradeoff [JYC^{+}24].

Traffic [BTR^{+}18, CS23, DSA^{+}20, HF17, HSL^{+}15, IIPK20, LHX16, PSR^{+}22, SMR^{+}14, SYOY12, ZZM^{+}22, WECC07]. Traffic-Adaptive [HF17, LHX16]. Traffic-Aware [CS23]. Trail [KASD09]. Train [LXYT24]. Training [FWF^{+}23]. Trajectories [SDZZ24]. Trajectory [SLC^{+}22, WLW^{+}23]. Transceiver [KGDC22]. Transfer [BASM16, LDC^{+}19, LYST23, MLX^{+}24, SZX17, SMZ^{+}17, WLZ23, ZZC^{+}23, GCRB12]. Transferable [AAHS18]. Transit [MCLW23]. Transition [SLC^{+}22]. Transmission [KLC^{+}16, KPCB20, LMZ^{+}16, LCH^{+}20, MDC17, MGS^{+}19, WXL^{+}19, ZCZL22, ZLW^{+}24, GCBL06, PR10, WWXY13]. Transmission-Based [MDC17]. Transmissions [BBEM^{+}24, XHZG22, YYXL22]. Transmit [KR18]. transport [HR13, PG10]. Transportation [BDP24, RMB^{+}10]. trap [CLH^{+}13]. Travel [FLCH23, Gel07]. Tree [JJ15, SB16, AH20, GFJ^{+}13, JKS^{+}10]. Trees [CHSA18, SCL^{+}14]. Trends [AAJ^{+}23, AMTH^{+}17]. triangle [YJWL13]. Triggered [SDBT19]. Tropical [LWL^{+}21]. Troubleshooting [KLA^{+}14]. True [CA22]. Trust [BJW^{+}22, LSX24, RBS16, SBCF20, TBS^{+}24, LYG^{+}13, YBY^{+}24]. Trust-based [BJW^{+}22]. trusted [HTC^{+}10]. Trustworthy [HWT^{+}22]. Truth [MJS^{+}19, WJY^{+}24, ZGH^{+}21]. Truthful [YHC^{+}24]. TSCH [TDD^{+}19]. TSDroid [ZLB^{+}23]. tunnels [MPC^{+}10]. Turf [WWB^{+}19]. TV [BAP^{+}17]. Twin [GXQ^{+}22, ZLX^{+}24]. Twin-enabled [GXQ^{+}22]. Twins [LCF^{+}22]. Two [DGS16, GCAK17, LHZZ20, WLZ23, WHST16]. Two-Connected [GCAK17]. Two-Hop [DGS16]. Two-phased [WLZ23]. Two-stream [LHZZ20]. Two-Tiered [WHST16]. Type [MGS^{+}19]. types [NRC^{+}09].

UAV [HZX^{+}24, HWF^{+}24, TZZ22, WLW^{+}23, XXW^{+}24, XQL^{+}24]. UAV-Aided [WLW^{+}23, XQL^{+}24]. UAV-Assisted [TZZ22]. UAVs [KVI^{+}13, ZHT^{+}23]. Ubiquitous [LWLT24, TGG^{+}19, ZZZ^{+}22]. Ultra [CP20, MDC^{+}09, PKC^{+}18]. Ultra-low [MDC^{+}09]. Ultra-wideband [CP20]. Ultrasonic [LDL^{+}24]. Ultrasound [ZJZ^{+}24a]. unattended [PMST12]. Uncertainty [TFL^{+}24]. Uncontrollable [RD16]. Underground [LL09, PGY^{+}24]. Understanding [BBEM^{+}24, XTXW22, YCL^{+}19]. Undervolting [KBW16]. Underwater [ELR^{+}22, EY14, GAMW22, HF17, KGDC22, LCF^{+}22, MGN22, PCA^{+}23, PSR^{+}22, RHS20, SNK^{+}22, XWW^{+}23, XWC^{+}23, SHY13]. Unfolding [CS18]. Unit [FLCH23, IHGS15, FKMS06]. Units [XYW^{+}22]. Unknown [LGTL19]. Unlabeled [ALS23]. Unmanned [HWS^{+}20]. Unobtrusive [CPX^{+}20]. unreliability [ZK07]. Unreliable [WKYH17]. Unrestricted [XLG^{+}22]. Unsupervised [HWT^{+}22, SLC^{+}22, TPM^{+}17]. Update [DCBL15, SNY^{+}24, PBM11]. Uplink [YYXL22]. upper [ZH05]. Urban [CTWG24, DXL^{+}15, MCLM20, MCLW23, YJL^{+}22, ZZX^{+}20, ZWWZ20, LNV^{+}05]. URLLC [SE23]. usable [VG10]. Usage [Pha16, TPM^{+}17]. Useful [SCLG24]. User [CYD^{+}24, CLJ^{+}23, LZGX23, WSC^{+}23, WLW^{+}23, WHQ^{+}23, XDX^{+}14, XLO^{+}23, YYXL22, YYSL08]. User-Centric [XDX^{+}14]. User-independent [WHQ^{+}23]. user-trace [YYSL08]. User/Device [LZGX23]. Users [CJL^{+}20, LLZ^{+}20]. Using [AMTH^{+}17, BQB^{+}11, CHPP23, CC23, DSA^{+}20, DML^{+}16, GYG^{+}23, HZX^{+}24, HLZ^{+}24, JGK^{+}23, KVS23, KR18, LTDZ22, LLDZ23, LYY24, LDGG21, LGLD23, LZN19, MDC17, PHKK17, PSR^{+}22, PCPK14, RKRP17, RMB^{+}10, RKLM23, SZX17, SYX^{+}23, SMZ^{+}17, SZG^{+}15, TPM^{+}17, TAT14, TCC^{+}23, WSC^{+}23, WTX^{+}16, WB17, WHYC19, WXG^{+}24, WWL15, WTH^{+}23, WHQ^{+}23, WNM^{+}24, XYJ^{+}23, XAKV15, YPZ^{+}17, YB17, ZZH^{+}23, ZZY^{+}23, ZYC^{+}23, ZJZ^{+}24a, ZGH^{+}21, BNPR20, CHSA18, CRY^{+}10, DLD09, EGG13, FLJ^{+}13, HR13, JYB^{+}21, KCPC13, KLA^{+}14, KVI^{+}13, KNSM14, LCC^{+}13, LK09, LFS09, LC14a, MS12, ORRJ12, RR09, SZG13, SPK14, SYOY12, WL14, XRS10, ZBA07, ZGT11, KAH^{+}10]. Utility [EMBP12, SJH^{+}18, PDMJ10]. Utility-based [EMBP12, PDMJ10]. Utilization [VPB^{+}20]. Utilizing [QM13]. UWB [CWY24, WCLD23].

validity [FLFW13]. value [BKS13, VG10]. value-based [VG10]. Valued [WHYC19]. Variability [MG24]. Variable [ZDG09, PR10]. variant [TTBH14]. Variation [GWS^{+}24, KR18]. Varying [GM14, VMS10]. Vehicles [LXR^{+}16, MDB^{+}23]. Vehicular [HKW^{+}24]. Verification [LJW^{+}21]. versatile [DDHC^{+}12]. versus [LP08]. via [CJL^{+}20, CG18, HPS^{+}18, HKG^{+}19, JZX^{+}20, KLJ12, LKA10, LJW^{+}21, LYL^{+}24, LXR^{+}16, MYW^{+}24, NXW^{+}22, SBSD18, SMS22, TLRE13, TGG^{+}17, WZZ^{+}23, WLX^{+}23, XXHL16, YA24, YYSL08, ZWL^{+}24]. Vibration [DBC^{+}24, WHQ^{+}23, ZDS^{+}21, KPS12]. Vibration-based [DBC^{+}24, ZDS^{+}21, KPS12]. Video [MYH^{+}24, NJL24, XKW^{+}22, ZZM^{+}22, DVS^{+}14, dLM14]. Videos [ZTZX23]. View [DSZ^{+}24, JM16, MCT14, WC13]. views [KNSM14]. VigilNet [HKL^{+}06, VHC^{+}09]. VILL [NZH^{+}23]. Virtual [LDGG21, DABNR10]. Vision [CZC^{+}24, ELYR14, IW14]. Visitor [KSR^{+}20]. ViST [LWLT24]. Visual [CYD^{+}24, NZH^{+}23, SYT22, XDM^{+}21, YYL^{+}23, DVS^{+}14, KQ12, KQ14, MAG13]. Vital [WNM^{+}24, YPZ^{+}17]. VLSI [GAJ^{+}06]. VNF [XZL^{+}20]. volcanic [TXC^{+}13]. Volumetric [WWL^{+}16]. Voronoi [MLZ^{+}24]. VSSB [TBS^{+}24]. VSSB-Raft [TBS^{+}24]. Vulnerabilities [HAH22, SWH^{+}24].

W3W [ZLYW19]. Wait [WTX^{+}23]. Wait-for [WTX^{+}23]. Wake [CWY^{+}15, NK15, GAJ^{+}06, ODCP13]. Wake-Up [CWY^{+}15, GAJ^{+}06, ODCP13]. wakeup [SHY13]. Walking [KGBS18, WTC22]. WAN [GSM^{+}22]. warfare [LNV^{+}05]. Water [AMTH^{+}17, DXL^{+}15, KYM17, PK19, KPS12, LCC^{+}13]. Wave [BY19, TYD^{+}07, YPZ^{+}17, ZCZ^{+}23]. Wavelengths [BNN^{+}20]. Waves [LYL^{+}24]. Waving [LJLW19]. way [SAZ10]. Weak [HXZ23a]. Wearable [XJR^{+}17]. Wearables [CLL^{+}23, JLZL19]. weighted [CPH06]. weighted-multidimensional [CPH06]. where [SYOY12]. while [GPL^{+}12]. Whisper [BLGS19]. Who [GLL^{+}24, SYOY12]. Wi [XYJ^{+}23, ZZZ^{+}22]. Wi-Fi [XYJ^{+}23, ZZZ^{+}22]. WIB [ZYC^{+}23]. Wide [LWKZ22, LCD22, SBK22, WQH^{+}22, KNSM14, WJ21, YSM08]. Wide-area [LCD22, KNSM14]. Wide-Area-Networks [SBK22]. Wideband [PKC^{+}18, CP20]. WiFi [GYG^{+}23, LCH^{+}19a, LWJ^{+}23, ZWG24]. WiFi-based [ZWG24]. WiFine [XYJ^{+}23]. Wild [DML^{+}16, SWH^{+}24]. wildlife [DEM^{+}12]. WILDSENSING [DEM^{+}12]. will [SYOY12]. Wind [DXL^{+}15]. Wireless [AMTH^{+}17, AMAT^{+}18, AKSM15, Amm16, Amm23, AH14, BAHS24, BYD^{+}15, BGMP15, BDO14, BAP^{+}17, BCMY22, BIMD19, BASM16, BLGS19, BSI^{+}15, BGP^{+}23, CBSA18, CKHP19, CWY^{+}15, CS23, DPB19, DRW^{+}14, DRC17, DDA11, DSH16, DGS16, DML^{+}16, EA15, GZK^{+}23, GLS^{+}14, GSGA23, GCAK17, GTL19, GZZ^{+}14, HBKP14, HCL15, HLL^{+}23, IPMGL18, JM16, KJD^{+}23, KOD^{+}14, KKRR15, KK15, KBW16, KRP15, LL16, LCC^{+}17, LDC^{+}19, LXY^{+}22, LXYT24, LZAH^{+}15, LMZ^{+}16, LWM^{+}21, LGLD23, LWCJ14, LHX16, LFL^{+}19, LFW^{+}19, LCH^{+}20, LCLY22, MCGZ21, MB16, MSB17, MPRS16, MSAJ18, NGBB14, NK15, NK14, PGY^{+}24, PPM15, PDP^{+}17, PTDD16, Pha16, PNL^{+}22, PSB^{+}14, PCPK14, QNN^{+}22, RFB^{+}14, RBS16, SSL^{+}19, SCL^{+}14, SCG^{+}15, SXD^{+}15, SGB15, SZG^{+}15, SDBT19, TCN^{+}17, TPM^{+}17, TFL^{+}24, TNBG18, WWFX11, WPL^{+}16, WKYH17, WZLM21, WS14, WBS14, WLS^{+}16, WHST16, WXD^{+}23, XDX^{+}14, XXHL16, YM14, YRM^{+}24, YTB^{+}14]. Wireless [YB17, ZHCA17, ZLW^{+}15, ZZZ^{+}20, ZLZ21, ZGCL23, ZWY21, ZZW^{+}23b, ZZC^{+}23, ZLGL19, ZLGL20, dOEC^{+}23, ADF12, BKM^{+}12, BHA^{+}13, BNG12, CJS11, CA06, CDGC12, CYS^{+}10, CCMT09, CC11, CLSW12, CNMH08, CLX09, CLH^{+}13, CVY09, CGD12, DLD09, Den09, DD09, DABNR10, DIE14, DDHC^{+}12, ENPNF13, EMBP12, FLJ^{+}13, FT06, GFJ^{+}13, HM07a, HWT^{+}11, HTC^{+}10, HLTC06, HTW07, HCXT09, HR13, IV12, JHU^{+}13, JLYG13, KBD14, KXTZ09, KCPC13, KC14, KPK12, KLJ12, KLA^{+}14, KRJ09, KSMH13, LDH06, LPV^{+}09, LP05, LPR09, LKA10, LSW06, LL09, LDZ13, LYG^{+}13, LCC10, LWH^{+}06, LND08, LFS09, MZWT10, MPS10, MS12, MKK^{+}13, MPC^{+}10, NZR10, NLD08, NC10, OBB^{+}13, ODCP13, PDMJ10, PG10, PEFSV13, PKG08, PMST12, PCR13, QM13, RBLP09, RBD13, RJL^{+}10, RR09, SYL09, SAZ10, SZG13, SSGM10, SPK^{+}10, SCWC13, SH09, SPK14, SZZC08, SDTL10, Su07, SEZA13, TP07]. wireless [TXC^{+}13, TXY^{+}13, TBL07, VAC13, WZL07, WLD10, WWLX13, XBWX13, XLZ^{+}07, XTZ08, XRH^{+}13, YS07, YVS07, ZK07, ZSKH08, ZJX10, ZJZ12, ZCLJ14, ZHKS06, ZDW^{+}10]. Wireless-Charging-Based [CKHP19]. Wireless-Sensor-Network-Enabled [KOD^{+}14]. without [LHX^{+}21, SSGM10]. Workloads [LDG^{+}21]. World [GKRW17, SGG10, YSM08]. Worn [SDX^{+}20]. worst [JKS^{+}10]. worst-case [JKS^{+}10]. WPANs [LED20]. Wrist [SDX^{+}20]. Wrist-Worn [SDX^{+}20]. Writing [YXG^{+}19]. WSN [JAC19]. WSNs [AMAT^{+}18, ABM13, AH20, ARWK19, KLC13, WWL^{+}16, WJD16, WLW^{+}23, WTX^{+}23, WYD^{+}22, XAKV15, YLSZ19, Yan22, ZGX^{+}16]. WUGS [RRA22]. WVC [ZYL^{+}24]. Wyner [DVS^{+}14].

X [CC23]. X-ray [CC23]. XNAS [Kun22].

Y-Networks [JJ15].

Zero [LSX24, TBS^{+}24, VRSR15, WL23, YBY^{+}24]. Zero-Delay [VRSR15]. Zero-Shot [WL23]. Zero-trust [YBY^{+}24]. Zig-Bee [AH20, SMS22, SCS22]. ZigBee-like [AH20]. Ziv [DVS^{+}14].

References

Arici:2006:PEB

[AAA06] Tarik Arici, Toygar Akgun, and Yucel Altunbasak. A prediction error-based hypothesis testing method for sensor data acquisition. ACM Transactions on Sensor Networks, 2(4):529–556, November 2006. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Arief-Ang:2018:SRO

[AAHS18] Irvan B. Arief-Ang, Margaret Hamilton, and Flora D. Salim.

A scalable room occupancy prediction with transferable time series decomposition of CO_{2} sensor data. ACM Transactions on Sensor Networks, 14(3–4):21:1– 21:??, December 2018. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Adil:2023:CSH

[AAJ^{+}23] Muhammad Adil, Jehad Ali, Muhammad Mohsin Jadoon, Sattam Rabia Alotaibi, Neeraj Kumar, Ahmed Farouk, and Houbing Song. COVID-19: Secure healthcare Internet of Things networks, current trends and challenges with future research directions. ACM Transactions on Sensor Networks, 19(3):54:1–54:??, August 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3558519.

Abbas:2018:IHG

[ABC^{+}18] Samar Abbas, Abu Bakar, Yasra Chandio, Khadija Hafeez, Ayesha Ali, Tariq M. Jadoon, and Muhammad Hamad Alizai. Inverted HVAC: Greenifying older buildings, one room at a time. ACM Transactions on Sensor Networks, 14(3–4):26:1– 26:??, December 2018. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Agarwal:2006:SOS

[ABM06] Pankaj K. Agarwal, David Brady, and Jiří Matoušek. Segmenting object space by geometric reference structures. ACM Transactions on Sensor Networks, 2(4):455–465, November 2006. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Abrardo:2013:GTD

[ABM13] Andrea Abrardo, Lapo Balucanti, and Alessandro Mecocci. A game theory distributed approach for energy optimization in WSNs. ACM Transactions on Sensor Networks, 9(4):44:1– 44:??, July 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Alippi:2013:HFS

[ACG^{+}13] Cesare Alippi, Romolo Camplani, Cristian Galperti, Antonio Marullo, and Manuel Roveri. A high-frequency sampling monitoring system for environmental and structural applications. ACM Transactions on Sensor Networks, 9(4):41:1–41:??, July 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ayday:2012:DAA

[ADF12] Erman Ayday, Farshid Delgosha, and Faramarz Fekri. Data authenticity and availability in multihop wireless sensor networks. ACM Transactions on Sensor Networks, 8(2):10:1– 10:??, March 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Anagnostopoulos:2014:APC

[AH14] Christos Anagnostopoulos and Stathes Hadjiefthymiades. Advanced principal componentbased compression schemes for wireless sensor networks. ACM Transactions on Sensor Networks, 11(1):7:1–7:??, August 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ahmad:2020:EED

[AH20] Aasem Ahmad and Zdenek Hanzalek. An energy-efficient distributed TDMA scheduling algorithm for ZigBee-like clustertree WSNs. ACM Transactions on Sensor Networks, 16(1): 3:1–3:41, February 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3360722.

Anagnostopoulos:2016:ADD

[AHK16] Christos Anagnostopoulos, Stathes Hadjiefthymiades, and Kostas Kolomvatsos. Accurate, dynamic, and distributed localization of phenomena for mobile sensor networks. ACM Transactions on Sensor Networks, 12 (2):9:1–9:??, May 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Alaziz:2020:BBM

[AJH^{+}20] Musaab Alaziz, Zhenhua Jia, Richard Howard, Xiaodong Lin, and Yanyong Zhang. In-bed body motion detection and classification system. ACM Transactions on Sensor Networks, 16 (2):13:1–13:26, April 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3372023.

Andersen:2018:DAB

[AKC^{+}18] Michael P. Andersen, John Kolb, Kaifei Chen, Gabe Fierro, David E. Culler, and Randy Katz. Democratizing authority in the built environment. ACM Transactions on Sensor Networks, 14(3–4):17:1–17:??, December 2018. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ali:2015:AHC

[AKSM15] Azad Ali, Abdelmajid Khelil, Neeraj Suri, and Mohammadreza Mahmudimanesh. Adaptive hybrid compression for wireless sensor networks. ACM Transactions on Sensor Networks, 11(4):53:1–53:??, December 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Alamos:2022:DLS

[ÁKSW22] José Álamos, Peter Kietzmann, Thomas C. Schmidt, and Matthias Wählisch. DSME-LoRa: Seamless long-range communication between arbitrary nodes in the constrained IoT. ACM Transactions on Sensor Networks, 18(4):69:1– 69:??, November 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3552432.

Aula:2022:ELC

[ALNT22] Kasimir Aula, Eemil Lagerspetz, Petteri Nurmi, and Sasu Tarkoma. Evaluation of lowcost air quality sensor calibration models. ACM Transactions on Sensor Networks, 18 (4):72:1–72:??, November 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3512889.

Ahmed:2023:DHA

[ALS23] Usman Ahmed, Jerry Chun-Wei Lin, and Gautam Srivastava. Deep hierarchical attention active learning for mental disorder unlabeled data in AIoMT. ACM Transactions on Sensor Networks, 19(3):49:1–49:??, August 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3519304.

An:2023:LRP

[ALY^{+}23] Zhenlin An, Qiongzheng Lin, Lei Yang, Yi Guo, and Ping Li. Localizing RFIDs in pixel dimensions. ACM Transactions on Sensor Networks, 19(1): 1:1–1:??, February 2023. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3517012.

Adu-Manu:2018:EHW

[AMAT^{+}18] Kofi Sarpong Adu-Manu, Nadir Adam, Cristiano Tapparello, Hoda Ayatollahi, and Wendi Heinzelman. Energy-harvesting wireless sensor networks (EH-WSNs): a review. ACM Transactions on Sensor Networks, 14 (2):10:1–10:??, July 2018. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ammari:2013:JCD

[Amm13] Habib M. Ammari. Joint kcoverage and data gathering in sparsely deployed sensor networks — impact of purposeful mobility and heterogeneity. ACM Transactions on Sensor Networks, 10(1):8:1–8:??, November 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ammari:2016:KCC

[Amm16] Habib M. Ammari. 3D-k Cov-ComFor: an energy-efficient framework for composite forwarding in three-dimensional duty-cycled k-covered wireless sensor networks. ACM Transactions on Sensor Networks, 12 (4):35:1–35:??, November 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ammari:2023:CGB

[Amm23] Habib M. Ammari. A computational geometry-based approach for planar k-coverage in wireless sensor networks. ACM Transactions on Sensor Networks, 19 (2):35:1–35:??, May 2023. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3564272.

Adu-Manu:2017:WQM

[AMTH^{+}17] Kofi Sarpong Adu-Manu, Cristiano Tapparello, Wendi Heinzelman, Ferdinand Apietu Katsriku, and Jamal-Deen Abdulai. Water quality monitoring using wireless sensor networks: Current trends and future research directions. ACM Transactions on Sensor Networks, 13(1):4:1– 4:??, February 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Alghamdi:2019:RAM

[ARWK19] Wael Alghamdi, Mohsen Rezvani, Hui Wu, and Salil S. Kanhere. Routing-aware and malicious node detection in a concealed data aggregation for WSNs. ACM Transactions on Sensor Networks, 15 (2):18:1–18:??, April 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3293537.

Babatunde:2024:GRE

[BAHS24] Simeon Babatunde, Arwa Alsubhi, Josiah Hester, and Jacob Sorber. Greentooth: Robust and energy efficient wireless networking for batteryless devices. ACM Transactions on Sensor Networks, 20(3):66:1–66:??, May 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3649221.

Bedogni:2017:PAF

[BAP^{+}17] Luca Bedogni, Andreas Achtzehn, Marina Petrova, Petri Mähönen, and Luciano Bononi. Performance assessment and feasibility analysis of IEEE 802.15.4m wireless sensor networks in TV grayspaces. ACM Transactions on Sensor Networks, 13(1):8:1– 8:??, February 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Bhatti:2016:EHW

[BASM16] Naveed Anwar Bhatti, Muhammad Hamad Alizai, Affan A. Syed, and Luca Mottola. Energy harvesting and wireless transfer in sensor network applications: Concepts and experiences. ACM Transactions on Sensor Networks, 12(3):24:1–24:??, August 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Baset:2023:ART

[BBD^{+}23] Aniqua Baset, Christopher Becker, Kurt Derr, Shamik Sarkar, and Sneha Kumar Kasera. AviSense: a real-time system for detection, classification, and analysis of aviation signals. ACM Transactions on Sensor Networks, 19(1):8:1– 8:35, February 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3526089.

Baddeley:2024:UCT

[BBEM^{+}24] Michael Baddeley, Carlo Alberto Boano, Antonio Escobar-Molero, Ye Liu, Xiaoyuan Ma, Victor Marot, Usman Raza, Kay Römer, Markus Schuss, and Aleksandar Stanoev. Understanding concurrent transmissions: The impact of carrier frequency offset and RF interference on physical layer performance. ACM Transactions on Sensor Networks, 20(1): 2:1–2:??, January 2024. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3604430.

Bartolini:2012:SAR

[BCL^{+}12] Novella Bartolini, Tiziana Calamoneri, Tom La Porta, Chiara Petrioli, and Simone Silvestri. Sensor activation and radius adaptation (SARA) in heterogeneous sensor networks. ACM Transactions on Sensor Networks, 8(3):24:1–24:??, July 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Bezerra:2022:AMP

[BCMY22] Pamela Bezerra, Po-Yu Chen, Julie A. McCann, and Weiren Yu. Adaptive monitor placement for near real-time node failure localisation in wireless sensor networks. ACM Transactions on Sensor Networks, 18(1):

2:1–2:41, February 2022. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3466639.

Barenboim:2014:DEO

[BDO14] Leonid Barenboim, Shlomi Dolev, and Rafail Ostrovsky. Deterministic and energy-optimal wireless synchronization. ACM Transactions on Sensor Networks, 11(1):13:1–13:??, August 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Biegelmeyer:2024:ESB

[BDP24] Anderson Biegelmeyer, Alexandre Dos Santos Roque, and Edison Pignaton de Freitas. An experimental study on BLE 5 mesh applied to public transportation. ACM Transactions on Sensor Networks, 20(3):59:1–59:??, May 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3647641.

Bruck:2009:LRS

[BGJ09] Jehoshua Bruck, Jie Gao, and Anxiao (Andrew) Jiang. Localization and routing in sensor networks by local angle information. ACM Transactions on Sensor Networks, 5(1):7:1–7:??, February 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Bagchi:2015:ORC

[BGMP15] Amitabha Bagchi, Sainyam Galhotra, Tarun Mangla, and Cristina M. Pinotti. Optimal radius for connectivity in duty-cycled wireless sensor networks. ACM Transactions on Sensor Networks, 11(2):36:1– 36:??, February 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Burman:2023:TFW

[BGP^{+}23] Scott G. Burman, Jingya Gao, Gregory B. Pasternack, Nann A. Fangue, Paul Cadrett, Elizabeth Campbell, and Dipak Ghosal. TempMesh — a flexible wireless sensor network for monitoring river temperatures. ACM Transactions on Sensor Networks, 19 (1):15:1–15:28, February 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3542697.

Bessos:2021:ISN

[BH21] Mai Ben Adar Bessos and Amir Herzberg. Intercepting a stealthy network. ACM Transactions on Sensor Networks, 17(2): 10:1–10:39, June 2021. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3431223.

Blumrosen:2013:ERB

[BHA^{+}13] Gaddi Blumrosen, Bracha Hod, Tal Anker, Danny Dolev, and Boris Rubinsky. Enhancing RSSI-based tracking accuracy in wireless sensor networks. ACM Transactions on Sensor Networks, 9(3):29:1–29:??, May 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Bhargava:2019:LFA

[BIMD19] Kriti Bhargava, Stepan Ivanov, Diarmuid McSweeney, and William Donnelly. Leveraging fog analytics for context-aware sensing in cooperative wireless sensor networks. ACM Transactions on Sensor Networks, 15 (2):23:1–23:??, April 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3306147.

Bashir:2018:MPC

[BIST18] Noman Bashir, David Irwin, Prashant Shenoy, and Jay Taneja. Mechanisms and policies for controlling distributed solar capacity. ACM Transactions on Sensor Networks, 14(3–4):25:1– 25:??, December 2018. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Basha:2015:NDS

[BJR15] Elizabeth Basha, Raja Jurdak, and Daniela Rus. In-network distributed solar current prediction. ACM Transactions on Sensor Networks, 11(2):23:1–23:??, February 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Bradbury:2022:TMG

[BJW^{+}22] Matthew Bradbury, Arshad Jhumka, Tim Watson, Denys Flores, Jonathan Burton, and Matthew Butler. Threatmodeling-guided trust-based task offloading for resourceconstrained Internet of Things. ACM Transactions on Sensor Networks, 18(2):29:1–29:41, May 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3510424.

Baccour:2012:RLQ

[BKM^{+}12] Nouha Baccour, Anis Koubâa, Luca Mottola, Marco Antonio Zúñiga, Habib Youssef, Carlo Alberto Boano, and Mário Alves. Radio link quality estimation in wireless sensor networks: a survey. ACM Transactions on Sensor Networks, 8 (4):34:1–34:??, September 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Bisdikian:2013:QVI

[BKS13] Chatschik Bisdikian, Lance M. Kaplan, and Mani B. Srivastava. On the quality and value of information in sensor networks. ACM Transactions on Sensor Networks, 9(4):48:1–48:??, July 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Brachmann:2019:WFF

[BLGS19] Martina Brachmann, Olaf Landsiedel, Diana Göhringer, and Silvia Santini. Whisper: Fast flooding for low-power wireless networks. ACM Transactions on Sensor Networks, 15(4):47:1– 47:??, October 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3356341.

Biswas:2006:SPB

[BLWY06] Pratik Biswas, Tzu-Chen Lian, Ta-Chung Wang, and Yinyu Ye. Semidefinite programming based algorithms for sensor network localization. ACM Transactions on Sensor Networks, 2(2):188– 220, May 2006. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Boers:2012:SCI

[BNG12] Nicholas M. Boers, Ioanis Nikolaidis, and Pawel Gburzynski. Sampling and classifying interference patterns in a wireless sensor network. ACM Transactions on Sensor Networks, 9(1): 2:1–2:??, November 2012. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Bui:2020:SBS

[BNN^{+}20] Nam Bui, Anh Nguyen, Phuc Nguyen, Hoang Truong, Ashwin Ashok, Thang Dinh, Robin Deterding, and Tam Vu. Smartphonebased SpO 2 measurement by exploiting wavelengths separation and chromophore compensation. ACM Transactions on Sensor Networks, 16(1):9:1– 9:30, February 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3360725.

Bhandari:2020:DLD

[BNPR20] Ravi Bhandari, Akshay Uttama Nambi, Venkata N. Padmanabhan, and Bhaskaran Raman. Driving lane detection on smartphones using deep neural networks. ACM Transactions on Sensor Networks, 16(1): 2:1–2:22, February 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3358797.

Busnel:2011:ADT

[BQB^{+}11] Yann Busnel, Leonardo Querzoni, Roberto Baldoni, Marin Bertier, and Anne-Marie Kermarrec. Analysis of deterministic tracking of multiple objects using a binary sensor network. ACM Transactions on Sensor Networks, 8(1):8:1–8:??, August 2011. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Bui:2015:SAS

[BR15] Nicola Bui and Michele Rossi. Staying alive: System design for self-sufficient sensor networks. ACM Transactions on Sensor Networks, 11(3):40:1– 40:??, February 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Brass:2007:BCT

[Bra07] Peter Brass. Bounds on coverage and target detection capabilities for models of networks of mobile sensors. ACM Transactions on Sensor Networks, 3(2): 9:1–9:??, June 2007. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Bhandari:2018:CCF

[BRR^{+}18] Ravi Bhandari, Bhaskaran Raman, K. K. Ramakrishnan, Deepthi Chander, Naveen Aggarwal, Divya Bansal, Mahima Choudhary, Nisha Moond, Aneesh Bansal, and Megha Chaudhary. CrowdLoc: Cellular fingerprinting for crowds by crowds. ACM Transactions on Sensor Networks, 14(1):4:1– 4:??, March 2018. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Branco:2015:TFS

[BSI^{+}15] Adriano Branco, Francisco Sant’anna, Roberto Ierusalimschy, Noemi Rodriguez, and Silvana Rossetto. Terra: Flexibility and safety in wireless sensor networks. ACM Transactions on Sensor Networks, 11(4):59:1– 59:??, December 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Bhotto:2018:NBS

[BT18] MD. Zulfiquar Ali Bhotto and Wee Peng Tay. Non-Bayesian social learning with observation reuse and soft switching. ACM Transactions on Sensor Networks, 14(2):14:1–14:??, July 2018. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Bhardwaj:2018:AAT

[BTR^{+}18] Romil Bhardwaj, Gopi Krishna Tummala, Ganesan Ramalingam, Ramachandran Ramjee, and Prasun Sinha. AutoCalib: Automatic traffic camera calibration at scale. ACM Transactions on Sensor Networks, 14(3– 4):19:1–19:??, December 2018. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Bhuiyan:2014:SPM

[BWCW14] Md Zakirul Alam Bhuiyan, Guojun Wang, Jiannong Cao, and Jie Wu. Sensor placement with multiple objectives for structural health monitoring. ACM Transactions on Sensor Networks, 10 (4):68:1–68:??, June 2014. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Bai:2024:IHB

[BWP^{+}24] Xuewei Bai, Yongcai Wang, Haodi Ping, Xiaojia Xu, Deying Li, and Shuo Wang. InferLoc: Hypothesis-based joint edge inference and localization in sparse sensor networks. ACM Transactions on Sensor Networks, 20(1):8:1–8:??, January 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3608477.

Buiquang:2019:BJD

[BY19] Chung Buiquang and Zhongfu Ye. Blind joint 2-D DOA/ symbols estimation for 3-D millimeter wave massive MIMO communication systems. ACM Transactions on Sensor Networks, 15(4):46:1–46:??, October 2019. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic).  URL https://dl.
acm.org/ft_gateway.cfm?id=
3352487.

Bagaa:2015:DLL

[BYD^{+}15] Miloud Bagaa, Mohamed Younis, Djamel Djenouri, Abdelouahid Derhab, and Nadjib Badache. Distributed low-latency data aggregation scheduling in wireless sensor networks. ACM Transactions on Sensor Networks, 11(3):49:1– 49:??, May 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Cao:2006:SLC

[CA06] Qing Cao and Tarek Abdelzaher. Scalable logical coordinates framework for routing in wireless sensor networks. ACM Transactions on Sensor Networks, 2(4): 557–593, November 2006. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Cui:2022:SES

[CA22] Hang Cui and Tarek Abdelzaher. SenseLens: an efficient social signal conditioning system for true event detection. ACM Transactions on Sensor Networks, 18(2):16:1–16:27, May 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3485047.

Chandio:2018:NWE

[CBSA18] Yasra Chandio, Jó Ágila Bitsch, Affan A. Syed, and Muhammad Hamad Alizai. Networking wireless energy in embedded networks. ACM Transactions on Sensor Networks, 14(2): 9:1–9:??, July 2018. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Chan:2011:SFP

[CC11] Aldar C-F. Chan and Claude Castelluccia. A security framework for privacy-preserving data aggregation in wireless sensor networks. ACM Transactions on Sensor Networks, 7(4):29:1– 29:??, February 2011. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Chen:2023:CDS

[CC23] Mu-Yen Chen and Po-Ru Chiang. COVID-19 diagnosis system based on chest X-ray images using optimized convolutional neural network. ACM Transactions on Sensor Networks, 19(3):53:1–53:??, August 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3558098.

Chen:2021:ECA

[CCC^{+}21] Quan Chen, Zhipeng Cai, Lianglun Cheng, Hong Gao, and Jianzhong Li. Energy-collision-aware minimum latency aggregation scheduling for energyharvesting sensor networks. ACM Transactions on Sensor Networks, 17(4):40:1–40:34, July 2021. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3461013.

Crespi:2008:TTA

[CCJ08] Valentino Crespi, George Cybenko, and Guofei Jiang. The theory of trackability with applications to sensor networks. ACM Transactions on Sensor Networks, 4(3):16:1–16:??, May 2008. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Castelluccia:2009:EPS

[CCMT09] Claude Castelluccia, Aldar C-F. Chan, Einar Mykletun, and Gene Tsudik. Efficient and provably secure aggregation of encrypted data in wireless sensor networks. ACM Transactions on Sensor Networks, 5(3):20:1– 20:??, May 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Chen:2021:RJA

[CD21] Gonglong Chen and Wei Dong.

Reactive jamming and attack mitigation over cross-technology communication links. ACM Transactions on Sensor Networks, 17(1):4:1–4:25, January 2021. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3418210.

Cao:2012:ITM

[CDGC12] Zhen Cao, Hui Deng, Zhi Guan, and Zhong Chen. Informationtheoretic modeling of false data filtering schemes in wireless sensor networks. ACM Transactions on Sensor Networks, 8(2):14:1– 14:??, March 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Chitnis:2008:AML

[CDR08] Laukik Chitnis, Alin Dobra, and Sanjay Ranka. Aggregation methods for large-scale sensor networks. ACM Transactions on Sensor Networks, 4(2):9:1– 9:??, March 2008. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Cichon:2018:ACA

[CG18] Jacek Cichoń and Karol Gotfryd. Average counting via approximate histograms. ACM Transactions on Sensor Networks, 14 (2):8:1–8:??, July 2018. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Chen:2019:TPO

[CGB^{+}19] Yueyue Chen, Deke Guo, MD Zakirul Alam Bhuiyan, Ming Xu, Guojun Wang, and Pin Lv. Towards profit optimization during online participant selection in compressive mobile crowdsensing. ACM Transactions on Sensor Networks, 15(4): 38:1–38:??, October 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3342515.

Choi:2012:NFE

[CGD12] Wook Choi, Giacomo Ghidini, and Sajal K. Das. A novel framework for energy-efficient data gathering with random coverage in wireless sensor networks. ACM Transactions on Sensor Networks, 8(4):36:1–36:??, September 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Carbunar:2006:RCD

[CGVC06] Bogdan Cărbunar, Ananth Grama, Jan Vitek, and Octavian Cărbunar. Redundancy and coverage detection in sensor networks. ACM Transactions on Sensor Networks, 2(1):94–128, February 2006. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Chen:2013:LBC

[CHN^{+}13] Phoebus Chen, Kirak Hong, Nikhil Naikal, S. Shankar Sastry, Doug Tygar, Posu Yan, Allen Y. Yang, Lung-Chung Chang, Leon Lin, Simon Wang, Edgar Lobatón, Songhwai Oh, and Parvez Ahammad. A low-bandwidth camera sensor platform with applications in smart camera networks. ACM Transactions on Sensor Networks, 9(2):21:1– 21:??, March 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Chakraborty:2023:FRS

[CHPP23] Roshni Chakraborty, Josefine Holm, Torben Bach Pedersen, and Petar Popovski. Finding representative sampling subsets in sensor graphs using timeseries similarities. ACM Transactions on Sensor Networks, 19 (4):89:1–89:32, November 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3595181.

Cauchi:2018:MSB

[CHSA18] Nathalie Cauchi, Khaza Anuarul Hoque, Marielle Stoelinga, and Alessandro Abate. Maintenance of smart buildings using fault trees. ACM Transactions on Sensor Networks, 14(3–4):28:1– 28:??, December 2018. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Cheng:2020:ICU

[CJL^{+}20] Yushi Cheng, Xiaoyu Ji, Xiaopeng Li, Tianchen Zhang, Sharaf Malebary, Xianshan Qu, and Wenyuan Xu. Identifying child users via touchscreen interactions. ACM Transactions on Sensor Networks, 16 (4):35:1–35:25, October 2020. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3403574.

Cai:2011:CSD

[CJS11] Haiyan Cai, Xiaohua Jia, and Mo Sha. Critical sensor density for partial connectivity in large area wireless sensor networks. ACM Transactions on Sensor Networks, 7(4):35:1– 35:??, February 2011. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Cevher:2009:ASN

[CK09] Volkan Cevher and Lance M.

Kaplan. Acoustic sensor network design for position estimation. ACM Transactions on Sensor Networks, 5(3):21:1–21:??, May 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Chong:2013:SLP

[CK13] Poh Kit Chong and Daeyoung Kim. Surface-level path loss modeling for sensor networks in flat and irregular terrain. ACM Transactions on Sensor Networks, 9(2):15:1–15:??, March 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Chang:2019:PPN

[CKHP19] Sang-Yoon Chang, Sristi Lakshmi Sravana Kumar, Yih-Chun Hu, and Younghee Park. Powerpositive networking: Wirelesscharging-based networking to protect energy against battery DoS attacks. ACM Transactions on Sensor Networks, 15(3): 27:1–27:??, August 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3317686.

Chen:2009:SRP

[CKL^{+}09] Yingying Chen, Konstantinos Kleisouris, Xiaoyan Li, Wade Trappe, and Richard P. Martin. A security and robustness performance analysis of localization algorithms to signal strength attacks. ACM Transactions on Sensor Networks, 5(1):2:1–2:??, February 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Chen:2013:EET

[CLH^{+}13] Jiming Chen, Junkun Li, Shibo He, Tian He, Yu Gu, and Youxian Sun. On energy-efficient trap coverage in wireless sensor networks. ACM Transactions on Sensor Networks, 10(1):2:1–2:??, November 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Cao:2023:LHD

[CLJ^{+}23] Hangcheng Cao, Daibo Liu, Hongbo Jiang, Ruize Wang, Zhe Chen, and Jie Xiong. LIPAuth: Hand-dependent light intensity patterns for resilient user authentication. ACM Transactions on Sensor Networks, 19(3): 64:1–64:??, August 2023. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3572909.

Cao:2023:TRD

[CLL^{+}23] Yetong Cao, Fan Li, Xiaochen Liu, Song Yang, and Yu Wang. Towards reliable driver drowsiness detection leveraging wearables. ACM Transactions on Sensor Networks, 19(2):39:1– 39:??, May 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3560821.

Cucuringu:2012:SNL

[CLS12] Mihai Cucuringu, Yaron Lipman, and Amit Singer. Sensor network localization by eigenvector synchronization over the Euclidean group. ACM Transactions on Sensor Networks, 8(3): 19:1–19:??, July 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Chang:2012:PRS

[CLSW12] Shih-Ying Chang, Yue-Hsun Lin, Hung-Min Sun, and Mu-En Wu. Practical RSA signature scheme based on periodical rekeying for wireless sensor networks. ACM Transactions on Sensor Networks, 8(2):13:1– 13:??, March 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Chen:2009:MGQ

[CLX09] Ai Chen, Ten H. Lai, and Dong Xuan. Measuring and guaranteeing quality of barrier coverage for general belts with wireless sensors. ACM Transactions on Sensor Networks, 6(1):2:1– 2:??, December 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Chang:2021:DDL

[CLX^{+}21] Xiangmao Chang, Gangkai Li, Guoliang Xing, Kun Zhu, and Linlin Tu. DeepHeart: a deep learning approach for accurate heart rate estimation from PPG signals. ACM Transactions on Sensor Networks, 17 (2):14:1–14:18, June 2021. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3441626.

Concas:2021:LCO

[CML^{+}21] Francesco Concas, Julien Mineraud, Eemil Lagerspetz, Samu Varjonen, Xiaoli Liu, Kai Puolamäki, Petteri Nurmi, and Sasu Tarkoma. Low-cost outdoor air quality monitoring and sensor calibration: a survey and critical analysis. ACM Transactions on Sensor Networks, 17 (2):20:1–20:44, June 2021. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3446005.

Chatterjea:2008:DSO

[CNMH08] Supriyo Chatterjea, Tim Nieberg, Nirvana Meratnia, and Paul Havinga. A distributed and self-organizing scheduling algorithm for energy-efficient data aggregation in wireless sensor networks. ACM Transactions on Sensor Networks, 4(4):20:1– 20:??, August 2008. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Choi:2016:DIM

[COP^{+}16] Woohyeok Choi, Jeungmin Oh, Taiwoo Park, Seongjun Kang, Miri Moon, Uichin Lee, Inseok Hwang, Darren Edge, and Junehwa Song. Designing interactive multiswimmer exergames: a case study. ACM Transactions on Sensor Networks, 12(3):17:1– 17:??, August 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Cardell-Oliver:2019:BAC

[COS19] Rachel Cardell-Oliver and Chayan Sarkar. BuildSense: Accurate, cost-aware, fault-tolerant monitoring with minimal sensor infrastructure. ACM Transactions on Sensor Networks, 15(3): 36:1–36:??, August 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3341171.

Corbalan:2020:UWC

[CP20] Pablo Corbalán and Gian Pietro Picco. Ultra-wideband concurrent ranging. ACM Transactions on Sensor Networks, 16 (4):41:1–41:41, October 2020. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3409477.

Costa:2006:DWM

[CPH06] Jose A. Costa, Neal Patwari, and Alfred O. Hero III. Distributed weighted-multidimensional scaling for node localization in sensor networks. ACM Transactions on Sensor Networks, 2(1): 39–64, February 2006. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Chiariotti:2020:GTA

[CPL^{+}20] Federico Chiariotti, Chiara Pielli, Nicola Laurenti, Andrea Zanella, and Michele Zorzi. A game-theoretic analysis of energy-depleting jamming attacks with a learning counterstrategy. ACM Transactions on Sensor Networks, 16(1):6:1– 6:25, February 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3365838.

Chen:2017:DEM

[CPP^{+}17] Xinlei Chen, Aveek Purohit, Shijia Pan, Carlos Ruiz, Jun Han, Zheng Sun, Frank Mokaya, Patric Tague, and Pei Zhang. Design experiences in minimalistic flying sensor node platform through SensorFly. ACM Transactions on Sensor Networks, 13 (4):33:1–33:??, December 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Chen:2023:ISS

[CPSS23] Mu-Yen Chen, Vincenzo Piuri, Alireza Souri, and Mohammad Shojafar. Introduction to the special section on Internet of behavior for emerging technologies. ACM Transactions on Sensor Networks, 19(2):23:1–23:3, May 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3589021.

Chang:2020:ISS

[CPX^{+}20] Xiangmao Chang, Cheng Peng, Guoliang Xing, Tian Hao, and Gang Zhou. iSleep: a smartphone system for unobtrusive sleep quality monitoring. ACM Transactions on Sensor Networks, 16(3):27:1– 27:32, August 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3392049.

Chen:2021:RND

[CQDW21] Ning Chen, Tie Qiu, Mahmoud Daneshmand, and Dapeng Oliver Wu. Robust networking: Dynamic topology evolution learning for Internet of Things. ACM Transactions on Sensor Networks, 17(3):28:1–28:23, June 2021. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3446937.

Cheng:2007:CBP

[CRW07] Maggie X. Cheng, Lu Ruan, and Weili Wu. Coverage breach problems in bandwidth-constrained sensor networks. ACM Transactions on Sensor Networks, 3(2): 12:1–12:??, June 2007. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Chin:2010:ILL

[CRY^{+}10] Jren-Chit Chin, Nageswara S. V.

Rao, David K. Y. Yau, Mallikarjun Shankar, Yong Yang, Jennifer C. Hou, Srinivasagopalan Srivathsan, and Sitharama Iyengar. Identification of low-level point radioactive sources using a sensor network. ACM Transactions on Sensor Networks, 7 (3):21:1–21:??, September 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Chen:2020:HDC

[CRZ^{+}20] Xinlei Chen, Carlos Ruiz, Sihan Zeng, Liyao Gao, Aveek Purohit, Stefano Carpin, and Pei Zhang. H-DrunkWalk: Collaborative and adaptive navigation for heterogeneous MAV swarm. ACM Transactions on Sensor Networks, 16(2): 20:1–20:27, April 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3382094.

Cheong:2017:AKK

[CS17] Se-Hang Cheong and Yain-Whar Si. Accelerating the Kamada– Kawai algorithm for boundary detection in a mobile ad hoc network. ACM Transactions on Sensor Networks, 13(1):3:1–3:??, February 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Cheong:2018:BND

[CS18] Se-Hang Cheong and Yain-Whar Si. Boundary node detection and unfolding of complex non-convex ad hoc networks. ACM Transactions on Sensor Networks, 14(1): 1:1–1:??, March 2018. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Cheng:2023:ATA

[CS23] Xia Cheng and Mo Sha. Autonomous traffic-aware scheduling for industrial wireless sensoractuator networks. ACM Transactions on Sensor Networks, 19 (2):38:1–38:??, May 2023. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3561056.

Chakrabarti:2006:CPO

[CSA06] Arnab Chakrabarti, Ashutosh Sabharwal, and Behnaam Aazhang. Communication power optimization in a sensor network with a path-constrained mobile observer. ACM Transactions on Sensor Networks, 2(3):297–324, August 2006. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Chen:2023:DIS

[CSLJ23] Tao Chen, Longfei Shangguan, Zhenjiang Li, and Kyle Jamieson. The design and implementation of a steganographic communication system over in-band acoustical channels. ACM Transactions on Sensor Networks, 19(4):90:1– 90:25, November 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3587162.

Chen:2019:BLS

[CT19] Kongyang Chen and Guang Tan. BikeGPS: Localizing shared bikes in street canyons with low-level GPS cooperation. ACM Transactions on Sensor Networks, 15(4):45:1–45:??, October 2019. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL https://dl.

                                        acm.org/ft_gateway.cfm?id=
                                        3343857.

Chen:2015:SSH

[CTW^{+}15] Jinzhu Chen, Rui Tan, Yu Wang, Guoliang Xing, Xiaorui Wang, Xiaodong Wang, Bill Punch, and Dirk Colbry. A sensor system for high-fidelity temperature distribution forecasting in data centers. ACM Transactions on Sensor Networks, 11(2):30:1–30:??, February 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Chang:2024:DCS

[CTWG24] Qingyi Chang, Dan Tao, Jiangtao Wang, and Ruipeng Gao. Deep compressed sensing based data imputation for urban environmental monitoring. ACM Transactions on Sensor Networks, 20(1):17:1–17:??, January 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3599236.

Cuevas:2013:SDS

[CUdVY13] Ángel Cuevas, Manuel Urueña, Gustavo de Veciana, and Aditya Yadav. STARR-DCS: Spatiotemporal adaptation of random replication for data-centric storage. ACM Transactions on Sensor Networks, 10(1):14:1–14:??, November 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Cheng:2009:DAN

[CVY09] Bing Hwa Cheng, Lieven Vandenberghe, and Kung Yao. Distributed algorithm for node localization in wireless ad-hoc networks. ACM Transactions on Sensor Networks, 6(1):8:1–8:??, December 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Cui:2022:ISS

[CWK^{+}22] Laizhong Cui, Yulei Wu, Ryan Ko, Alex Ladur, and Jianping Wu. Introduction to the special section on resiliency for AIenabled smart critical infrastructures for 5G and beyond. ACM Transactions on Sensor Networks, 18(3):40:1–40:??, August 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3538515.

Cai:2022:WTR

[CWS^{+}22] Haofan Cai, Ge Wang, Xiaofeng Shi, Junjie Xie, Minmei Wang, Chen Qian, and Shigang Chen. When tags ‘read’ each other: Enabling low-cost and convenient tag mutual identification. ACM Transactions on Sensor Networks, 18(2):22:1–22:22, May 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3494541.

Chen:2015:RMR

[CWY^{+}15] Li Chen, Jeremy Warner, Pak Lam Yung, Dawei Zhou, Wendi Heinzelman, Ilker Demirkol, Ufuk Muncuk, Kaushik Chowdhury, and Stefano Basagni. REACH 2-Mote: a rangeextending passive wake-up wireless sensor node. ACM Transactions on Sensor Networks, 11 (4):64:1–64:??, December 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Chen:2024:EAL

[CWY24] Yijie Chen, Jiliang Wang, and Jing Yang. Exploiting anchor links for NLOS combating in UWB localization. ACM Transactions on Sensor Networks, 20 (3):72:1–72:??, May 2024. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3657639.

Cai:2024:MUM

[CYD^{+}24] Xinjun Cai, Zheng Yang, Liang Dong, Qiang Ma, Xin Miao, and Zhuo Liu. Multi-user mobile augmented reality with ID-Aware visual interaction. ACM Transactions on Sensor Networks, 20(1):20:1–20:??, January 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3623638.

Carbunar:2010:QPW

[CYS^{+}10] Bogdan Carbunar, Yang Yu, Weidong Shi, Michael Pearce, and Venu Vasudevan. Query privacy in wireless sensor networks. ACM Transactions on Sensor Networks, 6(2):14:1– 14:??, February 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Cheng:2024:ECS

[CZC^{+}24] Yushi Cheng, Boyang Zhou, Yanjiao Chen, Yi-Chao Chen, Xiaoyu Ji, and Wenyuan Xu. Evaluating compressive sensing on the security of computer vision systems. ACM Transactions on Sensor Networks, 20 (3):56:1–56:??, May 2024. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3645093.

Cao:2023:CCO

[CZMM23] Zhichao Cao, Xiaolong Zheng, Qiang Ma, and Xin Miao. COFlood: Concurrent opportunistic flooding in asynchronous duty cycle networks. ACM Transactions on Sensor Networks, 19(3):58:1–58:??, August 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3570163.

Chang:2022:MBO

[CZX^{+}22] Xiangmao Chang, Jun Zhan, Guoliang Xing, Jun Huang, Bing Chen, and Lu Zhou.

Measurement-based optimization of cell selection in NB-IoT networks. ACM Transactions on Sensor Networks, 18(4):65:1– 65:??, November 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3544017.

Dong:2010:SRV

[DABNR10] Jing Dong, Kurt E. Ackermann, Brett Bavar, and Cristina Nita-Rotaru. Secure and robust virtual coordinate system in wireless sensor networks. ACM Transactions on Sensor Networks, 6(4):29:1–29:??, July 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Dong:2024:PSV

[DBC^{+}24] Yiwen Dong, Amelie Bonde, Jesse R. Codling, Adeola Bannis, Jinpu Cao, Asya Macon, Gary Rohrer, Jeremy Miles, Sudhendu Sharma, Tami Brown-Brandl, Akkarit Sangpetch, Orathai Sangpetch, Pei Zhang, and Hae Young Noh. PigSense: Structural vibration-based activity and health monitoring system for pigs. ACM Transactions on Sensor Networks, 20(1):1:1– 1:??, January 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3604806.

Doudou:2016:GTF

[DBOD^{+}16] Messaoud Doudou, Jose M. Barcelo-Ordinas, Djamel Djenouri, Jorge Garcia-Vidal, Abdelmadjid Bouabdallah, and Nadjib Badache. Game theory framework for MAC parameter optimization in energydelay constrained sensor networks. ACM Transactions on Sensor Networks, 12(2):10:1– 10:??, May 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Dong:2015:ORC

[DCBL15] Wei Dong, Chun Chen, Jiajun Bu, and Wen Liu. Optimizing relocatable code for efficient software update in networked embedded systems. ACM Transactions on Sensor Networks, 11(2): 22:1–22:??, February 2015. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ding:2024:EDR

[DCD24] Xianzhong Ding, Alberto Cerpa, and Wan Du. Exploring deep reinforcement learning for holistic smart building control. ACM Transactions on Sensor Networks, 20(3):70:1–70:??, May 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3656043.

Dietrich:2009:LWS

[DD09] Isabel Dietrich and Falko Dressler. On the lifetime of wireless sensor networks. ACM Transactions on Sensor Networks, 5(1):5:1–5:??, February 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Dereszynski:2011:SMD

[DD11] Ethan W. Dereszynski and Thomas G. Dietterich. Spatiotemporal models for dataanomaly detection in dynamic environmental monitoring campaigns. ACM Transactions on Sensor Networks, 8(1):3:1–3:??, August 2011. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

DiFrancesco:2011:DCW

[DDA11] Mario Di Francesco, Sajal K. Das, and Giuseppe Anastasi. Data collection in wireless sensor networks with mobile elements: a survey. ACM Transactions on Sensor Networks, 8(1): 7:1–7:??, August 2011. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Dutta:2012:MVE

[DDHC^{+}12] Prabal Dutta, Stephen Dawson-Haggerty, Yin Chen, Chieh-Jan Mike Liang, and Andreas Terzis. A-MAC: a versatile and efficient receiver-initiated link layer for low-power wireless. ACM Transactions on Sensor Networks, 8(4):30:1–30:??, September 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Dyo:2012:WDD

[DEM^{+}12] Vladimir Dyo, Stephen A. Ellwood, David W. Macdonald, Andrew Markham, Niki Trigoni, Ricklef Wohlers, Cecilia Mascolo, Bence Pásztor, Salvatore Scellato, and Kharsim Yousef. WILDSENSING: Design and deployment of a sustainable sensor network for wildlife monitoring. ACM Transactions on Sensor Networks, 8(4):29:1–29:??, September 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Deng:2009:MDF

[Den09] Jing Deng. Multihop/Direct Forwarding (MDF) for static wireless sensor networks. ACM Transactions on Sensor Networks, 5(4):35:1–35:??, November 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Dong:2016:THR

[DGS16] Jie Dong, Yu Ge, and David B.

Smith. Two-hop relay-assisted cooperative communication in wireless body area networks: an empirical study. ACM Transactions on Sensor Networks, 12 (4):32:1–32:??, November 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Donmez:2014:APC

[DIE14] Mehmet Yunus Donmez, Sinan Isik, and Cem Ersoy. Analysis of a prioritized contention model for multimedia wireless sensor networks. ACM Transactions on Sensor Networks, 10(2):36:1– 36:??, January 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Djidjev:2010:AAC

[Dji10] Hristo N. Djidjev. Approximation algorithms for computing minimum exposure paths in a sensor field. ACM Transactions on Sensor Networks, 7(3):23:1– 23:??, September 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

De:2009:DAM

[DLD09] Pradip De, Yonghe Liu, and Sajal K. Das. Deploymentaware modeling of node compromise spread in wireless sensor networks using epidemic theory. ACM Transactions on Sensor Networks, 5(3):23:1–23:??, May 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Dai:2023:OOP

[DLD^{+}23] Zhigang Dai, Wenjun Lyu, Yi Ding, Yiwei Song, and Yunhuai Liu. OPTI: Order preparation time inference for ondemand delivery. ACM Transactions on Sensor Networks, 19 (4):97:1–97:18, November 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3592610.

Dong:2021:THS

[DLG^{+}21] Wei Dong, Borui Li, Gaoyang Guan, Zhihao Cheng, Jiadong Zhang, and Yi Gao. TinyLink: a holistic system for rapid development of IoT applications. ACM Transactions on Sensor Networks, 17(1):2:1–2:29, January 2021. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3412366.

deLeo:2014:MVS

[dLM14] Carter de Leo and B. S. Manjunath. Multicamera video summarization and anomaly detection from activity motifs. ACM Transactions on Sensor Networks, 10(2):27:1–27:??, January 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Dressler:2016:MBW

[DML^{+}16] Falko Dressler, Margit Mutschlechner, Bijun Li, Rüdiger Kapitza, Simon Ripperger, Christopher Eibel, Benedict Herzog, Timo Hönig, and Wolfgang Schröder-Preikschat. Monitoring bats in the wild: On using erasure codes for energy-efficient wireless sensor networks. ACM Transactions on Sensor Networks, 12(1): 7:1–7:??, March 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

deOliveira:2023:MLT

[dOEC^{+}23] Leonardo L. de Oliveira, Gabriel H.

Eisenkraemer, Everton A. Carara, João B. Martins, and Jose Monteiro. Mobile localization techniques for wireless sensor networks: Survey and recommendations. ACM Transactions on Sensor Networks, 19(2):36:1– 36:??, May 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3561512.

Demetri:2019:LLA

[DPB19] Silvia Demetri, Gian Pietro Picco, and Lorenzo Bruzzone. LaPS: LiDAR-assisted placement of wireless sensor networks in forests. ACM Transactions on Sensor Networks, 15 (2):17:1–17:??, April 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

                                        URL https://dl.acm.org/ft_
                                        gateway.cfm?id=3293500.

Devarajan:2006:DMC

[DRC06] Dhanya Devarajan, Richard J. Radke, and Haeyong Chung. Distributed metric calibration of ad hoc camera networks. ACM Transactions on Sensor Networks, 2(3):380–403, August 2006. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Dezfouli:2017:RRT

[DRC17] Behnam Dezfouli, Marjan Radi, and Octav Chipara. REWIMO: a real-time and reliable lowpower wireless mobile network. ACM Transactions on Sensor Networks, 13(3):17:1–17:??, September 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Dezfouli:2014:CEM

[DRW^{+}14] Behnam Dezfouli, Marjan Radi, Kamin Whitehouse, Shukor Abd Razak, and Hwee-Pink Tan. CAMA: Efficient modeling of the capture effect for low-power wireless networks. ACM Transactions on Sensor Networks, 11(1):20:1–20:??, August 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Dimri:2020:BUB

[DSA^{+}20] Anuj Dimri, Harsimran Singh, Naveen Aggarwal, Bhaskaran Raman, K. K. Ramakrishnan, and Divya Bansal. BaroSense: Using barometer for road traffic congestion detection and path estimation with crowdsourcing. ACM Transactions on Sensor Networks, 16(1):4:1– 4:24, February 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3364697.

Dong:2016:SOC

[DSH16] Jie Dong, David B. Smith, and Leif W. Hanlen. Socially optimal coexistence of wireless body area networks enabled by a noncooperative game. ACM Transactions on Sensor Networks, 12 (4):26:1–26:??, November 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Du:2024:FVM

[DSZ^{+}24] Hongwei Du, Jingfang Su, Zhao Zhang, Zhenhua Duan, Cong Tian, and Ding-Zhu Du. Full view maximum coverage of camera sensors: Moving object monitoring. ACM Transactions on Sensor Networks, 20 (3):63:1–63:??, May 2024. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3649314.

Dong:2023:GNN

[DTW^{+}23] Guimin Dong, Mingyue Tang, Zhiyuan Wang, Jiechao Gao, Sikun Guo, Lihua Cai, Robert Gutierrez, Bradford Campbel, Laura E. Barnes, and Mehdi Boukhechba. Graph neural networks in IoT: a survey. ACM Transactions on Sensor Networks, 19(2):47:1–47:??, May 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3565973.

Deng:2022:RDS

[DTY^{+}22] Xianjun Deng, Yuan Tian, Lingzhi Yi, Laurence Tianruo Yang, Yunzhi Xia, Xiao Tang, and Chenlu Zhu. Resilient deployment of smart nodes for improving confident information coverage in 5G IoT. ACM Transactions on Sensor Networks, 18(3):44:1–44:??, August 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3526196.

Deligiannis:2014:PRW

[DVS^{+}14] Nikos Deligiannis, Frederik Verbist, Jürgen Slowack, Rik van de Walle, Peter Schelkens, and Adrian Munteanu. Progressively refined Wyner–Ziv video coding for visual sensors. ACM Transactions on Sensor Networks, 10(2):21:1–21:??, January 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Dai:2023:MMR

[DWF^{+}23] Yuanchao Dai, Jing Wu, Yuanzhao Fan, Jin Wang, Jianwei Niu, Fei Gu, and Shigen Shen. MSEva: a musculoskeletal rehabilitation evaluation system based on EMG signals. ACM Transactions on Sensor Networks, 19(1):6:1–6:23, February 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522739.

Dong:2021:ESC

[DXC^{+}21] Liang Dong, Jingao Xu, Guoxuan Chi, Danyang Li, Xinglin Zhang, Jianbo Li, Qiang Ma, and Zheng Yang. Enabling surveillance cameras to navigate. ACM Transactions on Sensor Networks, 17(4):35:1– 35:20, November 2021. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3446633.

Du:2015:SPM

[DXL^{+}15] Wan Du, Zikun Xing, Mo Li, Bingsheng He, Lloyd Hock Chye Chua, and Haiyan Miao. Sensor placement and measurement of wind for water quality studies in urban reservoirs. ACM Transactions on Sensor Networks, 11(3): 41:1–41:??, February 2015. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ebrahimi:2015:NCA

[EA15] Dariush Ebrahimi and Chadi Assi. Network coding-aware compressive data gathering for energy-efficient wireless sensor networks. ACM Transactions on Sensor Networks, 11(4):61:1– 61:??, December 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Erickson:2014:OMP

[ECPC14] Varick L. Erickson, Miguel Á. Carreira-Perpiñán, and Alberto E. Cerpa. Occupancy modeling and prediction for building energy management. ACM Transactions on Sensor Networks, 10(3):42:1–42:??, April 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Efrat:2010:FDA

[EFI^{+}10] Alon Efrat, David Forrester, Anand Iyer, Stephen G. Kobourov, Cesim Erten, and Ozan Kilic. Force-directed approaches to sensor localization. ACM Transactions on Sensor Networks, 7 (3):27:1–27:??, September 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ercan:2013:OTP

[EGG13] Ali O. Ercan, Abbas El Gamal, and Leonidas J. Guibas. Object tracking in the presence of occlusions using multiple cameras: a sensor network approach. ACM Transactions on Sensor Networks, 9(2):16:1–16:??, March 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Edara:2008:ANP

[ELR08] Pavan Edara, Ashwin Limaye, and Krithi Ramamritham. Asynchronous in-network prediction: Efficient aggregation in sensor networks. ACM Transactions on Sensor Networks, 4(4): 25:1–25:??, August 2008. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Elhoseny:2022:USM

[ELR^{+}22] Mohamed Elhoseny, Abdullah Lakhan, Ahmed Rashid, Mazin Mohammed, and Karrar Abdulkareem. Underwater sensor multi-parameter scheduling for heterogeneous computing nodes. ACM Transactions on Sensor Networks, 18(3):35:1– 35:??, August 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3476513.

Esterle:2014:SEV

[ELYR14] Lukas Esterle, Peter R. Lewis, Xin Yao, and Bernhard Rinner. Socio-economic vision graph generation and handover in distributed smart camera networks. ACM Transactions on Sensor Networks, 10(2):20:1–20:??, January 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Eswaran:2012:UBB

[EMBP12] Sharanya Eswaran, Archan Misra, Flavio Bergamaschi, and Thomas La Porta. Utilitybased bandwidth adaptation in mission-oriented wireless sensor networks. ACM Transactions on Sensor Networks, 8(2):17:1– 17:??, March 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Eisenman:2009:BMS

[EML^{+}09] Shane B. Eisenman, Emiliano Miluzzo, Nicholas D. Lane, Ronald A. Peterson, Gahng-Seop Ahn, and Andrew T. Campbell. BikeNet: a mobile sensing system for cyclist experience mapping. ACM Transactions on Sensor Networks, 6(1): 6:1–6:??, December 2009. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Eslami:2013:RFW

[ENPNF13] Ali Eslami, Mohammad Nekoui, Hossein Pishro-Nik, and Faramarz Fekri. Results on finite wireless sensor networks: Connectivity and coverage. ACM Transactions on Sensor Networks, 9(4):51:1–51:??, July 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Erdem:2012:EPH

[ES12] Uğur Murat Erdem and Stan Sclaroff. Event prediction in a hybrid camera network. ACM Transactions on Sensor Networks, 8(2):16:1–16:??, March 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Emokpae:2014:SRB

[EY14] Lloyd Emokpae and Mohamed Younis. Surface-reflection-based communication and localization in underwater sensor networks. ACM Transactions on Sensor Networks, 10(3):50:1– 50:??, April 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Fraternali:2020:AAC

[FBAG20] Francesco Fraternali, Bharathan Balaji, Yuvraj Agarwal, and Rajesh K. Gupta. ACES: Automatic configuration of energy harvesting sensors with reinforcement learning. ACM Transactions on Sensor Networks, 16 (4):36:1–36:31, October 2020. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3404191.

Fierro:2018:DAQ

[FC18] Gabe Fierro and David E.

Culler. Design and analysis of a query processor for Brick. ACM Transactions on Sensor Networks, 14(3–4):18:1–18:??, December 2018. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Fan:2022:IBM

[FHST22] Chun-I Fan, Ya-Wen Hsu, Cheng-Han Shie, and Yi-Fan Tseng. ID-based multireceiver homomorphic proxy reencryption in federated learning. ACM Transactions on Sensor Networks, 18(4):55:1– 55:??, November 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3540199.

Funke:2006:SID

[FKMS06] Stefan Funke, Alexander Kesselman, Ulrich Meyer, and Michael Segal. A simple improved distributed algorithm for minimum CDS in unit disk graphs. ACM Transactions on Sensor Networks, 2(3):444–453, August 2006. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Fang:2023:TTP

[FLCH23] Hao Fang, Yiwei Liu, Chi-Hua Chen, and Feng-Jang Hwang. Travel time prediction method based on spatial-feature-based hierarchical clustering and deep multi-input gated recurrent unit. ACM Transactions on Sensor Networks, 19(2):26:1–26:21, May 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3544976.

Fu:2013:TBE

[FLFW13] Huai-Lei Fu, Phone Lin, Yuguang Fang, and Ting-Yu Wang. Trade-off between energy efficiency and report validity for mobile sensor networks. ACM Transactions on Sensor Networks, 9(4):49:1–49:??, July 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Feng:2013:EED

[FLJ^{+}13] Jing Feng, Yung-Hsiang Lu, Byunghoo Jung, Dimitrios Peroulis, and Y. Charlie Hu. Energy-efficient data dissemination using beamforming in wireless sensor networks. ACM Transactions on Sensor Networks, 9(3):31:1–31:??, May 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Fan:2014:BCS

[FLS^{+}14] Haosheng Fan, Minming Li, Xianwei Sun, Peng-Jun Wan, and Yingchao Zhao. Barrier coverage by sensors with adjustable ranges. ACM Transactions on Sensor Networks, 11(1):14:1– 14:??, August 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Fortuna:2015:FDC

[FM15] Carolina Fortuna and Mihael Mohorcic. A framework for dynamic composition of communication services. ACM Transactions on Sensor Networks, 11(2): 32:1–32:??, February 2015. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Fierro:2020:MOT

[FPA^{+}20] Gabe Fierro, Marco Pritoni, Moustafa Abdelbaky, Daniel Lengyel, John Leyden, Anand Prakash, Pranav Gupta, Paul Raftery, Therese Peffer, Greg Thomson, and David E. Culler. Mortar: an open testbed for portable building analytics. ACM Transactions on Sensor Networks, 16(1):7:1– 7:31, February 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3366375.

Forte:2013:TAS

[FS13] Domenic Forte and Ankur Srivastava. Thermal-aware sensor scheduling for distributed estimation. ACM Transactions on Sensor Networks, 9(4):53:1– 53:??, July 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Feldman:2015:IGS

[FSSR15] Dan Feldman, Cynthia Sung, Andrew Sugaya, and Daniela Rus. iDiary: From GPS signals to a text-searchable diary. ACM Transactions on Sensor Networks, 11(4):60:1–60:??, December 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Fan:2023:EDP

[FSTH23] Chun-I Fan, Cheng-Han Shie, Yi-Fan Tseng, and Hui-Chun Huang. An efficient data protection scheme based on hierarchical ID-based encryption for MQTT. ACM Transactions on Sensor Networks, 19(3):

61:1–61:??, August 2023. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3570506.

Fan:2024:BDW

[FSTH24] Boyu Fan, Xiang Su, Sasu Tarkoma, and Pan Hui. Behave differently when clustering: a semi-asynchronous federated learning approach for IoT. ACM Transactions on Sensor Networks, 20(3):51:1–51:??, May 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3639825.

Fragouli:2006:CCT

[FT06] Christina Fragouli and Tarik Tabet. On conditions for constant throughput in wireless networks. ACM Transactions on Sensor Networks, 2(3):359–379, August 2006. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Fan:2023:NPT

[FWF^{+}23] Jinxiao Fan, Pengfei Wang, Yu Fan, Liang Liu, and Huadong Ma. Num2vec: Pre-training numeric representations for time series forecasting in the sensing system. ACM Transactions on Sensor Networks, 19(4):94:1– 94:23, November 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3599728.

Goldberg:2006:VIE

[GAJ^{+}06] David H. Goldberg, Andreas G. Andreou, Pedro Julián, Philippe O. Pouliquen, Laurence Riddle, and Rich Rosasco. VLSI implementation of an energy-aware wake-up detector for an acoustic surveillance sensor network. ACM Transactions on Sensor Networks, 2(4):594–611, November 2006. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Gazi:2022:RLB

[GAMW22] Firoj Gazi, Nurzaman Ahmed, Sudip Misra, and Wei Wei. Reinforcement learning-based MAC protocol for underwater multimedia sensor networks. ACM Transactions on Sensor Networks, 18(3):37:1–37:??, August 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3484201.

Ganeriwal:2008:RBF

[GBS08] Saurabh Ganeriwal, Laura K. Balzano, and Mani B. Srivastava. Reputation-based framework for high integrity sensor networks. ACM Transactions on Sensor Networks, 4(3):15:1– 15:??, May 2008. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ghosh:2017:MBY

[GCAK17] Avishek Ghosh, Arpan Chattopadhyay, Anish Arora, and Anurag Kumar. Measurement based as-you-go deployment of two-connected wireless relay networks. ACM Transactions on Sensor Networks, 13(3):23:1– 23:??, September 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ganesan:2006:PES

[GCBL06] Deepak Ganesan, Razvan Cristescu, and Baltasar Beferull-Lozano. Power-efficient sensor placement and transmission structure for data gathering under distortion constraints. ACM Transactions on Sensor Networks, 2(2):155– 181, May 2006. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Gabale:2012:PMT

[GCRB12] Vijay Gabale, Kameswari Chebrolu, Bhaskaran Raman, and Sagar Bijwe. PIP: a multichannel, TDMA-based MAC for efficient and scalable bulk transfer in sensor networks. ACM Transactions on Sensor Networks, 8 (4):28:1–28:??, September 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Giaretta:2022:SCF

[GDM22] Alberto Giaretta, Nicola Dragoni, and Fabio Massacci. S×C4IoT: a security-by-contract framework for dynamic evolving IoT devices. ACM Transactions on Sensor Networks, 18(1):12:1– 12:51, February 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3480462.

Gelenbe:2007:DMP

[Gel07]  Erol Gelenbe.  A diffusion

model for packet travel time in a random multihop medium. ACM Transactions on Sensor Networks, 3(2):10:1–10:??, June 2007. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Gnawali:2013:CER

[GFJ^{+}13] Omprakash Gnawali, Rodrigo Fonseca, Kyle Jamieson, Maria Kazandjieva, David Moss, and Philip Levis. CTP: an efficient, robust, and reliable collection tree protocol for wireless sensor networks. ACM Transactions on Sensor Networks, 10(1):16:1– 16:??, November 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Guo:2022:TEC

[GHZ^{+}22] Xiuzhen Guo, Yuan He, Jia Zhang, Haotian Jiang, Zihao Yu, and Xin Na. Taming the errors in cross-technology communication: a probabilistic approach. ACM Transactions on Sensor Networks, 18(1): 3:1–3:20, February 2022. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3469031.

Gruenwedel:2014:LCS

[GJNC^{+}14] Sebastian Gruenwedel, Vedran Jelaca, Jorge Oswaldo Nino-Castaneda, Peter van Hese, Dimitri van Cauwelaert, Dirk van Haerenborgh, Peter Veelaert, and Wilfried Philips. Lowcomplexity scalable distributed multicamera tracking of humans. ACM Transactions on Sensor Networks, 10(2):24:1–24:??, January 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Gu:2022:AAS

[GJT^{+}22] Chaojie Gu, Linshan Jiang, Rui Tan, Mo Li, and Jun Huang. Attack-aware synchronizationfree data timestamping in Lo-RaWAN. ACM Transactions on Sensor Networks, 18(1):10:1– 10:31, February 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3474368.

Griffiths:2017:EDS

[GKRW17] Erin Griffiths, Avinash Kalyanaraman, Juhi Ranjan, and Kamin Whitehouse. An empirical design space analysis of doorway tracking systems for real-world environments. ACM Transactions on Sensor Networks, 13(4):26:1–26:??, December 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Gamage:2023:LEC

[GLG^{+}23] Amalinda Gamage, Jansen Liando, Chaojie Gu, Rui Tan, Mo Li, and Olivier Seller. LMAC: Efficient carrier-sense multiple access for LoRa. ACM Transactions on Sensor Networks, 19(2):44:1–44:??, May 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3564530.

Gong:2024:WSW

[GLL^{+}24] Liangyi Gong, Hao Lin, Daibo Liu, Lanqi Yang, Hongyi Wang, Jiaxing Qiu, Zhenhua Li, and Feng Qian. Who should we blame for Android app crashes? An in-depth study at scale and practical resolutions. ACM Transactions on Sensor Networks, 20(3):62:1–62:??, May 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3649895.

Ghosh:2022:SSE

[GLQ^{+}22] Pradipta Ghosh, Xiaochen Liu, Hang Qiu, Marcos A. M. Vieira, Gaurav S. Sukhatme, and Ramesh Govindan. Sensing the sensor: Estimating camera properties with minimal information. ACM Transactions on Sensor Networks, 18(2):28:1– 28:26, May 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3508393.

Ghadimi:2014:ORL

[GLS^{+}14] Euhanna Ghadimi, Olaf Landsiedel, Pablo Soldati, Simon Duquennoy, and Mikael Johansson. Opportunistic routing in low duty-cycle wireless sensor networks. ACM Transactions on Sensor Networks, 10(4):67:1– 67:??, June 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ghaffarkhah:2014:DNC

[GM14] Alireza Ghaffarkhah and Yasamin Mostofi. Dynamic networked coverage of time-varying environments in the presence of fading communication channels. ACM Transactions on Sensor Networks, 10(3):45:1– 45:??, April 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Gupta:2008:EGC

[GNDC08] Himanshu Gupta, Vishnu Navda, Samir Das, and Vishal Chowdhary. Efficient gathering of correlated data in sensor networks. ACM Transactions on Sensor Networks, 4(1):4:1–4:??, January 2008. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Guha:2012:ALT

[GPL^{+}12] Santanu Guha, Kurt Plarre, Daniel Lissner, Somnath Mitra, Bhagavathy Krishna, Prabal Dutta, and Santosh Kumar. AutoWitness: Locating and tracking stolen property while tolerating GPS and radio outages. ACM Transactions on Sensor Networks, 8(4):31:1– 31:??, September 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Girod:2007:ESE

[GRE^{+}07] Lewis Girod, Nithya Ramanathan, Jeremy Elson, Thanos Stathopoulos, Martin Lukac, and Deborah Estrin. Emstar: a software environment for developing and deploying heterogeneous sensor-actuator networks. ACM Transactions on Sensor Networks, 3(3):13:1–13:??, August 2007. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ghahroudi:2023:DND

[GSGA23] Mahsa Sadeghi Ghahroudi, Alireza Shahrabi, Seyed Mohammad Ghoreyshi, and Faisal Abdulaziz Alfouzan. Distributed node deployment algorithms in mobile wireless sensor networks: Survey and challenges. ACM Transactions on Sensor Networks, 19(4):91:1–91:26, November 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3579034.

Gao:2010:CLC

[GSL10] Jie Gao, Radu Sion, and Sol Lederer. Collaborative location certification for sensor networks. ACM Transactions on Sensor Networks, 6(4):30:1–30:??, July 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Grochla:2022:EAA

[GSM^{+}22] Krzysztof Grochla, Anna Strzoda, Rafał Marjasz, Przemysław Głomb, Kamil Ksiazek, and Zbigniew Laskarzewski. Energy-aware algorithm for assignment of relays in LP WAN. ACM Transactions on Sensor Networks, 18(4):60:1–60:??, November 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3544561.

Gandhi:2009:CEM

[GSW09] Sorabh Gandhi, Subhash Suri, and Emo Welzl. Catching elephants with mice: Sparse sampling for monitoring sensor networks. ACM Transactions on Sensor Networks, 6(1):1:1–1:??, December 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Gu:2019:OHB

[GTL19] Chaojie Gu, Rui Tan, and Xin Lou. One-hop out-of-band control planes for multi-hop wireless sensor networks. ACM Transactions on Sensor Networks, 15(4): 40:1–40:??, October 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3342100.

Gu:2024:RTR

[GWS^{+}24] Xiaolin Gu, Wenjia Wu, Aibo Song, Ming Yang, Zhen Ling, and Junzhou Luo. RF-TESI: Radio frequency fingerprint-based smartphone identification under temperature variation. ACM Transactions on Sensor Networks, 20(2):41:1–41:??, March 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3636462.

Gai:2022:DTE

[GXQ^{+}22] Keke Gai, Qiang Xiao, Meikang Qiu, Guolei Zhang, Jianyu Chen, Yihang Wei, and Yue Zhang. Digital twin-enabled AI enhancement in smart critical infrastructures for 5G. ACM Transactions on Sensor Networks, 18(3):45:1–45:??, August 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3526195.

Guo:2023:BFG

[GYG^{+}23] Zhengxin Guo, Wenyang Yuan, Linqing Gui, Biyun Sheng, and Fu Xiao. BreatheBand: a fine-grained and robust respiration monitor system using WiFi signals. ACM Transactions on Sensor Networks, 19 (4):82:1–82:18, November 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3582079.

Gao:2016:NSS

[GYNY16] Mingjie Gao, Ka-Fai Cedric Yiu, Sven Nordholm, and Yinyu Ye. On a new SDP–SOCP method for acoustic source localization problem. ACM Transactions on Sensor Networks, 12(4):36:1– 36:??, November 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Gunia:2023:ADM

[GZJE23] Marco Gunia, Adrian Zinke, Niko Joram, and Frank Ellinger. Analysis and design of a MuSiCbased angle of arrival positioning system. ACM Transactions on Sensor Networks, 19(3): 66:1–66:??, August 2023. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3577927.

Gao:2023:APM

[GZK^{+}23] Honghao Gao, Lin Zhou, Jung Yoon Kim, Ying Li, and Wanqiu Huang. Applying probabilistic model checking to the behavior guidance and abnormality detection for A-MCI patients under wireless sensor network. ACM Transactions on Sensor Networks, 19(3):48:1– 48:??, August 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3499426.

Guo:2014:DFN

[GZZ^{+}14] Shuo Guo, Heng Zhang, Ziguo Zhong, Jiming Chen, Qing Cao, and Tian He. Detecting faulty nodes with data errors for wireless sensor networks. ACM Transactions on Sensor Networks, 10(3):40:1–40:??, April 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Guo:2023:LCD

[GZZ^{+}23] Kehua Guo, Feihong Zhu, Xiaokang Zhou, Lingyan Zhang, Yifei Wang, and Jian Kang. LesionTalk: Core data extraction and multi-class lesion detection in IoT-based intelligent healthcare. ACM Transactions on Sensor Networks, 19(3): 50:1–50:??, August 2023. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3526194.

Hessel:2022:LSE

[HAH22] Frank Hessel, Lars Almon, and Matthias Hollick. LoRaWAN security: an evolvable survey on vulnerabilities, attacks and their systematic mitigation. ACM Transactions on Sensor Networks, 18(4):70:1– 70:??, November 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3561973.

Hauer:2014:LHM

[Hau14] Jan-Hinrich Hauer. Leveraging human mobility for communication in body area networks. ACM Transactions on Sensor Networks, 10(3):39:1– 39:??, April 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Hu:2009:DEH

[HBC^{+}09] Wen Hu, Nirupama Bulusu, Chun Tung Chou, Sanjay Jha, Andrew Taylor, and Van Nghia Tran. Design and evaluation of a hybrid sensor network for cane toad monitoring. ACM Transactions on Sensor Networks, 5(1): 4:1–4:??, February 2009. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Hariharan:2014:ESF

[HBKP14] Srikanth Hariharan, Chatschik Bisdikian, Lance M. Kaplan, and Tien Pham. Efficient solutions framework for optimal multitask resource assignments for data fusion in wireless sensor networks. ACM Transactions on Sensor Networks, 10(3): 48:1–48:??, April 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Huang:2005:FFA

[HBLR05] Qingfeng Huang, Sangeeta Bhattacharya, Chenyang Lu, and Gruia-Catalin Roman. FAR: Face-Aware Routing for mobicast in large-scale sensor networks. ACM Transactions on Sensor Networks, 1(2):240–271, November 2005. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Hu:2018:SIC

[HBW^{+}18] Chuang Hu, Wei Bao, Dan Wang, Yi Qian, Muqiao Zheng, and Shi Wang. sTube+: an IoT communication sharing architecture for smart after-sales maintenance in buildings. ACM Transactions on Sensor Networks, 14(3–4):29:1–29:??, December 2018. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Hsieh:2015:EBC

[HCL15] Hung-Yun Hsieh, Chih-Hua Chang, and Wei-Chih Liao. Not every bit counts: Data-centric resource allocation for correlated data gathering in machine-to-machine wireless networks. ACM Transactions on Sensor Networks, 11(2):38:1–38:??, February 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Huang:2009:SSF

[HCXT09] Pei Huang, Hongyang Chen, Guoliang Xing, and Yongdong Tan. SGF: a state-free gradientbased forwarding protocol for wireless sensor networks. ACM Transactions on Sensor Networks, 5(2):14:1–14:??, March 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Han:2017:TTA

[HF17] Yu Han and Yunsi Fei. TARS: a traffic-adaptive receiver-synchronized MAC protocol for underwater sensor networks. ACM Transactions on Sensor Networks, 13 (4):27:1–27:??, December 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

He:2019:EBS

[HKG^{+}19] Liang He, Linghe Kong, Yu Gu, Cong Liu, Tian He, and Kang G.

Shin. Extending battery system operation via adaptive reconfiguration. ACM Transactions on Sensor Networks, 15(1): 11:1–11:??, February 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3284556.

He:2006:VIS

[HKL^{+}06] Tian He, Sudha Krishnamurthy, Liqian Luo, Ting Yan, Lin Gu, Radu Stoleru, Gang Zhou, Qing Cao, Pascal Vicaire, John A. Stankovic, Tarek F. Abdelzaher, Jonathan Hui, and Bruce Krogh. VigilNet: an integrated sensor network system for energy-efficient surveillance. ACM Transactions on Sensor Networks, 2(1):1–38, February 2006. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Huang:2024:SDS

[HKW^{+}24] Junqin Huang, Linghe Kong, Jingwei Wang, Guihai Chen, Jianhua Gao, Gang Huang, and Muhammad Khurram Khan. Secure data sharing over vehicular networks based on multisharding blockchain. ACM Transactions on Sensor Networks, 20(2):31:1–31:??, March 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3579035.

He:2017:ISA

[HL17] Bin He and Gang Li. Intelligent self-adaptation data behavior control inspired by speech acts. ACM Transactions on Sensor Networks, 13(2):13:1–13:??, June 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Huang:2023:RPR

[HLL^{+}23] Qianyi Huang, Youjing Lu, Zhicheng Luo, Hao Wang, Fan Wu, Guihai Chen, and Qian Zhang. Rethinking privacy risks from wireless surveillance camera. ACM Transactions on Sensor Networks, 19(3):60:1– 60:??, August 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3570504.

He:2011:PPP

[HLN^{+}11] Wenbo He, Xue Liu, Hoang Viet Nguyen, Klara Nahrstedt, and Tarek Abdelzaher. PDA: Privacy-preserving data aggregation for information collection. ACM Transactions on Sensor Networks, 8(1):6:1–6:??, August 2011. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Huang:2006:DEC

[HLTC06] Chi-Fu Huang, Li-Chu Lo, Yu-Chee Tseng, and Wen-Tsuen Chen. Decentralized energy-conserving and coveragepreserving protocols for wireless sensor networks. ACM Transactions on Sensor Networks, 2 (2):182–187, May 2006. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Hong:2024:DIA

[HLZ^{+}24] Zhen Hong, Lingling Lu, Dehua Zheng, Jiahui Suo, Peng Sun, Raheem Beyah, and Zhenyu Wen. Detect insider attacks in industrial cyberphysical systems using multiphysical features-based fingerprinting. ACM Transactions on Sensor Networks, 20(2):29:1– 29:??, March 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3582691.

Hoang:2007:CBC

[HM07a] Anh Tuan Hoang and Mehul Motani. Collaborative broadcasting and compression in cluster-based wireless sensor networks. ACM Transactions on Sensor Networks, 3(3):17:1– 17:??, August 2007. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Huang:2007:SPK

[HM07b] Dijiang Huang and Deep Medhi.

Secure pairwise key establishment in large-scale sensor networks: an area partitioning and multigroup key predistribution approach. ACM Transactions on Sensor Networks, 3(3):16:1– 16:??, August 2007. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Huang:2024:SSE

[HMG^{+}24] Haiyang Huang, Tianhui Meng, Jianxiong Guo, Xuekai Wei, and Weijia Jia. SecEG: a secure and efficient strategy against DDoS attacks in mobile edge computing. ACM Transactions on Sensor Networks, 20 (3):55:1–55:??, May 2024. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3641106.

Harb:2017:DBD

[HMLJ17] Hassan Harb, Abdallah Makhoul, David Laiymani, and Ali Jaber. A distance-based data aggregation technique for periodic sensor networks. ACM Transactions on Sensor Networks, 13(4):32:1– 32:??, December 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Han:2018:SHO

[HPS^{+}18] Jun Han, Shijia Pan, Manal Kumar Sinha, Hae Young Noh, Pei Zhang, and Patrick Tague. Smart home occupant identification via sensor fusion across onobject devices. ACM Transactions on Sensor Networks, 14(3– 4):23:1–23:??, December 2018. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Huang:2013:CEA

[HR13] Xiaolong Huang and Izhak Rubin. Capacity- and energy-aware activation of sensor nodes for area phenomenon reproduction using wireless network transport. ACM Transactions on Sensor Networks, 9(4):52:1–52:??, July 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Hossain:2016:NDM

[HSD16] A. K. M. Mahtab Hossain, Cormac J. Sreenan, and Rodolfo De Paz Alberola. Neighbourdisjoint multipath for low-power and lossy networks. ACM Transactions on Sensor Networks, 12(3):23:1–23:??, August 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Hare:2021:PRP

[HSGW21] James Z. Hare, Junnan Song, Shalabh Gupta, and Thomas A. Wettergren. POSE.R: Predictionbased opportunistic sensing for resilient and efficient sensor networks. ACM Transactions on Sensor Networks, 17(1):5:1– 5:41, January 2021. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3419755.

Hu:2015:SSB

[HSL^{+}15] Shaohan Hu, Lu Su, Hengchang Liu, Hongyan Wang, and Tarek F. Abdelzaher. SmartRoad: Smartphone-based crowd sensing for traffic regulator detection and identification. ACM Transactions on Sensor Networks, 11 (4):55:1–55:??, December 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Hester:2017:RRE

[HSSS17] Josiah Hester, Lanny Sitanayah, Timothy Scott, and Jacob Sorber. Realistic and repeatable emulation of energy harvesting environments. ACM Transactions on Sensor Networks, 13(2):16:1– 16:??, June 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Hu:2010:TTW

[HTC^{+}10] Wen Hu, Hailun Tan, Peter Corke, Wen Chan Shih, and Sanjay Jha. Toward trusted wireless sensor networks. ACM Transactions on Sensor Networks, 7(1): 5:1–5:??, August 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Huang:2007:DPE

[HTW07] Chi-Fu Huang, Yu-Chee Tseng, and Hsiao-Lu Wu. Distributed protocols for ensuring both coverage and connectivity of a wireless sensor network. ACM Transactions on Sensor Networks, 3 (1):??, March 2007. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Huang:2024:CTP

[HWF^{+}24] Ziyao Huang, Weiwei Wu, Chenchen Fu, Xiang Liu, Feng Shan, Jianping Wang, and Xueyong Xu. Communicationtopology-preserving motion planning: Enabling static routing in UAV networks. ACM Transactions on Sensor Networks, 20 (1):24:1–24:??, January 2024.

CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3631530.

Huang:2020:CEC

[HWS^{+}20] Ziyao Huang, Weiwei Wu, Feng Shan, Yuxin Bian, Kejie Lu, Zhenjiang Li, Jianping Wang, and Jin Wang. CoUAS: Enable cooperation for unmanned aerial systems. ACM Transactions on Sensor Networks, 16(3): 24:1–24:19, August 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3388323.

Holland:2011:OPL

[HWT^{+}11] Matthew Holland, Tianqi Wang, Bulent Tavli, Alireza Seyedi, and Wendi Heinzelman. Optimizing physical-layer parameters for wireless sensor networks. ACM Transactions on Sensor Networks, 7(4):28:1–28:??, February 2011. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Huang:2022:EET

[HWT^{+}22] Zijie Huang, Yulei Wu, Niccolò Tempini, Hui Lin, and Hao Yin. An energy-efficient and trustworthy unsupervised anomaly detection framework (EATU) for IIoT. ACM Transactions on Sensor Networks, 18 (4):56:1–56:??, November 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3543855.

Hou:2023:DMW

[HXZ23a] Ningning Hou, Xianjin Xia, and Yuanqing Zheng. Don’t miss weak packets: Boosting LoRa reception with antenna diversities. ACM Transactions on Sensor Networks, 19(2):41:1–41:??, May 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3563698.

Hou:2023:JLP

[HXZ23b] Ningning Hou, Xianjin Xia, and Yuanqing Zheng. Jamming of LoRa PHY and countermeasure. ACM Transactions on Sensor Networks, 19 (4):80:1–80:27, November 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3583137.

Hua:2007:ARS

[HY07] Cunqing Hua and Tak-Shing Peter Yum. Asynchronous random sleeping for sensor networks. ACM Transactions on Sensor Networks, 3(3):15:1–15:??, August 2007. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

He:2005:FTI

[HZGS05] Guanghui He, Rong Zheng, Indranil Gupta, and Lui Sha. A framework for time indexing in sensor networks. ACM Transactions on Sensor Networks, 1(1):

101–133, August 2005. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

He:2024:DIN

[HZX^{+}24] Yuan He, Jia Zhang, Rui Xi, Xin Na, Yimiao Sun, and Beibei Li. Detection and identification of non-cooperative UAV using a COTS mmWave radar. ACM Transactions on Sensor Networks, 20(2):44:1–44:??, March 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3638767.

Ingelrest:2010:SAS

[IBS^{+}10] François Ingelrest, Guillermo Barrenetxea, Gunnar Schaefer, Martin Vetterli, Olivier Couach, and Marc Parlange. SensorScope: Application-specific sensor network for environmental monitoring. ACM Transactions on Sensor Networks, 6(2): 17:1–17:??, February 2010. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Iwanicki:2015:BMU

[IHGS15] Konrad Iwanicki, Przemyslaw Horban, Piotr Glazar, and Karol Strzelecki. Bringing modern unit testing techniques to sensornets. ACM Transactions on Sensor Networks, 11(2):25:1–25:??, February 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Istomin:2020:RFR

[IIPK20] Timofei Istomin, Oana Iova, Gian Pietro Picco, and Csaba Kiraly. Route or flood? Reliable and efficient support for downward traffic in RPL. ACM Transactions on Sensor Networks, 16(1):1:1– 1:41, February 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3355997.

Illiano:2018:DRG

[IPMGL18] Vittorio P. Illiano, Andrea Paudice, Luis Muñoz-González, and Emil C. Lupu. Determining resilience gains from anomaly detection for event integrity in wireless sensor networks. ACM Transactions on Sensor Networks, 14(1):5:1–5:??, March 2018. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ilyas:2012:DPA

[IR12] Muhammad U. Ilyas and Hayder Radha. A dynamic programming approach to maximizing a statistical measure of the lifetime of sensor networks. ACM Transactions on Sensor Networks, 8 (2):18:1–18:??, March 2012. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Iwanicki:2012:CHR

[IV12] Konrad Iwanicki and Maarten Van Steen. A case for hierarchical routing in low-power wireless embedded networks. ACM Transactions on Sensor Networks, 8(3):25:1–25:??, July 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ilie:2014:OCA

[IW14] Adrian Ilie and Greg Welch.

Online control of active camera networks for computer vision tasks. ACM Transactions on Sensor Networks, 10(2):25:1– 25:??, January 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Jellali:2019:BDS

[JAC19] Zakia Jellali, Leïla Najjar Atallah, and Sofiane Cherif. Bidimensional signal compression based on linear prediction coding: Application to WSN. ACM Transactions on Sensor Networks, 15(3):29:1–29:??, August 2019. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic).  URL https://dl.
acm.org/ft_gateway.cfm?id=
3317688.

Jeong:2012:PTM

[JC12] Jaein Jeong and David Culler.

A practical theory of micro-solar power sensor networks. ACM Transactions on Sensor Networks, 9(1):9:1–9:??, November 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Jurdak:2013:EEL

[JCC^{+}13] Raja Jurdak, Peter Corke, Alban Cotillon, Dhinesh Dharman, Chris Crossman, and Guillaume Salagnac. Energy-efficient localization: GPS duty cycling with radio ranging. ACM Transactions on Sensor Networks, 9(2): 23:1–23:??, March 2013. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ji:2022:DFM

[JCZ^{+}22] Xiaoyu Ji, Yushi Cheng, Juchuan Zhang, Yuehan Chi, Wenyuan Xu, and Yi-Chao Chen. Device fingerprinting with magnetic induction signals radiated by CPU modules. ACM Transactions on Sensor Networks, 18(2):23:1– 23:28, May 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3495158.

Jiang:2023:ESS

[JGK^{+}23] Jielin Jiang, Jiajie Guo, Maqbool Khan, Yan Cui, and Wenmin Lin. Energy-saving service offloading for the Internet of Medical Things using deep reinforcement learning. ACM Transactions on Sensor Networks, 19(3):55:1–55:??, August 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3560265.

Ji:2013:CBS

[JHU^{+}13] Shouling Ji, Jing (Selena) He, A. Selcuk Uluagac, Raheem Beyah, and Yingshu Li. Cellbased snapshot and continuous data collection in wireless sensor networks. ACM Transactions on Sensor Networks, 9(4): 47:1–47:??, July 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Jafarizadeh:2015:ADL

[JJ15] Saber Jafarizadeh and Abbas Jamalipour. Adapting distributed LT codes to Y-networks: an abstraction of collection tree in sensor networks. ACM Transactions on Sensor Networks, 11(4):54:1– 54:??, December 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Jaggi:2008:NOA

[JKK08] Neeraj Jaggi, Koushik Kar, and Ananth Krishnamurthy. Nearoptimal activation policies in rechargeable sensor networks under spatial correlations. ACM Transactions on Sensor Networks, 4(3):17:1–17:??, May 2008. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Jurcik:2010:DWC

[JKS^{+}10] Petr Jurcik, Anis Koubâa, Ricardo Severino, Mário Alves, and Eduardo Tovar. Dimensioning and worst-case analysis of cluster-tree sensor networks. ACM Transactions on Sensor Networks, 7(2):14:1–14:??, August 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Jiang:2013:PMW

[JLYG13] Xiaoye Jiang, Mo Li, Yuan Yao, and Leonidas Guibas. Property management in wireless sensor networks with overcomplete radon bases. ACM Transactions on Sensor Networks, 9(3):36:1– 36:??, May 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Jiang:2019:MED

[JLZL19] Shiqi Jiang, Zhenjiang Li, Pengfei Zhou, and Mo Li. Memento: an emotion-driven lifelogging system with wearables. ACM Transactions on Sensor Networks, 15(1):8:1–8:??, February 2019. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL https://dl.

                                        acm.org/ft_gateway.cfm?id=
                                        3281630.

Jhumka:2016:NVC

[JM16] Arshad Jhumka and Luca Mottola. Neighborhood view consistency in wireless sensor networks. ACM Transactions on Sensor Networks, 12(3):19:1– 19:??, August 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Jindal:2006:MSC

[JP06] Apoorva Jindal and Konstantinos Psounis. Modeling spatially correlated data in sensor networks. ACM Transactions on Sensor Networks, 2(4):466–499, November 2006. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Jourdan:2008:OSP

[JR08] Damien B. Jourdan and Nicholas Roy. Optimal sensor placement for agent localization. ACM Transactions on Sensor Networks, 4(3):13:1–13:??, May 2008. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Jurdak:2009:DBO

[JROH09] Raja Jurdak, Antonio G.

Ruzzelli, Gregory M. P. O’hare, and Russell Higgs. Directed broadcast with overhearing for sensor networks. ACM Transactions on Sensor Networks, 6(1): 3:1–3:??, December 2009. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Johnson:2012:MMB

[JSBN^{+}12] Matthew P. Johnson, Deniz Sariöz, Amotz Bar-Noy, Theodore Brown, Dinesh Verma, and Chai W. Wu. More is more: The benefits of denser sensor deployment. ACM Transactions on Sensor Networks, 8(3):22:1– 22:??, July 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Jiang:2020:RBN

[JTE20] Linshan Jiang, Rui Tan, and Arvind Easwaran. Resilience bounds of network clock synchronization with fault correction. ACM Transactions on Sensor Networks, 16(4):38:1– 38:30, October 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3409804.

Jung:2009:SNL

[JTS09] Deokwoo Jung, Thiago Teixeira, and Andreas Savvides. Sensor node lifetime analysis: Models and tools. ACM Transactions on Sensor Networks, 5(1):3:1– 3:??, February 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ju:2021:ESD

[JYB^{+}21] Weiyu Ju, Dong Yuan, Wei Bao, Liming Ge, and Bing Bing Zhou. eDeepSave: Saving DNN inference using early exit during handovers in mobile edge environment. ACM Transactions on Sensor Networks, 17 (3):30:1–30:28, June 2021. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3447267.

Ju:2024:RST

[JYC^{+}24] Ying Ju, Mingjie Yang, Chinmay Chakraborty, Lei Liu, Qingqi Pei, Ming Xiao, and Keping Yu. Reliability-security tradeoff analysis in mmWave ad hocbased CPS. ACM Transactions on Sensor Networks, 20(2): 28:1–28:??, March 2024. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3582556.

Jia:2019:ORC

[JZL^{+}19] Riheng Jia, Jinbei Zhang, Xiao-Yang Liu, Peng Liu, Luoyi Fu, and Xinbing Wang. Optimal rate control for energy-harvesting systems with random data and energy arrivals. ACM Transactions on Sensor Networks, 15 (1):13:1–13:??, February 2019. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3293535.

Ji:2020:OOP

[JZX^{+}20] Xiaoyu Ji, Xinyan Zhou, Miao Xu, Wenyuan Xu, and Yabo Dong. OPCIO: Optimizing power consumption for embedded devices via GPIO configuration. ACM Transactions on Sensor Networks, 16(2): 16:1–16:28, April 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3373417.

Kwon:2013:PES

[KA13] Youngmin Kwon and Gul Agha.

Performance evaluation of sensor networks by statistical modeling and Euclidean model checking. ACM Transactions on Sensor Networks, 9(4):39:1–39:??, July 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Keller:2013:SNC

[KAAF13] Lorenzo Keller, Emre Atsan, Katerina Argyraki, and Christina Fragouli. SenseCode: Network coding for reliable sensor networks. ACM Transactions on Sensor Networks, 9(2):25:1– 25:??, March 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ko:2010:HNU

[KAH^{+}10] Teresa Ko, Shaun Ahmadian, John Hicks, Mohammad Rahimi, Deborah Estrin, Stefano Soatto, Sharon Coe, and Michael P. Hamilton. Heartbeat of a nest: Using imagers as biological sensors. ACM Transactions on Sensor Networks, 6(3):19:1– 19:??, June 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Kalpakis:2010:ESA

[Kal10] Konstantinos Kalpakis. Everywhere sparse approximately optimal minimum energy data gathering and aggregation in sensor networks. ACM Transactions on Sensor Networks, 7(1): 9:1–9:??, August 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Kusy:2014:RDR

[KAR^{+}14] Branislav Kusy, David Abbott, Christian Richter, Cong Huynh, Mikhail Afanasyev, Wen Hu, Michael Brünig, Diethelm Ostry, and Raja Jurdak. Radio diversity for reliable communication in sensor networks. ACM Transactions on Sensor Networks, 10(2):32:1–32:??, January 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Kusy:2010:RDS

[KAS^{+}10] Branislav Kusý, Isaac Amundson, Janos Sallai, Peter Völgyesi, Akos Lédeczi, and Xenofon Koutsoukos. RF Doppler shiftbased mobile sensor tracking and navigation. ACM Transactions on Sensor Networks, 7(1):1:1– 1:??, August 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Kulathumani:2009:TDS

[KASD09] Vinodkrishnan Kulathumani, Anish Arora, Mukundan Sridharan, and Murat Demirbas. Trail: a distance-sensitive sensor network service for distributed object tracking. ACM Transactions on Sensor Networks, 5(2):15:1– 15:??, March 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Kamal:2013:PLA

[KBD13] Abu Raihan M. Kamal, Chris Bleakley, and Simon Dobson. Packet-Level Attestation (PLA): a framework for in-network sensor data reliability. ACM Transactions on Sensor Networks, 9 (2):19:1–19:??, March 2013. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Kamal:2014:FDW

[KBD14] Abu Raihan M. Kamal, Chris J.

Bleakley, and Simon Dobson. Failure detection in wireless sensor networks: a sequence-based dynamic approach. ACM Transactions on Sensor Networks, 10 (2):35:1–35:??, January 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Kulau:2016:IRU

[KBW16] Ulf Kulau, Felix Büsching, and Lars Wolf. IdealVolting: Reliable undervolting on wireless sensor nodes. ACM Transactions on Sensor Networks, 12(2): 11:1–11:??, May 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Kapnadak:2014:OND

[KC14] Vibhav Kapnadak and Edward J. Coyle. Optimal nonuniform deployment of sensors for distributed detection in wireless sensor networks. ACM Transactions on Sensor Networks, 10(2): 29:1–29:??, January 2014. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Karapetyan:2020:MAC

[KCE^{+}20] Areg Karapetyan, Sid Chi-Kin Chau, Khaled Elbassioni, Syafiq Kamarul Azman, and Majid Khonji. Multisensor adaptive control system for IoT-empowered smart lighting with oblivious mobile sensors. ACM Transactions on Sensor Networks, 16(1):11:1– 11:21, February 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3369392.

Kamthe:2013:IWL

[KCPC13] Ankur Kamthe, Miguel Á.

Carreira-Perpiñán, and Alberto E. Cerpa. Improving wireless link simulation using multilevel Markov models. ACM Transactions on Sensor Networks, 10(1):17:1–17:??, November 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Khalil:2018:SPI

[KGBS18] Nacer Khalil, Omprakash Gnawali, Driss Benhaddou, and Jaspal Subhlok. SonicDoor: a person identification system based on modeling of shape, behavior, and walking patterns. ACM Transactions on Sensor Networks, 14(3– 4):27:1–27:??, December 2018. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Khosravy:2022:UIN

[KGDC22] Mahdi Khosravy, Neeraj Gupta, Nilanjan Dey, and Rubén González Crespo. Underwater IoT network by blind MIMO OFDM transceiver based on probabilistic Stone’s blind source separation. ACM Transactions on Sensor Networks, 18(3):32:1– 32:??, August 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3462674.

Krause:2011:RSP

[KGGK11] Andreas Krause, Carlos Guestrin, Anupam Gupta, and Jon Kleinberg. Robust sensor placements at informative and communication-efficient locations. ACM Transactions on Sensor Networks, 7(4):31:1– 31:??, February 2011. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Kaswan:2023:DDM

[KJD^{+}23] Amar Kaswan, Prasanta K. Jana, Madhusmita Dash, Anupam Kumar, and Bhabani P. Sinha. DMCP: a distributed mobile charging protocol in wireless rechargeable sensor networks. ACM Transactions on Sensor Networks, 19(1):7:1– 7:29, February 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3526090.

Ko:2015:DRS

[KJP^{+}15] Jeonggil Ko, Jongsoo Jeong, Jongjun Park, Jong Arm Jun, Omprakash Gnawali, and Jeongyeup Paek. DualMOP– RPL: Supporting multiple modes of downward routing in a single RPL network. ACM Transactions on Sensor Networks, 11 (2):39:1–39:??, February 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Knox:2015:WFI

[KK15] D. A. Knox and T. Kunz. Wireless fingerprints inside a wireless sensor network. ACM Transactions on Sensor Networks, 11(2): 37:1–37:??, February 2015. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Karenos:2008:CBC

[KKK08] Kyriakos Karenos, Vana Kalogeraki, and Srikanth V. Krishnamurthy. Cluster-based congestion control for sensor networks. ACM Transactions on Sensor Networks, 4(1):5:1–5:??, January 2008. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Kansal:2007:RMM

[KKP^{+}07] Aman Kansal, William Kaiser, Gregory Pottie, Mani Srivastava, and Gaurav Sukhatme. Reconfiguration methods for mobile sensor networks. ACM Transactions on Sensor Networks, 3(4): 22:1–22:??, October 2007. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Kuppannagari:2018:ODN

[KKP18] Sanmukh R. Kuppannagari, Rajgopal Kannan, and Viktor K. Prasanna. Optimal discrete netload balancing in smart grids with high PV penetration. ACM Transactions on Sensor Networks, 14(3–4):24:1–24:??, December 2018. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Klonowski:2015:MRD

[KKRR15] Marek Klonowski, Miroslaw Kutylowski, Michal Ren, and Katarzyna Rybarczyk. Mixing in random digraphs with application to the forward-secure key evolution in wireless sensor networks. ACM Transactions on Sensor Networks, 11(2):29:1– 29:??, February 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Khan:2014:TIC

[KLA^{+}14] Mohammad Maifi Hasan Khan, Hieu Khac Le, Hossein Ahmadi, Tarek F. Abdelzaher, and Jiawei Han. Troubleshooting interactive complexity bugs in wireless sensor networks using data mining techniques. ACM Transactions on Sensor Networks, 10(2): 31:1–31:??, January 2014. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ko:2013:GSC

[KLC13] Ren-Song Ko, Po-Liang Lin, and Pei-Yu Chiang. Gauss–Seidel correction algorithm: a macroscopic model-derived routing algorithm for WSNs. ACM Transactions on Sensor Networks, 10(1):9:1–9:??, November 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Kim:2016:REE

[KLC^{+}16] Hyung-Sin Kim, Myung-Sup Lee, Young-June Choi, Jeonggil Ko, and Saewoong Bahk. Reliable and energy-efficient downward packet delivery in asymmetric transmission power-based networks. ACM Transactions on Sensor Networks, 12(4):34:1– 34:??, November 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Kasirajan:2012:NDA

[KLJ12] Priya Kasirajan, Carl Larsen, and S. Jagannathan. A new data aggregation scheme via adaptive compression for wireless sensor networks. ACM Transactions on Sensor Networks, 9(1):5:1–5:??, November 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Kwon:2010:RLS

[KMS^{+}10] Youngmin Kwon, Kirill Mechitov, Sameer Sundresh, Wooyoung Kim, and Gul Agha. Resilient localization for sensor networks in outdoor environments. ACM Transactions on Sensor Networks, 7(1):3:1–3:??, August 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Kuo:2014:CWA

[KNSM14] Thomas Kuo, Zefeng Ni, Santhoshkumar Sunderrajan, and B. S. Manjunath. Calibrating a wide-area camera network with non-overlapping views using mobile devices. ACM Transactions on Sensor Networks, 10(2):26:1– 26:??, January 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Kazmi:2014:RWS

[KOD^{+}14] Aqeel H. Kazmi, Michael J.

O’Grady, Declan T. Delaney, Antonio G. Ruzzelli, and Gregory M. P. O’Hare. A review of wireless-sensor-network-enabled building energy management systems. ACM Transactions on Sensor Networks, 10(4): 66:1–66:??, June 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Koukoutsidis:2018:ESA

[Kou18] Ioannis Koukoutsidis. Estimating spatial averages of environmental parameters based on mobile crowdsensing. ACM Transactions on Sensor Networks, 14 (1):2:1–2:??, March 2018. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Krasniewski:2008:EED

[KPB^{+}08] Mark D. Krasniewski, Rajesh Krishna Panta, Saurabh Bagchi, Chin-Lung Yang, and William J. Chappell. Energyefficient on-demand reprogramming of large-scale sensor networks. ACM Transactions on Sensor Networks, 4(1):2:1–2:??, January 2008. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Kim:2020:PRJ

[KPCB20] Hyung-Sin Kim, Jeongyeup Paek, David E. Culler, and Saewoong Bahk. PC-RPL: Joint control of routing topology and transmission power in real low-power and lossy networks. ACM Transactions on Sensor Networks, 16(2): 14:1–14:32, April 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3372026.

Karumbu:2012:DOE

[KPK12] Premkumar Karumbu, Venkata K.

Prasanthi, and Anurag Kumar. Delay optimal event detection on ad hoc wireless sensor networks. ACM Transactions on Sensor Networks, 8(2):12:1– 12:??, March 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Karvonen:2014:CLO

[KPRH14] Heikki Karvonen, Carlos Pomalaza-Ráez, and Matti Hämäläinen. A cross-layer optimization approach for lower layers of the protocol stack in sensor networks. ACM Transactions on Sensor Networks, 11(1):16:1– 16:??, August 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Kim:2012:LSV

[KPS12] Younghun Kim, Heemin Park, and Mani B. Srivastava. A longitudinal study of vibrationbased water flow sensing. ACM Transactions on Sensor Networks, 9(1):8:1–8:??, November 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Karakaya:2012:CEC

[KQ12] Mahmut Karakaya and Hairong Qi. Coverage estimation for crowded targets in visual sensor networks. ACM Transactions on Sensor Networks, 8(3): 26:1–26:??, July 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Karakaya:2014:CLV

[KQ14] Mahmut Karakaya and Hairong Qi. Collaborative localization in visual sensor networks. ACM Transactions on Sensor Networks, 10(2):18:1–18:??, January 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

King:2018:DCC

[KR18] Alex King and Utz Roedig. Differentiating clear channel assessment using transmit power variation. ACM Transactions on Sensor Networks, 14(2):15:1– 15:??, July 2018. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Kho:2009:DCA

[KRJ09] Johnsen Kho, Alex Rogers, and Nicholas R. Jennings. Decentralized control of adaptive sampling in wireless sensor networks. ACM Transactions on Sensor Networks, 5(3):19:1–19:??, May 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Kumar:2015:GEB

[KRP15] Dheeraj Kumar, Sutharshan Rajasegarar, and Marimuthu Palaniswami. Geospatial estimationbased auto drift correction in wireless sensor networks. ACM Transactions on Sensor Networks, 11(3):50:1–50:??, May 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Kominami:2013:CSO

[KSMH13] Daichi Kominami, Masashi Sugano, Masayuki Murata, and Takaaki Hatauchi. Controlled and self-organized routing for large-scale wireless sensor networks. ACM Transactions on Sensor Networks, 10(1):13:1– 13:??, November 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Kaur:2020:JMC

[KSR^{+}20] Manpreet Kaur, Flora D. Salim, Yongli Ren, Jeffrey Chan, Martin Tomko, and Mark Sanderson. Joint modelling of cyber activities and physical context to improve prediction of visitor behaviors. ACM Transactions on Sensor Networks, 16(3): 28:1–28:25, August 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3393692.

Keeler:2011:MFG

[KT11] Holger P. Keeler and Peter G.

Taylor. A model framework for greedy routing in a sensor network with a stochastic power scheme. ACM Transactions on Sensor Networks, 7(4): 34:1–34:??, February 2011. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Kung:2022:XRP

[Kun22] S. Y. Kung. XNAS: a regressive/progressive NAS for deep learning. ACM Transactions on Sensor Networks, 18(4):57:1– 57:??, November 2022. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3543669.

Klein:2013:LSA

[KVI^{+}13] Daniel J. Klein, Sriram Venkateswaran, Jason T. Isaacs, Jerry Burman, Tien Pham, João Hespanha, and Upamanyu Madhow. Localization with sparse acoustic sensor network using UAVs as information-seeking data mules. ACM Transactions on Sensor Networks, 9(3):30:1–30:??, May 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Khan:2023:URS

[KVS23] Usman Mahmood Khan, Raghav H. Venkatnarayan, and Muhammd Shahzad. Using RF signals to generate indoor maps. ACM Transactions on Sensor Networks, 19(1):12:1–12:30, February 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3534121.

Kulkarni:2009:EEM

[KW09] Sandeep Kulkarni and Limin Wang. Energy-efficient multihop reprogramming for sensor networks. ACM Transactions on Sensor Networks, 5(2):16:1– 16:??, March 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Kamat:2009:TPW

[KXTZ09] Pandurang Kamat, Wenyuan Xu, Wade Trappe, and Yanyong Zhang. Temporal privacy in wireless sensor networks: Theory and practice. ACM Transactions on Sensor Networks, 5 (4):28:1–28:??, November 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Kartakis:2017:RSO

[KYM17] Sokratis Kartakis, Shusen Yang, and Julie A. Mccann. Reliability or sustainability: Optimal data stream estimation and scheduling in smart water networks. ACM Transactions on Sensor Networks, 13(3):18:1– 18:??, September 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Lambrou:2015:OCD

[Lam15] Theofanis P. Lambrou. Optimized cooperative dynamic coverage in mixed sensor networks. ACM Transactions on Sensor Networks, 11(3):46:1– 46:??, February 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Liu:2014:DDL

[LC14a] Tao Liu and Alberto E. Cerpa.

Data-driven link quality prediction using link features. ACM Transactions on Sensor Networks, 10(2):37:1–37:??, January 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Liu:2014:TAL

[LC14b] Tao Liu and Alberto E. Cerpa. Temporal adaptive link quality prediction with online learning. ACM Transactions on Sensor Networks, 10(3):46:1–46:??, April 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ling:2010:APA

[LCC10] Yibei Ling, Chung-Min Chen, and Shigang Chen. Analysis of power-aware buffering schemes in wireless sensor networks. ACM Transactions on Sensor Networks, 7(3):26:1– 26:??, September 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Lai:2013:MHW

[LCC^{+}13] Ted Tsung-Te Lai, Wei-Ju Chen, Yu-Han Tiffany Chen, Polly Huang, and Hao-Hau Chu. Mapping hidden water pipelines using a mobile sensor droplet. ACM Transactions on Sensor Networks, 9(2):20:1–20:??, March 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Li:2017:AHA

[LCC^{+}17] Ji Li, Siyao Cheng, Zhipeng Cai, Jiguo Yu, Chaokun Wang, and Yingshu Li. Approximate holistic aggregation in wireless sensor networks. ACM Transactions on Sensor Networks, 13(2): 11:1–11:??, June 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Lv:2022:ERC

[LCD22] Jiamei Lv, Gonglong Chen, and Wei Dong. Exploiting rateless codes and cross-layer optimization for low-power widearea networks. ACM Transactions on Sensor Networks, 18 (4):62:1–62:??, November 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3544560.

Lv:2022:AIU

[LCF^{+}22] Zhihan Lv, Dongliang Chen, Hailin Feng, Wei Wei, and Haibin Lv. Artificial intelligence in underwater digital twins sensor networks. ACM Transactions on Sensor Networks, 18(3): 39:1–39:??, August 2022. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3519301.

Luo:2009:DIE

[LCH^{+}09] Liqian Luo, Qing Cao, Chengdu Huang, Lili Wang, Tarek F. Abdelzaher, John A. Stankovic, and Michael Ward. Design, implementation, and evaluation of EnviroMic: a storage-centric audio sensor network. ACM Transactions on Sensor Networks, 5(3): 22:1–22:??, May 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Leng:2019:NMM

[LCH^{+}19a] Quan Leng, Wei-Ju Chen, Pei-Chi Huang, Yi-Hung Wei, Aloysius K. Mok, and Song Han. Network management of multicluster RT-WiFi networks. ACM Transactions on Sensor Networks, 15(1):12:1–12:??, February 2019. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=
3283451.

Liu:2019:ECO

[LCH^{+}19b] Daibo Liu, Zhichao Cao, Yuan He, Xiaoyu Ji, Mengshu Hou, and Hongbo Jiang. Exploiting concurrency for opportunistic forwarding in duty-cycled IoT networks. ACM Transactions on Sensor Networks, 15(3): 31:1–31:??, August 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3322496.

Liu:2020:PLT

[LCH^{+}20] Daibo Liu, Zhichao Cao, Mengshu Hou, Huigui Rong, and Hongbo Jiang. Pushing the limits of transmission concurrency for low power wireless networks. ACM Transactions on Sensor Networks, 16(4):40:1– 40:29, October 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3406834.

Liu:2023:CLP

[LCJ^{+}23] Daibo Liu, Zhichao Cao, Hongbo Jiang, Siwang Zhou, Zhu Xiao, and Fanzi Zeng. Concurrent low-power listening: a new design paradigm for duty-cycling communication. ACM Transactions on Sensor Networks, 19(1): 4:1–4:24, February 2023. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3517013.

Liu:2022:ORP

[LCLY22] Yu Liu, Joshua Comden, Zhenhua Liu, and Yuanyuan Yang. Online resource provisioning for wireless data collection. ACM Transactions on Sensor Networks, 18(1):7:1–7:2, February 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3470648.

Liu:2021:SSD

[LCM21] Liang Liu, Bo Chen, and Huadong Ma. SDCN: Sensory data-centric networking for building the sensing layer of IoT. ACM Transactions on Sensor Networks, 17(1):6:1– 6:25, January 2021. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3402452.

Li:2019:RCF

[LDC^{+}19] Lanlan Li, Haipeng Dai, Guihai Chen, Jiaqi Zheng, Wanchun Dou, and Xiaobing Wu. Radiation constrained fair charging for wireless power transfer. ACM Transactions on Sensor Networks, 15(2):15:1–15:??, April 2019. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL https://dl.

                                        acm.org/ft_gateway.cfm?id=
                                        3289182.

Laidi:2024:TST

[LDDL24] Roufaida Laidi, Djamel Djenouri, Youcef Djenouri, and Jerry Chun-Wei Lin. TG-SPRED: Temporal graph for sensorial data PREDiction. ACM Transactions on Sensor Networks, 20(3):64:1–64:??, May 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3649892.

Li:2021:QQA

[LDG^{+}21] Borui Li, Wei Dong, Gaoyang Guan, Jiadong Zhang, Tao Gu, Jiajun Bu, and Yi Gao. Queec: QoE-aware edge computing for IoT devices under dynamic workloads. ACM Transactions on Sensor Networks, 17(3): 27:1–27:23, June 2021. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3442363.

Lin:2021:SVF

[LDGG21] Yuxiang Lin, Wei Dong, Yi Gao, and Tao Gu. SateLoc: a virtual fingerprinting approach to outdoor LoRa localization using satellite images. ACM Transactions on Sensor Networks, 17 (4):43:1–43:28, July 2021. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3461012.

Law:2006:SBB

[LDH06] Yee Wei Law, Jeroen Doumen, and Pieter Hartel. Survey and benchmark of block ciphers for wireless sensor networks. ACM Transactions on Sensor Networks, 2(1):65–93, February 2006. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Lian:2024:EFG

[LDL^{+}24] Jie Lian, Changlai Du, Jiadong Lou, Li Chen, and Xu Yuan. EchoSensor: Fine-grained ultrasonic sensing for smart home intrusion detection. ACM Transactions on Sensor Networks, 20 (1):10:1–10:??, January 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3615658.

Li:2022:ROC

[LDS^{+}22] Jie Li, Yuxing Deng, Wei Sun, Weitao Li, Ruidong Li, Qiyue Li, and Zhi Liu. Resource orchestration of cloud-edge-based smart grid fault detection. ACM Transactions on Sensor Networks, 18(3):46:1–46:??, August 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3529509.

Li:2013:AEE

[LDZ13] Wei Li, Flávia C. Delicato, and Albert Y. Zomaya. Adaptive energy-efficient scheduling for hierarchical wireless sensor networks. ACM Transactions on Sensor Networks, 9(3):33:1– 33:??, May 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Lee:2020:DRE

[LED20] Tim Van Der Lee, Georgios Exarchakos, and Sonia Heemstra De Groot. Distributed reliable and energy-efficient scheduling for LR-WPANs. ACM Transactions on Sensor Networks, 16 (4):32:1–32:20, October 2020. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3399805.

Liu:2019:LCR

[LFL^{+}19] Chen Liu, Dingyi Fang, Xinyan Liu, Dan Xu, Xiaojiang Chen, Chieh-Jan Mike Liang, Baoying Liu, and Zhanyong Tang. Lowcost and robust geographic opportunistic routing in a strip topology wireless network. ACM Transactions on Sensor Networks, 15(2):24:1–24:??, April 2019. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic).  URL https://dl.
acm.org/ft_gateway.cfm?id=
3309701.

Li:2014:PSA

[LFNS14] Xu Li, Greg Fletcher, Amiya Nayak, and Ivan Stojmenovic. Placing sensors for area coverage in a complex environment by a team of robots. ACM Transactions on Sensor Networks, 11(1):

3:1–3:??, August 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Liu:2009:CEE

[LFS09] Sha Liu, Kai-Wei Fan, and Prasun Sinha. CMAC: an energy-efficient MAC layer protocol using convergent packet forwarding for wireless sensor networks. ACM Transactions on Sensor Networks, 5(4):29:1– 29:??, November 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Liu:2019:MSC

[LFW^{+}19] Xuecheng Liu, Luoyi Fu, Jiliang Wang, Xinbing Wang, and Guihai Chen. Multicast scaling of capacity and energy efficiency in heterogeneous wireless sensor networks. ACM Transactions on Sensor Networks, 15(3): 33:1–33:??, August 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3322497.

Lin:2023:DRA

[LGLD23] Yuxiang Lin, Yi Gao, Bingji Li, and Wei Dong. Detecting rogue access points using client-agnostic wireless fingerprints. ACM Transactions on Sensor Networks, 19(1):14:1– 14:25, February 2023. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3536423.

Liando:2019:KUF

[LGTL19] Jansen C. Liando, Amalinda Gamage, Agustinus W. Tengourtius, and Mo Li. Known and unknown facts of LoRa: Experiences from a large-scale measurement study. ACM Transactions on Sensor Networks, 15 (2):16:1–16:??, April 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

                                        URL https://dl.acm.org/ft_
                                        gateway.cfm?id=3293534.

Li:2023:AAP

[LGXC23] Jinxi Li, Deke Guo, Junjie Xie, and Sheng Chen. Availabilityaware provision of service function chains in mobile edge computing. ACM Transactions on Sensor Networks, 19(3):57:1– 57:??, August 2023. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3565483.

Lim:2009:DLA

[LH09] Hyuk Lim and Jennifer C. Hou. Distributed localization for anisotropic sensor networks. ACM Transactions on Sensor Networks, 5(2):11:1–11:??, March 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Liao:2024:ESB

[LHHW24] Qianru Liao, Yongzhi Huang, Yandao Huang, and Kaishun Wu. An eavesdropping system based on magnetic side-channel signals leaked by speakers. ACM Transactions on Sensor Networks, 20(2):39:1–39:??, March 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3637063.

Lachenmann:2009:MLG

[LHRM09] Andreas Lachenmann, Klaus Herrmann, Kurt Rothermel, and Pedro José Marrón. On meeting lifetime goals and providing constant application quality. ACM Transactions on Sensor Networks, 5(4):36:1–36:??, November 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Liu:2016:TMT

[LHX16] Chin-Jung Liu, Pei Huang, and Li Xiao. TAS-MAC: a trafficadaptive synchronous MAC protocol for wireless sensor networks. ACM Transactions on Sensor Networks, 12(1):1:1–1:??, March 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Liu:2021:CBE

[LHX^{+}21] Meng Liu, Hongsheng Hu, Haolong Xiang, Chi Yang, Lingjuan Lyu, and Xuyun Zhang. Clustering-based efficient privacy-preserving face recognition scheme without compromising accuracy. ACM Transactions on Sensor Networks, 17(3): 31:1–31:27, June 2021. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3448414.

Li:2020:SSB

[LHZZ20] Yantao Li, Hailong Hu, Zhangqian Zhu, and Gang Zhou. SCANet: Sensor-based continuous authentication with two-stream convolutional neural networks. ACM Transactions on Sensor Networks, 16(3):29:1–29:27, August 2020. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/abs/10.1145/3397179.

Liu:2021:EEC

[Liu21] Yunhao Liu. Editorial from the Editor-in-Chief. ACM Transactions on Sensor Networks, 17(2): 10e:1–10e:2, June 2021. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3448130.

Liu:2019:ROA

[LJLW19] Yang Liu, Yonghang Jiang, Zhenjiang Li, and Jianping Wang. Rulers on our arms: Waving to measure object size through contactless sensing. ACM Transactions on Sensor Networks, 15(1):14:1– 14:??, February 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

                                        URL https://dl.acm.org/ft_
                                        gateway.cfm?id=3289183.

Li:2021:SER

[LJW^{+}21] Chaohao Li, Xiaoyu Ji, Bin Wang, Kai Wang, and Wenyuan Xu. SenCS: Enabling real-time indoor proximity verification via contextual similarity. ACM Transactions on Sensor Networks, 17(2):19:1–19:22, June 2021. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3449071.

Lim:2010:RRP

[LJY^{+}10] Jun Bum Lim, Beakcheol Jang, Suyoung Yoon, Mihail L. Sichitiu, and Alexander G. Dean. RaPTEX: Rapid prototyping tool for embedded communication systems. ACM Transactions on Sensor Networks, 7(1): 7:1–7:??, August 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Li:2009:CNL

[LK09] Li Li and Thomas Kunz. Cooperative node localization using nonlinear data projection. ACM Transactions on Sensor Networks, 5(1):1:1–1:??, February 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Lee:2010:NLO

[LKA10] Huang Lee, Abtin Keshavarzian, and Hamid Aghajan. Nearlifetime-optimal data collection in wireless sensor networks via spatio-temporal load balancing. ACM Transactions on Sensor Networks, 6(3):26:1–26:??, June 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Li:2009:UCM

[LL09] Mo Li and Yunhao Liu. Underground coal mine monitoring with wireless sensor networks. ACM Transactions on Sensor Networks, 5(2):10:1– 10:??, March 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Li:2016:TLL

[LL16] Yimei Li and Yao Liang. Temporal lossless and lossy compression in wireless sensor networks. ACM Transactions on Sensor Networks, 12(4):37:1– 37:??, November 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Li:2021:GGB

[LL21] Xin Li and Dawei Li. GPFS: a graph-based human pose forecasting system for smart home with online learning. ACM Transactions on Sensor Networks, 17(3):34:1–34:19, June 2021. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3460199.

Li:2023:SNA

[LLDZ23] Yantao Li, Jiaxing Luo, Shaojiang Deng, and Gang Zhou. SearchAuth: Neural architecture search-based continuous authentication using auto augmentation search. ACM Transactions on Sensor Networks, 19 (4):92:1–92:23, November 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3599727.

Lin:2022:ABM

[LLH22] Kai Lin, Jiayi Liu, and Guangjie Han. AI-Based mean field game against resource-consuming attacks in edge computing. ACM Transactions on Sensor Networks, 18(4):52:1–52:??, November 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3519303.

Li:2014:TEF

[LLL14] Zhenjiang Li, Mo Li, and Yunhao Liu. Towards energyfairness in asynchronous dutycycling sensor networks. ACM Transactions on Sensor Networks, 10(3):38:1–38:??, April 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Li:2023:CEI

[LLW^{+}23] Yangfan Li, Kenli Li, Wei Wei, Tianyi Zhou, and Cen Chen. CoRec: an efficient Internet behavior-based recommendation framework with edge-cloud collaboration on deep convolution neural networks. ACM Transactions on Sensor Networks, 19 (2):24:1–24:28, May 2023. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3526191.

Li:2014:FOT

[LLX^{+}14] Huan Li, Dong Liang, Lihui Xie, Gong Zhang, and Krithi Ramamritham. Flash-optimized temporal indexing for time-series data storage on sensor platforms. ACM Transactions on Sensor Networks, 10(4):62:1– 62:??, June 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Li:2022:SPM

[LLX^{+}22] Jing Li, Weifa Liang, Zichuan Xu, Xiaohua Jia, and Wanlei Zhou. Service provisioning for multi-source IoT applications in mobile edge computing. ACM Transactions on Sensor Networks, 18(2):17:1–17:25, May 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3484200.

Liu:2020:IMC

[LLZ^{+}20] Zhao Liu, Kenli Li, Xu Zhou, Ningbo Zhu, and Keqin Li. Incentive mechanisms for crowdsensing: Motivating users to preprocess data for the crowdsourcer. ACM Transactions on Sensor Networks, 16(4):39:1– 39:24, October 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3409475.

Lan:2022:EAC

[LLZ^{+}22] Guohao Lan, Zida Liu, Yunfan Zhang, Tim Scargill, Jovan Stojkovic, Carlee Joe-Wong, and Maria Gorlatova. Edge-assisted collaborative image recognition for mobile augmented reality. ACM Transactions on Sensor Networks, 18(1):9:1– 9:31, February 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3469033.

Langendoen:2010:AMPa

[LM10a] Koen Langendoen and Andreas Meier. Analyzing MAC protocols for low data-rate applications. ACM Transactions on Sensor Networks, 7(1):10:1– 10:??, August 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Langendoen:2010:AMPb

[LM10b] Koen Langendoen and Andreas Meier. Analyzing MAC protocols for low data-rate applications. ACM Transactions on Sensor Networks, 7(2):19:1– 19:??, August 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Laoudias:2014:FFT

[LMP14] Christos Laoudias, Michalis P.

Michaelides, and Christos G. Panayiotou. ftTRACK: Faulttolerant target tracking in binary sensor networks. ACM Transactions on Sensor Networks, 10(4): 64:1–64:??, June 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Lin:2016:AAT

[LMZ^{+}16] Shan Lin, Fei Miao, Jingbin Zhang, Gang Zhou, Lin Gu, Tian He, John A. Stankovic, Sang Son, and George J. Pappas. ATPC: Adaptive transmission power control for wireless sensor networks. ACM Transactions on Sensor Networks, 12(1): 6:1–6:??, March 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Liu:2005:IKP

[LN05] Donggang Liu and Peng Ning. Improving key predistribution with deployment knowledge in static sensor networks. ACM Transactions on Sensor Networks, 1(2):204–239, November 2005. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Liu:2008:GBK

[LND08] Donggang Liu, Peng Ning, and Wenliang Du. Group-based key predistribution for wireless sensor networks. ACM Transactions on Sensor Networks, 4(2):11:1– 11:??, March 2008. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ledeczi:2005:CSU

[LNV^{+}05] Ákos Lédeczi, András Nádas, Péter Völgyesi, György Balogh, Branislav Kusy, János Sallai, Gábor Pap, Sebestyén Dóra, Károly Molnár, Miklós Maróti, and Gyula Simon. Countersniper system for urban warfare. ACM Transactions on Sensor Networks, 1(2):153–177, November 2005. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Lazos:2005:SRL

[LP05] Loukas Lazos and Radha Poovendran. SeRLoc: Robust localization for wireless sensor networks. ACM Transactions on Sensor Networks, 1(1):73–100, August 2005. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Lazos:2006:SCH

[LP06] Loukas Lazos and Radha Poovendran. Stochastic coverage in heterogeneous sensor networks. ACM Transactions on Sensor Networks, 2(3):325–358, August 2006. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Lai:2008:OBE

[LP08] Wei Lai and Ioannis C. Paschalidis. Optimally balancing energy consumption versus latency in sensor network routing. ACM Transactions on Sensor Networks, 4(4):21:1–21:??, August 2008. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Lazos:2009:AET

[LPR09] Loukas Lazos, Radha Poovendran, and James A. Ritcey. Analytic evaluation of target detection in heterogeneous wireless sensor networks. ACM Transactions on Sensor Networks, 5 (2):18:1–18:??, March 2009. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Law:2009:EEL

[LPV^{+}09] Yee Wei Law, Marimuthu Palaniswami, Lodewijk Van Hoesel, Jeroen Doumen, Pieter Hartel, and Paul Havinga. Energy-efficient link-layer jamming attacks against wireless sensor network MAC protocols. ACM Transactions on Sensor Networks, 5(1):6:1–6:??, February 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Lin:2023:SAL

[LPW^{+}23] Qi Lin, Shuhua Peng, Yuezhong Wu, Jun Liu, Hong Jia, Wen Hu, Mahbub Hassan, Aruna Seneviratne, and Chun H. Wang. Subject-adaptive loosefitting smart garment platform for human activity recognition. ACM Transactions on Sensor Networks, 19(4):84:1– 84:23, November 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3584986.

Li:2005:NPS

[LR05] Qun Li and Daniela Rus. Navigation protocols in sensor networks. ACM Transactions on Sensor Networks, 1(1):3–35, August 2005. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Liaskovitis:2010:LRS

[LS10] Periklis G. Liaskovitis and Curt Schurgers. Leveraging redundancy in sampling-interpolation applications for sensor networks: a spectral approach. ACM Transactions on Sensor Networks, 7(2):12:1–12:??, August 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Li:2006:LTC

[LSW06] Xiang-Yang Li, Wen-Zhan Song, and Yu Wang. Localized topology control for heterogeneous wireless sensor networks. ACM Transactions on Sensor Networks, 2(1):129–153, February 2006. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Lu:2014:SBH

[LSW14] Jiakang Lu, Yamina Taskin Shams, and Kamin Whitehouse. Smart blueprints: How simple sensors can collaboratively map out their own locations in the home. ACM Transactions on Sensor Networks, 11(1):19:1– 19:??, August 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Li:2024:ESI

[LSX24] Fangyu Li, WenZhan Song, and Xiaohua Xu. Editorial: Special issue on cyber-physical security and zero trust. ACM Transactions on Sensor Networks, 20(2):26:1–26:??, March 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3634700.

Li:2022:DCB

[LTDZ22] Yantao Li, Peng Tao, Shaojiang Deng, and Gang Zhou. DeF-Fusion: CNN-based continuous authentication using deep feature fusion. ACM Transactions on Sensor Networks, 18 (2):18:1–18:20, May 2022. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3485060.

Li:2024:DSM

[LTL^{+}24] Guopeng Li, Haisheng Tan, Liuyan Liu, Hao Zhou, Shaofeng H.-C. Jiang, Zhenhua Han, Xiang-Yang Li, and Guoliang Chen. DAG scheduling in mobile edge computing. ACM Transactions on Sensor Networks, 20 (1):12:1–12:??, January 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3616374.

Li:2018:NTP

[LTY18] Yang Li, Rui Tan, and David K. Y. Yau. Natural timestamps in powerline electromagnetic radiation. ACM Transactions on Sensor Networks, 14(2): 13:1–13:??, July 2018. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Liu:2024:CAD

[LTZ^{+}24] Junyi Liu, Yifu Tang, Haimeng Zhao, Xieheng Wang, Fangyu Li, and Jingyi Zhang. CPS attack detection under limited local information in cyber security: an ensemble multi-node multi-class classification approach. ACM Transactions on Sensor Networks, 20(2):33:1–33:??, March 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3585520.

Liu:2014:DAF

[LWCJ14] Hongbo Liu, Hui Wang, Yingying Chen, and Dayong Jia. Defending against frequency-based attacks on distributed data storage in wireless networks. ACM Transactions on Sensor Networks, 10(3):49:1–49:??, April 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Lederer:2009:CBL

[LWG09] Sol Lederer, Yue Wang, and Jie Gao. Connectivity-based localization of large-scale sensor networks with complex shape. ACM Transactions on Sensor Networks, 5(4):31:1–31:??, November 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Liu:2006:ORT

[LWH^{+}06] Xue Liu, Qixin Wang, Wenbo He, Marco Caccamo, and Lui Sha. Optimal real-time sampling rate assignment for wireless sensor networks. ACM Transactions on Sensor Networks, 2(2): 263–295, May 2006. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Lu:2022:SSK

[LWH^{+}22] Youjing Lu, Fan Wu, Qianyi Huang, Shaojie Tang, Linghe Kong, and Guihai Chen. Shared secret key generation by exploiting inaudible acoustic channels. ACM Transactions on Sensor Networks, 18(1):13:1– 13:26, February 2022. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3480461.

Lin:2023:CAS

[LWJ^{+}23] Chi Lin, Pengfei Wang, Chuanying Ji, Mohammad S. Obaidat, Lei Wang, Guowei Wu, and Qiang Zhang. A contactless authentication system based on WiFi CSI. ACM Transactions on Sensor Networks, 19 (2):29:1–29:20, May 2023. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3532095.

Li:2022:ISI

[LWKZ22] Mo Li, Jiliang Wang, Swarun Kumar, and Yuanqing Zheng. Introduction to the special issue on low power wide area networks. ACM Transactions on Sensor Networks, 18(4):58:1– 58:??, November 2022. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3586058.

Le:2021:DRL

[LWL^{+}21] Duc Van Le, Rongrong Wang, Yingbo Liu, Rui Tan, Yew-Wah Wong, and Yonggang Wen. Deep reinforcement learning for tropical air free-cooled data center control. ACM Transactions on Sensor Networks, 17 (3):24:1–24:28, June 2021. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3439332.

Li:2024:VUM

[LWLT24] Junsheng Li, Ling Wang, Jie Liu, and Jinshan Tang. ViST: a ubiquitous model with multimodal fusion for crop growth prediction. ACM Transactions on Sensor Networks, 20(1):23:1– 23:??, January 2024. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3627707.

Lin:2021:SEE

[LWM^{+}21] Deyu Lin, Quan Wang, Weidong Min, Jianfeng Xu, and Zhiqiang Zhang. A survey on energyefficient strategies in static wireless sensor networks. ACM Transactions on Sensor Networks, 17(1):3:1–3:48, January 2021. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3414315.

Liang:2012:DSE

[LWSL12] Jinling Liang, Zidong Wang, Bo Shen, and Xiaohui Liu. Distributed state estimation in sensor networks with randomly occurring nonlinearities subject to time delays. ACM Transactions on Sensor Networks, 9(1):4:1– 4:??, November 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Liu:2021:RRL

[LWX^{+}21] Tang Liu, Baijun Wu, Wenzheng Xu, Xianbo Cao, Jian Peng, and Hongyi Wu. RLC: a reinforcement learning-based charging algorithm for mobile devices. ACM Transactions on Sensor Networks, 17(4):36:1–36:23, July 2021. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3453682.

Liang:2021:EEC

[LWY^{+}21] Yunji Liang, Xin Wang, Zhiwen Yu, Bin Guo, Xiaolong Zheng, and Sagar Samtani. Energyefficient collaborative sensing: Learning the latent correlations of heterogeneous sensors. ACM Transactions on Sensor Networks, 17(3):33:1–33:28, June 2021. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3448416.

Li:2024:TED

[LWZ24] Mingzhe Li, Wei Wang, and Jin Zhang. Towards efficient and deposit-free blockchain-based spatial crowdsourcing. ACM Transactions on Sensor Networks, 20(3):73:1–73:??, May 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3656343.

Liang:2016:MLS

[LXR^{+}16] Weifa Liang, Wenzheng Xu, Xiaojiang Ren, Xiaohua Jia, and Xiaola Lin. Maintaining largescale rechargeable sensor networks perpetually via multiple mobile charging vehicles. ACM Transactions on Sensor Networks, 12(2):14:1–14:??, May 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Li:2022:WLS

[LXY^{+}22] Danyang Li, Jingao Xu, Zheng Yang, Chenshu Wu, Jianbo Li, and Nicholas D. Lane. Wireless localization with spatialtemporal robust fingerprints. ACM Transactions on Sensor Networks, 18(1):15:1–15:23, February 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3488281.

Li:2024:TOL

[LXYT24] Danyang Li, Jingao Xu, Zheng Yang, and Chengpei Tang. Train once, locate anytime for anyone:

Adversarial learning-based wireless localization. ACM Transactions on Sensor Networks, 20 (2):37:1–37:??, March 2024. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3614095.

Liu:2023:DMA

[LYF^{+}23] Miaomiao Liu, Kang Yang, Yanjie Fu, Dapeng Wu, and Wan Du. Driving maneuver anomaly detection based on deep autoencoder and geographical partitioning. ACM Transactions on Sensor Networks, 19(2):37:1– 37:??, May 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563217.

Li:2013:SAH

[LYG^{+}13] Ming Li, Shucheng Yu, Joshua D. Guttman, Wenjing Lou, and Kui Ren. Secure ad hoc trust initialization and key management in wireless body area networks. ACM Transactions on Sensor Networks, 9(2):18:1– 18:??, March 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Lian:2024:RSL

[LYL^{+}24] Jie Lian, Xu Yuan, Jiadong Lou, Li Chen, Hao Wang, and Nianfeng Tzeng. Room-scale location trace tracking via continuous acoustic waves. ACM Transactions on Sensor Networks, 20 (3):61:1–61:??, May 2024. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3649136.

Luo:2023:PDD

[LYST23] Wenjie Luo, Zhenyu Yan, Qun Song, and Rui Tan. Physicsdirected data augmentation for deep model transfer to specific sensor. ACM Transactions on Sensor Networks, 19 (1):21:1–21:30, February 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3549076.

Li:2024:EAO

[LYY24] Jianbo Li, Genji Yuan, and Zheng Yang. Edge-assisted object segmentation using multimodal feature aggregation and learning. ACM Transactions on Sensor Networks, 20(1):9:1– 9:??, January 2024. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3612922.

Le:2024:DDE

[LYZ^{+}24] Duc Van Le, Joy Qiping Yang, Siyuan Zhou, Daren Ho, and Rui Tan. Design, deployment, and evaluation of an industrial AIoT system for quality control at HP factories. ACM Transactions on Sensor Networks, 20(1):18:1–18:??, January 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3618300.

Lin:2015:TSN

[LZAH^{+}15] Shan Lin, Gang Zhou, Mo’taz Al-Hami, Kamin Whitehouse, Yafeng Wu, John A. Stankovic, Tian He, Xiaobing Wu, and Hengchang Liu. Toward stable network performance in wireless sensor networks: a multilevel perspective. ACM Transactions on Sensor Networks, 11(3):42:1– 42:??, February 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Li:2023:DPE

[LZGX23] Shancang Li, Shanshan Zhao, Prosanta Gope, and Li Da Xu. Data privacy enhancing in the IoT user/device behavior analytics. ACM Transactions on Sensor Networks, 19 (2):32:1–32:??, May 2023. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3534648.

Liu:2019:GFS

[LZN19] Yunhuai Liu, Qian Zhang, and Lionel Ni. A general framework for spectrum sensing using dedicated spectrum sensor networks. ACM Transactions on Sensor Networks, 15(1):7:1– 7:??, February 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

                                        URL https://dl.acm.org/ft_
                                        gateway.cfm?id=3275244.

Li:2015:IGS

[LZZ^{+}15] Mo Li, Pengfei Zhou, Yuanqing Zheng, Zhenjiang Li, and Guobin Shen. IODetector: a generic service for indoor/ outdoor detection. ACM Transactions on Sensor Networks, 11 (2):28:1–28:??, February 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Munishwar:2013:CAV

[MAG13] Vikram P. Munishwar and Nael B. Abu-Ghazaleh. Coverage algorithms for visual sensor networks. ACM Transactions on Sensor Networks, 9(4): 45:1–45:??, July 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Maierbacher:2009:LCC

[MB09] Gerhard Maierbacher and João Barros. Low-complexity coding and source-optimized clustering for large-scale sensor networks. ACM Transactions on Sensor Networks, 5(3):24:1–24:??, May 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Midi:2016:NLF

[MB16] Daniele Midi and Elisa Bertino.

Node or link? Fine-grained analysis of packet-loss attacks in wireless sensor networks. ACM Transactions on Sensor Networks, 12(2):8:1–8:??, May 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ma:2021:BEH

[MCGZ21] Qiang Ma, Zhichao Cao, Wei Gong, and Xiaolong Zheng. BOND: Exploring hidden bottleneck nodes in large-scale wireless sensor networks. ACM Transactions on Sensor Networks, 17 (2):13:1–13:21, June 2021. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3439956.

Ma:2020:QST

[MCLM20] Qiang Ma, Zhichao Cao, Kebin Liu, and Xin Miao. QA-Share: Toward an efficient QoSaware dispatching approach for urban taxi-sharing. ACM Transactions on Sensor Networks, 16 (2):17:1–17:21, April 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3375406.

Mo:2023:PID

[MCLW23] Xiaoyun Mo, Chu Cao, Mo Li, and David Z. W. Wang. Predicting the impact of disruptions to urban rail transit systems. ACM Transactions on Sensor Networks, 19(1):2:1–2:??, February 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3517015.

Mavrinac:2014:CQS

[MCT14] Aaron Mavrinac, Xiang Chen, and Yonghong Tan. Coverage quality and smoothness criteria for online view selection in a multi-camera network. ACM Transactions on Sensor Networks, 10(2):33:1–33:??, January 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Moran:2016:BMS

[MCW^{+}16] Bill Moran, Fred Cohen, Zengfu Wang, Sofia Suvorova, Douglas Cochran, Tom Taylor, Peter Farrell, and Stephen Howard. Bounds on multiple sensor fusion. ACM Transactions on Sensor Networks, 12(2):16:1– 16:??, May 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Mezair:2023:TAD

[MDB^{+}23] Tinhinane Mezair, Youcef Djenouri, Asma Belhadi, Gautam Srivastava, and Jerry Chun-Wei Lin. Towards an advanced deep learning for the Internet of behaviors: Application to connected vehicles. ACM Transactions on Sensor Networks, 19 (2):30:1–30:??, May 2023. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3526192.

Mathur:2009:ULP

[MDC^{+}09] Gaurav Mathur, Peter Desnoyers, Paul Chukiu, Deepak Ganesan, and Prashant Shenoy. Ultra-low power data storage for sensor networks. ACM Transactions on Sensor Networks, 5 (4):33:1–33:??, November 2009.

CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Mohammad:2017:IPS

[MDC17] Mobashir Mohammad, Manjunath Doddavenkatappa, and Mun Choon Chan. Improving performance of synchronous transmission-based protocols using capture effect over multichannels. ACM Transactions on Sensor Networks, 13(2):10:1– 10:??, June 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Majid:2020:DTB

[MDM^{+}20] Amjad Yousef Majid, Carlo Delle Donne, Kiwan Maeng, Alexei Colin, Kasim Sinan Yildirim, Brandon Lucia, and Przemysław Pawełczak. Dynamic task-based intermittent execution for energy-harvesting devices. ACM Transactions on Sensor Networks, 16(1):5:1– 5:24, February 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3360285.

Muzaffar:2021:DEC

[ME21] Shahzad Muzaffar and Ibrahim (Abe) M. Elfadel. Dynamic edge-coded protocols for lowpower, device-to-device communication. ACM Transactions on Sensor Networks, 17(1):8:1– 8:24, January 2021. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3426181.

Mishra:2024:MPB

[MG24] Rahul Mishra and Hari Prabhat Gupta. A model personalizationbased federated learning approach for heterogeneous participants with variability in the dataset. ACM Transactions on Sensor Networks, 20(1):22:1– 22:??, January 2024. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3629978.

MontenegroMarin:2022:ISI

[MGN22] Carlos Enrique Montenegro Marin, Paulo Alonso Gaona Garcia, and Edward Rolando Nuñez Valdez. Introduction to the special issue on artificial intelligence for underwater sensor networks. ACM Transactions on Sensor Networks, 18(3):30:1– 30:??, August 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3557051.

Margolies:2015:EHA

[MGS^{+}15] Robert Margolies, Maria Gorlatova, John Sarik, Gerald Stanje, Jianxun Zhu, Paul Miller, Marcin Szczodrak, Baradwaj Vigraham, Luca Carloni, Peter Kinget, Ioannis Kymissis, and Gil Zussman. Energy-Harvesting Active Networked Tags (En-HANTs): Prototyping and experimentation. ACM Transactions on Sensor Networks, 11 (4):62:1–62:??, December 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Mu:2019:ROS

[MGS^{+}19] Di Mu, Yunpeng Ge, Mo Sha, Steve Paul, Niranjan Ravichandra, and Souma Chowdhury. Robust optimal selection of radio type and transmission power for Internet of Things. ACM Transactions on Sensor Networks, 15 (4):39:1–39:??, October 2019. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

                                        URL https://dl.acm.org/ft_
                                        gateway.cfm?id=3342516.

Miao:2019:PPT

[MJS^{+}19] Chenglin Miao, Wenjun Jiang, Lu Su, Yaliang Li, Suxin Guo, Zhan Qin, Houping Xiao, Jing Gao, and Kui Ren. Privacypreserving truth discovery in crowd sensing systems. ACM Transactions on Sensor Networks, 15(1):9:1–9:??, February 2019. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic).  URL https://dl.
                                        acm.org/ft_gateway.cfm?id=
                                        3277505.

Mohammadi:2023:RDI

[MKFD^{+}23] Mojtaba Mohammadi, Abdollah Kavousi-Fard, Moslem Dehghani, Mazaher Karimi, Vincenzo Loia, Hassan Haes Alhelou, and Pierluigi Siano. Reinforcing data integrity in renewable hybrid AC-DC microgrids from social-economic perspectives. ACM Transactions on Sensor Networks, 19(2):25:1– 25:19, May 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3512891.

Misra:2013:ART

[MKK^{+}13] Prasant Misra, Navinda Kottege, Branislav Kusy, Diethelm Ostry, and Sanjay Jha. Acoustical ranging techniques in embedded wireless sensor networked devices. ACM Transactions on Sensor Networks, 10(1):15:1– 15:??, November 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Miao:2020:QAO

[MKM^{+}20] Xin Miao, Yanrong Kang, Qiang Ma, Kebin Liu, and Lei Chen. Quality-aware online task assignment in mobile crowdsourcing. ACM Transactions on Sensor Networks, 16(3):30:1– 30:21, August 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3397180.

Ming:2022:SCD

[MLS^{+}22] Zhao Ming, Xiuhua Li, Chuan Sun, Qilin Fan, Xiaofei Wang, and Victor C. M. Leung. Sleeping cell detection for resiliency enhancements in 5G/B5G mobile edge-cloud computing networks. ACM Transactions on Sensor Networks, 18(3):42:1– 42:??, August 2022. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3512893.

Meng:2024:AAG

[MLX^{+}24] Xiangwei Meng, Wei Liang, Zisang Xu, Kuanching Li, Muhammad Khurram Khan, and Xiaoyan Kui. An anonymous authenticated group key agreement scheme for transfer learning edge services systems. ACM Transactions on Sensor Networks, 20(3):75:1–75:??, May 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3657292.

Ma:2024:VDQ

[MLZ^{+}24] Chaofan Ma, Wei Liang, Meng Zheng, Xiaofang Xia, and Lin Chen. A Voronoi diagram and Qlearning based relay node placement method subject to radio irregularity. ACM Transactions on Sensor Networks, 20 (1):13:1–13:??, January 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3617124.

Mokaya:2018:MEI

[MNLZ18] Frank Mokaya, Hae Young Noh, Roland Lucas, and Pei Zhang. MyoVibe: Enabling inertial sensor-based muscle activation detection in high-mobility exercise environments. ACM Transactions on Sensor Networks, 14 (1):6:1–6:??, March 2018. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Miller:2010:RER

[MP10] Chris Miller and Christian Poellabauer. Reliable and efficient reprogramming in sensor networks. ACM Transactions on Sensor Networks, 7(1):6:1– 6:??, August 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Mottola:2010:AWS

[MPC^{+}10] Luca Mottola, Gian Pietro Picco, Matteo Ceriotti, Ştefan Gună, and Amy L. Murphy. Not all wireless sensor networks are created equal: a comparative study on tunnels. ACM Transactions on Sensor Networks, 7(2): 15:1–15:??, August 2010. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Minakov:2016:CSR

[MPRS16] Ivan Minakov, Roberto Passerone, Alessandra Rizzardi, and Sabrina Sicari. A comparative study of recent wireless sensor network simulators. ACM Transactions on Sensor Networks, 12(3):20:1– 20:??, August 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Martin:2010:KPH

[MPS10] Keith M. Martin, Maura B. Paterson, and Douglas R. Stinson. Key predistribution for homogeneous wireless sensor networks with group deployment of nodes. ACM Transactions on Sensor Networks, 7(2):11:1– 11:??, August 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Manohar:2009:PCS

[MRM09] Pallavi Manohar, S. Sundhar Ram, and D. Manjunath. Path coverage by a sensor field: The nonhomogeneous case. ACM Transactions on Sensor Networks, 5(2):17:1–17:??, March 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Manulis:2009:SMF

[MS09] Mark Manulis and Jörg Schwenk. Security model and framework for information aggregation in sensor networks. ACM Transactions on Sensor Networks, 5 (2):13:1–13:??, March 2009. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Misra:2012:LPB

[MS12] Sudip Misra and Sweta Singh. Localized policy-based target tracking using wireless sensor networks. ACM Transactions on Sensor Networks, 8(3):27:1– 27:??, July 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Movassaghi:2018:OSA

[MSAJ18] Samaneh Movassaghi, David B. Smith, Mehran Abolhasan, and Abbas Jamalipour. Opportunistic spectrum allocation for interference mitigation amongst coexisting wireless body area networks. ACM Transactions on Sensor Networks, 14(2):7:1– 7:??, July 2018. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Midi:2017:SRP

[MSB17] Daniele Midi, Salmin Sultana, and Elisa Bertino. A system for response and prevention of security incidents in wireless sensor networks. ACM Transactions on Sensor Networks, 13(1):1:1– 1:??, February 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Mishra:2023:HMO

[MSK^{+}23] Alekha Kumar Mishra, Osho Singh, Abhay Kumar, Deepak Puthal, Pradip Kumar Sharma, and Biswajeet Pradhan. Hybrid mode of operation schemes for P2P communication to analyze end-point individual behaviour in IoT. ACM Transactions on Sensor Networks, 19 (2):31:1–31:??, May 2023. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3548686.

Malan:2008:IPK

[MWS08] David J. Malan, Matt Welsh, and Michael D. Smith. Implementing public-key infrastructure for sensor networks. ACM Transactions on Sensor Networks, 4(4):22:1–22:??, August 2008. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Miao:2024:RLF

[MY24] Zhuoyi Miao and Jun Yu. A robust learning framework for smart grids in defense against false-data injection attacks. ACM Transactions on Sensor Networks, 20(2):30:1– 30:??, March 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3588439.

Ma:2024:LAS

[MYH^{+}24] Qiang Ma, Hao Yuan, Zhe Hu, Xu Wang, and Zheng Yang. A liquidity analysis system for large-scale video streams in the oilfield. ACM Transactions on Sensor Networks, 20 (3):65:1–65:??, May 2024. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3649222.

Mei:2024:ELP

[MYW^{+}24] Luoyu Mei, Zhimeng Yin, Shuai Wang, Xiaolei Zhou, Taiwei Ling, and Tian He. ECR-LoRa: LoRa packet recovery under low SNR via edge-cloud collaboration. ACM Transactions on Sensor Networks, 20(2): 40:1–40:??, March 2024. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3604936.

Mo:2023:EOT

[MZKC23] Lei Mo, Qi Zhou, Angeliki Kritikakou, and Xianghui Cao. Energy optimized task mapping for reliable and real-time networked systems. ACM Transactions on Sensor Networks, 19 (4):76:1–76:26, November 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3584985.

Ma:2019:FCS

[MZW^{+}19] Zhi Ma, Sheng Zhang, Jie Wu, Zhuzhong Qian, Yanchao Zhao, and Sanglu Lu. Fast charging scheduling under the nonlinear superposition model with adjustable phases. ACM Transactions on Sensor Networks, 15(4): 48:1–48:??, October 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3356342.

Machado:2010:CPC

[MZWT10] Renita Machado, Wensheng Zhang, Guiling Wang, and Sirin Tekinay. Coverage properties of clustered wireless sensor networks. ACM Transactions on Sensor Networks, 7(2):13:1– 13:??, August 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ning:2010:DST

[NC10] Xu Ning and Christos G. Cassandras. Dynamic sleep time control in wireless sensor networks. ACM Transactions on Sensor Networks, 6(3):21:1– 21:??, June 2010. CODEN ????

ISSN 1550-4859 (print), 1550-4867 (electronic).

Nordio:2010:IQE

[NCV10] Alessandro Nordio, Carla-Fabiana Chiasserini, and Emanuele Viterbo. The impact of quasiequally spaced sensor topologies on signal reconstruction. ACM Transactions on Sensor Networks, 6(2):11:1–11:??, February 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Noshadi:2013:BOD

[NDM^{+}13] Hyduke Noshadi, Foad Dabiri, Saro Meguerdichian, Miodrag Potkonjak, and Majid Sarrafzadeh. Behavior-oriented data resource management in medical sensing systems. ACM Transactions on Sensor Networks, 9(2):12:1–12:??, March 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Nath:2012:TAH

[NEKK12] Swaprava Nath, Venkatesan N. Ekambaram, Anurag Kumar, and P. Vijay Kumar. Theory and algorithms for hop-count-based localization with random geometric graph models of dense sensor networks. ACM Transactions on Sensor Networks, 8 (4):35:1–35:??, September 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Nabi:2014:ECM

[NGBB14] Majid Nabi, Marc Geilen, Twan Basten, and Milos Blagojevic. Efficient cluster mobility support for TDMA-based MAC protocols in wireless sensor networks. ACM Transactions on Sensor Networks, 10(4):65:1– 65:??, June 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Nath:2008:SDR

[NGSA08] Suman Nath, Phillip B. Gibbons, Srinivasan Seshan, and Zachary Anderson. Synopsis diffusion for robust aggregation in sensor networks. ACM Transactions on Sensor Networks, 4(2): 7:1–7:??, March 2008. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Nan:2024:LSV

[NJL24] Ya Nan, Shiqi Jiang, and Mo Li.

Large-scale video analytics with cloud-edge collaborative continuous learning. ACM Transactions on Sensor Networks, 20 (1):14:1–14:??, January 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3624478.

Nguyen:2005:KBL

[NJS05] Xuanlong Nguyen, Michael I.

Jordan, and Bruno Sinopoli. A kernel-based learning approach to ad hoc sensor network localization. ACM Transactions on Sensor Networks, 1(1):134– 152, August 2005. CODEN ????

ISSN 1550-4859 (print), 1550-4867 (electronic).

Noh:2018:ISI

[NJZ18] Hae Young Noh, Xiaofan (Fred) Jiang, and Pei Zhang. Introduction to the special issue on BuildSys’17. ACM Transactions on Sensor Networks, 14(3–4): 16:1–16:??, December 2018. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Nguyen:2014:CMF

[NK14] Diep N. Nguyen and Marwan Krunz. A cooperative MIMO framework for wireless sensor networks. ACM Transactions on Sensor Networks, 10(3):43:1– 43:??, April 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Naveen:2015:RSC

[NK15] K. P. Naveen and Anurag Kumar. Relay selection with channel probing in sleep-wake cycling wireless sensor networks. ACM Transactions on Sensor Networks, 11(3):52:1–52:??, May 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ning:2008:MAA

[NLD08] Peng Ning, An Liu, and Wenliang Du. Mitigating DoS attacks against broadcast authentication in wireless sensor networks. ACM Transactions on Sensor Networks, 4(1):1:1–1:??, January 2008. CODEN ????

ISSN 1550-4859 (print), 1550-4867 (electronic).

Niu:2019:REA

[NLH^{+}19] Qun Niu, Mingkuan Li, Suining He, Chengying Gao, S.- H. Gary Chan, and Xiaonan Luo. Resource-efficient and automated image-based indoor localization. ACM Transactions on Sensor Networks, 15 (2):19:1–19:??, April 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3284555.

Ni:2012:SND

[NP12] Kevin Ni and Greg Pottie. Sensor network data fault detection with maximum a posteriori selection and Bayesian modeling. ACM Transactions on Sensor Networks, 8(3):23:1–23:??, July 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ni:2009:SND

[NRC^{+}09] Kevin Ni, Nithya Ramanathan, Mohamed Nabil Hajj Chehade, Laura Balzano, Sheela Nair, Sadaf Zahedi, Eddie Kohler, Greg Pottie, Mark Hansen, and Mani Srivastava. Sensor network data fault types. ACM Transactions on Sensor Networks, 5(3): 25:1–25:??, May 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ning:2022:RST

[NXW^{+}22] Jingyi Ning, Lei Xie, Chuyu Wang, Yanling Bu, Fu Xiao, Baoliu Ye, and Sanglu Lu. Revolving scanning on tagged objects: 3D structure detection of logistics packages via RFID systems. ACM Transactions on Sensor Networks, 18 (2):20:1–20:29, May 2022. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3490171.

Niu:2023:VTE

[NZH^{+}23] Qun Niu, Kunxin Zhu, Suining He, Shaoqi Cen, S.-H. Gary Chan, and Ning Liu. VILL: Toward efficient and automatic visual landmark labeling. ACM Transactions on Sensor Networks, 19(4):74:1– 74:25, November 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3580497.

Nguyen:2015:GEE

[NZLH15] Nam Tuan Nguyen, Rong Zheng, Jie Liu, and Zhu Han. Green-Locs: an energy-efficient indoor place identification framework. ACM Transactions on Sensor Networks, 11(3):43:1– 43:??, February 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Nguyen:2021:SSI

[NZM21] Vanh Khuyen Nguyen, Wei Emma Zhang, and Adnan Mahmood.

Semi-supervised intrusive appliance load monitoring in smart energy monitoring system. ACM Transactions on Sensor Networks, 17(3):32:1–32:20, June 2021. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3448415.

Ni:2010:DRS

[NZR10] Jinfeng Ni, Li Zhou, and Chinya V. Ravishankar. Dealing with random and selective attacks in wireless sensor systems. ACM Transactions on Sensor Networks, 6(2):15:1– 15:??, February 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Odonovan:2013:GSW

[OBB^{+}13] Tony O’donovan, James Brown, Felix Büsching, Alberto Cardoso, José Cecílio, Jose Do Ó, Pedro Furtado, Paulo Gil, Anja Jugel, Wolf-Bastian Pöttner, Utz Roedig, Jorge Sá Silva, Ricardo Silva, Cormac J. Sreenan, Vasos Vassiliou, Thiemo Voigt, Lars Wolf, and Zinon Zinonos. The GINSENG system for wireless monitoring and control: Design and deployment experiences. ACM Transactions on Sensor Networks, 10(1):4:1–4:??, November 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Oller:2013:DDP

[ODCP13] Joaquim Oller, Ilker Demirkol, Jordi Casademont, and Josep Paradells. Design, development, and performance evaluation of a low-cost, low-power wake-up radio system for wireless sensor networks. ACM Transactions on Sensor Networks, 10(1):11:1– 11:??, November 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Osborne:2012:RTI

[ORRJ12] Michael A. Osborne, Stephen J. Roberts, Alex Rogers, and Nicholas R. Jennings. Real-time information processing of environmental sensor network data using Bayesian Gaussian processes. ACM Transactions on Sensor Networks, 9(1):1:1–1:??, November 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ouyang:2023:CCB

[OXZ^{+}23] Xiaomin Ouyang, Zhiyuan Xie, Jiayu Zhou, Guoliang Xing, and Jianwei Huang. ClusterFL: a clustering-based federated learning system for human activity recognition. ACM Transactions on Sensor Networks, 19 (1):17:1–17:32, February 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3554980.

Prabh:2005:ECD

[PA05] K. Shashi Prabh and Tarek F. Abdelzaher. Energy-conserving data cache placement in sensor networks. ACM Transactions on Sensor Networks, 1(2):178–203, November 2005. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Pan:2022:LES

[PAYL22] Qingrui Pan, Zhenlin An, Lei Yang, and Qiongzheng Lin. LSAB: Enhancing spatiotemporal efficiency of AoA tracking systems. ACM Transactions on Sensor Networks, 18 (4):58:1–58:??, November 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3534123.

Panta:2011:EIC

[PBM11] Rajesh Krishna Panta, Saurabh Bagchi, and Samuel P. Midkiff. Efficient incremental code update for sensor networks. ACM Transactions on Sensor Networks, 7(4):30:1–30:??, February 2011. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Paschalidis:2010:SAD

[PC10] Ioannis Ch. Paschalidis and Yin Chen. Statistical anomaly detection with sensor networks. ACM Transactions on Sensor Networks, 7(2):17:1–17:??, August 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Pal:2023:CUS

[PCA^{+}23] Amitangshu Pal, Filippo Campagnaro, Khadija Ashraf, Md Rashed Rahman, Ashwin Ashok, and Hongzhi Guo. Communication for underwater sensor networks: a comprehensive summary. ACM Transactions on Sensor Networks, 19(1):22:1– 22:44, February 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3546827.

Premnath:2014:EHR

[PCPK14] Sriram Nandha Premnath, Jessica Croft, Neal Patwari, and Sneha Kumar Kasera. Efficient high-rate secret key extraction in wireless sensor networks using collaboration. ACM Transactions on Sensor Networks, 11(1): 2:1–2:??, August 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Porter:2013:MSE

[PCR13] Barry Porter, Geoff Coulson, and Utz Roedig. Managing software evolution in large-scale wireless sensor and actuator networks. ACM Transactions on Sensor Networks, 9(4):54:1– 54:??, July 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Padhy:2010:UBA

[PDMJ10] Paritosh Padhy, Rajdeep K. Dash, Kirk Martinez, and Nicholas R. Jennings. A utilitybased adaptive sensing and multihop communication protocol for wireless sensor networks. ACM Transactions on Sensor Networks, 6(3):27:1–27:??, June 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Penil:2017:HLD

[PDP^{+}17] Pablo Peñil, Alvaro Díaz, Hector Posadas, Julio Medina, and Pablo Sánchez. High-level design of wireless sensor networks for performance optimization under security hazards. ACM Transactions on Sensor Networks, 13 (3):19:1–19:??, September 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Park:2013:DCO

[PEFSV13] Pangun Park, Sinem Coleri Ergen, Carlo Fischione, and Alberto Sangiovanni-Vincentelli. Duty-cycle optimization for IEEE 802.15.4 wireless sensor networks. ACM Transactions on Sensor Networks, 10(1):12:1– 12:??, November 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Park:2013:MSA

[PFJ13] Pangun Park, Carlo Fischione, and Karl Henrik Johansson. Modeling and stability analysis of hybrid multiple access in the IEEE 802.15.4 protocol. ACM Transactions on Sensor Networks, 9(2):13:1–13:??, March 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Paschalidis:2009:RDS

[PG09] Ioannis Ch. Paschalidis and Dong Guo. Robust and distributed stochastic localization in sensor networks: Theory and experimental results. ACM Transactions on Sensor Networks, 5(4):34:1–34:??, November 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Paek:2010:RRC

[PG10] Jeongyeup Paek and Ramesh Govindan. RCRT: Ratecontrolled reliable transport protocol for wireless sensor networks. ACM Transactions on Sensor Networks, 7(3):20:1– 20:??, September 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Paek:2010:TAT

[PGG^{+}10] Jeongyeup Paek, Ben Greenstein, Omprakash Gnawali, Ki-Young Jang, August Joki, Marcos Vieira, John Hicks, Deborah Estrin, Ramesh Govindan, and Eddie Kohler. The Tenet architecture for tiered sensor networks. ACM Transactions on Sensor Networks, 6(4):34:1– 34:??, July 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Pal:2024:TWU

[PGY^{+}24] Amitangshu Pal, Hongzhi Guo, Sijung Yang, Mustafa Alper Akkas, and Xufeng Zhang. Taking wireless underground: a comprehensive summary. ACM Transactions on Sensor Networks, 20(1):19:1–19:??, January 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3587934.

Puccinelli:2010:RDD

[PH10] Daniele Puccinelli and Martin Haenggi. Reliable data delivery in large-scale low-power sensor networks. ACM Transactions on Sensor Networks, 6(4): 28:1–28:??, July 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Pham:2016:QLR

[Pha16]  Congduc Pham.  QoS for

long-range wireless sensors under duty-cycle regulations with shared activity time usage. ACM Transactions on Sensor Networks, 12(4):33:1–33:??, November 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Park:2017:ESN

[PHKK17] Yongtae Park, Jihun Ha, Hyogon Kim, and Jeonggil Ko. Enabling sensor network to Smartphone interaction using software radios. ACM Transactions on Sensor Networks, 13(1):2:1–2:??, February 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Pal:2019:WFD

[PK19] Amitangshu Pal and Krishna Kant. Water flow driven sensor networks for leakage and contamination monitoring in distribution pipelines. ACM Transactions on Sensor Networks, 15(4):

37:1–37:??, October 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3342513.

Pal:2020:SSC

[PK20] Amitangshu Pal and Krishna Kant. Smart sensing, communication, and control in perishable food supply chain. ACM Transactions on Sensor Networks, 16 (1):12:1–12:41, February 2020. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3360726.

Pannuto:2018:HUW

[PKC^{+}18] Pat Pannuto, Benjamin Kempke, Li-Xuan Chuo, David Blaauw, and Prabal Dutta. Harmonium: Ultra wideband pulse generation with bandstitched recovery for fast, accurate, and robust indoor localization. ACM Transactions on Sensor Networks, 14(2):11:1– 11:??, July 2018. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Pattem:2008:ISC

[PKG08] Sundeep Pattem, Bhaskar Krishnamachari, and Ramesh Govindan. The impact of spatial correlation on routing with compression in wireless sensor networks. ACM Transactions on Sensor Networks, 4(4):24:1– 24:??, August 2008. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Pramanik:2023:ALL

[PKS^{+}23] Prithviraj Pramanik, Prasenjit Karmakar, Praveen Kumar Sharma, Soumyajit Chatterjee, Abhijit Roy, Santanu Mandal, Subrata Nandi, Sandip Chakraborty, Mousumi Saha, and Sujoy Saha. AQuaMoHo: Localized low-cost outdoor air quality sensing over a thermohygrometer. ACM Transactions on Sensor Networks, 19(3):69:1– 69:??, August 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3580279.

Pang:2024:ATC

[PLW^{+}24] Bowen Pang, Sicong Liu, Hongli Wang, Bin Guo, Yuzhan Wang, Hao Wang, Zhenli Sheng, Zhongyi Wang, and Zhiwen Yu. AdaMEC: Towards a contextadaptive and dynamically combinable DNN deployment framework for mobile edge computing. ACM Transactions on Sensor Networks, 20(1):21:1– 21:??, January 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3630098.

Pietro:2012:SHU

[PMST12] Roberto Di Pietro, Di Ma, Claudio Soriente, and Gene Tsudik. Self-healing in unattended wireless sensor networks. ACM Transactions on Sensor Networks, 9(1):7:1–7:??, November 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Pham:2022:MLD

[PNL^{+}22] Van-Trung Pham, Tu N. Nguyen, Bing-Hong Liu, My T. Thai, Braulio Dumba, and Tong Lin. Minimizing latency for data aggregation in wireless sensor networks: an algorithm approach. ACM Transactions on Sensor Networks, 18(3):30:1– 30:??, August 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3450350.

Panigrahi:2015:ESN

[PPM15] Trilochan Panigrahi, Ganapati Panda, and Bernard Mulgrew. Error saturation nonlinearities for robust incremental LMS over wireless sensor networks. ACM Transactions on Sensor Networks, 11(2):27:1– 27:??, February 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Peleg:2010:LSC

[PR10] David Peleg and Liam Roditty. Localized spanner construction for ad hoc networks with variable transmission range. ACM Transactions on Sensor Networks, 7 (3):25:1–25:??, September 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Peyravi:2017:LMD

[PS17] Hassan Peyravi and Rahul Sehgal. Link modeling and delay analysis in networks with disruptive links. ACM Transactions on Sensor Networks, 13(4):31:1– 31:??, December 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Pottner:2014:CST

[PSB^{+}14] Wolf-Bastian Pöttner, Hans Seidel, James Brown, Utz Roedig, and Lars Wolf. Constructing schedules for time-critical data delivery in wireless sensor networks. ACM Transactions on Sensor Networks, 10(3):44:1– 44:??, April 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Pradhan:2022:TCU

[PSR^{+}22] B. Pradhan, Gautam Srivastava, D. S. Roy, K. H. K. Reddy, and Jerry Chun-Wei Lin. Traffic classification in underwater networks using SDN and data-driven hybrid metaheuristics. ACM Transactions on Sensor Networks, 18(3):34:1– 34:??, August 2022. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3474556.

Perazzo:2016:SPW

[PTDD16] Pericle Perazzo, Lorenzo Taponecco, Antonio A. D’amico, and Gianluca Dini. Secure positioning in wireless sensor networks through enlargement miscontrol detection. ACM Transactions on Sensor Networks, 12(4):27:1– 27:??, November 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ping:2023:NLI

[PWS^{+}23] Haodi Ping, Yongcai Wang, Xingfa Shen, Deying Li, and Wenping Chen. On node localizability identification in barycentric linear localization. ACM Transactions on Sensor Networks, 19(1):19:1–19:26, February 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3547143.

Pongaliur:2013:SNS

[PX13] Kanthakumar Pongaliur and Li Xiao. Sensor node source privacy and packet recovery under eavesdropping and node compromise attacks. ACM Transactions on Sensor Networks, 9(4):50:1– 50:??, July 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Peng:2021:TPC

[PZOZ21] Chaoqun Peng, Xinglin Zhang, Zhaojing Ou, and Junna Zhang. Task planning considering location familiarity in spatial crowdsourcing. ACM Transactions on Sensor Networks, 17(2):16:1– 16:24, June 2021. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3442698.

Qin:2013:MUA

[QM13] Fei Qin and John E. Mitchell. AS-MAC: Utilizing the adaptive spreading code length for wireless sensor networks. ACM Transactions on Sensor Networks, 10(1):1:1–1:??, November 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Quan:2022:GMN

[QNN^{+}22] La Van Quan, Minh Hieu Nguyen, Thanh Hung Nguyen, Kien Nguyen, and Phi Le Nguyen. On the global maximization of network lifetime in wireless rechargeable sensor networks. ACM Transactions on Sensor Networks, 18(4):71:1– 71:??, November 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3510423.

Qi:2022:SEI

[QWC^{+}22] Saiyu Qi, Wei Wei, Jingxian Cheng, Yuanqing Zheng, Zhou Su, Jingning Zhang, and Yong Qi. Secure and efficient item traceability for cloudaided IIoT. ACM Transactions on Sensor Networks, 18 (4):54:1–54:??, November 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522740.

Qiu:2022:ISS

[QXZZ22] Meikang Qiu, Ke Xu, Cheng Zhang, and Tianwei Zhang. Introduction to the special section on energy-efficient and secure computing for artificial intelligence and beyond. ACM Transactions on Sensor Networks, 18 (4):51:1–51:??, November 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3558553.

Razzaque:2013:CWS

[RBD13] M. A. Razzaque, Chris Bleakley, and Simon Dobson. Compression in wireless sensor networks: a survey and comparative evaluation. ACM Transactions on Sensor Networks, 10(1):5:1– 5:??, November 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Rajasegarar:2009:EAW

[RBLP09] Sutharshan Rajasegarar, James C. Bezdek, Christopher Leckie, and Marimuthu Palaniswami. Elliptical anomalies in wireless sensor networks. ACM Transactions on Sensor Networks, 6(1):7:1–7:??, December 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Rathore:2016:CAS

[RBS16] Heena Rathore, Venkataramana Badarla, and Supratim Shit. Consensus-aware sociopsychological trust model for wireless sensor networks. ACM Transactions on Sensor Networks, 12(3): 21:1–21:??, August 2016. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Restuccia:2016:OLS

[RD16] Francesco Restuccia and Sajal K. Das. Optimizing the lifetime of sensor networks with uncontrollable mobile sinks and QoS constraints. ACM Transactions on Sensor Networks, 12(1): 2:1–2:??, March 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Restuccia:2016:IMP

[RDP16] Francesco Restuccia, Sajal K.

Das, and Jamie Payton. Incentive mechanisms for participatory sensing: Survey and research challenges. ACM Transactions on Sensor Networks, 12 (2):13:1–13:??, May 2016. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ramachandran:2007:ACA

[RDR07] Iyappan Ramachandran, Arindam K.

Das, and Sumit Roy. Analysis of the contention access period of IEEE 802.15.4 MAC. ACM Transactions on Sensor Networks, 3(1):??, March 2007. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ramos:2014:TRM

[RFB^{+}14] Heitor S. Ramos, Alejandro C.

Frery, Azzedine Boukerche, Eduardo M. R. Oliveira, and Antonio A. F. Loureiro. Topologyrelated metrics and applications for the design and operation of wireless sensor networks. ACM Transactions on Sensor Networks, 10(3):53:1–53:??, April 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Restuccia:2019:FFO

[RFS^{+}19] Francesco Restuccia, Pierluca Ferraro, Timothy S. Sanders, Simone Silvestri, Sajal K. Das, and Giuseppe Lo Re. FIRST: a framework for optimizing information quality in mobile crowdsensing systems. ACM Transactions on Sensor Networks, 15(1): 5:1–5:??, February 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3267105.

Restuccia:2017:QIM

[RGB^{+}17] Francesco Restuccia, Nirnay Ghosh, Shameek Bhattacharjee, Sajal K. Das, and Tommaso Melodia. Quality of information in mobile crowdsensing: Survey and research challenges. ACM Transactions on Sensor Networks, 13(4):34:1–34:??, December 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Razzaque:2017:QBA

[RHD17] M. A. Razzaque, Muta Tah Hira, and Mukta Dira. QoS in body area networks: a survey. ACM Transactions on Sensor Networks, 13(3):25:1–25:??, September 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Renner:2020:AIL

[RHS20] Bernd-Christian Renner, Jan Heitmann, and Fabian Steinmetz. ahoi: Inexpensive, lowpower communication and localization for underwater sensor networks and μAUVs. ACM Transactions on Sensor Networks, 16(2):18:1–18:46, April 2020. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/abs/10.1145/3376921.

Rowaihy:2010:SMA

[RJL^{+}10] Hosam Rowaihy, Matthew P.

Johnson, Ou Liu, Amotz Bar-Noy, Theodore Brown, and Thomas La Porta. Sensormission assignment in wireless sensor networks. ACM Transactions on Sensor Networks, 6(4): 36:1–36:??, July 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Rajamani:2009:IGA

[RKJ09] Vasanth Rajamani, Sanem Kabadayi, and Christine Julien. An interrelational grouping abstraction for heterogeneous sensors. ACM Transactions on Sensor Networks, 5(3):27:1–27:??, May 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Rezaei:2023:SPP

[RKLM23] Yoones Rezaei, Talha Khan, Stephen Lee, and Daniel Mossé. Solar-powered parking analytics system using deep reinforcement learning. ACM Transactions on Sensor Networks, 19 (4):75:1–75:27, November 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3584949.

Rathore:2017:MEB

[RKRP17] Punit Rathore, Dheeraj Kumar, Sutharshan Rajasegarar, and Marimuthu Palaniswami. Maximum entropy-based auto drift correction using high- and lowprecision sensors. ACM Transactions on Sensor Networks, 13 (3):24:1–24:??, September 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Ramachandran:2006:DDF

[RKW^{+}06] Umakishore Ramachandran, Rajnish Kumar, Matthew Wolenetz, Brian Cooper, Bikash Agarwalla, Junsuk Shin, Phillip Hutto, and Arnab Paul. Dynamic data fusion for future sensor networks. ACM Transactions on Sensor Networks, 2(3): 404–443, August 2006. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Reddy:2010:UMP

[RMB^{+}10] Sasank Reddy, Min Mun, Jeff Burke, Deborah Estrin, Mark Hansen, and Mani Srivastava. Using mobile phones to determine transportation modes. ACM Transactions on Sensor Networks, 6(2):13:1–13:??, February 2010. CODEN ????

ISSN 1550-4859 (print), 1550-4867 (electronic).

Ruj:2009:KPU

[RR09] Sushmita Ruj and Bimal Roy.

Key predistribution using combinatorial designs for grid-group deployment scheme in wireless sensor networks. ACM Transactions on Sensor Networks, 6(1): 4:1–4:??, December 2009. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

R:2022:ABE

[RRA22] Kanthavel R., Dhaya R., and Ahilan A. AI-based efficient WUGS network channel modeling and clustered cooperative communication. ACM Transactions on Sensor Networks, 18(3): 33:1–33:??, August 2022. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3469034.

Reijers:2019:IAT

[RS19] Niels Reijers and Chi-Sheng Shih. Improved ahead-of-time compilation of stack-based JVM Bytecode on resourceconstrained devices. ACM Transactions on Sensor Networks, 15(3):34:1–34:??, August 2019. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic).  URL https://dl.
acm.org/ft_gateway.cfm?id=
3341170.

Roy:2021:OSD

[RSK^{+}21] Dhrubojyoti Roy, Sangeeta Srivastava, Aditya Kusupati, Pranshu Jain, Manik Varma, and Anish Arora. One size does not fit all: Multi-scale, cascaded RNNs for radar classification. ACM Transactions on Sensor Networks, 17(2):12:1–12:27, June 2021. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3439957.

Saeed:2019:RTC

[SAK^{+}19] Ahmed Saeed, Ahmed Abdelkader, Mouhyemen Khan, Azin Neishaboori, Khaled A. Harras, and Amr Mohamed. On realistic target coverage by autonomous drones. ACM Transactions on Sensor Networks, 15(3):32:1–32:??, August 2019. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3325512.

Sang:2010:LAO

[SAZ10] Lifeng Sang, Anish Arora, and Hongwei Zhang. On link asymmetry and one-way estimation in wireless sensor networks. ACM Transactions on Sensor Networks, 6(2):12:1–12:??, February 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Sharma:2016:NOD

[SB16] Gokarna Sharma and Costas Busch. Near-optimal deterministic Steiner tree maintenance in sensor networks. ACM Transactions on Sensor Networks, 12(1): 4:1–4:??, March 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Salimitari:2020:PTA

[SBCF20] Mehrdad Salimitari, Shameek Bhattacharjee, Mainak Chatterjee, and Yaser P. Fallah. A prospect theoretic approach for trust management in IoT networks under manipulation attacks. ACM Transactions on Sensor Networks, 16(3):26:1– 26:26, August 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3392058.

Sangar:2022:NTI

[SBK22] Yaman Sangar, Yoganand Biradavolu, and Bhuvana Krishnaswamy. A novel timeinterval based modulation for large-scale, low-power, widearea-networks. ACM Transactions on Sensor Networks, 18 (4):68:1–68:??, November 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3549543.

Shah:2018:DGC

[SBSD18] Vijay K. Shah, Shameek Bhattacharjee, Simone Silvestri, and Sajal K. Das. Designing green communication systems for smart and connected communities via dynamic spectrum access. ACM Transactions on Sensor Networks, 14(3–4):31:1– 31:??, December 2018. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Sun:2012:QCC

[SC12] Xusheng Sun and Edward J. Coyle. Quantization, channel compensation, and optimal energy allocation for estimation in sensor networks. ACM Transactions on Sensor Networks, 8 (2):15:1–15:??, March 2012. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Salmani:2015:RRR

[SC15] Vahid Salmani and Pai H. Chou. Resilient round robin: a lightweight deterministic MAC primitive. ACM Transactions on Sensor Networks, 11(2):31:1– 31:??, February 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Sorbelli:2024:DBB

[SCD^{+}24] Francesco Betti Sorbelli, Federico Coró, Sajal K. Das, Lorenzo Palazzetti, and Cristina M. Pinotti. Drone-based bug detection in orchards with nets: a novel orienteering approach. ACM Transactions on Sensor Networks, 20(3):68:1–68:??, May 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3653713.

Shu:2015:TLW

[SCG^{+}15] Yuanchao Shu, Peng Cheng, Yu Gu, Jiming Chen, and Tian He. TOC: Localizing wireless rechargeable sensors with time of charge. ACM Transactions on Sensor Networks, 11(3):44:1– 44:??, February 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Schieferdecker:2015:LFD

[Sch15] Dennis Schieferdecker. Locationfree detection of network boundaries. ACM Transactions on Sensor Networks, 11(4):58:1– 58:??, December 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Shan:2014:BML

[SCL^{+}14] Mengfan Shan, Guihai Chen, Dijun Luo, Xiaojun Zhu, and Xiaobing Wu. Building maximum lifetime shortest path data aggregation trees in wireless sensor networks. ACM Transactions on Sensor Networks, 11(1):11:1– 11:??, August 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Shi:2019:DSC

[SCL^{+}19] Tuo Shi, Siyao Cheng, Jianzhong Li, Hong Gao, and Zhipeng Cai. Dominating sets construction in RF-based battery-free sensor networks with full coverage guarantee. ACM Transactions on Sensor Networks, 15(4): 43:1–43:??, October 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3352486.

Shi:2024:OAU

[SCLG24] Tuo Shi, Zhipeng Cai, Jianzhong Li, and Hong Gao. Optimize the age of useful information in edgeassisted energy-harvesting sensor networks. ACM Transactions on Sensor Networks, 20(2): 49:1–49:??, March 2024. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3640342.

Shi:2022:ECTb

[SCS22] Junyang Shi, Xingjian Chen, and Mo Sha. Enabling crosstechnology communication from LoRa to ZigBee in the 2.4 GHz band. ACM Transactions on Sensor Networks, 18 (2):21:1–21:23, May 2022. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3491222.

Sheu:2013:ACC

[SCWC13] Jang-Ping Sheu, Guey-Yun Chang, Shan-Hung Wu, and Yen-Ting Chen. Adaptive kcoverage contour evaluation and deployment in wireless sensor networks. ACM Transactions on Sensor Networks, 9(4):40:1– 40:??, July 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Sutton:2019:BLL

[SDBT19] Felix Sutton, Reto Da Forno, Jan Beutel, and Lothar Thiele. BLITZ: Low latency and energyefficient communication for event-triggered wireless sensing systems. ACM Transactions on Sensor Networks, 15 (2):25:1–25:??, April 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3309702.

Strasser:2010:DRJ

[SDČ10] Mario Strasser, Boris Danev, and Srdjan Čapkun. Detection of reactive jamming in sensor networks. ACM Transactions on Sensor Networks, 7(2):16:1– 16:??, August 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Srinivasan:2010:ESL

[SDTL10] Kannan Srinivasan, Prabal Dutta, Arsalan Tavakoli, and Philip Levis. An empirical study of low-power wireless. ACM Transactions on Sensor Networks, 6(2):16:1–16:??, February 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Sun:2023:ALA

[SDW^{+}23] Xue Sun, Wenwen Deng, Xudong Wei, Dingyi Fang, Baochun Li, and Xiaojiang Chen. Akte-Liquid: Acousticbased liquid identification with smartphones. ACM Transactions on Sensor Networks, 19 (1):18:1–18:24, February 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3551640.

Shen:2020:SCP

[SDX^{+}20] Yiran Shen, Bowen Du, Weitao Xu, Chengwen Luo, Bo Wei, Lizhen Cui, and Hongkai Wen. Securing cyber-physical social interactions on wrist-worn devices. ACM Transactions on Sensor Networks, 16(2): 19:1–19:22, April 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3378669.

Shi:2022:EEP

[SDYC22] Yimin Shi, Haihan Duan, Lei Yang, and Wei Cai. An energy-efficient and privacyaware decomposition framework for edge-assisted federated learning. ACM Transactions on Sensor Networks, 18(4):53:1– 53:??, November 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3522741.

Shen:2024:RST

[SDZZ24] Zhihao Shen, Wan Du, Xi Zhao, and Jianhua Zou. Retrieving similar trajectories from cellular data of multiple carriers at city scale. ACM Transactions on Sensor Networks, 20(2): 47:1–47:??, March 2024. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3613245.

Samaddar:2023:OSR

[SE23] Ankita Samaddar and Arvind Easwaran. Online schedule randomization to mitigate timing attacks in 5G periodic URLLC communications. ACM Transactions on Sensor Networks, 19 (4):93:1–93:26, November 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3600093.

Sundaram:2013:DTW

[SEZA13] Vinaitheerthan Sundaram, Patrick Eugster, Xiangyu Zhang, and Vamsidhar Addanki. Diagnostic tracing for wireless sensor networks. ACM Transactions on Sensor Networks, 9(4):38:1– 38:??, July 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Sugihara:2008:PMS

[SG08] Ryo Sugihara and Rajesh K.

Gupta. Programming models for sensor networks: a survey. ACM Transactions on Sensor Networks, 4(2):8:1–8:??, March 2008. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Sugihara:2010:SCS

[SG10] Ryo Sugihara and Rajesh K.

Gupta. Speed control and scheduling of data mules in sensor networks. ACM Transactions on Sensor Networks, 7(1): 4:1–4:??, August 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Sugihara:2011:PPD

[SG11] Ryo Sugihara and Rajesh K. Gupta. Path planning of data mules in sensor networks. ACM Transactions on Sensor Networks, 8(1):1:1–1:??, August 2011. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Steine:2015:DRA

[SGB15] Marcel Steine, Marc Geilen, and Twan Basten. A distributed reconfiguration approach for quality-of-service provisioning in dynamic heterogeneous wireless sensor networks. ACM Transactions on Sensor Networks, 11 (2):34:1–34:??, February 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Sharma:2010:SFD

[SGG10] Abhishek B. Sharma, Leana Golubchik, and Ramesh Govindan. Sensor faults: Detection methods and prevalence in real-world datasets. ACM Transactions on Sensor Networks, 6(3):23:1– 23:??, June 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Sengul:2008:APB

[SGM08] Cigdem Sengul, Indranil Gupta, and Matthew J. Miller. Adaptive probability-based broadcast forwarding in energy-saving sensor networks. ACM Transactions on Sensor Networks, 4(2): 6:1–6:??, March 2008. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Shi:2009:OBS

[SH09] Yi Shi and Y. Thomas Hou. Optimal base station placement in wireless sensor networks. ACM Transactions on Sensor Networks, 5(4):32:1–32:??, November 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Shen:2020:COM

[SHWW20] Shihao Shen, Yiwen Han, Xiaofei Wang, and Yan Wang. Computation offloading with multiple agents in edge-computing-supported IoT. ACM Transactions on Sensor Networks, 16(1): 8:1–8:27, February 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3372025.

Syed:2013:TRM

[SHY13] Affan A. Syed, John Heidemann, and Wei Ye. Tones for real: Managing multipath in underwater acoustic wakeup. ACM Transactions on Sensor Networks, 9 (2):27:1–27:??, March 2013. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Sangogboye:2018:FPP

[SJH^{+}18] Fisayo Caleb Sangogboye, Ruoxi Jia, Tianzhen Hong, Costas Spanos, and Mikkel Baun Kjærgaard. A framework for privacypreserving data publishing with enhanced utility for cyberphysical systems. ACM Transactions on Sensor Networks, 14(3– 4):30:1–30:??, December 2018.

CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Sangaiah:2022:IQS

[SJP^{+}22] Arun Kumar Sangaiah, Amir Javadpour, Pedro Pinto, Forough Ja’fari, and Weizhe Zhang. Improving quality of service in 5G resilient communication with the cellular structure of smartphones. ACM Transactions on Sensor Networks, 18(3): 43:1–43:??, August 2022. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3512890.

Singh:2011:MTT

[SKM^{+}11] Jaspreet Singh, Rajesh Kumar, Upamanyu Madhow, Subhash Suri, and Richard Cagley. Multiple-target tracking with binary proximity sensors. ACM Transactions on Sensor Networks, 8(1):5:1–5:??, August 2011. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Shen:2022:TMD

[SLC^{+}22] Xingfa Shen, Chuang Li, Weijie Chen, Yongcai Wang, and Quanbo Ge. Transition modeldriven unsupervised localization framework based on crowdsensed trajectory data. ACM Transactions on Sensor Networks, 18(2):26:1–26:21, May 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3499425.

Sun:2022:SAC

[SLS^{+}22] Qindong Sun, Kai Lin, Chengxiang Si, Yanyue Xu, Shancang Li, and Prosanta Gope. A secure and anonymous communicate scheme over the Internet of Things. ACM Transactions on Sensor Networks, 18(3): 40:1–40:??, August 2022. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3508392.

Silva:2018:FPD

[SML18] Nuno Silva, Eduardo R. B. Marques, and Luís M. B. Lopes. Flux: a platform for dynamically reconfigurable mobile crowdsensing. ACM Transactions on Sensor Networks, 14(3–4):20:1– 20:??, December 2018. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Shrivastava:2009:TTB

[SMMS09] Nisheeth Shrivastava, Raghuraman Mudumbai, Upamanyu Madhow, and Subhash Suri. Target tracking with binary proximity sensors. ACM Transactions on Sensor Networks, 5 (4):30:1–30:??, November 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Sen:2014:RRP

[SMR^{+}14] Rijurekha Sen, Abhinav Maurya, Bhaskaran Raman, Rupesh Mehta, Ramkrishnan Kalyanaraman, and Amarjeet Singh. Road-RFSense: a practical RF sensing-based road traffic estimation system for developing regions. ACM Transactions on Sensor Networks, 11(1):4:1–4:??, August 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Shi:2022:ECTa

[SMS22] Junyang Shi, Di Mu, and Mo Sha. Enabling crosstechnology communication from LoRa to ZigBee via payload encoding in sub-1 GHz bands. ACM Transactions on Sensor Networks, 18(1):6:1–6:26, February 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3470452.

Shah:2023:ISI

[SMW23] Syed Hassan A. Shah, Shahid Mumtaz, and Wei Wei. Introduction to the special issue on cognitive computing for Internet of Medical Things in smart healthcare. ACM Transactions on Sensor Networks, 19(3):48:1–48:??, August 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3584742.

Sun:2017:ITC

[SMZ^{+}17] Boyuan Sun, Qiang Ma, Shanfeng Zhang, Kebin Liu, and Yunhao Liu. iSelf: Towards cold-start emotion labeling using transfer learning with Smartphones. ACM Transactions on Sensor Networks, 13(4):30:1– 30:??, December 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Sharma:2023:BBP

[SNC^{+}23] Pratima Sharma, Suyel Namasudra, Naveen Chilamkurti, Byung-Gyu Kim, and Ruben Gonzalez Crespo. Blockchainbased privacy preservation for IoT-Enabled healthcare system. ACM Transactions on Sensor Networks, 19(3):56:1– 56:??, August 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3577926.

Sah:2022:LEM

[SNK^{+}22] Dinesh Kumar Sah, Tu N.

Nguyen, Manjusha Kandulna, Korhan Cengiz, and Tarachand Amgoth. 3D localization and error minimization in underwater sensor networks. ACM Transactions on Sensor Networks, 18(3): 31:1–31:??, August 2022. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3460435.

Sun:2024:FEE

[SNY^{+}24] Zehua Sun, Tao Ni, Huanqi Yang, Kai Liu, Yu Zhang, Tao Gu, and Weitao Xu. FLoRa+: Energy-efficient, reliable, beamforming-assisted, and secure over-the-air firmware update in LoRa networks. ACM Transactions on Sensor Networks, 20(3):54:1–54:??, May 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3641548.

Saxena:2024:HEA

[SPI^{+}24] Ravi Raj Saxena, Joydeep Pal, Srinivasan Iyengar, Bhawana Chhaglani, Anurag Ghosh, Venkata N. Padmanabhan, and Prabhakar T. Venkata. Holistic energy awareness and robustness for intelligent drones. ACM Transactions on Sensor Networks, 20(3):57:1–57:??, May 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3641855.

Shen:2010:EDD

[SPK^{+}10] Chung-Ching Shen, William L. Plishker, Dong-Ik Ko, Shuvra S. Bhattacharyya, and Neil Goldsman. Energy-driven distribution of signal processing applications across wireless sensor networks. ACM Transactions on Sensor Networks, 6(3):24:1–24:??, June 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Shin:2014:PDC

[SPK14] Paul J. Shin, Johnny Park, and Avinash C. Kak. A predictive duty cycle adaptation framework using augmented sensing for wireless camera networks. ACM Transactions on Sensor Networks, 10(2):22:1–22:??, January 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Shpungin:2013:IMS

[SS13] Hanan Shpungin and Michael Segal. Improved multicriteria spanners for ad-hoc networks under energy and distance metrics. ACM Transactions on Sensor Networks, 9(4):37:1–37:??, July 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Schmid:2010:ICP

[SSC^{+}10] Thomas Schmid, Roy Shea, Zainul Charbiwala, Jonathan Friedman, Mani B. Srivastava, and Young H. Cho. On the interaction of clocks, power, and synchronization in duty-cycled embedded sensor nodes. ACM Transactions on Sensor Networks, 7(3):24:1–24:??, September 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Saukh:2010:BRL

[SSGM10] Olga Saukh, Robert Sauter, Matthias Gauger, and Pedro José Marrón. On boundary recognition without location information in wireless sensor networks. ACM Transactions on Sensor Networks, 6(3):20:1– 20:??, June 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Saifullah:2019:CEW

[SSL^{+}19] Abusayeed Saifullah, Sriram Sankar, Jie Liu, Chenyang Lu, Ranveer Chandra, and Bodhi Priyantha. CapNet: Exploiting wireless sensor networks for data center power capping. ACM Transactions on Sensor Networks, 15(1):6:1–6:??, February 2019. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic).  URL https://dl.
acm.org/ft_gateway.cfm?id=
3278624.

Song:2022:CEL

[SSL^{+}22] Yihang Song, Chao Song, Li Lu, Shen Yang, Songfan Li, Chong Zhang, Qianhe Meng, Xiandong Shao, and Haili Wang. Chipnet: Enabling large-scale backscatter network with processor-free devices. ACM Transactions on Sensor Networks, 18(4):61:1– 61:??, November 2022. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3544492.

Shrivastava:2008:DCS

[SST08] Nisheeth Shrivastava, Subhash Suri, and Csaba D. Tóth. Detecting cuts in sensor networks. ACM Transactions on Sensor Networks, 4(2):10:1–10:??, March 2008. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Shirmohammadi:2012:SLS

[ST12] Babak Shirmohammadi and Camillo J. Taylor. Self-localizing smart camera networks. ACM Transactions on Sensor Networks, 8(2):11:1–11:??, March 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Su:2007:CAA

[Su07] Xun Su. A combinatorial algorithmic approach to energy efficient information collection in wireless sensor networks. ACM Transactions on Sensor Networks, 3(1):??, March 2007. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Sardar:2023:SFR

[SUR^{+}23] Alamgir Sardar, Saiyed Umer, Ranjeet Kr. Rout, Shui-Hua Wang, and M. Tanveer. A secure face recognition for IoTenabled healthcare system. ACM Transactions on Sensor Networks, 19(3):52:1–52:??, August 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3534122.

Silvestri:2019:FIS

[SUZK19] Simone Silvestri, Rahul Urgaonkar, Murtaza Zafer, and Bong Jun Ko. A framework for the inference of sensing measurements based on correlation. ACM Transactions on Sensor Networks, 15(1):4:1– 4:??, February 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3272035.

Schroder:2022:IPB

[SW22] Yannic Schröder and Lars Wolf. InPhase: Phase-based ranging and localization. ACM Transactions on Sensor Networks, 18 (2):24:1–24:39, May 2022. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3494542.

Song:2024:TAC

[SWH^{+}24] Jinke Song, Shangfeng Wan, Min Huang, Jiqiang Liu, Limin Sun, and Qiang Li. Toward automatically connecting IoT devices with vulnerabilities in the wild. ACM Transactions on Sensor Networks, 20(1): 6:1–6:??, January 2024. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3608951.

Sun:2021:IDC

[SWYW21] Danfeng Sun, Jia Wu, Jian Yang, and Huifeng Wu. Intelligent data collaboration in heterogeneous-device IoT platforms. ACM Transactions on Sensor Networks, 17(3):22:1– 22:17, June 2021. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3427912.

Song:2015:ETP

[SXD^{+}15] Wen-Zhan Song, Mingsen Xu, Debraj De, Deukhyoun Heo, Jong-Hoon Kim, and Byeong-Sam Kim. ECPC: Toward preserving downtime data persistence in disruptive wireless sensor networks. ACM Transactions on Sensor Networks, 11(2):24:1– 24:??, February 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Sadek:2009:EEC

[SYL09] Ahmed K. Sadek, Wei Yu, and K. J. Ray Liu. On the energy efficiency of cooperative communications in wireless sensor networks. ACM Transactions on Sensor Networks, 6(1):5:1–5:??, December 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Sun:2022:RAL

[SYL^{+}22] Zehua Sun, Huanqi Yang, Kai Liu, Zhimeng Yin, Zhenjiang Li, and Weitao Xu. Recent advances in LoRa: a comprehensive survey. ACM Transactions on Sensor Networks, 18 (4):67:1–67:??, November 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3543856.

Shuai:2012:TMP

[SYOY12] Zaihong Shuai, Sangseok Yoon, Songhwai Oh, and Ming-Hsuan Yang. Traffic modeling and prediction using sensor networks: Who will go where and when? ACM Transactions on Sensor Networks, 9(1):6:1–6:??, November 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Song:2022:DMT

[SYT22] Qun Song, Zhenyu Yan, and Rui Tan. DeepMTD: Moving target defense for deep visual sensing against adversarial examples. ACM Transactions on Sensor Networks, 18(1): 5:1–5:32, February 2022. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3469032.

Shang:2023:TCC

[SYX^{+}23] Fei Shang, Panlong Yang, Jie Xiong, Yuanhao Feng, and Xiangyang Li. Tamera: Contactless commodity tracking, material and shopping behavior recognition using COTS RFIDs. ACM Transactions on Sensor Networks, 19(2):43:1–43:??, May 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3563777.

Shaabana:2019:CPH

[SZ19] Ala Shaabana and Rong Zheng. CRONOS: a post-hoc data driven multi-sensor synchronization approach. ACM Transactions on Sensor Networks, 15(3): 26:1–26:??, August 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

                                        URL https://dl.acm.org/ft_
                                        gateway.cfm?id=3309703.

Sarkar:2011:HSG

[SZG11] Rik Sarkar, Xianjin Zhu, and Jie Gao. Hierarchical spatial gossip for multiresolution representations in sensor networks. ACM Transactions on Sensor Networks, 8(1):4:1–4:??, August 2011. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Sarkar:2013:DCR

[SZG13] Rik Sarkar, Xianjin Zhu, and Jie Gao. Distributed and compact routing using spatial distributions in wireless sensor networks. ACM Transactions on Sensor Networks, 9(3):32:1–32:??, May 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Suresh:2015:TOM

[SZG^{+}15] Mahima Agumbe Suresh, Wei Zhang, Weijiao Gong, Radu Stoleru, Amin Rasekh, and M. Katherine Banks. Toward optimal monitoring of flow-based systems using mobile wireless sensor networks. ACM Transactions on Sensor Networks, 11 (3):48:1–48:??, February 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Shaabana:2017:ICI

[SZX17] Ala Shaabana, Rong Zheng, and Zhipeng Xu. Inferring clothing insulation levels using mechanisms of heat transfer. ACM Transactions on Sensor Networks, 13(4):28:1–28:??, December 2017. CODEN ????

ISSN 1550-4859 (print), 1550-4867 (electronic).

Song:2008:LPP

[SZZC08] Hui Song, Sencun Zhu, Wensheng Zhang, and Guohong Cao. Least privilege and privilege deprivation: Toward tolerating mobile sink compromises in wireless sensor networks. ACM Transactions on Sensor Networks, 4(4):23:1–23:??, August 2008. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Tas:2014:LCI

[TAT14] Baris Tas, Nihat Altiparmak, and Ali Saman Tosun. Low-cost indoor location management for robots using IR leds and an IR camera. ACM Transactions on Sensor Networks, 10(4):63:1– 63:??, June 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Tiwari:2007:EEW

[TBL07] Ankit Tiwari, Prasanna Ballal, and Frank L. Lewis. Energyefficient wireless sensor network design and implementation for condition-based maintenance. ACM Transactions on Sensor Networks, 3(1):??, March 2007. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Tian:2024:VRS

[TBS^{+}24] Siben Tian, Fenhua Bai, Tao Shen, Chi Zhang, and Bei Gong.

VSSB-Raft: a secure and efficient zero trust consensus algorithm for blockchain. ACM Transactions on Sensor Networks, 20(2):34:1–34:??, March 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3611308.

Tovar:2014:CFS

[TCB^{+}14] Benjamin Tovar, Fred Cohen, Leonardo Bobadilla, Justin Czarnowski, and Steven M. Lavalle. Combinatorial filters: Sensor beams, obstacles, and possible paths. ACM Transactions on Sensor Networks, 10(3): 47:1–47:??, April 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Teng:2023:PID

[TCC^{+}23] Fei Teng, Yanjiao Chen, Yushi Cheng, Xiaoyu Ji, Boyang Zhou, and Wenyuan Xu. PDGes: an interpretable detection model for Parkinson’s disease using smartphones. ACM Transactions on Sensor Networks, 19 (4):77:1–77:21, November 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3585314.

Tan:2017:JDC

[TCN^{+}17] Rui Tan, Sheng-Yuan Chiu, Hoang Hai Nguyen, David K. Y. Yau, and Deokwoo Jung. A joint data compression and encryption approach for wireless energy auditing networks. ACM Transactions on Sensor Networks, 13 (2):9:1–9:??, June 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Tiloca:2019:DDS

[TDD^{+}19] Marco Tiloca, Domenico De Guglielmo, Gianluca Dini, Giuseppe Anastasi, and Sajal K. Das. DISH: DIstributed SHuffling against selective jamming attack in IEEE 802.15.4e TSCH networks. ACM Transactions on Sensor Networks, 15(1):3:1– 3:??, February 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

                                        URL https://dl.acm.org/ft_
                                        gateway.cfm?id=3241052.

Tan:2022:BCI

[TDZ^{+}22] Zhaowei Tan, Boyan Ding, Jinghao Zhao, Yunqi Guo, and Songwu Lu. Breaking cellular IoT with forged data-plane signaling: Attacks and countermeasure. ACM Transactions on Sensor Networks, 18 (4):59:1–59:??, November 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3534124.

Tang:2024:FET

[TFL^{+}24] Jianzhi Tang, Luoyi Fu, Shiyu Liang, Fei Long, Lei Zhou, Xinbing Wang, and Chenghu Zhou. FlowerCast: Efficient time-sensitive multicast in wireless sensor networks with link uncertainty. ACM Transactions on Sensor Networks, 20(1):

3:1–3:??, January 2024. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3605551.

Teng:2017:IIO

[TGG^{+}17] Xiaoqiang Teng, Deke Guo, Yulan Guo, Xiaolei Zhou, Zeliu Ding, and Zhong Liu. IONavi: an indoor-outdoor navigation service via mobile crowdsensing. ACM Transactions on Sensor Networks, 13(2):12:1– 12:??, June 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Teng:2019:CTU

[TGG^{+}19] Xiaoqiang Teng, Deke Guo, Yulan Guo, Xiaolei Zhou, and Zhong Liu. CloudNavi: Toward ubiquitous indoor navigation service with 3D point clouds. ACM Transactions on Sensor Networks, 15(1):1:1– 1:??, February 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3216722.

Tan:2014:CPL

[TJLK14] Guang Tan, Hongbo Jiang, Jun Liu, and Anne-Marie Kermarrec. Convex partitioning of largescale sensor networks in complex fields: Algorithms and applications. ACM Transactions on Sensor Networks, 10(3):41:1– 41:??, April 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Tang:2013:EED

[TJWK13] Bin Tang, Neeraj Jaggi, Haijie Wu, and Rohini Kurkal. Energyefficient data redistribution in sensor networks. ACM Transactions on Sensor Networks, 9 (2):11:1–11:??, March 2013. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Tan:2013:CBA

[TJZ^{+}13] Guang Tan, Hongbo Jiang, Shengkai Zhang, Zhimeng Yin, and Anne-Marie Kermarrec. Connectivity-based and anchorfree localization in large-scale 2D/3D sensor networks. ACM Transactions on Sensor Networks, 10(1):6:1–6:??, November 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Taherkordi:2013:OSN

[TLRE13] Amir Taherkordi, Frederic Loiret, Romain Rouvoy, and Frank Eliassen. Optimizing sensor network reprogramming via in situ reconfigurable components. ACM Transactions on Sensor Networks, 9(2):14:1–14:??, March 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Tessens:2014:CST

[TMAP14] Linda Tessens, Marleen Morbee, Hamid Aghajan, and Wilfried Philips. Camera selection for tracking in distributed smart camera networks. ACM Transactions on Sensor Networks, 10(2):23:1–23:??, January 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Tavakoli:2018:DIA

[TNBG18] Rasool Tavakoli, Majid Nabi, Twan Basten, and Kees Goossens. Dependable interference-aware time-slotted channel hopping for wireless sensor networks. ACM Transactions on Sensor Networks, 14(1):3:1–3:??, March 2018. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Tague:2007:CSA

[TP07] Patrick Tague and Radha Poovendran. A canonical seed assignment model for key predistribution in wireless sensor networks. ACM Transactions on Sensor Networks, 3(4):19:1– 19:??, October 2007. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Tan:2017:URP

[TPM^{+}17] Rui Tan, Dennis E. Phillips, Mohammad-Mahdi Moazzami, Guoliang Xing, and Jinzhu Chen. Unsupervised residential power usage monitoring using a wireless sensor network. ACM Transactions on Sensor Networks, 13(3):20:1–20:??, September 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Thai:2014:DTV

[TTBH14] My T. Thai, Ravi Tiwari, Raja Bose, and Abdelsalam Helal. On detection and tracking of variant phenomena clouds. ACM Transactions on Sensor Networks, 10(2):34:1–34:??, January 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Tan:2013:FBV

[TXC^{+}13] Rui Tan, Guoliang Xing, Jinzhu Chen, Wen-Zhan Song, and Renjie Huang. Fusion-based volcanic earthquake detection and timing in wireless sensor networks. ACM Transactions on Sensor Networks, 9(2):17:1– 17:??, March 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Tan:2013:SLC

[TXY^{+}13] Rui Tan, Guoliang Xing, Zhaohui Yuan, Xue Liu, and Jianguo Yao. System-level calibration for data fusion in wireless sensor networks. ACM Transactions on Sensor Networks, 9(3): 28:1–28:??, May 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Trigoni:2007:WSR

[TYD^{+}07] Niki Trigoni, Yong Yao, Alan Demers, Johannes Gehrke, and Rajmohan Rajaraman. Wave scheduling and routing in sensor networks. ACM Transactions on Sensor Networks, 3(1): ??, March 2007. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Tian:2015:SSH

[TYGW15] Jie Tian, Tan Yan, Xin Gao, and Guiling Wang. Scheduling survivability-heterogeneous sensor networks for critical location surveillance. ACM Transactions on Sensor Networks, 11(4):56:1– 56:??, December 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Tan:2022:JOR

[TZZ22] Tiao Tan, Ming Zhao, and Zhiwen Zeng. Joint offloading and resource allocation based on UAV-assisted mobile edge computing. ACM Transactions on Sensor Networks, 18(3): 36:1–36:??, August 2022. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3476512.

Voulkidis:2013:EEW

[VAC13] Artemis C. Voulkidis, Markos P.

Anastasopoulos, and Panayotis G. Cottis. Energy efficiency in wireless sensor networks: a game-theoretic approach based on coalition formation. ACM Transactions on Sensor Networks, 9(4):43:1– 43:??, July 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Voulgaris:2016:DNL

[VDV16] Spyros Voulgaris, Matthew Dobson, and Maarten Van Steen. Decentralized network-level synchronization in mobile ad hoc networks. ACM Transactions on Sensor Networks, 12(1):5:1– 5:??, March 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Venkatasubramanian:2010:PVB

[VG10] Krishna K. Venkatasubramanian and Sandeep K. S. Gupta. Physiological value-based efficient usable security solutions for body sensor networks. ACM Transactions on Sensor Networks, 6(4): 31:1–31:??, July 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Vicaire:2009:ALT

[VHC^{+}09] Pascal Vicaire, Tian He, Qing Cao, Ting Yan, Gang Zhou, Lin Gu, Liqian Luo, Radu Stoleru, John A. Stankovic, and Tarek F. Abdelzaher. Achieving long-term surveillance in VigilNet. ACM Transactions on Sensor Networks, 5(1):9:1–9:??, February 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Vedantam:2010:ADE

[VMS10] Satish Vedantam, Urbashi Mitra, and Ashutosh Sabharwal. Asymptotic distortion exponents for the estimation of timevarying channels in multihop sensor networks. ACM Transactions on Sensor Networks, 6(4): 33:1–33:??, July 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Verma:2020:QPF

[VPB^{+}20] Rahul Kumar Verma, K. K. Pattanaik, Sourabh Bharti, Divya Saxena, and Jiannong Cao. A query processing framework for efficient network resource utilization in shared sensor networks. ACM Transactions on Sensor Networks, 16(4):31:1– 31:28, October 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3397809.

Viswanatha:2015:EER

[VRSR15] Kumar Viswanatha, Sharadh Ramaswamy, Ankur Saxena, and Kenneth Rose. Error/ erasure-resilient and complexityconstrained zero-delay distributed coding for large-scale sensor networks. ACM Transactions on Sensor Networks, 11(2):35:1– 35:??, February 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Viswanathan:2018:EEG

[VTY18] Sreejaya Viswanathan, Rui Tan, and David K. Y. Yau. Exploiting electrical grid for accurate and secure clock synchronization. ACM Transactions on Sensor Networks, 14(2):12:1– 12:??, July 2018. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wang:2017:SNP

[WB17] Changda Wang and Elisa Bertino. Sensor network provenance compression using dynamic Bayesian networks. ACM Transactions on Sensor Networks, 13(1):5:1–5:??, February 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wang:2010:DEE

[WBS10] Zijian Wang, Eyuphan Bulut, and Boleslaw K. Szymanski. Distributed energy-efficient target tracking with binary sensor networks. ACM Transactions on Sensor Networks, 6(4):32:1– 32:??, July 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wu:2014:DPF

[WBS14] Xiuchao Wu, Kenneth N. Brown, and Cormac J. Sreenan. Data pre-forwarding for opportunistic data collection in wireless sensor networks. ACM Transactions on Sensor Networks, 11(1): 8:1–8:??, August 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wettergren:2009:OPD

[WC09] Thomas A. Wettergren and Russell Costa. Optimal placement of distributed sensors against moving targets. ACM Transactions on Sensor Networks, 5(3):26:1– 26:??, May 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wettergren:2012:OMP

[WC12] Thomas A. Wettergren and Russell Costa. Optimal multiobjective placement of distributed sensors against moving targets. ACM Transactions on Sensor Networks, 8(3):21:1–21:??, July 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wang:2013:AFV

[WC13] Yi Wang and Guohong Cao.

Achieving full-view coverage in camera sensor networks. ACM Transactions on Sensor Networks, 10(1):3:1–3:??, November 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wang:2023:ESS

[WCLD23] Ju Wang, Xi Chen, Xue Liu, and Gregory Dudek. Eliminating space scanning: Fast mmWave beam alignment with UWB radios. ACM Transactions on Sensor Networks, 19 (4):79:1–79:20, November 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3588438.

Winkler:2020:OOI

[WCPC20] Daniel A. Winkler, Miguel Á.

Carreira-Perpiñán, and Alberto E. Cerpa. OPTICS: OP-Timizing Irrigation Control at Scale. ACM Transactions on Sensor Networks, 16(3):22:1– 22:38, August 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3372024.

Wei:2018:SSA

[WCV^{+}18] Peter Wei, Xiaoqi Chen, Jordan Vega, Stephen Xia, Rishikanth Chandrasekaran, and Xiaofan Jiang. A scalable system for apportionment and tracking of energy footprints in commercial buildings. ACM Transactions on Sensor Networks, 14(3–4):22:1– 22:??, December 2018. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wang:2023:FDD

[WCW^{+}23] Guang Wang, Yuefei Chen, Shuai Wang, Fan Zhang, and Desheng Zhang. ForETaxi: Data-driven fleet-oriented charging resource allocation in largescale electric taxi networks. ACM Transactions on Sensor Networks, 19(3):63:1–63:??, August 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3570958.

Wang:2009:SST

[WDLN09] Ronghua Wang, Wenliang Du, Xiaogang Liu, and Peng Ning. ShortPK: a short-term public key scheme for broadcast authentication in sensor networks. ACM Transactions on Sensor Networks, 6(1):9:1–9:??, December 2009. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wan:2011:EEC

[WEC11] Chieh-Yih Wan, Shane B. Eisenman, and Andrew T. Campbell. Energy-efficient congestion detection and avoidance in sensor networks. ACM Transactions on Sensor Networks, 7(4):32:1– 32:??, February 2011. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wan:2007:OTM

[WECC07] Chieh-Yih Wan, Shane B. Eisenman, Andrew T. Campbell, and Jon Crowcroft. Overload traffic management for sensor networks. ACM Transactions on Sensor Networks, 3(4):18:1– 18:??, October 2007. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wu:2023:TDF

[WHQ^{+}23] Kaishun Wu, Yandao Huang, Minghui Qiu, Zhenkan Peng, and Lu Wang. Toward devicefree and user-independent fall detection using floor vibration. ACM Transactions on Sensor Networks, 19(1):5:1– 5:20, February 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3519302.

Wu:2016:RFM

[WHST16] Fang-Jing Wu, Hsiu-Chi Hsu, Chien-Chung Shen, and Yu-Chee Tseng. Range-free mobile actor relocation in a twotiered wireless sensor and actor network. ACM Transactions on Sensor Networks, 12(2):15:1– 15:??, May 2016. CODEN ????

ISSN 1550-4859 (print), 1550-4867 (electronic).

Wei:2019:RCE

[WHYC19] Bo Wei, Wen Hu, Mingrui Yang, and Chun Tung Chou. From real to complex: Enhancing radiobased activity recognition using complex-valued CSI. ACM Transactions on Sensor Networks, 15(3):35:1–35:??, August 2019. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic).  URL https://dl.
                                        acm.org/ft_gateway.cfm?id=
                                        3338026.

Wang:2011:OSM

[WIF^{+}11] Guiling Wang, Mary Jane Irwin, Haoying Fu, Piotr Berman, Wensheng Zhang, and Tom La Porta. Optimizing sensor movement planning for energy efficiency. ACM Transactions on Sensor Networks, 7(4):33:1– 33:??, February 2011. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wei:2021:DDS

[WJ21] Peter Wei and Xiaofan Jiang. A data-driven system for citywide energy footprinting and apportionment. ACM Transactions on Sensor Networks, 17 (2):11:1–11:24, June 2021. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3433639.

Wang:2016:CBS

[WJD16] Chen Wang, Hongbo Jiang, and Yan Dong. Connectivity-based space filling curve construction algorithms in high genus 3D surface WSNs. ACM Transactions on Sensor Networks, 12(3):22:1– 22:??, August 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wang:2024:HBT

[WJY^{+}24] Pengfei Wang, Dian Jiao, Leyou Yang, Bin Wang, and Ruiyun Yu. Hypergraph-based truth discovery for sparse data in mobile crowdsensing. ACM Transactions on Sensor Networks, 20 (3):69:1–69:??, May 2024. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3649894.

Weiss:2021:DBS

[WJZ21] Wolfgang Weiss, Víctor J. Expósito Jiménez, and Herwig Zeiner. Dynamic buffer sizing for outof-order event compensation for time-sensitive applications. ACM Transactions on Sensor Networks, 17(1):1:1–1:23, January 2021. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3410403.

Wang:2014:MLA

[WKA14] Dong Wang, Lance Kaplan, and Tarek F. Abdelzaher. Maximum likelihood analysis of conflicting observations in social sensing. ACM Transactions on Sensor Networks, 10(2):30:1–30:??, January 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wang:2017:EWN

[WKYH17] Shuai Wang, Song Min Kim, Zhimeng Yin, and Tian He. Encode when necessary: Correlated network coding under unreliable wireless links. ACM Transactions on Sensor Networks, 13(1): 7:1–7:??, February 2017. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wan:2014:DDA

[WL14] Jiuqing Wan and Li Liu. Distributed data association in smart camera networks using belief propagation. ACM Transactions on Sensor Networks, 10(2): 19:1–19:??, January 2014. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wang:2023:GZS

[WL23] Wei Wang and Qingzhong Li. Generalized zero-shot activity recognition with embeddingbased method. ACM Transactions on Sensor Networks, 19(3):72:1–72:25, August 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3582690.

Wang:2010:EED

[WLD10] Jing Wang, Yonghe Liu, and Sajal K. Das. Energy-efficient data gathering in wireless sensor networks with asynchronous sampling. ACM Transactions on Sensor Networks, 6(3):22:1– 22:??, June 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wang:2024:RTC

[WLLZ24] Di Wang, Fangyu Li, Kaibo Liu, and Xi Zhang. Realtime cyber-physical security solution leveraging an integrated learning-based approach. ACM Transactions on Sensor Networks, 20(2):27:1–27:??, March 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3582009.

Wu:2016:EMC

[WLS^{+}16] Yafeng Wu, Kin Sum Liu, John A. Stankovic, Tian He, and Shan Lin. Efficient multichannel communications in wireless sensor networks. ACM Transactions on Sensor Networks, 12(1): 3:1–3:??, March 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wu:2012:SSM

[WLW12] Xiaopei Wu, Mingyan Liu, and Yue Wu. In-situ soil moisture sensing: Optimal sensor placement and field estimation. ACM Transactions on Sensor Networks, 8(4):33:1–33:??, September 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wang:2020:TEM

[WLW^{+}20] Yanyan Wang, Jia Liu, Xia Wang, Xingyu Chen, Yingli Yan, and Lijun Chen. Time-efficient missing tag identification in an open RFID system. ACM Transactions on Sensor Networks, 16(3):21:1–21:27, August 2020. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3386242.

Wang:2023:JUS

[WLW^{+}23] Xindi Wang, Xinyu Liu, Jianjian Wu, Wei Ju, Xiaojing Chen, and Ling Shen. Joint user scheduling, power configuration and trajectory planning strategy for UAVaided WSNs. ACM Transactions on Sensor Networks, 19 (1):10:1–10:27, February 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3529508.

Wu:2023:SDR

[WLX^{+}23] Yue Wu, Fan Li, Yadong Xie, Yu Wang, and Zheng Yang. SymListener: Detecting respiratory symptoms via acoustic sensing in driving environments. ACM Transactions on Sensor Networks, 19(1):3:1– 3:21, February 2023. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3517014.

Wang:2013:MSA

[WLZ13] Dan Wang, Jiangchuan Liu, and Qian Zhang. On mobile sensor assisted field coverage. ACM Transactions on Sensor Networks, 9(2):22:1–22:??, March 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wang:2023:TPP

[WLZ23] Jian Wang, Jiaxin Liu, and Guosheng Zhao. Two-phased participant selection method based on partial transfer learning in mobile crowdsensing. ACM Transactions on Sensor Networks, 19(2):42:1–42:??, May 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3563776.

Wu:2019:EIL

[WMT^{+}19] Hang Wu, Ziliang Mo, Jiajie Tan, Suining He, and S.- H. Gary Chan. Efficient indoor localization based on geomagnetism. ACM Transactions on Sensor Networks, 15(4): 42:1–42:??, October 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3342517.

Wu:2024:NIH

[WNM^{+}24] Yingxiao Wu, Haocheng Ni, Changlin Mao, Jianping Han, and Wenyao Xu. Non-intrusive human vital sign detection using mmWave sensing technologies: a review. ACM Transactions on Sensor Networks, 20 (1):16:1–16:??, January 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3627161.

Wang:2016:FTM

[WPL^{+}16] Tian Wang, Zhen Peng, Junbin Liang, Sheng Wen, Md Zakirul Alam Bhuiyan, Yiqiao Cai, and Jiannong Cao. Following targets for mobile tracking in wireless sensor networks. ACM Transactions on Sensor Networks, 12(4):31:1–31:??, November 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wang:2022:EPO

[WQH^{+}22] Lu Wang, Xiaoke Qi, Ruifeng Huang, Kaishun Wu, and Qian Zhang. Exploring partially overlapping channels for low-power wide area networks. ACM Transactions on Sensor Networks, 18 (4):63:1–63:??, November 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3546075.

Wang:2010:MLL

[WRS10] Chao Wang, Parameswaran Ramanathan, and Kewal K. Saluja. Modeling latency — lifetime trade-off for target detection in mobile sensor networks. ACM Transactions on Sensor Networks, 7(1):8:1–8:??, August 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wang:2011:DSS

[WRYL11] Qian Wang, Kui Ren, Shucheng Yu, and Wenjing Lou. Dependable and secure sensor data storage with dynamic integrity assurance. ACM Transactions on Sensor Networks, 8(1):9:1–9:??, August 2011. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Won:2014:LSG

[WS14] Myounggyu Won and Radu Stoleru. A low-stretch-guaranteed and lightweight geographic routing protocol for large-scale wireless sensor networks. ACM Transactions on Sensor Networks, 11(1):18:1–18:??, August 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wan:2023:MUR

[WSC^{+}23] Haoran Wan, Shuyu Shi, Wenyu Cao, Wei Wang, and Guihai Chen. Multi-user room-scale respiration tracking using COTS acoustic devices. ACM Transactions on Sensor Networks, 19 (4):85:1–85:28, November 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3594220.

Wu:2022:PFG

[WTC22] Hang Wu, Jiajie Tan, and S.- H. Gary Chan. Pedometer-free geomagnetic fingerprinting with casual walking speed. ACM Transactions on Sensor Networks, 18(1):8:1–8:21, February 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3470850.

Wu:2023:PFI

[WTH^{+}23] Jimmy Ming-Tai Wu, Qian Teng, Shamsul Huda, Yeh-Cheng Chen, and Chien-Ming Chen. A privacy frequent itemsets mining framework for collaboration in IoT using federated learning. ACM Transactions on Sensor Networks, 19(2): 27:1–27:15, May 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3532090.

Wang:2016:EEA

[WTX^{+}16] Yu Wang, Rui Tan, Guoliang Xing, Jianxun Wang, Xiaobo Tan, and Xiaoming Liu. Energyefficient aquatic environment monitoring using Smartphonebased robots. ACM Transactions on Sensor Networks, 12(3):25:1– 25:??, August 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wang:2023:TMW

[WTX^{+}23] Zuyan Wang, Jun Tao, Yifan Xu, Yang Gao, and Dikai Zou. Toward the minimal wait-for delay for rechargeable WSNs with multiple mobile chargers. ACM Transactions on Sensor Networks, 19(4):78:1– 78:24, November 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3579093.

Winkler:2019:DDI

[WWB^{+}19] Daniel A. Winkler, Robert Wang, François Blanchette, Miguel Á. Carreira-Perpiñán, and Alberto E. Cerpa. DIC-TUM: Distributed Irrigation aCtuation with Turf hUmidity Modeling. ACM Transactions on Sensor Networks, 15(4): 41:1–41:??, October 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3342514.

Wang:2011:MMR

[WWFX11] Xiaorui Wang, Xiaodong Wang, Xing Fu, and Guoliang Xing. MCRT: Multichannel real-time communications in wireless sensor networks. ACM Transactions on Sensor Networks, 8(1): 2:1–2:??, August 2011. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wu:2015:SSM

[WWL15] Xiaopei Wu, Qingsi Wang, and Mingyan Liu. In-situ soil moisture sensing: Measurement scheduling and estimation using sparse sampling. ACM Transactions on Sensor Networks, 11(2): 26:1–26:??, February 2015. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wang:2016:BTD

[WWL^{+}16] Chen Wang, Wei Wei, Hongzhi Lin, Hongbo Jiang, and John C. S. Lui. BLOW–UP: Toward distributed and scalable space filling curve construction in 3D volumetric WSNs. ACM Transactions on Sensor Networks, 12 (4):30:1–30:??, November 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wang:2013:DDD

[WWLX13] Xiaodong Wang, Xiaorui Wang, Liu Liu, and Guoliang Xing. DutyCon: a dynamic duty-cycle control approach to end-to-end delay guarantees in wireless sensor networks. ACM Transactions on Sensor Networks, 9(4): 42:1–42:??, July 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wang:2013:MTP

[WWXY13] Xiaodong Wang, Xiaorui Wang, Guoliang Xing, and Yanjun Yao. Minimum transmission power configuration in real-time sensor networks with overlapping channels. ACM Transactions on Sensor Networks, 9(2):10:1– 10:??, March 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wang:2021:PPD

[WWZ^{+}21] Jing Wang, Libing Wu, Sherali Zeadally, Muhammad Khurram Khan, and Debiao He. Privacy-preserving data aggregation against malicious data mining attack for IoT-enabled smart grid. ACM Transactions on Sensor Networks, 17 (3):25:1–25:25, June 2021. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3440249.

Wang:2024:TIC

[WWZ24] Weizheng Wang, Qing Wang, and Marco Zuniga. Taming irregular cardiac signals for biometric identification. ACM Transactions on Sensor Networks, 20(1):25:1–25:??, January 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3624570.

Wang:2008:SLC

[WX08] Chen Wang and Li Xiao. Sensor localization in concave environments. ACM Transactions on Sensor Networks, 4(1):3:1– 3:??, January 2008. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wu:2023:OCC

[WXD^{+}23] Sixu Wu, Lijie Xu, Haipeng Dai, Linfeng Liu, Fu Xiao, and Jia Xu. Optimizing comprehensive cost of charger deployment in multi-hop wireless charging. ACM Transactions on Sensor Networks, 19(4):83:1– 83:24, November 2023. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3584950.

Wei:2024:SBF

[WXG^{+}24] Bo Wei, Weitao Xu, Mingcen Gao, Guohao Lan, Kai Li, Chengwen Luo, and Jin Zhang. SolarKey: Battery-free key generation using solar cells. ACM Transactions on Sensor Networks, 20(1):7:1–7:??, January 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3605780.

Wang:2019:EEC

[WXL^{+}19] Wei Wang, Tiantian Xie, Xin Liu, Yao Yao, and Ting Zhu. ECT: Exploiting crosstechnology transmission for reducing packet delivery delay in IoT networks. ACM Transactions on Sensor Networks, 15 (2):20:1–20:??, April 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3293536.

Wang:2024:SSE

[WYC^{+}24] Shanyue Wang, Yubo Yan, Yujie Chen, Panlong Yang, and Xiang-Yang Li. Spray: a spectrumefficient and agile concurrent backscatter system. ACM Transactions on Sensor Networks, 20 (2):42:1–42:??, March 2024. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3638051.

Wu:2022:OCO

[WYD^{+}22] Tao Wu, Panlong Yang, Haipeng Dai, Chaocan Xiang, and Wanru Xu. Optimal charging oriented sensor placement and flexible scheduling in rechargeable WSNs. ACM Transactions on Sensor Networks, 18(3):50:1– 50:??, August 2022. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3512888.

Wang:2019:CMC

[WYY^{+}19] Liang Wang, Zhiwen Yu, Dingqi Yang, Tao Ku, Bin Guo, and Huadong Ma. Collaborative mobile crowdsensing in opportunistic D2D networks: a graphbased approach. ACM Transactions on Sensor Networks, 15(3): 30:1–30:??, August 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3317689.

Wang:2007:SPP

[WZL07] Dan Wang, Qian Zhang, and Jiangchuan Liu. The selfprotection problem in wireless sensor networks. ACM Transactions on Sensor Networks, 3(4): 20:1–20:??, October 2007. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wang:2008:PNC

[WZL08] Dan Wang, Qian Zhang, and Jiangchuan Liu. Partial network coding: Concept, performance, and application for continuous data collection in sensor networks. ACM Transactions on Sensor Networks, 4(3): 14:1–14:??, May 2008. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Wang:2021:CQC

[WZLM21] Yuting Wang, Xiaolong Zheng, Liang Liu, and Huadong Ma. CoHop: Quantitative correlationbased channel hopping for lowpower wireless networks. ACM Transactions on Sensor Networks, 17(2):15:1–15:29, June 2021. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3440248.

Wang:2021:SEM

[WZZ^{+}21] Beilun Wang, Jiaqi Zhang, Yan Zhang, Meng Wang, and Sen Wang. Scalable estimator for multi-task Gaussian graphical models based in an IoT network. ACM Transactions on Sensor Networks, 17(3):23:1– 23:33, June 2021. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3432312.

Wang:2023:DLC

[WZZ^{+}23] Yuting Wang, Fanhao Zhang, Xiaolong Zheng, Liang Liu, and Huadong Ma. Decoding LoRa collisions via parallel alignment. ACM Transactions on Sensor Networks, 19(3):62:1– 62:??, August 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3571586.

Xu:2015:HDA

[XAKV15] Xi Xu, Rashid Ansari, Ashfaq Khokhar, and Athanasios V.

Vasilakos. Hierarchical data aggregation using compressive sensing (HDACS) in WSNs. ACM Transactions on Sensor Networks, 11(3):45:1–45:??, February 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Xiao:2013:RLA

[XBWX13] Qingjun Xiao, Kai Bu, Zhijun Wang, and Bin Xiao. Robust localization against outliers in wireless sensor networks. ACM Transactions on Sensor Networks, 9(2):24:1–24:??, March 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Xu:2015:OEE

[XCC^{+}15] Lijie Xu, Guihai Chen, Jiannong Cao, Shan Lin, Haipeng Dai, Xiaobing Wu, and Fan Wu. Optimizing energy efficiency for minimum latency broadcast in low-duty-cycle sensor networks. ACM Transactions on Sensor Networks, 11(4):57:1–57:??, December 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Xie:2016:LLI

[XCT^{+}16] Bo Xie, Kongyang Chen, Guang Tan, Mingming Lu, Yunhuai Liu, Jie Wu, and Tian He. LIPS: a light intensity-based positioning system for indoor environments. ACM Transactions on Sensor Networks, 12(4):28:1– 28:??, November 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Xu:2024:BEE

[XDL^{+}24] Yifan Xu, Fan Dang, Kebin Liu, Zhui Zhu, Xinlei Chen, Xu Wang, Xin Miao, and Haitian Zhao. BEANet: an energy-efficient BLE solution for high-capacity equipment area network. ACM Transactions on Sensor Networks, 20(3):52:1– 52:??, May 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3641280.

Xu:2021:SBI

[XDM^{+}21] Jingao Xu, Erqun Dong, Qiang Ma, Chenshu Wu, and Zheng Yang. Smartphone-based indoor visual navigation with leaderfollower mode. ACM Transactions on Sensor Networks, 17 (2):18:1–18:22, June 2021. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3448417.

Xia:2014:MMU

[XDX^{+}14] Ming Xia, Yabo Dong, Wenyuan Xu, Xiangyang Li, and Dongming Lu. MC 2: Multimode user-centric design of wireless sensor networks for long-term monitoring. ACM Transactions on Sensor Networks, 10(3):52:1– 52:??, April 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Xu:2021:ECC

[XFZ^{+}21] Xiaolong Xu, Zijie Fang, Jie Zhang, Qiang He, Dongxiao Yu, Lianyong Qi, and Wanchun Dou. Edge content caching with deep spatiotemporal residual network for IoV in smart city. ACM Transactions on Sensor Networks, 17(3):29:1–29:33, June 2021. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3447032.

Xia:2022:PSL

[XHZG22] Xianjin Xia, Ningning Hou, Yuanqing Zheng, and Tao Gu. PCube: Scaling LoRa concurrent transmissions with reception diversities. ACM Transactions on Sensor Networks, 18 (4):66:1–66:??, November 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3545571.

Xia:2023:PAR

[XJL^{+}23] Ming Xia, Jiaquan Jin, Biqian Liu, Yu Hen Hu, Xiaoyan Wang, and Kaikai Chi. Physicalassisted routing for proactive avoidance of nomadic obstacles in IoT. ACM Transactions on Sensor Networks, 19 (2):45:1–45:??, May 2023. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3565021.

Xu:2017:GKG

[XJR^{+}17] Weitao Xu, Chitra Javali, Girish Revadigar, Chengwen Luo, Neil Bergmann, and Wen Hu. Gait-Key: a gait-based shared secret key generation protocol for wearable devices. ACM Transactions on Sensor Networks, 13(1):6:1– 6:??, February 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Xu:2022:ACC

[XKW^{+}22] Ran Xu, Rakesh Kumar, Pengcheng Wang, Peter Bai, Ganga Meghanath, Somali Chaterji, Subrata Mitra, and Saurabh Bagchi. ApproxNet: Content and contention-aware video object classification system for embedded clients. ACM Transactions on Sensor Networks, 18 (1):11:1–11:27, February 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3463530.

Xiang:2022:EEG

[XLG^{+}22] Tao Xiang, Hangcheng Liu, Shangwei Guo, Yan Gan, and Xiaofeng Liao. EGM: an efficient generative model for unrestricted adversarial examples. ACM Transactions on Sensor Networks, 18(4):51:1– 51:??, November 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3511893.

Xie:2023:NEM

[XLO^{+}23] Ying Xie, Xiaohui Liu, Mohammad S. Obaidat, Xiong Li, and Pandi Vijayakumar. Nondeterministic evaluation mechanism for user recruitment in mobile crowd-sensing. ACM Transactions on Sensor Networks, 19 (2):34:1–34:??, May 2023. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3546951.

Xing:2007:MPC

[XLZ^{+}07] Guoliang Xing, Chenyang Lu, Ying Zhang, Qingfeng Huang, and Robert Pless. Minimum power configuration for wireless communication in sensor networks. ACM Transactions on Sensor Networks, 3(2):11:1– 11:??, June 2007. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Xu:2024:FTM

[XQL^{+}24] Zichuan Xu, Haiyang Qiao, Weifa Liang, Zhou Xu, Qiufen Xia, Pan Zhou, Omer F. Rana, and Wenzheng Xu. Flow-time minimization for timely data stream processing in UAV-aided mobile edge computing. ACM Transactions on Sensor Networks, 20(3):58:1–58:??, May 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3643813.

Xu:2013:RTR

[XRH^{+}13] Yinsheng Xu, Fengyuan Ren, Tao He, Chuang Lin, Canfeng Chen, and Sajal K. Das. Realtime routing in wireless sensor networks: a potential field approach. ACM Transactions on Sensor Networks, 9(3):35:1– 35:??, May 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Xu:2010:CGM

[XRS10] Xiaochun Xu, Nageswara S. V. Rao, and Sartaj Sahni. A computational geometry method for localization using differences of distances. ACM Transactions on Sensor Networks, 6(2):10:1– 10:??, February 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Xu:2022:DDT

[XTXW22] Zhenqiang Xu, Shuai Tong, Pengjin Xie, and Jiliang Wang. From demodulation to decoding: Toward complete LoRa PHY understanding and implementation. ACM Transactions on Sensor Networks, 18(4):64:1– 64:??, November 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3546869.

Xu:2008:DWS

[XTZ08] Wenyuan Xu, Wade Trappe, and Yanyong Zhang. Defending wireless sensor networks from radio interference through channel adaptation. ACM Transactions on Sensor Networks, 4(4):18:1– 18:??, August 2008. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Xia:2023:IFC

[XWC^{+}23] Na Xia, Yin Wang, Bin Chen, Huazheng Du, Chaonong Xu, and Rong Zheng. IMF^{2}O^{2}: a fully connected sensor deployment algorithm for underwater sensor networks. ACM Transactions on Sensor Networks, 19(3): 67:1–67:??, August 2023. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3577201.

Xiong:2012:CBP

[XWDN12] Kaiqi Xiong, Ronghua Wang, Wenliang Du, and Peng Ning. Containing bogus packet insertion attacks for broadcast authentication in sensor networks. ACM Transactions on Sensor Networks, 8(3):20:1–20:??, July 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Xu:2024:RCP

[XWL24] Yaming Xu, Yan Wang, and Boliang Li. Robust classification and 6D pose estimation by sensor dual fusion of image and point cloud data. ACM Transactions on Sensor Networks, 20(2):46:1–46:??, March 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3639705.

Xing:2020:DRL

[XWW^{+}20] Tianzhang Xing, Qing Wang, Chase Q. Wu, Wei Xi, and Xiaojiang Chen. dWatch: a reliable and low-power drowsiness detection system for drivers based on mobile devices. ACM Transactions on Sensor Networks, 16 (4):37:1–37:22, October 2020. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3407899.

Xia:2023:HSD

[XWW^{+}23] Na Xia, Yin Wang, Qiong Wu, Chenguang Yuan, Xinyi Wen, Yue Wu, and Longya Lang. The hunting-style deployment of underwater sensor networks. ACM Transactions on Sensor Networks, 19 (4):96:1–96:22, November 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3604556.

Xing:2005:ICC

[XWZ^{+}05] Guoliang Xing, Xiaorui Wang, Yuanfang Zhang, Chenyang Lu, Robert Pless, and Christopher Gill. Integrated coverage and connectivity configuration for energy conservation in sensor networks. ACM Transactions on Sensor Networks, 1(1):36–72, August 2005. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Xu:2016:EET

[XXHL16] Miao Xu, Wenyuan Xu, Tingrui Han, and Zhiyun Lin. Energyefficient time synchronization in wireless sensor networks via temperature-aware compensation. ACM Transactions on Sensor Networks, 12(2):12:1– 12:??, May 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Xia:2024:AFG

[XXW^{+}24] Shuangqing Xia, Tianzhang Xing, Chase Q. Wu, Guoqing Liu, Jiadi Yang, and Kang Li. AQMon: a fine-grained air quality monitoring system based on UAV images for smart cities. ACM Transactions on Sensor Networks, 20(2):43:1– 43:??, March 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3638766.

Xing:2023:WRT

[XYJ^{+}23] Tianzhang Xing, Qing Yang, Zhiping Jiang, Xinhua Fu, Junfeng Wang, Chase Q. Wu, and Xiaojiang Chen. WiFine: Realtime gesture recognition using Wi-Fi with edge intelligence. ACM Transactions on Sensor Networks, 19(1):11:1– 11:24, February 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3532094.

Xie:2022:GST

[XYW^{+}22] Lei Xie, Peicheng Yang, Chuyu Wang, Tao Gu, Gaolei Duan, Xinran Lu, and Sanglu Lu. Gait-Tracker: 3D skeletal tracking for gait analysis based on inertial measurement units. ACM Transactions on Sensor Networks, 18 (2):27:1–27:27, May 2022. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3502722.

Xu:2020:QAV

[XZL^{+}20] Zichuan Xu, Zhiheng Zhang, Weifa Liang, Qiufen Xia, Omer Rana, and Guowei Wu. QoSaware VNF placement and service chaining for IoT applications in multi-tier mobile edge networks. ACM Transactions on Sensor Networks, 16(3):23:1– 23:27, August 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3387705.

Yang:2024:BEE

[YA24] Xin Yang and Omid Ardakanian. Blinder: End-to-end privacy protection in sensing systems via personalized federated learning. ACM Transactions on Sensor Networks, 20(1):15:1– 15:??, January 2024. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3623397.

Yang:2022:AAP

[Yan22] Weizhong Yang. Adversarial attack protection scalar multiplication for WSNs resistance machine-learning side-channel attack. ACM Transactions on Sensor Networks, 18(3):38:1– 38:??, August 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3486679.

Yu:2017:EEC

[YB17] Xiaohan Yu and Seung Jun Baek. Energy-efficient collection of sparse data in wireless sensor networks using sparse random matrices. ACM Transactions on Sensor Networks, 13 (3):22:1–22:??, September 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Yang:2024:ASC

[YBY^{+}24] Yinghong Yang, Fenhua Bai, Zhuo Yu, Tao Shen, Yingli Liu, and Bei Gong. An anonymous and supervisory cross-chain privacy protection protocol for zero-trust IoT application. ACM Transactions on Sensor Networks, 20(2):32:1–32:??, March 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3583073.

Yu:2019:PIM

[YCL^{+}19] Ruiyun Yu, Jiannong Cao, Rui Liu, Wenyu Gao, Xingwei Wang, and Junbin Liang. Participant incentive mechanism toward quality-oriented sensing: Understanding and application. ACM Transactions on Sensor Networks, 15(2):21:1–21:??, April 2019. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic).  URL https://dl.
acm.org/ft_gateway.cfm?id=
3303703.

Yang:2013:ASS

[YH13] Ou Yang and Wendi Heinzelman. An adaptive sensor sleeping solution based on sleeping multipath routing and dutycycled MAC protocols. ACM Transactions on Sensor Networks, 10(1):10:1–10:??, November 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Yu:2024:TOD

[YHC^{+}24] Yuning Yu, Shanglin Hsu, Andre Chen, Yutian Chen, and Bin Tang. Truthful and optimal data preservation in base station-less sensor networks: an integrated game theory and network flow approach. ACM Transactions on Sensor Networks, 20(1):5:1–5:??, January 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3606263.

Yau:2022:NIC

[YJL^{+}22] Cheuk-Wang Yau, Sukanya Jewsakul, Man-Ho Luk, Angela P. Y. Lee, Yun-Hin Chan, Edith C. H. Ngai, Philip W. T. Pong, King-Shan Lui, and Jiangchuan Liu. NB-IoT coverage and sensor node connectivity in dense urban environments: an empirical study. ACM Transactions on Sensor Networks, 18(3): 49:1–49:??, August 2022. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3536424.

Yang:2013:BTI

[YJWL13] Zheng Yang, Lirong Jian, Chenshu Wu, and Yunhao Liu. Beyond triangle inequality: Sifting noisy and outlier distance measurements for localization. ACM Transactions on Sensor Networks, 9(2):26:1–26:??, March 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Yen:2013:DLM

[YLL13] Li-Hsing Yen, Che-Ming Lin, and Victor C. M. Leung. Distributed lifetime-maximized target coverage game. ACM Transactions on Sensor Networks, 9 (4):46:1–46:??, July 2013. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Yang:2019:NAK

[YLSZ19] Zheng Yang, Junyu Lai, Yingbing Sun, and Jianying Zhou. A novel authenticated key agreement protocol with dynamic credential for WSNs. ACM Transactions on Sensor Networks, 15 (2):22:1–22:??, April 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

                                        URL https://dl.acm.org/ft_
                                        gateway.cfm?id=3303704.

Yang:2014:DOL

[YM14] Shusen Yang and Julie A. Mccann. Distributed optimal lexicographic max-min rate allocation in solar-powered wireless sensor networks. ACM Transactions on Sensor Networks, 11(1): 9:1–9:??, August 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Yin:2023:HIG

[YMY^{+}23] Xiaoyan Yin, Xiaoqian Mi, Sijia Yu, Yanjiao Chen, and Baochun Li. Harmony or involution: Game inspiring ageof-information optimization for edge data gathering in Internet of things. ACM Transactions on Sensor Networks, 19 (2):46:1–46:??, May 2023. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3565022.

Yuan:2013:STA

[YPW^{+}13] Yi Yuan, Dawei Pan, Dan Wang, Xiaohua Xu, Yu Peng, Xiyuan Peng, and Peng-Jun Wan. A study towards applying thermal inertia for energy conservation in rooms. ACM Transactions on Sensor Networks, 10(1):7:1– 7:??, November 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Yang:2017:VSS

[YPZ^{+}17] Zhicheng Yang, Parth H.

Pathak, Yunze Zeng, Xixi Liran, and Prasant Mohapatra. Vital sign and sleep monitoring using millimeter wave. ACM Transactions on Sensor Networks, 13(2): 14:1–14:??, June 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Yan:2022:PPP

[YQLD22] Zheng Yan, Xinren Qian, Shushu Liu, and Robert Deng. Privacy protection in 5G positioning and location-based services based on SGX. ACM Transactions on Sensor Networks, 18(3): 41:1–41:??, August 2022. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3512892.

Yoon:2017:FBC

[YRB^{+}17] Hee Jung Yoon, Ho-Kyeong Ra, Can Basaran, Sang Hyuk Son, Taejoon Park, and Jeonggil Ko. Fuzzy bin-based classification for detecting children’s presence with 3D depth cameras. ACM Transactions on Sensor Networks, 13(3):21:1–21:??, September 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

You:2024:PCP

[YRM^{+}24] Wei You, Meixuan Ren, Yuzhuo Ma, Die Wu, Jilin Yang, Xuxun Liu, and Tang Liu. Practical charger placement scheme for wireless rechargeable sensor networks with obstacles. ACM Transactions on Sensor Networks, 20(1):11:1–11:??, January 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3614431.

Yoon:2007:CAC

[YS07] Sunhee Yoon and Cyrus Shahabi. The Clustered AGgregation (CAG) technique leveraging spatial and temporal correlations in wireless sensor networks. ACM Transactions on Sensor Networks, 3(1):??, March 2007. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Yang:2015:PBD

[YSK^{+}15] Yong Yang, Lu Su, Mohammad Khan, Michael Lemay, Tarek Abdelzaher, and Jiawei Han. Power-based diagnosis of node silence in remote high-end sensing systems. ACM Transactions on Sensor Networks, 11(2):33:1– 33:??, February 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Yap:2008:MWA

[YSM08] Kok-Kiong Yap, Vikram Srinivasan, and Mehul Motani. MAX: Wide area human-centric search of the physical world. ACM Transactions on Sensor Networks, 4(4):26:1–26:??, August 2008. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Yang:2013:TSS

[YSZC13] Yi Yang, Min Shao, Sencun Zhu, and Guohong Cao. Towards statistically strong source anonymity for sensor networks. ACM Transactions on Sensor Networks, 9(3):34:1–34:??, May 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Yu:2014:CCW

[YTB^{+}14] Zuoming Yu, Jin Teng, Xiaole Bai, Dong Xuan, and Weijia Jia. Connected coverage in wireless networks with directional antennas. ACM Transactions on Sensor Networks, 10(3):51:1– 51:??, April 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Yang:2022:ASE

[YTR^{+}22] Fan Yang, Ashok Samraj Thangarajan, Gowri Sankar Ramachandran, Wouter Joosen, and Danny Hughes. AsTAR: Sustainable energy harvesting for the Internet of Things through adaptive task scheduling. ACM Transactions on Sensor Networks, 18(1):4:1– 4:34, February 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3467894.

Yuan:2023:MOS

[YTZ^{+}23] Xiaoming Yuan, Hansen Tian, Zedan Zhang, Zheyu Zhao, Lei Liu, Arun Kumar Sangaiah, and Keping Yu. A MEC offloading strategy based on improved DQN and simulated annealing for Internet of behavior. ACM Transactions on Sensor Networks, 19(2):28:1–28:20, May 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3532093.

Yoon:2007:TST

[YVS07] Suyoung Yoon, Chanchai Veerarittiphan, and Mihail L. Sichitiu. Tiny-sync: Tight time synchronization for wireless sensor networks. ACM Transactions on Sensor Networks, 3(2):8:1– 8:??, June 2007. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Yang:2021:MMN

[YWD^{+}21] Panlong Yang, Tao Wu, Haipeng Dai, Xunpeng Rao, Xiaoyu Wang, Peng-Jun Wan, and Xin He. MORE: Multi-node mobile charging scheduling for deadline constraints. ACM Transactions on Sensor Networks, 17 (1):7:1–7:21, January 2021. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3410454.

Yin:2017:THM

[YXFL17] Yafeng Yin, Lei Xie, Yuanyuan Fan, and Sanglu Lu. Tracking human motions in photographing: a context-aware energysaving scheme for smart phones. ACM Transactions on Sensor Networks, 13(4):29:1–29:??, December 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Yin:2019:ABC

[YXG^{+}19] Yafeng Yin, Lei Xie, Tao Gu, Yijia Lu, and Sanglu Lu. AirContour: Building contour-based model for in-air writing gesture recognition. ACM Transactions on Sensor Networks, 15(4): 44:1–44:??, October 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3343855.

Yin:2019:SBM

[YYC^{+}19] Junjie Yin, Zheng Yang, Hao Cao, Tongtong Liu, Zimu Zhou, and Chenshu Wu. A survey on Bluetooth 5.0 and Mesh: New milestones of IoT. ACM Transactions on Sensor Networks, 15(3):28:1–28:??, August 2019. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3317687.

Yin:2023:TTF

[YYL^{+}23] Junjie Yin, Zheng Yang, Sicong Liao, Chunhui Duan, Xuan Ding, and Li Zhang. Tag-Focus: Towards fine-grained multi-object identification in RFID-based systems with visual aids. ACM Transactions on Sensor Networks, 19(1):9:1– 9:22, February 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3526193.

Yau:2010:QMS

[YYM^{+}10] David K. Y. Yau, Nung Kwan Yip, Chris Y. T. Ma, Nageswara S. V. Rao, and Mallikarjun Shankar. Quality of monitoring of stochastic events by periodic and proportional-share scheduling of sensor coverage. ACM Transactions on Sensor Networks, 7(2):18:1–18:??, August 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Yin:2008:ARU

[YYSL08] Jie Yin, Qiang Yang, Dou Shen, and Ze-Nian Li. Activity recognition via user-trace segmentation. ACM Transactions on Sensor Networks, 4(4):19:1–19:??, August 2008. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Yan:2022:OBU

[YYXL22] Yubo Yan, Panlong Yang, Jie Xiong, and Xiang-Yang Li. OpenCarrier: Breaking the user limit for uplink MU-MIMO transmissions with coordinated APs. ACM Transactions on Sensor Networks, 18(2):19:1– 19:21, May 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3488382.

Yang:2023:AMA

[YZZD23] Kang Yang, Xi Zhao, Jianhua Zou, and Wan Du. ATPP: a mobile app prediction system based on deep marked temporal point processes. ACM Transactions on Sensor Networks, 19(3):71:1–71:??, August 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3582555.

Zheng:2007:LUB

[ZBA07] Yunhui Zheng, David J. Brady, and Pankaj K. Agarwal. Localization using boundary sensors: an analysis based on graph theory. ACM Transactions on Sensor Networks, 3(4):21:1–21:??, October 2007. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Zhang:2014:AIP

[ZCLJ14] Hongwei Zhang, Xin Che, Xiaohui Liu, and Xi Ju. Adaptive instantiation of the protocol interference model in wireless networked sensing and control. ACM Transactions on Sensor Networks, 10(2):28:1–28:??, January 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Zhang:2023:PLM

[ZCZ^{+}23] Guidong Zhang, Guoxuan Chi, Yi Zhang, Xuan Ding, and Zheng Yang. Push the limit of millimeter-wave radar localization. ACM Transactions on Sensor Networks, 19(3):59:1– 59:??, August 2023. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3570505.

Zhang:2022:TDT

[ZCZL22] Qingyang Zhang, Jie Cui, Hong Zhong, and Lu Liu. Toward data transmission security based on proxy broadcast re-encryption in edge collaboration. ACM Transactions on Sensor Networks, 18(3):48:1–48:??, August 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3529510.

Zhou:2009:VRC

[ZDG09] Zongheng Zhou, Samir R. Das, and Himanshu Gupta. Variable radii connected sensor cover in sensor networks. ACM Transactions on Sensor Networks, 5(1): 8:1–8:??, February 2009. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Zhao:2021:LLD

[ZDS^{+}21] Guangrong Zhao, Bowen Du, Yiran Shen, Zhenyu Lao, Lizhen Cui, and Hongkai Wen. LeaD: Learn to decode vibration-based communication for intelligent Internet of Things. ACM Transactions on Sensor Networks, 17 (3):26:1–26:25, June 2021. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3440250.

Zhu:2010:FTR

[ZDW^{+}10] Mengxia Zhu, Song Ding, Qishi Wu, R. R. Brooks, N. S. V. Rao, and S. S. Iyengar. Fusion of threshold rules for target detection in wireless sensor networks. ACM Transactions on Sensor Networks, 6(2):18:1– 18:??, February 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Zhang:2023:TOE

[ZGCL23] Jin Zhang, Hong Gao, Quan Chen, and Jianzhong Li. Taskoriented energy scheduling in wireless rechargeable sensor networks. ACM Transactions on Sensor Networks, 19(4):88:1– 88:32, November 2023. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3594874.

Zhu:2021:DBA

[ZGH^{+}21] Yi Zhu, Abhishek Gupta, Shaohan Hu, Weida Zhong, Lu Su, and Chunming Qiao. Driver behavior-aware parking availability crowdsensing system using truth discovery. ACM Transactions on Sensor Networks, 17 (4):41:1–41:26, July 2021. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3460200.

Zhu:2012:ALT

[ZGHZ12] Ting Zhu, Yu Gu, Tian He, and Zhi-Li Zhang. Achieving longterm operation with a capacitordriven energy storage and sharing network. ACM Transactions on Sensor Networks, 8(4):32:1– 32:??, September 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Zhang:2022:LQE

[ZGJ^{+}22] Jia Zhang, Xiuzhen Guo, Haotian Jiang, Xiaolong Zheng, and Yuan He. Link quality estimation of cross-technology communication: The case with physicallevel emulation. ACM Transactions on Sensor Networks, 18 (1):14:1–14:20, February 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3482527.

Zhu:2011:SNL

[ZGT11] Yuanchen Zhu, Steven J. Gortler, and Dylan Thurston. Sensor network localization using sensor perturbation. ACM Transactions on Sensor Networks, 7(4):36:1–36:??, February 2011. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Zhao:2016:CCA

[ZGX^{+}16] Yawei Zhao, Deke Guo, Jia Xu, Pin Lv, Tao Chen, and Jianping Yin. CATS: Cooperative allocation of tasks and scheduling of sampling intervals for maximizing data sharing in WSNs. ACM Transactions on Sensor Networks, 12(4):29:1– 29:??, November 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Zhang:2005:UBL

[ZH05] Honghai Zhang and Jennifer C.

Hou. On the upper bound of α-lifetime for large sensor networks. ACM Transactions on Sensor Networks, 1(2):272–300, November 2005. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Zhao:2005:I

[Zha05] Feng Zhao. Introduction. ACM Transactions on Sensor Networks, 1(1):1–2, August 2005. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Zarepour:2017:SSE

[ZHCA17] Eisa Zarepour, Mahbub Hassan, Chun Tung Chou, and Adesoji A. Adesina. SEMON: Sensorless event monitoring in self-powered wireless nanosensor networks. ACM Transactions on Sensor Networks, 13(2):15:1– 15:??, June 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Zhu:2020:RAH

[ZHJ^{+}20] Shaoyi Zhu, Weiqing Huang, Chenggang Jia, Siye Wang, Bowen Li, and Yanfang Zhang. RF-AMOC: Human-related RFID tag movement identification in access management of carries. ACM Transactions on Sensor Networks, 16(4):33:1– 33:23, October 2020. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3399678.

Zhou:2006:MSR

[ZHKS06] Gang Zhou, Tian He, Sudha Krishnamurthy, and John A. Stankovic. Models and solutions for radio irregularity in wireless sensor networks. ACM Transactions on Sensor Networks, 2 (2):221–262, May 2006. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Zhang:2015:GND

[ZHL^{+}15] Desheng Zhang, Tian He, Yunhuai Liu, Yu Gu, Fan Ye, Raghu K. Ganti, and Hui Lei. Generic neighbor discovery accelerations in mobile applications. ACM Transactions on Sensor Networks, 11(4):63:1– 63:??, December 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Zhu:2023:DMN

[ZHT^{+}23] Xiaojun Zhu, Zhouqing Han, Shaojie Tang, Lijie Xu, and Chao Dong. Deploying the minimum number of rechargeable UAVs for a quarantine barrier. ACM Transactions on Sensor Networks, 19(2):40:1–40:??, May 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3561303.

Zhang:2016:CSL

[ZHZ^{+}16] Desheng Zhang, Tian He, Fan Zhang, Mingming Lu, Yunhuai Liu, Haengju Lee, and Sang H. Son. Carpooling service for large-scale taxicab networks. ACM Transactions on Sensor Networks, 12(3):18:1– 18:??, August 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Zhang:2010:RTD

[ZJX10] Jun Zhang, Xiaohua Jia, and Guoliang Xing. Real-time data aggregation in contention-based wireless sensor networks. ACM Transactions on Sensor Networks, 7(1):2:1–2:??, August 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Zhang:2012:ACI

[ZJZ12] Jun Zhang, Xiaohua Jia, and Yuan Zhou. Analysis of capacity improvement by directional antennas in wireless sensor networks. ACM Transactions on Sensor Networks, 9(1):3:1–3:??, November 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Zhang:2024:UCU

[ZJZ^{+}24a] Guoming Zhang, Xiaoyu Ji, Xinyan Zhou, Donglian Qi, and Wenyuan Xu. Ultrasound communication using the nonlinearity effect of microphone circuits in smart devices. ACM Transactions on Sensor Networks, 20 (3):53:1–53:??, May 2024. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3631120.

Zhao:2024:FBR

[ZJZ24b] Ping Zhao, Jin Jiang, and Guanglin Zhang. FedSuper: a Byzantine-robust federated learning under supervision. ACM Transactions on Sensor Networks, 20(2):36:1– 36:??, March 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3630099.

Zamalloa:2007:AUA

[ZK07] Marco Zúñiga Zamalloa and Bhaskar Krishnamachari. An analysis of unreliability and asymmetry in low-power wireless links. ACM Transactions on Sensor Networks, 3(2):7:1– 7:??, June 2007. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Zhang:2010:DMM

[ZKS10] Zhiguo Zhang, Ajay D. Kshemkalyani, and Sol M. Shatz. Dynamic multiroot, multiquery processing based on data sharing in sensor networks. ACM Transactions on Sensor Networks, 6(3): 25:1–25:??, June 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Zhang:2023:TNA

[ZLB^{+}23] Gaofeng Zhang, Yu Li, Xudan Bao, Chinmay Chakarborty, Joel J. P. C. Rodrigues, Liping Zheng, Xuyun Zhang, Lianyong Qi, and Mohammad R. Khosravi. TSDroid: a novel Android malware detection framework based on temporal & spatial metrics in IoMT. ACM Transactions on Sensor Networks, 19(3):51:1–51:??, August 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3532091.

Zhang:2024:DRL

[ZLD^{+}24] Shun Zhang, Pengfei Lan, Benfei Duan, Zhili Chen, Hong Zhong, and Neal N. Xiong. DPIVE: a regionalized location obfuscation scheme with personalized privacy levels. ACM Transactions on Sensor Networks, 20(2): 35:1–35:??, March 2024. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3572029.

Zhang:2010:RPA

[ZLGG10] Lei Zhang, Ligang Liu, Craig Gotsman, and Steven J. Gortler. An as-rigid-as-possible approach to sensor network localization. ACM Transactions on Sensor Networks, 6(4):35:1–35:??, July 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Zhu:2019:BSB

[ZLGL19] Tongxin Zhu, Jianzhong Li, Hong Gao, and Yingshu Li. Broadcast scheduling in batteryfree wireless sensor networks. ACM Transactions on Sensor Networks, 15(4):49:1–49:??, October 2019. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=
3356472.

Zhu:2020:LED

[ZLGL20] Tongxin Zhu, Jianzhong Li, Hong Gao, and Yingshu Li. Latency-efficient data collection scheduling in battery-free wireless sensor networks. ACM Transactions on Sensor Networks, 16(3):25:1–25:21, August 2020. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/abs/10.1145/3390956.

Zhou:2022:CSB

[ZLL^{+}22] Siwang Zhou, Yi Lian, Daibo Liu, Hongbo Jiang, Yonghe Liu, and Keqin Li. Compressive sensing based distributed data storage for mobile crowdsensing. ACM Transactions on Sensor Networks, 18(2):25:1–25:21, May 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3498321.

Zhang:2015:ARF

[ZLW^{+}15] Shigeng Zhang, Xuan Liu, Jianxin Wang, Jiannong Cao, and Geyong Min. Accurate range-free localization for anisotropic wireless sensor networks. ACM Transactions on Sensor Networks, 11(3):51:1– 51:??, May 2015. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Zheng:2024:PDT

[ZLW^{+}24] Xiaolong Zheng, Ruinan Li, Yuting Wang, Liang Liu, and Huadong Ma. PolarScheduler: Dynamic transmission control for floating LoRa networks. ACM Transactions on Sensor Networks, 20(3):67:1–67:??, May 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3652856.

Zhang:2024:CMD

[ZLX^{+}24] Yuncan Zhang, Weifa Liang, Wenzheng Xu, Zichuan Xu, and Xiaohua Jia. Cost minimization of digital twin placements in mobile edge computing. ACM Transactions on Sensor Networks, 20(3):74:1–74:??, May 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3658449.

Zhang:2019:WEM

[ZLYW19] Qian Zhang, Fan Li, Song Yang, and Yu Wang. W3W: Energy management of hybrid energy supplied sensors for Internet of Things. ACM Transactions on Sensor Networks, 15(1): 10:1–10:??, February 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3280964.

Zhang:2021:EED

[ZLZ21] Yufan Zhang, Ertao Li, and Yi-Hua Zhu. Energy-efficient dual-codebook-based backscatter communications for wireless powered networks. ACM Transactions on Sensor Networks, 17 (1):9:1–9:20, January 2021. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3426885.

Zordan:2014:PLC

[ZMVR14] Davide Zordan, Borja Martinez, Ignasi Vilajosana, and Michele Rossi. On the performance of lossy compression schemes for energy constrained sensor networking. ACM Transactions on Sensor Networks, 11(1):15:1– 15:??, August 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Zhang:2024:EFG

[ZMXM24] Xiao Zhang, James Mariani, Li Xiao, and Matt W. Mutka. Exploiting fine-grained dimming with improved LiFi throughput. ACM Transactions on Sensor Networks, 20(3):60:1–60:??, May 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3643814.

Zhang:2024:INE

[ZPL^{+}24] Wen Zhang, Chen Pan, Tao Liu, Jeff (Jun) Zhang, Mehdi Sookhak, and Mimi Xie. Intelligent networking for energy harvesting powered IoT systems. ACM Transactions on Sensor Networks, 20(2):45:1– 45:??, March 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3638765.

Zhu:2009:SSF

[ZSG09] Xianjin Zhu, Rik Sarkar, and Jie Gao. Segmenting a sensor field: Algorithms and applications in network design. ACM Transactions on Sensor Networks, 5(2): 12:1–12:??, March 2009. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Zhu:2006:LES

[ZSJ06] Sencun Zhu, Sanjeev Setia, and Sushil Jajodia. LEAP+: Efficient security mechanisms for large-scale distributed sensor networks. ACM Transactions on Sensor Networks, 2(4):500–528, November 2006. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Zhu:2007:IHH

[ZSJN07] Sencun Zhu, Sanjeev Setia, Sushil Jajodia, and Peng Ning. Interleaved hop-by-hop authentication against false data injection attacks in sensor networks. ACM Transactions on Sensor Networks, 3(3):14:1–14:??, August 2007. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Zamalloa:2008:EGR

[ZSKH08] Marco Zúñiga Zamalloa, Karim Seada, Bhaskar Krishnamachari, and Ahmed Helmy. Efficient geographic routing over lossy links in wireless sensor networks. ACM Transactions on Sensor Networks, 4(3):12:1–12:??, May 2008. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Zhang:2023:NLO

[ZSLL23] Zengqi Zhang, Sheng Sun, Min Liu, and Zhongcheng Li. Network lifetime optimization in multi-hop industrial cognitive radio sensor networks. ACM Transactions on Sensor Networks, 19(1):20:1–20:22, February 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3549938.

Zhao:2020:DPS

[ZSZ20] Ping Zhao, Jiaxin Sun, and Guanglin Zhang. DAML: Practical secure protocol for data aggregation based on machine learning. ACM Transactions on Sensor Networks, 16(4):34:1– 34:18, October 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3404192.

Zhang:2023:FGC

[ZTZX23] Xinglin Zhang, Jiaqi Tian, Junna Zhang, and Chaocan Xiang. Fine-grained caching and resource scheduling for adaptive bitrate videos in edge networks. ACM Transactions on Sensor Networks, 19(4):95:1– 95:30, November 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3604555.

Zheng:2010:ODD

[ZVPS10] Rong Zheng, Khuong Vu, Amit Pendharkar, and Gangbing Song. Obstacle discovery in distributed actuator and sensor networks. ACM Transactions on Sensor Networks, 7(3):22:1– 22:??, September 2010. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Zhang:2005:ODS

[ZW05] Xin Zhang and Stephen B.

Wicker. On the optimal distribution of sensors in a random field. ACM Transactions on Sensor Networks, 1(2):301– 306, November 2005. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Zhang:2024:FTA

[ZW24] Jiarui Zhang and Jiliang Wang.

FusionTrack: Towards accurate device-free acoustic motion tracking with signal fusion. ACM Transactions on Sensor Networks, 20(3):71:1–71:??, May 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3654666.

Zhou:2024:SAD

[ZWG24] Zhipeng Zhou, Feng Wang, and Wei Gong. i-Sample: Augment domain adversarial adaptation models for WiFi-based HAR. ACM Transactions on Sensor Networks, 20(2):38:1– 38:??, March 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3616494.

Zhang:2024:AHF

[ZWL^{+}24] Jinghui Zhang, Jiawei Wang, Yaning Li, Fa Xin, Fang Dong, Junzhou Luo, and Zhihua Wu. Addressing heterogeneity in federated learning with client selection via submodular optimization. ACM Transactions on Sensor Networks, 20(2):48:1– 48:??, March 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3638052.

Zhang:2023:GPM

[ZWW^{+}23] Yu Zhang, Qinhan Wei, Yongcai Wang, Haodi Ping, and Deying Li. GPART: Partitioning maximal redundant rigid and maximal global rigid components in generic distance graphs. ACM Transactions on Sensor Networks, 19(4):86:1– 86:26, November 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3594668.

Zhang:2023:BBF

[ZWWL23] Xiangjun Zhang, Weiguo Wu, Jinyu Wang, and Song Liu. BiLSTM-based federated learning computation offloading and resource allocation algorithm in MEC. ACM Transactions on Sensor Networks, 19(3):68:1– 68:??, August 2023. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3579824.

Zheng:2020:UMM

[ZWWZ20] Zimu Zheng, Feng Wang, Dan Wang, and Liang Zhang. An urban mobility model with buildings involved: Bridging theory to practice. ACM Transactions on Sensor Networks, 16 (1):10:1–10:24, February 2020. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3366689.

Zhou:2021:DSS

[ZWY21] Pengzhan Zhou, Cong Wang, and Yuanyuan Yang. Design of self-sustainable wireless sensor networks with energy harvesting and wireless charging. ACM Transactions on Sensor Networks, 17(4):45:1–45:38, July 2021. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3459081.

Zhang:2023:WRT

[ZYC^{+}23] Jian Zhang, Wu Yuan, Yanjiao Chen, Mingxi Li, Junkongshuai Wang, and Qian Zhang. WIB: Real-time, non-intrusive blood pressure detection using smartphones. ACM Transactions on Sensor Networks, 19 (4):87:1–87:27, November 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3595182.

Zhang:2024:WTS

[ZYL^{+}24] Qian Zhang, Zheng Yang, Fan Li, Biaokai Zhu, and Pengpeng Chen. WVC: Towards secure device paring for mobile augmented reality. ACM Transactions on Sensor Networks, 20(1): 4:1–4:??, January 2024. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3600233.

Zhang:2019:DEF

[ZYZ^{+}19] Qingquan Zhang, Yao Yao, Ting Zhu, Ziqiao Zhou, Wei Xu, Ping Yi, and Sheng Xiao. Dynamic enhanced field division: an advanced localizing and tracking middleware. ACM Transactions on Sensor Networks, 15(1): 2:1–2:??, February 2019. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

                                        URL https://dl.acm.org/ft_
                                        gateway.cfm?id=3216721.

Zhang:2021:PLB

[ZZ21] Yifan Zhang and Xinglin Zhang. Price learning-based incentive mechanism for mobile crowd sensing. ACM Transactions on Sensor Networks, 17(2):17:1– 17:24, June 2021. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3447622.

Zhang:2023:IMT

[ZZ23] Yifan Zhang and Xinglin Zhang.

Incentive mechanism with task bundling for mobile crowd sensing. ACM Transactions on Sensor Networks, 19(3):70:1– 70:??, August 2023. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3581788.

Zhou:2023:IIM

[ZZC^{+}23] Wangqiu Zhou, Hao Zhou, Xiang Cui, Fengyu Zhou, Haisheng Tan, and Xiang-Yang Li. IMeP: Impedance matching enhanced power-delivered-to-load optimization for magnetic MIMO wireless power transfer system. ACM Transactions on Sensor Networks, 19(4):73:1– 73:25, November 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3582693.

Zhou:2024:MSD

[ZZG^{+}24] Zhiyuan Zhou, Xiaolei Zhou, Baoshen Guo, Shuai Wang, and Tian He. Multi-sensor data-driven route prediction in instant delivery with a 3conversion network. ACM Transactions on Sensor Networks, 20(2):50:1–50:??, March 2024. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3639405.

Zhang:2023:KEU

[ZZH^{+}23] Yuexin Zhang, Fengjuan Zhou, Xinyi Huang, Li Xu, and Ayong Ye. Key extraction using ambient sounds for smart devices. ACM Transactions on Sensor Networks, 19(1):16:1– 16:20, February 2023. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3544108.

Zhang:2022:IVI

[ZZM^{+}22] Xu Zhang, Yangchao Zhao, Geyong Min, Wang Miao, Haojun Huang, and Zhan Ma. Intelligent video ingestion for realtime traffic monitoring. ACM Transactions on Sensor Networks, 18(3):47:1–47:??, August 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3529511.

Zhang:2023:MOE

[ZZPW23] Xinglin Zhang, Jinyi Zhang, Chaoqun Peng, and Xiumin Wang. Multimodal optimization of edge server placement considering system response time. ACM Transactions on Sensor Networks, 19(1):13:1– 13:20, February 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3534649.

Zeng:2023:EBM

[ZZW^{+}23a] Yiming Zeng, Pengzhan Zhou, Cong Wang, Ji Liu, and Yuanyuan Yang. Economical behavior modeling and analyses for data collection in edge Internet of Things networks. ACM Transactions on Sensor Networks, 19(2):33:1–33:??, May 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3532092.

Zhou:2023:CAM

[ZZW^{+}23b] Wangqiu Zhou, Hao Zhou, Zhan Wang, Haisheng Tan, and Xiang-Yang Li. Context-aware magnetic MIMO wireless charging with parallel in-band communication. ACM Transactions on Sensor Networks, 19(4):81:1– 81:24, November 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3582692.

Zhao:2020:UST

[ZZX^{+}20] Yi Zhao, Zimu Zhou, Wang Xu, Tongtong Liu, and Zheng Yang. Urban scale trade area characterization for commercial districts with cellular footprints. ACM Transactions on Sensor Networks, 16(4):42:1– 42:20, October 2020. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

https://dl.acm.org/doi/10.
1145/3412372.

Zhang:2023:MCL

[ZZY^{+}23] Lan Zhang, Daren Zheng, Mu Yuan, Feng Han, Zhengtao Wu, Mengjing Liu, and Xiang-Yang Li. MultiSense: Crosslabelling and learning human activities using multimodal sensing data. ACM Transactions on Sensor Networks, 19(3):65:1– 65:??, August 2023. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3578267.

Zhang:2020:CAD

[ZZZ^{+}20] Jianhui Zhang, Siwen Zheng, Tianhao Zhang, Mengmeng Wang, and Zhi Li. Chargeaware duty cycling methods for wireless systems under energy harvesting heterogeneity. ACM Transactions on Sensor Networks, 16(2):15:1– 15:23, April 2020. CO-DEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3372800.

Zhang:2022:GTU

[ZZZ^{+}22] Yi Zhang, Yue Zheng, Guidong Zhang, Kun Qian, Chen Qian, and Zheng Yang. GaitSense: Towards ubiquitous gait-based human identification with Wi-Fi. ACM Transactions on Sensor Networks, 18(1):1:1– 1:24, February 2022. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3466638.